A Selected Bibliography of Publications by, and about, Edsger W. Dijkstra

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

25 August 2025 Version 1.44

Title word cross-reference

#1 [Dij96x].

(∃y: : (∀x: : p.x.y)) ⇒ (∀x(∃y: : p.x.y)) [Dij86b]. (n + 1) [Dij95c]. 2 [Dij80b, Dij99a, FCTB08]. 3 [Dij92f, LCZW14, VPM^{+}10]. 3 × 7 [Dij89-37]. 6 [Dij80b]. 9 [Dij83t]. [x] ≡ [ x] [Dij95m]. ↓ [Dij99p]. ∃∀ ⇒ ∀∃ [Dij95-28]. g1 [Dij82-30]. k [Dij99j]. L [MS14]. N [Dij98g, Dij95c, Dijxx-29]. N >> 0 [Dij98g]. sin.(α + β) [Dij96s, Dij96t]. ↑ [Dij99p]. x: [x] [Dij95f]. |x[n]| = x[n − 1] + x[n + 1] [Dij83a, Dij83t].

-coloured [Dij80b]. -colouring [Dij99a]. -concave [MS14]. -dimensional [Dij95c]. -function [Dij82-30]. -graph [Dij95c]. -graphs [Dij80b]. -Multiprogramming [Dij21c, Dij68p, Dij83v, Dij68q]. -Rekentechniek [Dij54a]. -th [Dij99j]. -valued [Dijxx-29].

0 [Dij85-51, Mil89]. 0-201-17536-3 [Mil89].

1 [Dij61a, Dij82o, Har99]. 10 [Dij81-29, MGH^{+}05]. 10.1.1986 [Dij86-44]. 10/6 [Dij79-30]. 10/8/81 [Dij81-29]. 11e [Dij70b]. 11th [Dij70b]. 1308 [Dij02d]. 13th [Dij73w]. 14 [Dij72n]. 141 [Dij96-36]. 15 [Dij85h]. 1609.4 [PS14]. 16th [Dij75-52, Dij82-106, Dij82-49]. 1862 [Dij80e]. 18th [Dij73v]. 1950s [Bac80, Day11b]. 1960s [Day11b]. 1962 [Dij62c]. 1965 [Dij65j]. 1966 [Dij66d]. 1967 [Dij67e, Dij67g]. 1968 [Dij68r, NR69]. 1969 [Dij69j, Dijxx-96]. 1971 [Dij71f, Dij71e]. 1972 [Dij72b, Dij72n, Dij72q, Dij72r]. 1973 [Dij73k, Dij73t, Dij73x, Dij82c, Dij82b, Dij82-54]. 1974 [Dij74g, Dij74u, Dij74x, Dij74w, Dij74t, Dij74v, Dij82-70, Dij82-109, Dij82-51]. 1975 [Dij75b, Dij75q, Dij75o, Dij75r, Dij75t, Dij75-34, Dij75-59, Dij75-50, Dij75-58, Dij75-54, Dij75-55, Dij75-57, Dij75-56, Dij75-62, Dij75-64, Dij75-65, Dij75-67, Dij82-110, Dij82-60]. 1976 [Dij75s, Dij76e, Dij76d, Dij76l, Dij76k, Dij76j, Dij76q, Dij76n, Dij76-31, Dij76-36, Dij76-81, Dij76-67, Dij76-79, Dij76-80, Dij76-73, Dij76-78, Dij76-75, Dij76-77, Dij76-71, Dij76-74, Dij76-76, Dij76-85, Dij76-87, Dij76-84, Dij76-86, Dij76-92, Dij76-93, Dij76-91, Dij77b, Dij82-112, Dij82-114, Dij82-113, Dij82-62, Dij82-63, Dij82-53, Maz76]. 1977 [Dij77a, Dij77e, Dij77f, Dij77-42, Dij77-39, Dij77-38, Dij77-37, Dij77-40, Dij77-41, Dij77-44, Dij77-45, Dij77-48, Dij77-49, Dij82-107, Dij82-111, Dij82-58, Dij82-61]. 1978 [Dij78-44, Dij78-49, Dij78-46, Dij78-47, Dij78-48, Dij78-45, Win78]. 1979 [Dij79-33, Dij79-30, Dij79-29, Dij79-31, Dij79-32]. 1980 [Dij80e, Dij80-47, Dij80-45, Dij80-46]. 1981 [Dij81-28, Dij81z, Dij81-27, Dij81-30, Dij81-31, Dij81-32, Dij81y, Dij81-33]. 1981/82 [Dij82o]. 1982 [Dij82-55, Dij82-52, Dij82-50, Dij82-56]. 1983 [Dij83x, Dij83-27, Dij83y, Dij83-28, Dij83z]. 1984 [Dij84u, Dij84t, Dij84w, Dij84y, Dij84z, Dij84s, Dij84x, Dij84v]. 1985 [Dij85-50, Dij85-49, Dij85-45, Dij85-46]. 1986 [Dij86-45]. 1987 [Dij87k, Dij87y]. 1988 [Dij88r, Dij88s]. 1989 [Dij89-38]. 1990 [TLR^{+}91]. 1994 [Bro96]. 1996 [Dij96i]. 1998 [Dij98d]. 1a [Dij57c]. 1st [Sam76].

2 [Kru03]. 2**N [Dij73g]. 2002 [Ano02b, Ano02c, Ano12]. 2004 [JSW04]. 2009 [Mos10]. 2012 [AMM^{+}12]. 2013 [Ano13]. 2018 [Sch18]. 20th [Dij73w]. 21 [Dij67e]. 21st [Dij73v, Dij73s]. 22 [Dij68r]. 233/AvG [Dij96-36]. 24 [Dij62c]. 25th [Dij74w, Dij74y]. 26/5 [Dij79-30]. 26/7 [Dij81-29]. 26th [Dij73s]. 27 [Dijxx-96]. 27th [BR70]. 28/29 [BD01b]. 28th [Dij74x]. 2de [Dij63d].

3 [Mil89]. 3-Layer [LWQ09]. 31st [BR70]. 360 [BMWC01]. 3de [Dij63e]. 3rd [Dij74x].

4.20 [Bar75]. 4th [IEE79b, IEE79a].

5 [Dij85-37]. 5th [Dij74w, Dij74y, Maz76, VU05].

6 [Dij67g]. 60 [Dij60b, Dij61c, Dij61a, Dij61b, Dij62a, Dij62e, Dij63a, Dij63f, Dij63i, Dij70b, Kru03, NBB^{+}60, RR64, Ran02, van19]. 60th [TLR^{+}91]. 68 [Dijxx-88, WTD^{+}69, vWMPK69].

7 [Dij73k]. ’79 [IEE79a]. 7th [Dij75-52, Dij82-106, Dij82-49, Win78].

8 [Dij72b, Dij74g]. 81 [Dij81-29]. 83g [Dij82o]. 8th [HW06].

= [Dij85b].

A* [DK13, YL11]. A. [DLP78, Dij78s, Dij80a, Dij86h, Dij86-39, DvG90a]. Aan [Dij66b, Dij67a, Dij68b, Dij76b, Dij76c, Dijxxa, Dijxxb, Dijxxd, Dijxxc, Dij65d, Dij65j, Dij65i, Dij66c, Dij67g, Dij69k, Dij71e, Dij72b, Dij72n, Dij73x, Dij80t, Dijxx-32, Dijxxt, Dijxx-76, Dijxx-77, Dijxx-97]. aanleiding [Dij77o, Dij80t]. aanschaffing [Dij64j]. Aanschaffingsbeleid [Dij68c]. aansluiting [Dij63m]. aanvang [Dijxxi]. Aarhus [Dij73k]. abolishment [Dij74k]. Abstract [WG01]. Abstraction [GHJV01, Dij73h, Dij76t]. Academic [Bar75, Dij94z, Dijxx-61, Dijxxp]. academisch [Dijxxp]. academische [Dijxx-61]. accelerating [ANGK09]. Acceptance [Dij74b, Dij82-70]. Access [Dij65d]. Acclaimed [Gol02]. accommodating [DFR74]. according [Boh08, Dij86-35, Dij89v]. ACE [Wil80b]. Achievements [Dij99b]. Achtste [Dij75b, Dij75c, Dij76e, Dij76d, Dij77a]. Acknowledgments [Ano80a]. ACM [Ash87, BSE^{+}90, Dij91n, Dijxxe, Gol02, WWSD90]. acquire [Dij64j]. Acta [Dij80-40, BB03]. action [ŁMB02]. Activity [Dij96-35, Dij65f, Dij65g]. Ada [Bis85]. adaptive [AAEG11]. addenda [Dij95g]. Addendum [Dij87a, SS04a]. Adding [BN94, SB04]. Addison [Mil89]. Addison-Wesley [Mil89]. Additional [Knu66a, Knu66b]. additive [Dij76-63]. Address [Dij86c, Dij86e, BDS82, Dij86d]. addressing [Bis85, Dijxx-37, Dij75-51, Dijxx-91]. Adds [Dij06]. adjusted [TZDM10]. administratie [Dijxx-62]. administration [Dij64b, Dij81a, Dijxx-65]. Adopted [Dij02e, Dij00b]. adressering [Dijxx-37]. Advanced [Big80, Bro96, Dij62f, Dijxx-83, VU05]. Advances [Fag01a]. Advent [Day11b]. Advice [Dij76b, Dij76c, Dij89-44]. Advies [Dij76b, Dij76c]. afdeling [Dij64j]. AFIPS [Dij74b, Dij82-70]. AFP [VU05]. afsluiting [Dij83w]. afstudeerder [Dijxx-91]. afstudeerders [Dij75-51]. afstudeerrichting [Dij72e, Dijxx-56]. After [Dij75d, Har99, Rei83b, Rei83a]. Afternoon [Dij78-42, Dij80-27, Dij80y, Dij82s, Dij85b]. against [Dij68d, Dijxx-51]. Agda [Swi11]. agency [Dij95p]. agent [BDBR01, OC10]. Ages [Dij77w, Dij78u]. agreement [vDO01]. AGV [LN14]. Aho [Dij76-40]. AHP [YL10]. AI [Dij84n]. aims [Dij64a, Dijxx-82]. airline [YJ10]. al [Tat23]. al-Khwarizmi [Tat23]. Alain [Dij78-27]. Albany [Dij79-30]. algebra [JH93, Tse95, Tse97]. Algebraic [BT79, Leh77, CRT78, Dij73f, DD98, GB96]. Algemeen [Dij56c, Dij57c].

ALGOL [Dij61a, Dijxx-88, DHPS59, Dij60b, Dij61a, Dij61b, Dij62a, Dij62e, Dij63b, Dij63f, Dij63i, Dij66f, Dij66e, Kru03, NBB^{+}60, RR64, vWMPK69, van19, Dij61c, Dij63a, Dij70b, Ran02, WTD^{+}69]. ALGOL-60 [Dij61a, Dij63f, Dij63i]. ALGOL-60-Übersetzer [Dij61a]. Algorithm [BMST03, BB99b, BB99a, CMMS98, CC12, DFvG83, DH01, FJ09, GLFM07, HCSW09, HZW09, Joh73, KLK08, KLMZ11, LX10, yLsLtZ10, LSMS12, LW11, LWQ09, MS01, MJC07, Nep96, ND00, NAGK09, Obr78, PMR11, PRC12, PPS^{+}13, SWW99, SWX10, Tat23, TS01, Uch14, UNTH14, WT79, WLXX09, XWZH12, ZJM12, AVL02, ANGK09, Anw03, Bar98, BDn^{+}10, BS99, BW00, CF04, Cha10, CH10, CSZ10, Dij76-59, DFvG82, DFvG85, Dij96e, Dijxxu, EKP96, EBR09, FS10, zFmLcWcJ10, FC14, GS97, HHM94, HKS93, ITOK04, JAE^{+}12, JHLC14, KYTY09, KGZ06, Knu77, Koh76, Kun80, KF08, Lee06, yLsLtZ09, LN14, LCL^{+}94, LCZW14, Mis90, MSS^{+}06, Mor90b, Muk12, MS14, NMM78, Nos85, NS00, OC10, PRM11, PRD14, PS14, PG13, Pet71, PRV09, SWW00, SZHT10, SS10, SxXq11, Sni06, SPPR92, SEEM01, Str72, SVSR04]. algorithm [SAID13, TZDM10, UOZ11, VPM^{+}10, WYY11, xXL10, XZY^{+}12, XLH^{+}07, YJ10, YL10, Yap83, ZWKT10, ZTZY12, ZLY13, Dij76s, Dij77n, Dij82-81, Dij93a, Dij98h, Dij98c, Dijxxo, GGAVGGRC15, Mis01, RPGS14, Du10]. algorithme [Dijxxu, Pet71]. Algorithmic [NBB^{+}60, Wel80, vWMPK69, Dij62e]. Algorithms [BMST03, DMS83, DvG82f, Dij82-83, GLFM07, JNK11, MCX13, MS01, WT79, CSZ10, Dij74h, DvG82g, DvG82d, DvG82e, DK13, ELCdCG14, FCTB08, FURD13, GL98, LvdS90, NA13, Nep00, Trä95, YL11]. alignment [vDO01]. allocation [Dij72d]. Allotted [Dij79-36, Dij79-37]. Alphern [Rem90]. alternate [Anw03]. Alternative [Dij82-43, Dij80v, Dij81b, Dij82, Dij81u, Dij81v, Dij82g, Dij82o, Dij82-42, Dij94a, Dij94w, Dij96a]. Alto [Dij87y]. ambacht [Dijxx-33]. amélioration [Bov86]. America [Bac80, Dij71h]. American [Dij94v, Dij80c, Dij81c, Dij95-31]. Amerika [Dij71h]. among [Dij74c, Dij75l, Dij91l]. Amsterdam [Dij81y, Sam76, Dij76-72, Dij82-108, Dij82-59]. an.’85 [Dij85-48]. analogy [Bro90a]. Analysis [DeM01, Har95, HCSW09, LWQ09, Dij73b, Dij81d, Dij01m, DK13, HKS93, JAE^{+}12, JM80, PS14, SAID13, YL10]. analytical [Dij83b]. Andrei [Dij89b]. Andrew [Dij96b]. Anecdotes [TLR^{+}91]. Angeles [Dij80-45]. Angelic [Hes95]. angelical [Dij86-32]. anger [Dij78-51]. angle [Dij86r, Dij97a]. angles [Dij94r]. Animadversions [Pad85]. Anniversary [Dij82-44, Dij74r]. annotated [Dij92l, Dij92w]. Announcing [AMM^{+}12]. anomaly [Dij90m]. anonymous [Dij85-44]. answer [Dij75e]. Answering [Dijxxh]. Answers [Dij00a]. anthropomorphism [Dij85z]. Antony [Dij77b, Dij99h]. Antwerp [Dij79-29]. Appalling [Dij98a]. apparatuur [Dijxxt]. Appendix [Dij63b]. Application [BBM07a, CC12, Nep96, WYY11, WLXX09, XWZH12, FS10, LSJT12, VPM^{+}10, XZY^{+}12, Pop95]. applications [Dij04, Nel92, PRV09, Dij90i]. applied [VPM^{+}10]. Appreciatie [Dijxxf]. Appreciation [Dijxxf].

Approach [Dij68h, Dij82-31, AAEG11, Bis85, Dij67b, Dij68g, Dij74i, Dij88d, Dij89g, Dij92u, ITOK04, LCL^{+}94, MHMO10]. April [Dij74u, Dij74x, Dij75-52, Dij78-44, Dij81-27, Dij82-106, Dij82-49, Dij84w, Dij84y, HW06, Dij66d, Dij66d, Dij95-29, Dij99m]. April/7th [Dij75-52, Dij82-49]. Apt [Dij91f]. Arbeitsblatt [DS76a]. Arbeitstagung [Dij65i]. arbitrary [Dij96-31]. Arbitration [DS80a, Dij80-35]. arc [KGZ06, Kun80]. Architectural [Bis85]. Architecture [Boa90, BDB93]. Archive [Ano06b]. arcs [Dij94r, Yap83]. Area [KP09, KP10]. Area-Efficient [KP09, KP10]. Aretz [Dij95s]. Argonne [Chu80]. Argument [DvG86c, DvG84a, DvG84b, DvG84c, DvG84d, DvG84e, DvG85b, Dij85-41, Dij85-42, Dij87n, Dij94c, Dij94q, Dij94b, Dij95a, Dij95z, Dij96c, Dij99o]. Arguments [Dij80-37, Dij78-39, Dij80z, Dij94u]. Aristotle [Dij77u, Dij82-95]. Arithmetic [D^{+}66a, Dij64d, Dij65c, Dij92a, Dij94c, Dij96c, Dij96d, Dij00f, Dij00g]. arithmetiek [Dij65c]. ARMAC [Dij56b, Dij56c, Dij56d, Dij56e, Dij56a, Dij57c, Dij57d]. ARRA [Dij53, Dij54a, Dij54b, Dij55c]. ARRA-subroutines [Dij54a]. Arrangement [Dij76p, Dij76o]. Array [Dij76f, Dijxxg, Dij85-34]. arrogance [Che16]. arrogant [DS72c]. Arsac [Dij77m]. ARSI [Dij75h]. Art [Dij83w, Pad85, Dij71g, Dij78v, Dijxx-54, Dijxx-79, PGB00]. Ascending [Dij88b, Dij76-52, Dij77-34, Dij93p]. aspect [SZHT10]. Assembly [Dij79e]. assertion [JM80]. assertional [Dij81e]. assertions [Gou90]. assignment [vT90]. Assisted [Dijxx-68, KKNI14]. Associates [Tom80]. Associative [Nep96, ND00, Dij92g, Nep00]. associativity [Dij86p]. Associons [DFR74, DFRxx]. asynchronous [Dij92u, Mar90]. ATAC [Dij85b, Dij86-35, Dij89c]. Atlantic [Dij77r, Dij82-91]. attempt [Dij62b, Dij72l]. Attribute [DS04]. attributed [Dij76u]. audience [Dij86-47]. Auer [Dij91k]. Aug [Dij81-32, Dij84v, Dij86-45]. August [Ano02b, Ano12, BS76, Bro96, Dij72q, Dij72r, Dij73t, Dij76-72, Dij77-37, Dij78-46, Dij80-45, Dij82b, Dij82-108, Dij82-54, Dij82-59, Sam76, VU05, Dij70d]. augustus [Dij72q, Dij72r]. aunt [Dij02f]. Austin [Dij78-45, Dij79-30, Dij79-31, Dij80-45, TLR^{+}91, Dij85b]. Australia [Dij77-38, Dij82-107, Dij82-58, Dij83x]. Austria [Dij85-47]. Ausübung [BS01]. automaat [Dijxxp]. Automata [Dij65i, Tur90b]. Automatentheorie [Dij65i]. Automatic [Dij59a, Dij59b, Flo77, Hus80, PPS^{+}13, BS01, DDvW57]. automation [Dij67a, LSJT12]. automatische [DDvW57]. automatischen [BS01]. Automatiseringswiskunde [Dij67a]. Automobiles [LSMS12]. Autorouting [SPPR92]. Autumn [Dij75b, Dij75q, Dij75o, Dij75r, Dij75t, Dij75-34, Dij75-59, Dij75-50, Dij75-58, Dij75-62, Dij75-64, Dij75-65, Dij75-67, Dij76d, Dij76j, Dij76n, Dij76-80, Dij76-84, Dij76-86, Dij76-91]. AvG [Dij96-36]. AvG16 [Dij82g]. AvG16/ [Dij82g]. AvG45 [Dij85-44]. AvG45/EWD901 [Dij85-44]. avoiding [Dijxx-48]. Award [Ash87, Dij78-37, Dij82-70, Ano13, Dij74b, Gol02]. Aware [SWX10]. Away [Coc02, Che16]. Axel [Dij77-28]. axial [DvG82b, DvG82c, Dij85o]. Axiomatic [Hoa01a, Dij73i, Dij73q, Dij73a, Dij74p]. axiomatization [Dij93p].

Babbage [Dij86-43]. Backus [Che16, Dij78-37]. Baden [Dij75-57]. Baffled [Dij86-34]. baffler [Dij87h, Dij85k, Dij85l, Dij85-51, Dij86l, Dij86m, Dij87i]. Bag [HB94]. bagatelle [Dij82h, Dij88c, Dij89d, Dij92v, Dij93a, Dij96e]. bags [Dij93j]. Bahamas [Dij78-44]. baked [Dijxx-40]. bakery [Lam22]. balance [Dij90b, Dij96f]. Band [Dij71f]. banker [Bro90a, Dij77n, Dij82-81, Dijxxo]. bankiersalgorithme [Dijxxo]. Banks [Cod01]. Barcelona [Dij83y]. bars [Dij92f]. base [Anw03, GA21]. Based [FJ09, HCSW09, HZW09, KP09, yLsLtZ10, LSMS12, LW11, LWQ09, PRC12, SWX10, Uch14, UNTH14, ZJM12, ZXZ^{+}09, AAEG11, BDB93, Du10, GS97, KP10, yLsLtZ09, LCL^{+}94, LCZW14, LvdS90, MTZ12, MJC07, NA13, PRM11, PRD14, TZDM10, WM14, xXL10, YJ10, YL10, ZTZY12, ZLY13]. basic [Dijxx-89, Dij69f]. basically [Dij77o]. Basis [Hoa01a, Dij73q, Dij73a, Dij74p, Gou90]. Bass [Dij75-44]. Batch [Bri01]. Batcher [Dij85k, Dij85l, Dij85-51, Dij86l, Dij86m, Dij87h, Dij87i]. Bauer [Dij86k, Dij71f, Dij72c, DS76a]. Be [Dij79z, Par01b, Dij79-40, DvG84h, DvG84i, PJ89]. beamforming [LCZW14]. Beantwoording [Dijxxh]. beast [Neu90]. beats [KYTY09]. Beautiful [Dij80-37, Dij82i, Dij82-93, Dij75g, Dij77t, Dij78-39, Dij91-33, DFxx]. Beautifying [Heh90]. Beauty [FvGGM90, Neu90]. been [Dij98e]. beginning [Dijxxi]. beginnings [Dij94-28]. behaviors [NMM78]. Behaviour [Dij79-27, Dij77-32, Kot90]. Behind [Alp18, Dij82-81, Mos10, Dij77n]. Being [IdKM06, Dij76-42, Dij82s, Dij88n, Dijxx-49]. Bekic [Dij74g]. Belated [Dij86f, Dij85c]. beledigend [Dij77o]. Bemerkung [Str72]. Ben [Dij91y]. Benders [Dij68f]. Bengt [Dij89-40]. berekeningen [Dij79-36, Dij79-37]. Berger [Dij88t, Dij99o]. Berlin [Bör94b, Dij77b]. beschrijving [Dij53, Dij56b, Dij65c]. best [Anw03]. betreffende [Dij72n]. Betrouwbaarheid [Dij73c]. better [BD89]. Between [Bau80, BDR85, Che16, Dij73g, Dij77s, Dij79b, Dij79c, Dij79v, Dij80l, Dijxx-41, Yap83, vDO01]. Beware [Dij96g]. Beyond [DM97]. Bezettingsadministratie [Dij64b]. bezighouden [Dij76-88]. Bezoek [Dij72b, Dij65d, Dij65j, Dij65i, Dij67g, Dij69k, Dij72n, Dij72q, Dij73x, Dijxx-76, Dijxx-77, Dijxx-97]. biased [Dij89m, Dij89w, Dij89x, Dij90n]. Bibliography [Ran80b]. bichrome [Dij89-37, Dij00h]. Bij [Dijxxi, Dijxx-92]. Bijdrage [Dij68c]. bijeenkomst [Dijxx-43]. Bill [Dij91h]. Billionth [Coc02]. Binary [Dij99c, Gri90, Dij79-34, Dij82-64, Dij87r, Dij88l, Dij95s, Dij01h]. Binding [Kot90]. binnen [Dij77c]. biobjective [SNR15]. Biographies [Wei88]. Bird [Dij90j, Dij94k]. Birgit [Dij95e]. birth [Dij88j, Dij88k, Dij88u].

Birthday [FvGGM90, TLR^{+}91]. birthdays [GA21]. bisectors [Dij97a]. bit [Dij98f, Dijxx-42]. bits [Dij92-31]. Black [Dij99k]. Blanchland [Dij73s]. Bletchley [Goo80]. blijven [Dijxx-99]. Block [MCX13, HHM94]. blocked [Dij56d]. Blokhuis [Dij86-39]. BLUE [Dij78z, Dij78c]. blunder [Dij88i]. Board [Dij95k]. boat [Dij96k]. Boldern [Dij74x]. bomen [Dij60c]. Bonn [BD01b]. Book [Bar75, Bar74, Bör94a, Bör94b, Flo77, Llo74, Mil89, Dij76g, Dij76-66, Dij78h, Dij78i, Dij79f, Dij82-39, Dij91w, Dij94t, Dij72c, Dij77b]. boolean [Bij90, Dij87b, Dij87c, DS90a, Dij90c, Dij94d]. Boston [Dij72r, Dij85-50, Dij72r]. bound [Dij91-28]. Bounded [Dij82-74, LSMS12, Bro90b, Dij72d, Dij76-42, Dij78l, Dij82n, Dijxx-49]. Boundedness [Dij82-99, Dij77-30]. Boute [Dij94w]. BP [Dij86-46]. Brahmagupta [Dij86-34]. brain [EBR09, VPM^{+}10]. Brasil [Dij83z]. breath [Che16]. brevity [Dij89o]. bridges [Dij00i]. Bruijn [Dijxx-75]. Bruijn-sequences [Dijxx-75]. buckets [GS97]. Budget [Dij87x]. buffer [Dij72i, Dij72j, Dij95q]. buffering [Dij65b]. buffers [Bro90b]. bug [Dij76g]. Bugs [Dij82-95, Dij77u]. build [Lam22]. Built [Dij82-86, Dij74j]. Bulterman [Dij92b, Dij96h]. bundle [Lee06]. Bureau [Hus80, Slu80]. Burroughs [Dij75-29]. Burstall [Dij76v, Dij82-76]. Business [FvGGM90]. Butterfly [Dij83b, Dij83m]. Buzzword [Par01a]. byte [Dij89s].

C [Bar75, Bar74, Bör94b, Dij88f, Dij89h, Dij90d, Llo74, Hoa90]. C-mos [Hoa90]. C. [Dij70d, Dij74a, Dij75-61, Dij77q, Dij82-89, Dij82s, Dij82-66, Dij86h, Dijxxz, Dijxx-29, Hoa82]. caching [Nel90]. Cactus [DvdS06]. CAD [ZXZ^{+}09]. calculating [BS01]. Calculation [DM97, Dij54a]. Calculational [Dij95b, Dij93k, Dij94i, Dij95j, Dij98b, DM00, DM01]. Calculations [vdW90, Dij79-36, Dij79-37]. calculators [DDvW57]. calculi [Dij92h, Dij92-29, Dij95n]. Calculus [Bör94a, BN94, Dij76-29, DS90g, Nel89, BBM07a, BBM07c, Ber99, BN98, Boh08, Bör94b, Dij74e, Dij75w, Dij83p, Dij84m, Dij85u, Dij87b, Dij87c, Dij89f, DS90a, Dij90c, Dij91q, Dij91-32, Dij92o, Dij99n, Hes95, ST99, SB04, vG90, vGDxx, Dij86-35, Dij89-38]. California [Dij81z, Dij70i]. calissons [Dij89-34]. called [Dij79-35]. Calls [Bri02]. Calvert [Dij91z]. Cambridge [Bir80, Dij99m, Wil80a]. Can [Dij85d, Dij89r, PJ89]. Canada [Dij74w, Dij75-52, Dij82-106, Dij82-49]. candy [Dij84a]. candy/sweet [Dij84a]. Canonical [Dij81f]. can’t [Dij86-48]. Canterbury [Dij73v]. Cantor [Dij99d, DM00, DM01]. Capacity [LSMS12]. Capita [vWBD56]. Capitalisation [vWBD56]. cards [Dijxxf]. Careful [Par18]. Carel [Bör94a, Dij92t]. Carlitz [Urb01]. Carpooling [SZHT10]. carries [Dij99f]. Carruth [Dij94q]. cascading [Dij99f]. Case [SWW99, BS99, Dij68d, Dij86-40, Dij01m, LCL^{+}94, SWW00]. case-based [LCL^{+}94]. cases [Kun80]. CAT [DvG82b, DvG82c, Dij85o]. Cathleen [Dij92x]. Cayley [Dij78y]. Central [Bau80, Zem80]. century [vdH17, MHR80]. ceremony [Dijxx-92]. certain [Kun80]. Certification [Dij75-47]. Ceva [Dij97l]. Ch [Dij85-37]. Ch.2 [Dij85u]. Ch.3 [Dij85-30]. Ch.5 [Dij85j]. Ch.6 [Dij85-33]. chair [Dij78r]. Chairman [Dij76-70, Dij82-105, Dij75-32, Dij95k, Dij82-82]. chairs [Dij80l]. Challenge [Coc02, Dij69g, Dij75f, Dij96b]. Challenges [Dij99b]. Chandy [Dij83e, Dij83f, Dij85i]. Chandy/Lamport/Misra [Dij83e]. Chang [Dij79n]. Change [LSMS12, ŁMB02]. channels [Dij56d, vL90]. Chap.4 [Dij85-36, Dij90r]. chapter [Dij91n]. characterization [Dij76h, Dij76-61, Dij76-65, Dij92m, Dijxxj, Rem90]. characterizations [BT79]. characters [Dij94f]. charlatans [Dijxx-30]. Charles [Dij86-43, Dij99h]. Chartreuse [Dij76-76, Dij82-113, Dij82-53]. Chechik [Ano13]. checkers [Dij92c]. Checking [Bai86, BDR85, Dij64e]. cheers [Dij82-65, Dij98g]. chemin [Bov86]. chemins [Pet71]. chessboard [Dij01b]. Chief [Dij80-40]. Choice [BN94, Dij63j, WM14, Dijxx-41]. chutspa [Dij80d]. chutzpah [Dij94e]. circles [Pos90, Wir90]. circuit [Hoa90]. circuits [Mar90]. circular [Dij79w]. city [GGAVGGRC15]. Claimed [Ano80g]. Class [DN01, DS82a, SD82, Dij72d, DS77, DS82d, Dij86v, Dij86w, Dij92y, Dijxxz]. classes [Dij76-59, Dij87v]. Classic [Bri01, Lew21, Hor90]. closure [Dij74d, DvG90b, Dij91q, Dij96-41, Dij00d, Dij00e, Leh77, vGD92]. closures [Dij98e]. Club [Dij78-42, Dij80-27, Dij80y, Dij82s, Dij85b]. Co [Dij68e, Dij68r, Dij69j, Dijxx-85]. Co-operating [Dij68e, Dij68r, Dij69j, Dijxx-85]. Coahuila [Dij91n]. Code [Fag01b, Har95, BD66, Dij56b, Dijxx-75]. coded [BS01]. codes [HHM94]. Coding [vL90]. coefficients [DvW55]. COGITANS [Dij75z]. coin [Dij89m, Dij89w, Dij89x, Dij90n]. coins [Dij90b, Dij96f, Dij97f]. Col [Dij78w]. Colchester [Dij84u]. Cold [Tat23]. Collaborative [BDBR01]. Colleagues [PSvE^{+}89]. Collection [Dij75-35, DLM^{+}76, DLM^{+}78, Dij82i, MHR80, Dij75g, Dij75-41, DLM^{+}75c, DLM^{+}75a, DLM^{+}75b, DLM^{+}77a, DLM^{+}77b, Dij81o]. college [Dijxx-54]. Collegebeschrijving [Dijxx-54]. Collocation [Dij02a]. Colloquium [Dij73v]. COLOSSUS [Ran80a]. coloured [Dij80b, Dij89y]. colouring [Dij99a]. Columbus [Dij77-39]. combination [AHHK93]. Combinatory [Dij80q]. combine [BD89, PRM11]. Combining [BDn^{+}10, Ber99, HSWW05]. come [DvG83b]. command [Dij56b, HSM97]. Commands [Bai86, Dij75y, Dij76-29, Dij96v, CRT78, Dij74e, Dij75x, Dij75w, Nel92]. Commemorating [Tat23]. comment [Dij79i, Dij87l, vDO01]. Commentaar [Dij68f, Dij75h, Dij77c]. Commentary [Dij78u, Van78, Dij68f]. Comments [Dij75i, Dij75j, DS76a, Dij80e, Dij82j, Dij89i, PSvE^{+}89, Dij64a, Dij89n, Dij89t, Dijxx-82, EM72, Knu66a, Knu66b, Rei83b, Rei83a, Dij75h, Dij77c]. commerce [JSW04]. commissie [Dij66b, Dij67a, Dij68b, Dijxxa]. commission [Dij66b]. Committee [BR70, DHPS59, Dij67a, NR69, Dij68b, Dijxxa, Dijxxb, Dijxx-96]. common [Dij90f, Dij96o, Dij97c]. Communicated [DS82i, DS76b, Dij91z, Dij96-34]. communicatie [Dijxxt]. communicatieapparaat [Dijxx-31]. communicatieapparatuur [Dij66c]. Communicatiebuffering [Dij65b]. communicatieprogramma [Dij55c]. Communicating [Dij82-71, Hoa78b, Hoa78a, Hoa85, Dij77d, Dij92u]. Communication [Dij59a, Dij59b, DS82a, Dij82c, SD82, Dij55c, Dij66c, DS77, DS82d, Dijxx-31, Dijxxt, Dij65b, Dij73u]. Communications [JNK11]. Company [Dij79g]. comparative [FCTB08]. Comparison [Nep00, Dij92d, ELCdCG14, Trä95]. compete [ITOK04]. Competition [Dij89q]. competitive [Nel90]. compilation [Har99]. compiled [BDR85]. compiler [Kru03]. Complete [Dij95d, Dij80-27, Dij95c, Dij00h, NMM78]. completeness [Tse95]. complex [Dij64d]. Complexity [CCD86, CMS92, Dij69c, Dij69d, Dij76i, Dijxxk, JM80, Bar98, Neu90, Nos85]. Components [Dij82q, Are90, Dij73d, Dij76w, Dijxx-27]. Composing [LX10]. Composite [SWX10]. Composition [Dij89f, CT90, Dij76-63]. Comput [Dij82o]. Computation [Dijxxl, Flo77, Dijxx-51, ITOK04, Mis01, WWZ05]. Computational [DK13]. Computations [DFvG83, DS78a, DS78b, DS79, DS80b, DFvG82, DFvG85, HSWW05]. Compute [Smi11]. Computer [Ano12, Big80, Bur80, CK02a, Chu80, Dij59a, Dij59b, Dij82-100, Dij89-32, Dijxxm, Hoa01a, Hur80, Lap96, Lav80, Lew21, Mar02, May80, SL98, Tat23, Tro80, Ano02a, Dij68l, Dij71d, Dij72e, Dij76x, Dij76-88, Dij78-40, Dij80t, Dij85d, Dij85e, Dij93e, Dijxxb, Dijxx-32, Maz76, RR64, Win78, Dij78t, Dijxx-61, Dijxxm, Dijxx-68, Dijxx-60, Hus80]. computerized [DvG82b, DvG82c, Dij85o]. Computers [Boo80, Bur80, Dij65d, Dij82c, Dij83c, Goo80, Raj80, Ran80b, Sti80, Svo80, Dij65d, Dij68c, Dij73u]. Computing [Ano13, Bau80, Bir80, Bul15, DM97, Der85, Dij72o, Dij79r, Dij82-69, Dij87d, Dij92e, Dij95-31, Dij99b, DLMB99, Gol02, MHR80, Sue80, TLR^{+}91, Ula80, Zem80, Zus80, Dij79q, Dij82-28, Dij82-29, Dij84k, Dij84r, Dij85d, Dij85-28, Dij86o, Dij86-36, Dij87-27, Dij88m, Dij88d, Dij89g, Dij91c, Dij91a, Dij01c, Dijxx-75, Lee06, SNR15, AMM^{+}12, Ano13, Dij89n, Dij90k, HTVW08]. concave [MS14]. concept [Dij68f, Dij68f]. Concepts [NRB76, Ban09, BN98, Dij62b]. Concern [Dij70a, Dij79-40]. Concerning [Dij85f]. concerns [Dij74c, Dij75l]. Concludes [Coc02]. conclusion [Dij83w]. Concurrency [Mal19, DFR74, Kot90]. Concurrent [Bri02, Dij65h, Dij75k, Dij78o, Dij79a, Dij83u, Dij96-37, Dij01j, Dij02i, Dij21b, Lam74, Ban07, Ban09, BBM07b, Dij77i, Dij77j, Dij77k, Dij78a, Dij78b, EM72, Knu66a, Knu66b, Lam19, Sch90a]. Concurrently [Pop95]. conditional [Dij77-33]. Conference [BD01b, BR70, Dij82-108, Dijxx-97, Dijxx-96, IEE79a, IEE79b, IdKM06, Sam76, BMW93, Dij76-72, Dij82-59, JSW04, NR69, Dij86-46, Dijxx-96]. Conferences [NRB76]. CONFIDENTIAL [Dij62d, Dij68n]. configuratie [Dij71b, Dij71c]. Configuration [Dij82u, Dij71b, Dij71c]. congres [Dij65j]. congress [Dij65j]. Conjecture [Dij86k]. connected [BBM07c]. connecting [Dij63m]. connection [Dij73g, Dij94p, NF95, Sni06]. connectives [Dij94d]. connexion [Dij59c]. conscience [Dij77v]. conscience-money [Dij77v]. consensus [Mis90]. consequence [Dij93s]. Consequences [Jon90, Dij76-62]. consideration [Dij76-62]. considerations [DS90f, Dij92o, Dijxxs]. Considered [Dij96u, Dij96-35, Dij01e, Dij02g, Dij21a, Dij65f, Dij65g, Dij68t, Dij68s, Dij69g, DS72c]. Consistency [BW00, Dij75-43]. conspiracy [Ban09]. constituent [Dij62b]. constraints [Dij78w, SPPR92]. Constructing [DR90a, Dij99c]. Construction [Dij82-45, Tse97, Wil80b, BB79, BMW93, Dij68o, Dij70a, Dij79o, DvG82b, DvG82c, Dij85o, Dij96-31]. constructions [AHHK93]. Constructive [Dij68h, Jac01, Dij67b, Dij68g, Dij73f]. constructs [Dij73q, Dij73a, Dij74p, Dij76-44]. containers [WWZ05]. content [Dij93e]. contents [Dij56d]. Context [Dijxxn, Sch18]. contexts [PG13]. contiguity [Dij78n]. Continuation [Dij63k, Dijxx-46]. continued [DFRxx]. Continuity [Dij82-74, DvG86c, Dij78l, Dij82n, DvG84b, DvG84c, DvG84d, DvG84e, DvG84h, DvG84i, DvG85b]. continuously [Dij79-35]. Contractie [Dij68i]. Contraction [Dij68i]. contribute [Dij86n]. contribution [Dij80-43, Dij81o, Dij68c]. Contributions [BD01b, BD02, BD01b]. Contributors [Ano80b, Ano80i]. Control [Dij65h, Dij74o, Dij82-40, Dij83u, Dij96-37, Dij01j, Dij02i, Dij21b, BGM90, Dij64k, Dij73l, Dij73o, Dij74n, Dij82e, Dij83q, Dij83r, Dij83s, EM72, Knu66a, Knu66b, KF08]. controlerende [Dijxx-31]. Controlled [Dij69c, Dij76i, Dij69d, Dijxxk]. controller [Dijxx-31]. controversy [Par78]. Conventions [Dij02e, Dij79e, Dij00b]. convergence [Dij99j]. Converse [DS90b, Dij96z, Dij97j]. Conversion [CCD86, Dij91d]. convex [Dij76-50]. Convocation [Dij96i]. Cooperating [Dij68j, Dij68k, Dij96j, Dij02b, Dij65a, Dij66d, Dij67e, Dijxx-86]. Cooperation [DLM^{+}76, DLM^{+}78, Sam76, DLM^{+}75c, DLM^{+}75a, DLM^{+}75b, DLM^{+}77a, DLM^{+}77b]. Copenhagen [Dij76-71, Dij82-50]. Copyright [Ano80c]. core [DD98]. correct [Dij77-28, Dijxx-93]. Correction [Dij85g, Dij78j, Dij83d, Dij85h, Dij91e, Dij97b]. correctly [TZDM10]. Correctness [Dij68h, Dij74c, Dij75l, Dij78o, Dij79a, Dij82-71, Hoa01b, BS99, Dij67b, Dij68g, Dij70a, Dij73f, Dij77d, Dij77i, Dij77j, Dij77k, Dij78-27, Dij78a, Dij78b, Dij79-40, Dij82f, GB96, Har81, MHMO10, Mau78]. Correlation [SWX10]. Correspondence [Gal73, Dij87q, Dij02f]. Corrigendum [Hoa78b]. corrupt [Dij99i]. cosine [Dij86g]. Cost [ZXZ^{+}09, HTVW08]. could [Dij84f, Dij98e]. Council [Dij76b, Dij76c, Dij87x]. Counting [Dij94f]. country [Dij98i]. couples [Dij91i, Dij96k, Dij96-43]. coupling [Dij89k, DvG90a]. course [Dij71d, Dij86q, Dij87k, Dij92h, Dij93h, Dij95n, Dij96y, DDvW57, Dij60b, Dij70b]. court [Bov86]. Courtesy [Dij86h, Dij91f, Dij95e, Dij96l, Dij99d]. courts [Pet71]. cover [Dij93o]. covered [Dij01b]. Covering [Dij92f, Dij89-29, Dij89-30]. Coxeter [Dij02c]. Cradle [Tat23]. craft [Dij76-58, Dij77-27, Dijxx-33]. Craftsman [Dij75n, Dij75m, Dij82-72]. Crane [WLXX09]. crawl [Dij79-38, Dij79-39]. Created [Lew21, Dij90k, Dij91a]. Crisis [Pay14, Dij82k, Dij82l]. Criteria [Par01b, Dij89-35]. criterion [Tse95]. critical [Dij77-33, SS10]. Crook [Dij88b]. crossflow [Dij76x]. Cruelty [Dij89-32, Dij88m]. Cruz [Dij79-31]. Cry [Day11a, Day11b, Day12b]. CSP [Mor90a]. cubic [Dij99j]. Cultural [Dij86x, Dij85w, Dij85x]. curious [Pos90]. curriculum [Dijxx-32]. Cursus [DDvW57, Dij60b, Dij70b]. curve [Dijxxa]. Curves [KM83]. cyber [SAID13]. cycle [vT90]. cycles [Dij80-27]. Cyclic [Dij82a, Dij73r, Dij82d]. cyclically [BBM07c]. Czechoslovakia [Svo80].

D [Dij76-70, Dij82-105, FCTB08, LCZW14, VPM^{+}10]. D. [Dij77q, Dij80f, Dij80-38, Dij82-89, Dij93b]. daarvan [Dijxxo]. DAGs [Dij95d]. Dahl [Ano02c, Bar75, Bar74, Llo74]. dance [Dij98i]. dans [Bov86, Pet71]. Danupon [Ano13]. dartboard [Dij89-39]. Data [Che01, Cod01, Dij63r, FJ09, Hoa01b, WG01, BS01, TZDM10]. Daten [BS01]. David [Dij75-36, Dij75-37, Dij76-38, Dij76-39, Dij96-38]. Dawn [Day12a]. Days [Bul15]. Deadlock [CMS92, BBM07c, Dij79n]. Deadlock-Free [CMS92]. deadly [Dijxxu]. deal [Dij76-88]. Dear [Dij70c, Dij79g, Dij99e, Dij63l]. dearest [Dij99e]. death [Dijxxw]. Debatable [BSE^{+}90, WWSD90]. debts [Dij80-33, Dij82-38]. Dec [Dij76-76, Dij78-45, Dij82-113, Dij82-53]. Dec.1979 [Dij80-44]. Dec.’84 [Dij85-48]. December [Dij74v, Dij82-109, Dij82-51, Dij67g, Dij72b, Dij73w, Dij96i]. Decimal [Gri90]. decision [TZDM10]. Declarations [DN01]. decoding [HHM94]. Decomposing [Dij88e, Par01b]. Deconstructing [Lam22]. dedicated [Dij77q, Dij82-89, Dij95s, Dij99h, Dij82s]. Dedication [Ano80d]. deduction [Boh08]. Deductive [Bro96, Bro96]. deep [Dij88d, Dij89g]. default [Dij56a]. Defense [Dij61b, Dij75-38]. defined [JHLC14]. Defining [Dij97c, Sch18, Dij00c]. definitely [Dijxx-79]. definition [Dij73i]. definitions [Dij80x, Dij91p]. Degree [Dij02a, Dij73b, Dij86u, Dijxx-92]. degrees [Dij87g]. delay [Lee06, Mar90]. delay-insensitivity [Mar90]. delays [Dij72d]. Delft [JSW04]. deliver [Dijxx-30]. DeMillo [DLP78, Dij78u, Van78, Kil79]. demon [Dij77v]. demonstranda [Dij91o]. dense [PG13]. Department [Ano12, Dij64j, Dij75-38]. dependent [Dijxxn]. Depot [KLMZ11]. Depth [Dij82u]. Derde [Dij75o, Dij75p, Dij76k, Dij76j, Dij77e]. Derivation [Dij75y, Dij76-29, DFvG82, DFvG83, DFvG85, Dij96v, Dij74e, Dij75x, Dij75w, Dij77-28, Dij84b, Dij88f, Dij89h, Dij90d, Dij93c, Dij93b, Hoa90]. derived [Dij87r, Dij88l, Dij96-42]. deriving [Ber99, Fei90]. Dertiende [Dij75q, Dij76l]. described [PJ89]. Description [Dij64c, Dijxxq, Dijxxr, Dij53, Dij56b, Dij65c, Dijxx-54]. descriptions [Dijxx-58]. Design [Dij63c, Dijxxs, Fag01b, GHJV01, Jac01, AHH^{+}95, Bis85, Bro96, Dij61d, Dij76-44, Dij78-27, Dij78s, Dij79h, Dij89-33, Dij91g, Dij92q, Dij92h, Dij93k, Dij93h, Dij94x, Dij95b, Dij01g, Dij05, Dijxx-47, Ter93, Dij62c]. Designed [Dij82-93, Dij77t, Dij79j, Dij86-39, Dij91r]. designer [Dij77-46, Dij78-33, Dij82-115]. designers [Dij75f]. Designing [DR90b, DM00, DM01, RD92]. designs [Hoa90]. desk [Dij95k]. detail [Dijxxs]. detailed [Dij93c]. Detection [DFvG83, DS78a, DS78b, DS79, DS80b, Dij82m, DFvG82, DFvG85, Dij87u, KP10, UOZ11, Dij79n]. Detector [KP09, MJC07]. determinacy [Dij74e, Dij75x, Dij75w, Dij76-29, Dij76-42, Dij77-30, Dijxx-49]. determinant [Dij97i]. determinants [Dij98g]. determining [Dij91-30]. Determinism [Dij74d]. deterministically [Lee06]. Detroit [Dij78-45, Dij78g]. DEUCE [Wil80b]. Developed [CH10, Cha10]. Development [Bau80, Big80, Chu80, Dij89p, ESB80, Fag01b, Hur80, JNK11, KP80, KM83, Lav80, WG01, Wir01a, Dij62c, Dij77c, Dij72k]. Developments [Bir80, Wil80a]. device [Dijxx-31]. d’Huez [Dij72b]. diabolical [Dij86-32]. diameter [Dij86v, Dij86w, Dij87g]. diamonds [Dij89-29, Dij89-30]. diatomic [Urb01]. Did [Dij82-99, Ham80, Dij77-30, Dij95x]. Dies [Gol02, Mar02]. differ [Dij86-42]. different [DS82e, FNS03, PS14]. Differential [Dij02a]. Difficult [Dij86i, Dij89-39]. difficulties [Dij79-36, Dij79-37]. diffusing [DS78a, DS78b, DS79, DS80b]. Digital [Bau80, Bul15, Ran80b, Dij64j]. digitale [Dij64j]. Dijkstra [AMM^{+}12, Ano02b, Ano02c, Ano06b, Ano12, Ano13, Bar75, Bar74, Bör94a, Bör94b, Du10, zFmLcWcJ10, GGAVGGRC15, Llo74, Mil89, Mos10, Sch18, TLR^{+}91, AVL02, Alp18, AHHK93, AHH^{+}95, ANGK09, Ano02a, Ano06a, Anw03, Apt02, AH22, AAEG11, Bai86, BBM07a, BBM07c, Bar98, BMST03, BB03, BDn^{+}10, BS99, BES95, Ber99, BN98, Boh08, BB99b, BB99a, Bov86, BFG^{+}02, BN94, BW00, CK02a, CK02b, CF04, CRT78, Cha10, CH10, Che02a, Che02b, Che16, CSZ10, Coc02, CMMS98, Cre02, CC12, Day11a, Day11b, Day12b, Day12a, Der85, DMS83, Dij71e, Dij71h, Dij72b, Dij72n, Dij72r, Dij73k, Dij75-53, Dij76-37, Dij80-31, Dij82-87, Dij82-57, Dij84l, Dij89i, Dij98h, Dijxx-76, DK13, DLMB99, DH01, DS72c, DCF^{+}02, EM72, EKP96, ELCdCG14, EBR09, FS10, FvGGM90, FC14, Flo77, FJ09]. Dijkstra [FNS03, FURD13, Gal73, GL98, GLFM07, Gaw64, Gol02, GS97, GA21, HHM94, Har98, HKS93, Hes95, Hoa03, HM12, HCSW09, HZW09, ITOK04, JAE^{+}12, Jef78, JHLC14, Joh73, JM80, JNK11, KYTY09, KLK08, KGZ06, Kil79, KP09, KP10, Knu77, Koh76, KLMZ11, Kos73, KKNI14, Kru03, Kun80, KF08, Lam74, Lam19, Lee06, yLsLtZ09, LX10, yLsLtZ10, LN14, LSMS12, LCL^{+}94, LW11, LSJT12, LCZW14, LWQ09, ŁMB02, MB97, MCX13, Mar02, MHMO10, MS01, MZS02, Mis10, Mis01, MR03, MSS^{+}06, Moo02, Muk12, MS14, MTZ12, MJC07, NA13, Nel89, Nel92, Nep96, ND00, Nep00, NB93, NBK95, NAGK09, NMM78, Nos85, NS00, OC10, PRM11, Par78, PSvE^{+}89, PMR11, PRC12, PRD14, PPS^{+}13, Pay14, PS14, PG13, Pet71, PRV09, Pop95, Ran03, Rei83b, Rei83a, Rei08, RPGS14, ST99, SL88]. Dijkstra [SWW99, SWW00, SNR15, SWX10, SZHT10, SS10, Sha81, She03b, She03a, SS04a, SS04b, SxXq11, SB04, Smi11, Sni06, SPPR92, SEEM01, Str72, SVSR04, SAID13, SWS^{+}13, TZDM10, Ter93, TS01, Tse95, Tse97, Uch14, UNTH14, Urb01, UOZ11, Van78, Vv72, VPM^{+}10, WYY11, WM14, WLXX09, xXL10, XZY^{+}12, XLH^{+}07, XWZH12, YJ10, YL10, YL11, ZWKT10, ZJM12, ZTZY12, ZLY13, ZXZ^{+}09, dB76, vDO01, vL04]. Dijkstra-Based [LSMS12, MTZ12, NA13]. Dijkstra-immune [AAEG11]. Dijkstra-Like [Der85, Kun80, SNR15]. Dijkstra-Verfahrens [Str72]. dilemma [Dij68m]. dimensional [Dij79l, Dij82-27, Dij95c]. dimensions [Dij76-50]. Dining [BBM07a]. diploma [Dijxx-92]. direct [AHHK93]. Directed [Dij82q, BDn^{+}10, Dij73d, Dij76w, Dijxx-27]. disappointing [Dij87q]. Disaster [JNK11]. Discipline [Dij74m, Dij75-31, Flo77, BD79, Dij73j, Dij76m, Dij76-58, Dij77-27, Dij89n, Dijxx-38]. disciplines [Dij86-27]. disconnected [Dij86u]. discontinuities [Bra90]. Discoveries [SL98]. Discrete [MMP87, Dij90i]. discriminated [Dij78-38]. disease [Dij81a, Rei83b]. disentangled [Dij92m]. disjoint [Dij01h]. disjunction [Dij91u]. Disputing [GA21]. Dissertation [Ano13]. Dissolving [vdH17]. distance [Dij83k, Dij83l, FCTB08]. Distances [DS81, MGH^{+}05]. Distributed [AMM^{+}12, Ano13, Bri01, Dij74o, DS80a, Dij82m, Dij82-40, DFvG83, BW00, Dij73l, Dij73o, Dij74n, Dij80-35, DFvG82, Dij82e, Dij83e, Dij83q, Dij83r, Dij83s, Dij85i, DFvG85, Lam22, SZHT10, Trä95, Dij83f]. distributedly [Rei08]. distributes [Dij99p]. Distribution [HCSW09, Dij86p, SVSR04]. distributivity [Dij92i]. Divide [Dij97g]. Division [Dijxxd]. divisor [Dij90f, Dij96o, Dij97c]. dMLP [Par78]. Do [Dij82-79, Dij75-27, Dij82t, Dij86n, Dij86t]. Doctoral [Ano13]. Documentatie [Dij66c, Dijxxt]. Documentation [Dij66c, Dijxxt]. DoD [Dij75j, Dij78z, Dij78c, Dij78k, Dij78-29, Dij78d, Dij78-30, Dij78e, Dij78-31, Dij78f]. DoD-I [Dij78k]. dodelijke [Dijxxu]. doelstelling [Dijxx-69]. Doing [Dij93l, Dij92r, Dij92s]. dominated [SNR15]. dominoes [Dij01b]. Don’t [Dij94g, Dij88p]. dood [Dijxxw]. door [Dij63n, Dij72n, Dijxx-96]. Doornbos [Dij91r]. double [Hab90]. Doubling [Dij02a]. Doug [Dij87p, Dij89-27]. Doyen [CK02a]. Dr [DS76a, Dij86y, Dij95e]. Dr. [Dij74g, Dij76v, Dij82-76, Dij78v]. Draft [Dijxx-54, Dijxx-55, Dij78t, WTD^{+}69, Dij76-62, DS80a, Dij80-43, Dij81-35, Dij82y, Dij85x, Dij85j, Dij85u, Dij85-30, Dij85-33, Dij85-37, Dij85-36, Dij90r, Dijxx-56]. Drawing [Wir90]. dream [Dij01n, Dij00k]. drijvende [Dij65c, Dij54a]. Drijvende-Komma [Dij54a]. driven [AHHK93, AHH^{+}95]. drum [Dij63q, Dij64b]. Dual [FNS03]. Duality [Tur90a]. due [JM80]. duikboten [Dij89r]. dummies [Dij89-36, Dij92n]. dummy [Dij91o]. d’un [Bov86]. duplicating [Dij79t]. Dutch [Dij76-51, DDvW57, Dij53, Dij54a, Dij54b, Dij55c, Dij55a, Dij55b, Dij56b, Dij56c, Dij56d, Dij56e, Dij56a, Dij57c, Dij57d, Dij60b, Dij60c, Dij62d, Dij62c, Dij63g, Dij63d, Dij63e, Dij63h, Dij63o, Dij63j, Dij63k, Dij63m, Dij63n, Dij63p, Dij63q, Dij64b, Dij64h, Dij64i, Dij64j, Dij65b, Dij65c, Dij65d, Dij65j, Dij65i, Dij66b, Dij66c, Dij66d, Dij67a, Dij67e, Dij67g, Dij68b, Dij68c, Dij68i, Dij68f, Dij68m, Dij68n, Dij68l, Dij68r, Dij69f, Dij69j, Dij69k, Dij70b, Dij70g, Dij70i, Dij71d, Dij71e, Dij71b, Dij71c, Dij71h, Dij72b, Dij72e, Dij72k, Dij72l, Dij72n, Dij72p, Dij72q, Dij72r, Dij73c, Dij73k, Dij73x, Dij75b, Dij75c, Dij75h, Dij75q, Dij75o, Dij75p, Dij75r, Dij75t, Dij75s, Dij75-34, Dij75-59, Dij75-50, Dij75-58, Dij75-51, Dij75-62, Dij75-64]. Dutch [Dij75-63, Dij75-65, Dij75-67, Dij75-66, Dij75-68, Dij76b, Dij76c, Dij76e, Dij76d, Dij76l, Dij76k, Dij76j, Dij76q, Dij76n, Dij76-36, Dij76-81, Dij76-67, Dij76-79, Dij76-80, Dij76-85, Dij76-87, Dij76-84, Dij76-86, Dij76-88, Dij76-92, Dij76-93, Dij76-91, Dij77a, Dij77c, Dij77e, Dij77g, Dij77f, Dij77o, Dij77-42, Dij77-44, Dij77-45, Dij77-48, Dij77-49, Dij79i, Dij79-36, Dij79-37, Dij80h, Dij80l, Dij80i, Dij80t, Dij82-73, Dij82v, Dij82k, Dij82l, Dij83j, Dij83w, Dij84a, DF84, Dij89r, Dij93e, Dij93t, Dij96m, Dijxxa, Dijxxb, Dijxxd, Dijxxc, Dijxx-61, Dijxxu, Dijxxf, Dijxxi, Dijxxh, Dijxxo, Dijxx-31, Dijxx-32, Dijxxm, Dijxxt, Dijxx-33, Dijxxw, Dijxx-37, Dijxx-30, Dijxx-36, Dijxx-57, Dijxx-39, Dijxx-38, Dijxx-41, Dijxxp, Dijxx-43, Dijxx-45, Dijxx-46, Dijxx-54, Dijxx-56, Dijxx-55, Dijxx-62, Dijxx-69, Dijxxv]. Dutch [Dijxx-76, Dijxx-77, Dijxx-65, Dijxx-59, Dijxx-81, Dijxx-84, Dijxx-58, Dijxx-63, Dijxx-64, Dijxx-85, Dijxx-86, Dijxx-89, Dijxx-90, Dijxx-91, Dijxx-92, Dijxx-95, Dijxx-60, Dijxx-34, Dijxx-98, Dijxx-97, Dijxx-96, Dijxx-100, Dijxx-99, Swi11, vWBD56]. dynamic [Dij04, EKP96, RPGS14, Sni06, UOZ11]. dynamically [Anw03]. dynamics [BDB93].

E. [Dij71h, Dij72b, Dij72n, Dij72r, Dij73k, Dij73t, Dij73s, Dij74u, Dij74x, Dij74w, Dij74t, Dij74v, Dij75-53, Dij75-52, Dij75-54, Dij75-55, Dij75-57, Dij75-56, Dij75-39, Dij76-37, Dij76-73, Dij76-78, Dij76-72, Dij76-75, Dij76-77, Dij76-71, Dij76-74, Dij76-76, Dij77-39, Dij77-38, Dij77-37, Dij77-40, Dij77-41, Dij78-44, Dij78-49, Dij78-46, Dij78-47, Dij78-48, Dij78-45, Dij79-33, Dij79-30, Dij79-29, Dij79-31, Dij79-32, Dij80-44, Dij80-47, Dij80-45, Dij80-46, Dij81-28, Dij81z, Dij81-27, Dij81-30, Dij81-31, Dij81-29, Dij81-32, Dij81y, Dij81-33, Dij82-87, Dij82b, Dij82-109, Dij82-106, Dij82-110, Dij82-108, Dij82-112, Dij82-114, Dij82-113, Dij82-107, Dij82-111, Dij82-54, Dij82-51, Dij82-49, Dij82-55, Dij82-52, Dij82-50, Dij82-56, Dij82-60, Dij82-59, Dij82-62, Dij82-63, Dij82-53, Dij82-58, Dij82-61, Dij82-88, Dij83x, Dij83-27, Dij83y, Dij83-28, Dij83z]. E. [Dij84u, Dij84t, Dij84w, Dij84y, Dij84z, Dij84s, Dij84x, Dij84v, Dij85-48, Dij85-47, Dij85-50, Dij85-49, Dij85-45, Dij85-46, Dij86-44, Dij86-45, Dij87y, Dij88r, Dij88s, Dijxx-76, SL88]. EAA [Dij77-30, Dij82-99]. Early [ANGK09, Bau80, Bul15, Day11b, Dij80-30, ESB80, KP80, Raj80, Sti80, Sue80, Wil80a, Dij76-57]. easy [Dij78r, Dij86i]. Ebergen [Dij83j]. ECI [Dij76-72, Dij82-108, Dij82-59, Sam76]. ECI-Conference [Dij82-108, Dij76-72, Dij82-59]. Economics [Boe01]. Economy [Dij92s, Dij93l, Dij92r]. Edger [Che02a]. edges [Dij80-27, NA13]. Edible [Dij90a]. Edinburgh [Dij74t]. edition [Dij70b]. EDITOR [Dijxx-88, BD01a, Dij61b, Dij68s, Dij80-40, Knu66b, RD68]. Editor-in-Chief [Dij80-40]. Editors [AG01]. Eds [Dij77b]. EDSAC [Dij79e, Har99]. Edsger [AMM^{+}12, Ano02c, Bör94a, Flo77, Mil89, Sch18, TLR^{+}91, Alp18, Ano02a, Ano02b, Ano06b, Ano12, Apt02, AH22, BB03, BFG^{+}02, CK02a, CK02b, Che02b, Che16, Coc02, Cre02, Dij71e, FvGGM90, Gol02, Hoa03, Hoa22, Mar02, Mis10, MR03, Moo02, Ran03, She03b, She03a, vL04]. educatief [Dij68m]. Education [Dij77b, Dij06, Dij83c]. educational [Dij68m, Dij80g, Dij80r]. Eerste [Dij75r, Dij75s, Dij76n, Dij77f, Dijxxw]. Effective [Dij76p, Dij82u, Dij76o, Dij96n, Dij96p]. Efficient [Bau90, Der85, KP09, BD89, KP10, WWZ05]. effort [DFR74, Dij80r, Dij85a, Dijxxx]. eigenlijk [Dij77o]. eight [Dij87z]. Eighth [Dij75b, Dij75c, Dij76e, Dij76d, Dij77a]. einde [Dijxx-33]. Eindhoven [Dij80t, IdKM06, Dij80t, Dij96-40]. Eisenstein [Urb01]. EL-X8 [Dij65b]. elected [Dij82s]. Electrical [Dij64j]. Electrologica [Kru03]. Electronic [Svo80, JSW04]. Elegance [Dij96n, Dij77h, Dij78m, Neu90]. elektrotechniek [Dij64j]. Elementary [DvG82a, Dij80-48]. Elementen [Dij72e]. Elements [Pad85, Dij73g, Dij72e]. Elephant [Dij82-73, Dij82-86, Dij74j, Dij77g, Dij77-31]. elephants [Neu90]. Eleventh [Dij75t, Dij76q]. Elfde [Dij75t, Dij76q]. Eliminating [Dij99f]. elimination [DvG82a, LH90]. ellipses [Wir90]. embedded [SPPR92, ZWKT10]. Embedded-GIS [ZWKT10]. Embedding [Dij64d]. Embellishments [Dij82-75, Dij75u]. embrace [Dijxxu]. emergence [Dij82k, Dij82l]. emergency [YL10]. Emerging [Dij82-101, Dij02k, Dij75-49]. Empirical [SWW99, HKS93, SWW00]. Employing [NAGK09]. empty [Dij96g]. encoding [BD89]. end [Dij01c, Dijxx-33]. ending [Dij82g]. ener [Dij72e, Dijxxh]. energy [Dij72e]. enforceable [Dij78w]. Engeland [Dij72q, Dijxx-76]. Engineer [Dijxx-100, Dij78-33, Dij83j, Dijxx-39]. Engineering [Boe01, BD01b, BD02, BR70, Day12a, Dij77b, Dij79z, Dij00a, Dijxxm, Dijxx-96, HSW08, IEE79a, NR69, NRB76, Ran02, Dij80t, Dij64j, Dijxxm, Dijxx-97, Dijxx-96, IEE79b, Tom80]. England [Dij76-73, Dij81-27, Dij72q, Dij73s, Dijxx-76]. enhanced [YL10]. ENIAC [Bur80, Eck80, Mau80]. enkel [Dijxxm]. enquete [Dijxxh]. enriched [Dij01f]. enter [Dij93g]. enterprise [Dij94z]. Entity [Che01]. entropy [Hel90]. environment [Dijxxp, Muk12]. Environments [DH01]. Eötvös [Dij89q]. Epp [Dij90i]. equality [Dij81q, DS90l, Dij91g, Dij91p, Dij91v, Dij01f]. Equation [Dij02a, Dij86-30, Dij95f]. equational [Boh08]. Equations [DS90c, Dij85j, Dij86j, Dij87e, Dij94s, Dij96-32]. Equilateral [Dij94h]. equipment [Dijxxt]. Equivalence [Dij82-74, Dij88g, Dij76-59, Dij78l, Dij82n, Dij82-65, Dij85f, DS86, Dij91-34, Wal90]. Era [Mar02, Tom80]. Ernest [Dij79n]. Errata [Dij82o, Knu66a]. Erratum [Dij75u, Dij82-75]. Error [Dij64e, Dij80j]. Errors [Fag01b, vL90]. Ershov [Dij89b]. Erster [Dij71f]. Essay [MR80, Dij76r, DvG82g, Dij84f]. Essays [Dij77h, Dij78m, MHR80, Pad85]. Estonia [VU05]. ETAC [Dij91f, Dij91i, Dij94a, Dij96a, Dij01p]. etc [Dij56b]. ETH [Dij75-53, Dij82-57]. Ethernet [MGH^{+}05]. Etude [SS04b, SS04a]. Euclid [Dij76s, Dij88h, Dij93a, Dij96e]. Euclidean [Dij94j, FCTB08]. Euler [Dij81p, DvG81c, Dij99l]. Europe [Dij86-44, Bau80]. European [Sam76, Zem80]. Evacuation [XWZH12]. Evaluation [VPM^{+}10]. Even [Der85]. event [Dij75-40]. events [BBM07b]. Everett [DvW55]. everywhere [Dij90e]. Evidence [Dij06]. Evolution [Wel80, Dij84o]. evolutionary [OC10]. EWD [Dij75-29, Dij02d]. EWD-series [Dij75-29]. EWD1002 [Dij87w]. EWD1016 [Dij96-39]. EWD102 [Dijxxr]. EWD1047 [Dij91t]. EWD1048 [Dij92z]. EWD1099 [Dij91h]. EWD1103 [Dij96-43].

EWD1107 [Dij92y]. EWD1112 [Dij91e]. EWD1115 [Dij96-43]. EWD1117 [Dij92n]. EWD1140 [Dij95y]. EWD1163 [Dij94a, Dij96a]. EWD1166 [Dij99g]. EWD1171 [Dij95y]. EWD1192 [Dij95g]. EWD1240 [Dij97b]. EWD1241 [Dij96-36]. EWD1247 [Dij96b]. EWD126 [Dijxx-80]. EWD1300 [Dij02e]. EWD140 [Dij66c]. EWD200 [Dij67c]. EWD201 [Dij67d]. EWD237 [Dijxx-44]. EWD239 [Dijxx-44]. EWD316 [Dijxx-79]. EWD386 [Dij82d]. EWD391 [Dij82e]. EWD398 [Dij73e]. EWD418 [Dij75-30]. EWD501 [Dij75-46, Dij82-36, Eds82]. EWD503 [Dij75u, Dij82-75]. EWD51 [Dijxx-46]. EWD535 [Dij75-48]. EWD54 [Dij63k]. EWD570 [Dij76-34, Dij82-84]. EWD591 [Dij77-34]. EWD592 [Dij77-29]. EWD619 [Dij78m]. EWD651 [Dij78j]. EWD653 [Dij78x]. EWD668 [Dij78s]. EWD708 [Dij79-39]. EWD709 [Dij79r]. EWD740 [Dij80-34]. EWD744 [Dij80j]. EWD755 [Dij80-28]. EWD764 [Dij89-37]. EWD771 [Dij80p]. EWD809 [Dij82g]. EWD822 [DS82f]. EWD842 [Dij82-41]. EWD851b [Dij83d]. EWD859 [DS83d]. EWD863 [Dij84p]. EWD866 [Dij83m]. EWD874 [DS84c]. EWD878 [Dij86-41]. EWD901 [Dij85-44]. EWD912 [Dij85p, Dij85-37]. EWD914 [Dij85h]. EWD914-15 [Dij85h]. EWD932b [Dij85g]. EWD935a [Dij85-51]. EWD935a-0 [Dij85-51]. EWD955 [Dij86-31]. EWD964 [Dij87l]. EWD969 [Dij87a]. EWD977 [Dij86-37]. EWD996 [Dij96x]. Exam [Dij66d, Dij67e, Dij68r, Dij69j, Dijxx-85, Dijxx-86]. examination [Dij79j]. Example [Dij79p, Sha81, Dij73f, Dij76-35, Dij80n, Dij89j, Dij89u, Dij89-45, Dij98b, vdW90]. Examples [CT90, Dij76-27, Dijxx-28]. Excerpt [Dij63g, Dij63h, Dij63d, Dij63e, Dij63g, Dij63d, Dij63e, Dij63h]. exclusion [BW00, CSZ10, Dij77-43, Dij83q, Dij83r, Dij83s, Dij85v, PJ89]. execution [Dij62b, JAE^{+}12]. Executional [Dij76t]. Exercise [DLM^{+}76, DLM^{+}78, Dij82-71, Dij82-76, Are90, Dij75-41, DLM^{+}75c, DLM^{+}75a, DLM^{+}75b, Dij76u, Dij76v, Dij77d, DLM^{+}77a, DLM^{+}77b, Dij79k, Dij79j, Dij81h, Dij82-67, Dij89-28, Dij00j, Fei90, Kes88, MP90]. Exercises [Dij75v, Dij82-77, Bir90]. existence [DvG81c]. Exit [Dijxxy]. expansie [Dij68i]. expansion [Dij68i]. expectation [Dij91-30]. expected [NMM78, Nos85]. expensive [Dij77-46, Dij82-115]. Experience [Har95, Dij62a, Dij75d]. experiences [ANGK09]. Experiment [Dijxxz, Dij80k, Dij95b, Dij99n]. Experimental [Dij80t, Dij86o, Trä95]. Experimentele [Dij80t]. Experimenting [vG90]. experiments [MGH^{+}05]. explanation [Dij77-46, Dij82-115, Dij86l, Dij86m, Dij87h, Dij87i]. exploitation [BDS82]. Exploiting [Dij78n]. exploration [Dij96p, Dijxxw]. Exploring [Bra90, Dij89k]. Export [Dij54b]. exposition [Dij79k, Dij80k]. expression [Mor90b]. expressions [Bij90, Dij94d]. extendable [Dij69h]. extended [Dij86-43, NS00]. Extending [JHLC14, KGZ06]. Extension [GGAVGGRC15, GLFM07, Dij85g, Vv72]. Extensions [DHPS59]. extract [Dij92y]. Extreme [Dij85j, Dij86j, Dij87e, DS83a, DS83b, DS84b, DS90c, Dij91s, Dij92m, Dij94s, Dij95i, Dij96-32, SNR15].

F [DS76a, Dij77q, Dij82-89]. F. [Dij80-39, Dij86k, Dij95s]. fabrikanten [Dij63g, Dij63d, Dij63e, Dij63h]. Fact [Dij82-91, Dij77r]. factor [Dij76-56]. factorial [Dij89l]. Factorizing [Dij89l]. factors [Dij95-27]. failed [Dij92-30]. Fair [BN94, Dij89m, Dij77v, Dij89w, Dij89x, Dij90n, Dij95q]. Fairness [Ban09, Ban07, CM88, Dij87t, Dij88o, SL88]. Fall [Dij87k, Dij89-38, Dij96n]. far [Dij76-62]. far-reaching [Dij76-62]. Fast [DvG82b, DvG82c, Dij82p]. fault [ZTZY12]. Feasibility [Dij75-47, Per90]. Feb [Dij80-47, Dij83x, Dij85-50]. februari [Dij80t]. February [Dij77-38, Dij80t, Dij82-57, Dij82-107, Dij82-58, Dij84t, Dij75-53]. feedback [YL11]. Feijen [Mil89, Dij79j, Dij79s, Dij79w, Dij86-30, Dij95-28]. fell [Dij80l]. Fellow [Dij82s]. Fermat [Dij80-36]. FERTA [Dij55a, Dij55b]. few [Dij82f, Dij89n, Dij95g, Dij01i, Yap83]. Fibonacci [Dij78p, Dij79m, Dij81g, Dij90f, Dij96o]. fiction [Dij86-36]. Fifth [Dij75-64, Dij76-87, Dij76-86, Dij77-45, Dijxx-98]. Fifty [DM97, Dij96-29, Dij96-30]. Fighting [Rei83b]. figure [Dij89-29, Dij89-30]. file [Dij76-83]. filler [Dij88a]. Fillers [Dij87f, Dij90g, Dij90h]. final [Dijxx-43, WTD^{+}69]. Finding [Dij73d, Dij76w, Dij77i, Dij77j, Dij77k, Dij78o, Dij78a, Dij78b, Dij79a, Dij82q, Dijxx-27, MP91, Anw03, LCL^{+}94, Pet71]. Fine [Dij82-89, Dij77q]. Fine-Grained [Dij82-89, Dij77q]. finite [Dij72i, Dij72j, DS86]. Fire [XWZH12]. firing [Pop95]. First [IdKM06, van19, Ash87, Dij76x, Dij78t, Dij89-45, Dij94-28, Dij96p, vGDxx, Dij75r, Dij75s, Dij76n, Dij77f, Dijxxw]. Firsts [Ano80g]. fix [DvG84b, DvG84c, DvG84d, DvG84e, DvG85b]. fix-point [DvG84b, DvG84c, DvG84d, DvG84e, DvG85b]. Fixed [dB76, Mis01]. fixed-point [Mis01]. Fixpoint [DvG86c, Dij96-27]. Flag [Dij76-51, Dij77g, Dij82-73, Swi11]. floating [Dij65c, Dij54a]. floating-point [Dij65c, Dij54a]. FLOPS [HW06]. flow [GGAVGGRC15, JM80]. flowchart [BT79]. Floyd [FURD13, DK13, HM12, LSJT12]. Floyd-Warshall [FURD13]. Fly [DLM^{+}76, DLM^{+}78, Dij75-41, DLM^{+}75c, DLM^{+}75a, DLM^{+}75b, DLM^{+}77a, DLM^{+}77b]. folklore [Dij80s]. Following [Dij77o, Dij80t]. Foolishness [Dij79u, Dij78-28]. football [Dij78-34]. forced [Dij95a]. Foreword [Dij64f, Dij96q, Dij96r, Dijxx-56, Dij93t]. Forge [Dij78-48]. form [Wal90]. Formal [BS99, DH96, Dij75y, Dij76y, Dij76z, Dij79p, Dij82r, Dij89p, Dij96v, BGM90, Dij64g, Dij75x, Dij76-27, Dij76-35, Dij84b, Dij89j, Dijxx-28, Dij77-28]. formalism [Wal90]. Formalizing [BES95, Har98, Hor90]. format [Dij87s, DS90d]. formed [CCD86]. Formelübersetzung. [SB01]. formula [Dij83i, Dij91r, Dij96s, Dij96t, Dij99h, SB01]. formulae [Dij82-46, DvG84f]. FORTRAN [CW75]. Forty [Hoa22, Dij89z]. Forum [BSE^{+}90, WWSD90]. foundations [Maz76, Win78]. founded [Dij84-28]. foundedness [DvG90a, DvG90b, Dij91-32, vGD92]. four [Dij71b, Dij71c, Dij73m]. four-state [Dij73m]. Fourier [Dij82p]. Fourth [Dij75-62, Dij75-63, Dij76-85, Dij76-84, Dij77-44]. Framework [DCF^{+}02]. France [Dij73w]. Free [CMS92, Dij77-43]. freedom [BBM07c, Dij89-36]. freely [TZDM10]. Freeman [Dij77b]. French [Dij93f, Pet71]. Friend [Dij82z, Dij75-28]. friendly [Dij84-27]. friends [Dij63l]. Front [Ano80e]. Frontiers [TLR^{+}91]. fruits [Dij83g, Dij83h]. Fuji [HW06]. Fuji-Susono [HW06]. Function [Dij69c, Dij76i, Dij82-84, HZW09, Dij69d, Dij76-34, Dij82-30, Dij86z, Dijxxk, Kes88, MS14, Urb01]. Functional [BMWC01, BS99, Tse95, VU05, Dij53, HW06]. Functionele [Dij53]. functions [Bac90, Dij76-63, Dij85n, Dij86-31, Dij88b, DS90l]. Fundamental [Dijxx-56, dB76]. Fundamentele [Dij69f, Dijxx-56, Dijxx-89]. Funeral [Moo02]. Further [EM72, HHM94]. fusc [Dij76-34, Dij82-84]. Future [Lew21, Pay14, Bra90, Dij92e, Dijxx-60, Dijxx-89]. fuzzy [Muk12].

G [Dij86h]. G. [Dij81e]. Galois [Dij94p]. gambling [Dij89m]. game [Dij81d, Rei08]. Gap [Dij86x, Dij85w, Dij85x]. Garbage [Dij75-35, DLM^{+}76, DLM^{+}78, Dij75-41, DLM^{+}75c, DLM^{+}75a, DLM^{+}75b, DLM^{+}77a, DLM^{+}77b]. Garmisch [NR69]. Gas [CC12]. Gasteren [Dij80-29, Dij86h, DvG90a]. Gatlinburg [Dijxxe]. Gauntlet [Dij76-39, Dij75-36, Dij75-37, Dij76-38]. Gauss [LH90]. GCD [Dij01d]. Gdansk [Maz76]. geblokkeerde [Dij56d]. Gebruiksvoorwerp [Dij68l]. gegenereerd [Dij63n]. gehalte [Dij93e]. General [Ano80f, Dij77-43, Dij80-42, DS82c, Dij56c, Dij57c, Dij65c, Dij83c]. Generalization [Day11a, Day11b, Nel89, AVL02, CF04, Day12b, Dij85n, Dij90j, Dijxx-29, Knu77, PRV09, ST99]. generalizations [Nel92]. generalized [Dij99l, HHM94, Luk94]. Generalizing [Dij83i]. generated [Dij63n]. Generation [KKNI14]. Generators [Dij95i]. Genetic [JNK11, FC14, KYTY09]. Genius [Apt02]. Gentleman [Dij83j, Dijxx-100]. gentlemen [Dijxx-39]. Geodesic [MMP87]. Geometric [WWZ05, Dij92a, Dij94c, Dij96c, Dij96d, Dij00f, Dij00g]. geometrical [Dij95x]. Geometry [Dij98d, Dij82-48]. Georg [Dij99d]. georganiseerd [Dijxx-96]. German [BS01, Dij61a, DF85, Dij88u, SB01, Str72]. Germany [Bro96, NR69, Zus80]. gewicht [Dij71d]. gewoon [Dijxx-99]. Ghosh [DS76b, DS82i]. Gigabit [MGH^{+}05]. GIS [KLMZ11, XZY^{+}12, YL10, ZWKT10]. GIS-based [YL10]. global [AHHK93, AAEG11]. Globale [Dij65c]. GMPLS [PMR11]. Go [Dij68t, Dij96u, Dij01e, Dij02g, Dij21a, Knu74, RD68, Dij68d, Dij68s]. goal [BDn^{+}10]. goal-directed [BDn^{+}10]. Gödel [Heh90]. Goes [KKNI14]. going [Dij93r]. Golden [Dij77-36, Dij82-103]. Good [Dij82-100, Dij78-40]. Goodbye [Sch90b]. Goode [Dij74b, Dij82-70]. Goos [Dij71f, Dij72c]. Goto [Che02b, Che02a, Gal73, Nag18]. Goto-less [Che02b, Che02a]. gotspe [Dij80h]. graduate [Dijxx-91]. graduations [Dij75-51]. Grained [Dij82-89, Dij77q]. Grammar [DS04]. Graph [Dij82q, KKNI14, Nep96, BT79, Dij60c, Dij60a, Dij73d, Dij76w, Dij95c, Dijxx-27, Pet71]. graphe [Pet71]. graphically [PJ89]. Graphs [Dij87g, LW11, AVL02, Bar98, Dij59c, Dij77p, Dij80b, Dij80-27, Dij86v, Dij86w, Dij91-28, Dij99k, Dij00h, HSW08, MSS^{+}06, NMM78]. gratitude [Dij77q, Dij82-89]. Great [Dij82-78, Lap96, SL98, Dij76-28]. greater [Dij86r]. greatest [Dij90f, Dij96o, Dij97c]. greedy [CF04]. GREEN [Dij78-29, Dij78d]. Grenoble [Dij67g]. Grid [Dij02a, Dij88i, Dij89-37, ELCdCG14, HTVW08]. grids [Dij94h]. Gries [Ber99, Dij75-36, Dij75-37, Dij76-38, Dij76-39, Dij76-48, Dij82-97, Dij96-38, NB93, NBK95, Ter93]. Gries/Dijkstra [Ter93]. Group [Dij73s]. Guarded [Dij74e, Dij75x, Dij75w, Dij75y, Dij76-29, Dij96v, CRT78, HSM97, Nel92, Bai86]. Guided [Dij91j]. guiding [Dij70a]. Gupta [Dij86-33]. Gutknecht [Dij89v]. guy [Che16].

H [Dij74g, Dij79j, Dij79s, Dij79w, Dij86-30, Mil89]. H. [Dij76-70, Dij82-105]. Haantjes [Dij86y]. habit [Dij91c]. half [Dijxx-40, vdH17]. half-baked [Dijxx-40]. Hall [Dij98j]. Hamming [Dij76u, Dij81h]. Han [Dij81-28]. Han-sur-Lesse [Dij81-28]. hand [Dij82h]. Handboek [Dij55a, Dij55b]. Handling [Bai86]. Hanover [Dij65i]. hard [Dij84a]. hardnekkigheid [Dij82k, Dij82l]. Hardware [TS01, Dij77-46, Dij82-115]. Harmful [Dij96u, Dij01e, Dij02g, Dij21a, Nag18, Dij68t, Dij68s, DS72c]. Harmony [Dij82-86, Dij74j]. Harry [Dij74b, Dij82-70]. Harvard [Dij95-29]. Haskell [DS04]. Hauck [Dij75-33, Dij75-39, Dij82-85, Dij82-88]. HCTNav [ELCdCG14]. head [Dij78-35, Dij85n]. Health [Dij06]. Heap [ZJM12]. Heapsort [DvGF81, Dij81b, Dij81l, Dij81m, Dij81n]. heavy [Dij92-31]. heavy-weight [Dij92-31]. Heel [Dij65d, Dij68n]. Heer [Dijxx-30, Dij83j, Dijxx-100]. Heidelberg [Dij77b]. Heidelberger [Dij71f]. held [BS76, Bro96]. hell [Dij89-43]. Helper [NAGK09]. Helsinki [Dij83-27]. Hendriks [Dijxx-39]. Henk [Dij91r]. heren [Dijxx-39]. Heuristic [BMST03, MS01, Dij86l, Dij86m, Dij87h, Dij87i]. Heuristics [Dij94i, Dij94j, Dij95j, Dij94c, Dij94k, Dij96c]. Hierarchical [Dij69c, Dij71a, Dij72f, Dij76i, Dij02h, Dijxx-35, Par01a, BDn^{+}10, Dij69d, Dijxxk]. hierarchies [BS76]. Hierarchy [Dij82-104, HB94, Dij74s, Dij76-68, Dij76-69, Dijxx-87, UOZ11]. High [KP09, KP10, DFR74, Dij00j, ZLY13, Dijxx-37]. High-order [Dijxx-37]. high-school [Dij00j]. High-Throughput [KP09, KP10]. Higher [LSMS12, SWS^{+}13]. higher-order [SWS^{+}13]. Higman [Dij91-29]. hint [Dij96w]. Historic [BD01b]. historical [Kru03]. Historiography [May80]. History [Leh80, MHR80, Sch18, Sue80, Tat23, Tro80, Zus80]. Hm [Dij95k]. Hoare [Bar75, Bar74, Dij70d, Dij74a, Dij75-61, Dij77q, Dij82-89, Dij82s, Dij82-66, Dij89-28, Dij99h, Dijxxz, Hoa82, Llo74]. Hoe [Dij89r, Dijxx-36]. Hoger [Dijxx-37]. HOL [Dij75j]. hole [Dij86-40, Dij91-31]. holonomic [PG13]. Home [Dij08]. Homework [Dij96x]. HOMO [Dij75z]. Honors [Dij96y]. Honour [Dij79m, Dij78p]. Honsberger [Dij75-45, Dij81o]. Hoogerwoord [Dij91f]. hoogleraar [Dijxx-99]. Hop [Smi11]. Hop-by-Hop [Smi11]. Hopcroft [Dij76-40]. Hopes [Dij79r, Dij79q]. Hopper [Pay14]. HOS [Dij83k, Dij83l]. hot [Dij62d]. hour(R) [SEEM01]. house [Dij73k]. Huge [FJ09]. Hugo [Dij96-42, Dij01a]. hull [Dij76-50]. Human [Dij96-35, IdKM06, Dij65f, Dij65g]. Humble [Dij68a, Dij72h, Dij72g, Dij87j].

Humming [Dij82-86, Dij74j]. hun [Dij79-36, Dij79-37]. Hungarian [Dij80m, Dij80a, Dij89q]. Hungry [Dij89s]. Hurt [Dij82-79, Dij75-27, Dij82t]. Hybrid [PG13, LSJT12, OC10, Yap83]. hybridized [FC14]. Hydrogenation [Dij90a]. hygiene [Dij89-31].

i.e [Dij75-42, Dij76-45, Dij85-37]. I.U.C.C. [Dij73v]. I/O [BD79]. i860 [BDB93]. IBM [Dij63r, Dij73u, Dij82c, Hur80, Dij79f]. IBM360 [Dijxx-57]. ICEC04 [JSW04]. ICSE [IEE79a]. Ideas [Lew21, Dijxx-40]. Idempotence [Dij86p]. identical [Dij93j]. identifiers [Dij89-43]. IEEE [PS14]. IFIP [Dij73s, Dij74v, Dij82-109, Dij82-51, Dij65j, Dijxx-76]. IFIP-congres [Dij65j]. IFIP-congress [Dij65j]. IFIP77 [Dij77z]. II [Dij61a, Dij64c, Dij79-39, Dijxx-65]. IJzeren [Dij02f]. ik [Dijxx-99]. ILLIAC [Rob80]. Image [BDBR01, Dij79l, Dij79o, DvG82b, DvG82c, Dij85o]. immediate [Dij73e]. immune [AAEG11]. impact [Dij77z]. Implementation [Dij82-83, FURD13, KP09, ND00, UOZ11, Dij74h, Dij77-33, Dij78q, Dij83q, Dij83r, Dij83s, RR64, RPGS14, SZHT10, vdH17]. Implementations [GS97, Dij76-41]. Implementing [LX10]. implication [Dij88c, Dij88g, Dij92-28, Dij97j]. import [Dij54b]. importance [DvG84h, DvG84i]. Imposed [Dij82-104, Dij74s, Dij76-68, Dij76-69, Dijxx-87]. Impressions [Bac80, Dij85-43]. Improved [CC12, KLK08, LWQ09, SWX10, ZJM12, AHHK93, AHH^{+}95, zFmLcWcJ10, KF08, SxXq11, WM14, XZY^{+}12, XLH^{+}07]. Improvement [Dij82-78, FS10, KLMZ11, Dij76-28, Dij81l, Dij81m, Dij81n, Dij85p, xXL10]. improving [Dij78v]. In- [Dij54b]. In-Depth [Dij82u]. in/one [CCD86]. Inc. [Dij74f, Dij82w, Dij82-82, Tom80]. included [Dij94c, Dij96c]. including [Dij94w]. Incremental [Dij85q, Dij85-39, Dij85-40]. incurable [Dij95-31]. Independent [Dij63c, Dij61d, LvdS90]. Index [Ano80f, Ano80g, Ano80h]. Indexes [BM01]. Indirect [Dij01f]. inducing [Dij72d]. Induction [Dij80-37, Dij82-90, Dij76-43, Dij78-39, Dij82-28, Dij82-29, Dij87m, Dij90o, Dij92k, Dij96-27, Hoo90]. inductive [JM80]. industry [Dij85d]. inequalities [Dij86-28]. infima [Dij00c]. infinite [Dij80y, Dij82-47, Dij91v, Dijxx-48]. Inform [San05]. Informatica [Dijxx-38, Dij71d, Dij72e, Dij76-88, Dij80t, Dij93e, Dijxx-32, BB03, Dij72k, Dij72l, Dij75h, Dij80-40]. informatica-ingenieur [Dij80t]. informatica-opleiding [Dij71d]. Informatics [Dij72k, Dij75h, Sam76, Dij72l, Dij80l, Dij88u, Dijxx-38]. Informatie [Dij70g]. informatieverwerkende [Dij64j]. Informatik [Dij71f, Dij72c, Dij88u]. Information [Dij72i, Dij72j, Dij64j, Dij79t, Dij70g]. Ingenieur [Dij83j, Dij80t, Dijxx-100]. Inhomogeneous [Uch14]. inhoud [Dij56d]. Inleiding [Dij82v, Dij83w, Dijxx-54]. input [Dijxx-74]. insensitivity [Mar90]. Insights [van19]. Inspections [Fag01a, Fag01b]. Inspired [Dij82-73, Dij91k, Dij95m, Dij77g]. instantiated [Dij91p]. Institut [Dij88u]. Institute [Bro96, Dij87f, Dij90g, Dij90h, Big80, Dij88a, Dij88u]. integer [Dij88e]. Integers [Dij79-28, Dij82-102, Dij77-35]. Integrating [Anw03, KYTY09, LCL^{+}94]. Integration [DvdS06]. intellectual [Dij69g, Dij89-31, Dij93e]. intellectually [Ano02a]. intellectuele [Dij93e]. Intelligent [XWZH12, NB93, NBK95]. Interaction [Ula80]. interactive [BD79]. Intercomputer [Pad85]. interdiction [KGZ06]. Interface [Dij77b]. interfaces [BS76, BDR85]. Interleaving [Dij92g, Gou90]. intern [Dij63n]. internal [Dij63n]. International [HW06, IEE79a, IEE79b, IdKM06, BS76, BB79, BMW93, JSW04, VU05]. Internationally [Gol02]. Internet [yLsLtZ09, yLsLtZ10]. Interplay [Dij79b, Dij79v, Dij77s, Dij79c]. interpolation [DvW55]. interpretatie [Dijxx-59, Dijxx-81]. interpretation [DS84b, Dijxx-59, Dijxx-81, MHMO10]. intersection [KYTY09]. intersection-turns [KYTY09]. interview [Mis10]. intriguing [Dij80n]. Introducing [Dij86q, Dij87k, Dij92h, Dij93h, Dij95n]. Introduction [Dij82-80, DvG82f, Dij83w, Dijxx-54, Dij71g, Dij77l, Dij78q, Dij82x, DvG82d, DvG82e, Dij89e, Dij91o, Dij92n, Dijxx-79, Dij82v]. Introductory [MR80, DvG82g, Dij89u, Dij89-45]. Intuitionistic [Boh08, Muk12]. Invariance [Dij84c, Dij84d, Dij84e, Dij85r, Hoo90, dB76]. invariants [Dij93i]. invented [Dij98e]. Inventions [Ano80g]. inverse [Bau90]. Inversion [Dij79x, Dij82-98, Dij78-36]. Investigate [Dij57a, Dij57b]. investigation [Dij75k, Dij76x, Dijxx-68]. investigations [BBM07c]. involving [Pri90]. IP [KF08]. IPW [Dij83n]. ir [Dijxx-92, Dijxx-92]. IR-degree [Dijxx-92]. ir-diploma [Dijxx-92]. Ir. [Dijxx-39]. ISBN [Mil89]. isles [Dij00i]. isn’t [Knu90]. isolated [Dij76-53, Dijxx-70]. Isomorphism [LW11]. issue [AG01]. issues [BBM07c, Dij78q, FURD13]. Italy [BR70]. Item [DDG^{+}70]. iteration [Dij82-34, Dij99j]. Ithaca [Dij79-30, Dij85-45]. IVUS [BDBR01].

J [Bar75, Bar74, DLP78, Dij78-27, Dij78s, Dij79j, Dij79s, Dij79w, Dij80a, Dij81w, Dij85-48, Dij86h, Dij86y, Dij86-30, Dij87l, DvG90a, Dij95s, Llo74, Mil89, Dij89v]. J. [Dij77m, Dij77q, Dij82-89, Dij86h, Dij88f, Dij89h, Dij90d, DR90b, Dij94q]. jaarboek [Dij93t]. Jack [Dij75e]. Jacobson [Her90]. Jahre [Dij88u]. Jan [Dij79-33, Dij80-47, Dij81-28, Dij83x, Dij84u, Dij69j, Dij73k, Dij94k, Dij94m, Dij94l, Dij95o]. Jan.1985 [Dij85-47]. Jansen [Dij63o]. January [Dij73k, Dij76-73, Dij77-39, Dij81z, Dij69j, Dij74y]. Japan [HW06, Sue80]. Jayadev [Dij99d]. jc [Dij80t]. Jill [Dij99e]. Johan [Ano02c]. John [Dij76-64, Che16, Dij78-37, Dij94-27]. Johnny [Dij86-48]. Jonathan [Dij75-28, Dij82z]. Jonkers [Dij82-35]. Jonsson [Dij89-40]. Jordan [LH90]. Joseph [Dij92u]. Joshi [Dij96l]. journal [Dij70g, Dijxx-40, Dij79f]. Judging [Dij83k, Dij83l]. judgment [Dij64j]. juli [Dij71e]. July [BS76, BMW93, Bro96, Dij71e, Dij73t, Dij76-31, Dij76-78, Dij77-37, Dij78-49, Dij78-46, Dij79-31, Dij81-31, Dij82b, Dij82-54, Dij83-28, Dij84v, Dij86-45, Dij74a]. Junctivity [DS83a, DS83b, Dij94n, Dij88g, Dij92-27, Sch90b]. June [BMW93, BD01b, Dij67e, Dij76-78, Dij79-29, Dij82-55, Dij83y, Dij83-28, Dij84s, Dij84x, Dij85-45, Dij85-46]. juni [Dij67e]. just [Dijxx-99].

Justification [Dij63j].

K&CK [Mos10]. K. [DvG81b, Dij83f, Dij85i]. Kac [Dij95v]. kanalen [Dij56d]. kanttekening [Dij80i]. Kaplansky [Dij92p]. Ken [Dij91z]. Kernel [Rua90]. Keuze [Dijxx-41, Dij63j]. keuzecommissie [Dijxxb]. Khalil [Dij85y]. Khwarizmi [Tat23]. kind [Dij96z]. Kingdom [Mil89, Dij87d]. Kleene [HM12]. Kleinrock [Dij81r]. KMP [Mor90b]. knight [Dij92j]. Know [Ham80, Dij95x]. knowledge [Anw03, LCL^{+}94]. knowledge-base [Anw03]. knowledge-based [LCL^{+}94]. Known [Gol02]. Knuth [Dij77q, Dij82-89]. kodierten [BS01]. Komma [Dij54a]. kort [Dij65d]. Korte [Dij56b]. Kortste [Dij60c]. Kraskal [Nep00]. Kristen [Ano02c]. Kroeger [Dij80-39]. Kruseman [Dij95s]. Kruskal [Dij98h]. Kuipers [Dij91y]. kunnen [Dij89r]. kunst [Dijxx-54, Dij83w].

L [DS76a, Dij81e, Dij84l, Dij86k, Dij94k, Dij94m, Dij94l]. L. [Dij75-44, Dij83f, Dij85i]. laatste [Dij75q, Dij76l, Dijxx-43]. Labeling [Der85]. laboratory [Dijxx-72, Chu80, Dijxxy, Wil80b]. ladder [Dij97d, Dij97e]. l’Alpe [Dij72b]. lambda [Dij89f]. lambda-calculus [Dij89f]. Lambek [Dij80w, Dij81i]. Lamport [Dij83e, Dij83f, Dij85i, Dij85v, Mal19]. Lamsweerde [Dij77-28]. Lancaster [Dij78-47]. landscape [JSW04]. Language [Dij79u, NBB^{+}60, Wel80, Wir01b, vWMPK69, Dij62e, Dij73q, Dij73a, Dij74p, Dij75-38, Dij76-61, Dij78c, Dij78-28, Dij78d, Dij78w, Dij78e, Dij78f, HSM97, JM80, Bai86, BS76, Dij78z, Dij78-29, Dij78-30, Dij78-31]. Languages [Dij63c, KP80, MB97, Dij61d, Dij76-60, ŁMB02]. Large [BM01, Cod01, Dij76-56, GL98]. large-scale [GL98]. Largely [Dij81j]. last [Dij75q, Dij76l]. Late [Day11b]. lattice [Dij94-28, Dij98f, Dijxx-42, Tse97]. Laurens [Dij96-34]. Law [Dij95-29]. Laws [HB94]. Layer [LWQ09]. leads [SB04, Tur90b]. leads-to [SB04]. lean [Du10]. learned [DvG81d, Dij86y]. Learning [DH96]. least [HTVW08]. least-cost-path [HTVW08]. leaves [Dij78x, Dij80-32]. Lecture [Dij81k, Dij82y, Dij75b, Dij75c, Dij75q, Dij75o, Dij75p, Dij75r, Dij75t, Dij75s, Dij75-34, Dij75-59, Dij75-50, Dij75-58, Dij75-62, Dij75-64, Dij75-63, Dij75-65, Dij75-67, Dij75-66, Dij75-68, Dij76e, Dij76d, Dij76l, Dij76k, Dij76j, Dij76q, Dij76n, Dij76-36, Dij76-81, Dij76-67, Dij76-79, Dij76-80, Dij76-85, Dij76-87, Dij76-84, Dij76-86, Dij76-92, Dij76-93, Dij76-91, Dij77a, Dij77e, Dij77f, Dij77-42, Dij77-44, Dij77-45, Dij77-48, Dij77-49, Dijxx-95, Dijxx-98, Dij78-37]. Lectures [Ash87, VU05]. Led [Dij02d, Dij01o]. leden [Dij66b, Dij68b, Dijxxd, Dijxxc]. left [Dij82h]. Legacy [AH22, Che02b, Che02a]. Leibniz [Dij90m, Dij90l, Dij94b, Dij94u, Dij96z, Dij00k, Dij01n]. Lemma [Dij91-29, Dij98h]. lemmata [Dij85-39]. Length [LSMS12, BD89, Dij76-52, Dij77-34, DvG83a, Dij92f, ELCdCG14, NA13, Udd90]. lengthen [Dij89j]. lengths [Kun80]. Leonardo [Dij81g]. Leron [Dij89-41]. Leslie [Mal19].

Less [Dij79p, Che02a, Che02b, Dij76-35, Dij94o]. Lesse [Dij81-28]. Letter [Dij61b, Dij68s, Dij70d, Dij75-29, Dij75-30, Dij77m, Dij82z, Dij82-82, Dij82-99, Hoa82, Kil79, Knu66b, RD68, Van78, Dij74a, Dij75-28, Dij75-32, Dij75-44, Dij75-45, Dij75-61, Dij76-31, Dij76-64, Dij77-30, Dij78v, Dij78w, Dij80-41, Dij80-40, Dij80-38, Dij80-39, Dij81w, Dij82-66, Dij85s, Dij89-41, Dij91y, Dij92x, Dij96-38, Mau78, Dij74g, Dijxx-55]. Letters [BD01a, Che16, DCF^{+}02]. Level [Dij82-104, Dij74s, Dij76-68, Dij76-69, Dijxx-87, GS97, Dijxx-77]. levels [CF04]. lex [Dij66e]. lexical [Dij89k, DvG90a]. lexicographic [Dij79w]. liberal [DS82h]. Liege [Dij80-44]. Life [AH22, Dij93d, Dij99g, LSJT12]. Lifting [Dij94p]. Like [Der85, Dij75a, Kun80, SNR15]. Limitation [HZW09]. Limitations [Kos73, Mar90]. limited [KGZ06]. Line [SWW99, Dij62d, DvG81c, Dij85-36, DS90i, Dij90r, SWW00]. Linear [Bri90, Dij76-32, DF89, Dij78n, Dij85-34, DF88b, HHM94, Dij85m]. Linearization [Dij82-27]. lines [Wir90]. Linial [Ano13]. Linkages [Svo80]. links [Lee06]. Lipton [DLP78, Dij78u, Kil79, Van78]. List [Ano80i, D^{+}66b, HB94, MJC07, Dij79w]. listener [Dijxxh]. lists [Tur90a]. Literate [Knu92]. little [Dij84f, Dij96k, Dij96-40, Dij98f, Dijxx-42, Fei90]. Liu [Dij92u]. Liveness [Sch18, Rem90]. Lives [SL98]. load [DK13]. localized [MTZ12, vL90]. locating [Dij72l]. location [ZTZY12]. locations [Dij64g]. Logic [HW06, BES95, Boh08, DD98, Dijxx-29, Dij80q]. Logical [Dij76p, Har95, Dij76o, Dij77v, Dijxxp]. logician [Dij90m]. logische [Dijxxp]. lollie [Dij84a]. lolly [Dij84a]. London [Bar75, Dij75-54, Dij78-47, Dij79-32, Dij84u, Dij84t, Dij84s, Dij85-48, Dij85-46, Boo80]. long [MGH^{+}05]. longer [Dij86r]. Longest [DMS83, Dij85t, Kal90, Dij85-31]. Look [Par18, Dij01a]. Lord [Dijxx-30]. losing [Dij80-49]. Lowest [ZXZ^{+}09]. Lt. [Dij78w]. Luxembourg [Dij74u].

M [Dij76v, Dij80-31, Dij80-41, Dij82-76, Dij83f, Dij85i, Dij86h, DvG90a]. M. [Dij77-28, Dij85y, Dij92c]. M.C. [Dij63n]. M.C.-translator [Dij63n]. M.C.-vertaler [Dij63n]. maart [Dij73x]. Machine [Dij63c, BS01, Dij61d, Dij64j, Dij67f, Dij79o, Dijxxp, Lam22]. machines [Dij73m, Dij73n]. macros [Dij91m]. made [Ano02a]. MAES [Dij77l, Dij82x]. MAES(R) [Dij82-80]. Mainly [Dij80o, Dij85u]. Maintenance [BM01, CCD86, Har95]. majority [Dij89v, Dij91l, Ein90]. Making [Dij63i, Dij82-77, Dij82-89, Dij89w, Dij89x, Dij90n, Dij75v, Dij77q, Dij61c]. Man [Alp18, OC10]. Management [Dij82u, Dij86s, BDS82, Dij63p, Dij64h, Dij04, Dijxx-62]. Manchester [Lav80]. MANIAC [Met80]. Manipulating [Dij91m]. Manna [Dij76-31]. Manual [Dij55a, Dij55b]. manuals [Dij76-41]. manufacturer [Dij85s]. manufacturers [Dij63g, Dij63d, Dij63e, Dij63h, Dijxx-55]. manuscript [Dij91n]. Many [BMST03, MS01, Dij75d]. maps [ELCdCG14]. March [Dij75-54, Dij77-38, Dij82-107, Dij82-58, Dij83-27, Dij85-49, Dij87y, Dij73x, Dij98d]. marked [Dij97f]. Marketing [Dij75-31]. Marktoberdorf [BS76, Bro96, Dij75-55, Dij78-46, Dij81-29, Dij84v, Dij86-45, Dij88r, BB03, Dij71e]. marriage [Dij95p]. Martin [Dij78v, Dij77t, Dij78-27, Dij78s, Dij80a, Dij82-93]. massaging [Dij94n]. massive [HTVW08]. Matching [BMST03, MS01, Dij76-47, Dijxx-67, TZDM10]. materials [Du10]. Mathai [Dij92u]. Mathematical [Dij74m, Dij80-37, Dij82-28, Dij82-29, Dij90o, Dij92k, Dij94v, Dij96-27, Dij97h, Dij72p, Dij73j, Dij77h, Dij78m, Dij78-39, Dij79-40, Dij80k, Dij80s, Dij80-43, Dij80z, Dij85-41, Dij85-42, Dij86q, Dij87o, Dij87m, Dij89-35, Dij90k, Dij91a, Dij91m, Dijxx-36, Dijxx-38, Hoo90, Dij87k, Dij89o, Dij89u, Dij89a, Dij89-35, Dij89-45, Dij90l, Dij91g, Dij91p, Dij91-31, Dij96y, Dij97g, Maz76, Win78]. mathematicians [Dij80-49, Dij88p]. Mathematics [Dij75-32, Dij78r, Dij79b, Dij79v, Dij82-31, Dij82-81, Dij90i, Dij92s, Dij93l, Tat23, Ula80, Dij67a, Dij74i, Dij76-88, Dij77s, Dij77n, Dij79c, Dij92r, Dij98k, vWBD56, BMW93, Dij74f, Dij80u, Dij80-43, Dij82w, Dij82-82, Dij84-27, Dij95k, Dijxxc]. Mathemetics [Dij86s]. matrices [Dij97i]. Matter [Ano80e, Dij95t, EBR09]. maturity [Dij88j, Dij88k, Dij88u]. Maximal [Are90, Dij76w, Dijxx-27]. maximization [MS14]. maximize [KGZ06]. maximizing [Dij83o]. Maximum [Dij82q, Dij73d, Dij76-52, Dij77-34, DvG83a, Udd90]. May [Ano02b, Ano12, Dij65j, Dij68r, Dij72n, Dij74x, Dij74w, Dij75-52, Dij80e, Dij81-30, Dij82-106, Dij82-49, Dij82-55, Dij84y, Dij84z, Dij85-45, IdKM06, TLR^{+}91, Dij79-38, Dij79-39]. Mazola [Dij75e]. McCarthy [Dij76-64, Dij82-30]. McDonald [Sha81]. McIlroy [Dij87p, Dij89-27]. me [Dij80-31, Dij80-29, Dij92c]. mean [Dij92a, Dij94c, Dij96c]. meaningful [Dij89-43]. means [Dij96d, Dij00f, Dij00g]. Measurement [CCD86]. Measuring [vDO01]. Mechanical [Svo80]. Mechanism [Dij90a]. Medema [Dijxx-39]. Meditations [Dij62f, Dijxx-83]. Medium [Uch14]. meet [HM12]. Meeting [Dij74x, Dij74v, Dij82-109, Dij82-51, Dijxxd, Dijxx-43]. mei [Dij65j, Dij68r, Dij72n]. member [Dij78-42, Dij70c]. members [Dij66b, Dij68b, Dij87x, Dijxxd, Dijxxc]. Memorial [Che02b, Dij74b, Dij82-70]. Memorials [Ano06b]. Memoriam [Dij94l, MR03, BFG^{+}02]. Memories [Dij80-30, Slu80, Dij76-57]. Memory [NAGK09, ANGK09, BDS82, BW00, Dij75f, Dijxx-34]. Menelaos [Dij90p]. Merging [Dij76-33]. mesh [DK13]. Message [D^{+}66b]. messages [Dij79-38, Dij79-39]. Method [Der85, Dij57a, Dij02a, Sha81, AAEG11, BS01, Ber99, Dij57b, DF84, DF85, DF88a, Mil89, NS00, OC10, SNR15, SS10, YL10]. methode [DF84, Bov86]. Methodical [Nel90]. Methodik [DF85]. methodological [Dij80p, Dij87m, Dij88i]. methodologies [Dij72m, Dijxx-73]. Methodology [Dij73s, Dij76-70, Dij82u, Dij82-105, Dij86q, Dijxx-47, Dijxx-90, NB93, NBK95, Dij87k, Dij89o, Dij89u, Dij89a, Dij89-35, Dij89-45, Dij90l, Dij91g, Dij91p, Dij91-31, Dij96y, Dij97h]. Methods [DH96, Jac01]. Michael [Dij66b]. microgrid [UOZ11]. microprocessor [Dij84a]. microprocessors [Dij77z]. Middle [Dij77w, Dij78u]. Might [Dij82-79, Dij75-27, Dij82t]. Mijn [Dijxx-43, Dij70i, Dij75c, Dij75p, Dij75s, Dij75-63, Dij75-66, Dij75-68, Dij76e, Dij76d, Dij76l, Dij76k, Dij76j, Dij76q, Dij76n, Dij76-36, Dij76-81, Dij76-67, Dij76-79, Dij76-80, Dij76-85, Dij76-87, Dij76-84, Dij76-86, Dij76-92, Dij76-93, Dij76-91, Dij77a, Dij77e, Dij77f, Dij77-42, Dij77-44, Dij77-45, Dij77-48, Dij77-49, Dijxx-95, Dijxx-98]. MIL [Dij80e]. MIL-STD-1862 [Dij80e]. mild [Dij80q]. milieu [Dijxxp]. Mills [Dij76-70, Dij82-105]. MIMO [KP09, KP10, MJC07]. Minds [SL98]. minimum [Dij79w, Dij01d]. minor [Dij81l, Dij81m, Dij81n]. Minority [DDG^{+}70, WTD^{+}69]. Minsegsumtwodim [Dij84g, Dij84h]. Mintree [Obr78, WT79]. MIRFAC [Dij64a, Dijxx-82, Gaw64]. miscellany [DS84a]. misconceptions [BN98]. misery [Dij95-32]. misguided [Dij80r, Dij85a]. Misra [Dij83e, Dij85v, Dij87l, Dij94q, Dij95q, Dij99d]. missed [Dij85-51]. Mission [Dij79-31]. mistake [Dij99m]. misunderstanding [Dij83g, Dij83h]. mix [Dij94g]. mixed [Dij79-35]. mobile [AAEG11, LSJT12]. mobility [PS14]. Modalities [Hes90]. Model [Che01, Cod01, LWQ09, PMR11, PRC12, Dij04, GB96, PRD14]. Modeling [Ban07, GLFM07, Dij72p]. models [DS86, DvdS06, GL98, HSM97, PS14]. modelvorming [Dij72p]. modest [Dij86v, Dij86w, Dij87g]. Modified [DMS83, FURD13, KP09, KP10, Lee06, SS10]. modularity [Bro90b]. Modules [Par01b, BDR85]. moeilijkheden [Dij79-36, Dij79-37]. Moerland [Dij92t]. moet [Dij76-88]. mogelijkheden [Dij79-36, Dij79-37]. Molecular [BDB93, ITOK04]. monad [SWS^{+}13]. money [Dij77v, Dij79y]. Monochrome [Dij89y]. Monotonic [Dij74h, Dij82-83, Dij91o, Bau90, Dij82-32, DvG83a, Dij92n]. Monotonicity [Dij84i, DvG84a, Dij86p, Dij87n, Dij88g, Dij96w]. Monthly [Dij98d, Dij94v]. moral [Dijxx-44]. Morawetz [Dij92x]. Morgan [Tur90a]. Morley [Dij89-33, Dij92q]. mos [Hoa90]. Moser [Dij80w, Dij81i]. Mosquitoes [Dij82-86]. mosquitos [Dij74j]. most [Dij76-53, Dij93m, Dij93n, Dij94o, Dij96-33, Dijxx-70, Dijxx-79]. mother [Dij81o, Dij83m]. motion [PG13]. Motivering [Dij63j]. MPLS [PMR11]. MPMA [DLMB99]. MR [Dij82o]. Mr. [Dij79g]. Ms. [OC10]. much [Dij87o, Dij91x, Dij91-27]. Multi [BDBR01, Dij66a, DS82b, LCZW14, DS72a, DS72b, EKP96, FC14, GS97, MP90, Dij65d]. Multi-agent [BDBR01]. multi-level [GS97]. multi-objective [FC14]. multi-process [MP90]. Multi-programming [Dij66a]. Multi-Record [DS82b, DS72a, DS72b]. multi-source [EKP96]. Multi-step [LCZW14]. Multicast [PS14, MTZ12]. Multidisciplinary [Dij82-31, Dij74i]. multilevel [HSW08]. multiple [PG13]. multiple-trailer [PG13]. multiplicity [Dij95r]. multiprocessing [Dij75-41]. multiprogrammed [Dij73b]. Multiprogrammering [Dij63k, Dijxx-45, Dijxx-46]. multiprogrammeringsproject [Dijxx-69]. Multiprogramming [Dij83v, Dij01k, Dij02j, Dij21c, Dij65e, Dij68p, Dij68q, Dijxx-69, Dij63k, Dijxx-45, Dijxx-46]. multiprograms [Fei90]. München [Dij81-30, Dij84w]. Munich [Dij73t, Dij74v, Dij76-74, Dij79-32, Dij82b, Dij82-109, Dij82-54, Dij82-51]. must [Dij76-88, Dij79-40]. mutual [BW00, CSZ10, Dij83q, Dij83r, Dij83s, Dij85v, Dij88g, PJ89]. My [Dij63l, Dij79q, Dij79r, Dij81o, Dij82z, Dij83m, Dij88i, Dij96-28, Dij01g, Dij05, Dij70i, Dij75c, Dij75p, Dij75s, Dij75-28, Dij75-63, Dij75-66, Dij75-68, Dij76e, Dij76d, Dij76g, Dij76l, Dij76k, Dij76j, Dij76q, Dij76n, Dij76-36, Dij76-81, Dij76-67, Dij76-79, Dij76-80, Dij76-85, Dij76-87, Dij76-84, Dij76-86, Dij76-92, Dij76-93, Dij76-91, Dij77a, Dij77e, Dij77f, Dij77-42, Dij77-44, Dij77-45, Dij77-48, Dij77-49, Dij78-35, Dij79y, Dij84l, Dij86c, Dij86d, Dij86e, Dij86t, Dij87k, Dij93d, Dij99g, Dij02f, Dijxx-95, Dijxx-98, Dijxx-43].

Naar [Dij77o, Dij80t, Dij70i, Dij71h, Dij72r]. Naive [Dij82-116, Dij77-47, Dij78-50]. najaar [Dij75b, Dij75q, Dij75o, Dij75r, Dij75t, Dij75-34, Dij75-59, Dij75-50, Dij75-58, Dij75-62, Dij75-64, Dij75-65, Dij75-67, Dij76d, Dij76j, Dij76n, Dij76-80, Dij76-84, Dij76-86, Dij76-91]. Names [Ano80h, Dijxxn]. naming [DvG86a]. Nanongkai [Ano13]. Nati [Ano13]. National [Chu80, Dij76-51, Dij77g, Dij82-73, Swi11, Hus80, Wil80b]. Native [MGH^{+}05]. NATO [Bro96, Dij75-55, BR70, Dijxx-97, Dijxx-96, NR69, NRB76]. natural [Dij78-28, Dij79u]. Nature [Dij74m, Dij72m, Dij73j, Dij77h, Dij78m, Dij78t, Dij84k, Dij85-28, Dij86t, Dijxx-73]. navigation [PPS^{+}13]. necessary [Dij91b]. necessity [Dij91j]. need [Dijxx-90]. Needs [Dij77b]. negative [Kun80, NA13, Yap83]. Negende [Dij75-34, Dij76-36]. neglected [Dij87o]. nephew [Dij84l]. net [Kot90]. Netherlands [Dij85-47, IdKM06, JSW04, Sam76]. nets [Kos73, Pop95]. Netty [Dij80-29, Dij88h, Dij01f]. Network [CC12, HCSW09, JNK11, ZTZY12, DK13, Ein90, GL98, OC10, SS10, XLH^{+}07, YJ10]. Network-based [ZTZY12]. Networking [Pad85, JHLC14]. Networks [Dij79-27, GLFM07, KKNI14, MTZ12, Dij77-32, KF08, Lee06, LCZW14, Muk12, SVSR04]. Neumann [Ula80]. neural [OC10]. Newcastle [Dij74t, Dij75-56, Dij77-40, Dij78-47, Dij80-46, Dij81-31, Dij81-32, Dij82c, Dij82-110, Dij82-111, Dij82-52, Dij82-60, Dij82-61, Dij73u]. Newcastle-upon [Dij82-52]. Newcastle-upon-Tyne [Dij77-40, Dij80-46, Dij82-111, Dij82-61]. Newport [Dij85-45]. News [Coc02, DDG^{+}70]. Next [DM97, Dij76-54, KKNI14, Dij89z, Dij96-29, Dij96-30]. nice [Dij82-32]. niet [Dijxx-99]. Niklaus [PGB00]. Nils [Dij81w]. Nilsson [Dij81w]. Ninth [Dij75-34, Dij76-36]. No [Bar75, Dij63h, Dij82o, Dij82-48, Dij92u]. Noble [Dij83j, Dijxx-39, Dijxx-100]. Node [CC12, Dij86u, Dij95r, KGZ06]. node-wise [KGZ06]. nodes [Dij78x, Dij99k, Yap83]. nomenclature [Dij81j]. Non [Dij76-29, Bau90, Dij73f, Dij74d, Dij74e, Dij75x, Dij75w, Dij76-42, Dij76-37, Dij77-30, Dij95r, Dijxx-49, PG13, SNR15, Tur90b, Dij82-87]. Non-determinacy [Dij76-29, Dij74e, Dij75x, Dij75w, Dij76-42, Dij77-30, Dijxx-49]. non-determinism [Dij74d]. non-dominated [SNR15]. non-holonomic [PG13]. non-monotonic [Bau90]. non-stepping [Tur90b]. non-unique [Dij95r]. nondegenerate [Dij88n]. Nondeterminacy [Dij75y, Dij82-74, Dij82-99, Dij78l, Dij82n, Dij84c, Dij84d, Dij84e, DvG84g, Dij85r, Hes90]. Nondeterminancy [Dij96v]. nondeterminism [Dij86-32]. Nondeterministic [Dij96-31, Har81]. nonlinear [GL98]. Nonoptimal [vDO01]. noodzaak [Dijxx-90]. Noord [Dij71h]. North [Dij71h]. notation [Dij78m, Dij81x, DvG82h, DvG85a, DvG86b, Dijxx-74, vGD86]. Notational [Dij92o, Dij02e, Dij80v, Dij89-36, Dij00b]. Note [BB99b, BB99a, Joh73, Bar98, BDR85, Dij59c, Dij76-63, Dij76-90, Dij80i, Dij80-35, Dij82-33, Dij86-38, Dij93i, Kru03, MZS02, Per90, Dij63m]. Notes [BB99b, Dij69a, Dij69e, Dij70e, Dij70f, Dij72a, Dij75-35, Dij01o, Dij81k, Dij82y, Dij02d]. Notice [Du10, zFmLcWcJ10]. notion [Dij76r, Dij86z, Dij87v]. notions [DS82e]. Notitie [Dij63m]. Nov [Dij76-71, Dij78-48, Dij78-45, Dij80-44, Dij81-33, Dij83z]. novel [AAEG11]. November [Dij76-74, Dij65d]. NPL [Dij65d]. Nuenen [Dij85-48, Dij89b]. number [Dij76-56, Dij76-89, Dij80-48, Dij94v]. numbering [Dij82-68]. numbers [Dij81g, Dij82-47, Dij88h, Dij01h]. numerical [vWBD56]. numerieke [vWBD56]. NumerMath59 [Dij82-37]. Nygaard [Ano02c].

O [Bar75, Dij92c, Llo74, BD79]. O. [Bar74, DS76a]. O.-J [Bar75, Llo74]. Oak [Hou80]. Obituaries [Ano06b]. Obituary [Ano02c, Ano12]. Object [GHJV01, Dij64c, Dij87o, Dijxxq, Dijxxr, Dij63n]. Object-Oriented [GHJV01]. objective [FC14]. Objectives [Dij77b, Dij72m, Dijxx-73]. Objectprogramma [Dij63n]. obtention [Pet71]. Occasion [Dij82-44, Dij74r, Dij82s, TLR^{+}91]. Occupation [Dij64b]. ocean [DvdS06, Dij77r, Dij82-91]. Oct [Dij76-77, Dij78-48, Dij81y, Dij82-114, Dij82-56, Dij82-63, Dij83z, Dij88s]. October [BR70, Dij62c, Dij73s, Dij74g, Dij77-41, Dijxx-96, JSW04, NR69, Dij94v]. Odd [Der85, Dij79-28, Dij82-102, Dij77-35]. Odd/Even [Der85]. OFDM [MJC07]. offensive [Dij77o]. offloading [LCZW14]. Oh [Wei88]. Oils [Dij90a]. oktober [Dij62c, Dijxx-96]. Old [Dij82z, van19, Dij75-28, Dij83i, vdH17]. Ole [Ano02c]. Ole-Johan [Ano02c]. omarming [Dijxxu]. omission [Dij80o, Dij94t]. On-Line [SWW99, SWW00]. On-the-Fly [DLM^{+}76, DLM^{+}78, Dij75-41, DLM^{+}75c, DLM^{+}75a, DLM^{+}75b, DLM^{+}77a, DLM^{+}77b]. onbelangkrijk [Dij89r]. Once [Dij89-37, Dij00h, Dij85q, Dij88d, Dij89g, Dij90e, Dij92z, Dij95v, Dij96b, Dij96d, Dij96-43, Dij97l, Dij99c]. once-deep [Dij88d, Dij89g]. onderafdeling [Dij76-88, Dijxxc]. onderafdelingsvergadering [Dijxxd]. onderwijs [Dij69f]. onderzoek [Dij69f]. One [Gri90, Dij75-40, Dij78y, Dij80-36, Dij01l, HKS93, CCD86]. One-in [CCD86]. One-in/one-out [CCD86]. one-out [CCD86]. one-sided [Dij75-40]. one-to-one [HKS93]. online [Rei08]. Only [Dij95t, Dij72d, Dijxxm]. ons [Dijxx-30]. ontology [TZDM10]. ontstaan [Dij82k, Dij82l]. Ontwerp [Dijxx-54, Dijxx-56, Dijxx-55]. Ontwerpschets [Dij62c]. ontwikkeling [Dij77c]. Ontwikkelingsplan [Dij72k, Dij62c]. oordeel [Dij64j]. op [Dij68f, Dij75h, Dij77c]. opdrachtencode [Dij56b]. Open [Dij82-99, Hoa82, Dij75-44, Dij75-45, Dij75-61, Dij76-64, Dij77-30, Dij78v, Dij78w, Dij80-41, Dij80-40, Dij80-38, Dij80-39, Dij81w, Dij82-66, Dij89-41, Dij91y, Dij92x, Dij96-38, Dij73k]. OpenFlow [FURD13]. Operating [Bri01, BD66, Dij62a, D^{+}66c, D^{+}66b, D^{+}66a, Dij66f, Dij66e, Dij68e, Dij69h, Dij01g, Dij05, Dijxxe, HP72, Dij68r, Dij69j, Dijxx-85]. operation [Dij85-29]. Operational [DS90f, Luk94, DS84b, JH93]. Operator [D^{+}66b, Dij90e, Dij99p, SB04]. operators [Dij87r, Dij88l, Dij94g]. opleiding [Dij71d, Dij80t, Dijxx-61]. opposite [Dij86r]. Optical [Uch14]. optimal [AAEG11, BDS82, KYTY09, OC10, WM14, YL11, ZLY13]. Optimization [HZW09, KLK08, ZXZ^{+}09, zFmLcWcJ10, YJ10]. Optimized [WLXX09, SZHT10, ZWKT10]. orde [Dijxx-37]. order [BES95, Dij92-27, Dijxx-37, SWS^{+}13]. Ordered [BM01]. Ordering [BBM07b, Dij69c, Dij71a, Dij72f, Dij76i, Dij02h, Dij69d, Dijxxk, Dijxx-35]. orders [Dij89d, Dij94p]. ORDVAC [Rob80]. Organisatie [Dij69f]. Organization [BM01, Dijxxv, Dij69f]. organized [Dijxx-96]. Oriented [GHJV01]. Origin [Bri02]. Original [BD01b]. Origins [Ran80b]. OS/360 [BMWC01]. Other [Pad85, Dij74c, Dij75l, Dij85t, Dij99e]. otherwise [Dij89-31]. Our [DvG81b, Dij87s, Dij94t, Dij80-33, Dij82-38, Dij85u, DS90d, Dij91w, FvGGM90, DvG83b]. outlook [Dijxx-89]. output [PRM11]. overlay [HSW08]. overrated [DvG84h, DvG84i]. Owicki [Dij82-97, Dij76-48]. Oxford [BMW93, Dij79-33, Dij91t].

P [Dij88f, Dij89b, Dij89h, Dij90d, Dij94-29, Dij77-33]. P/V [Dij77-33]. P/V-implementation [Dij77-33]. P1400 [Dij71b, Dij71c]. P880 [Dij71b, Dij71c]. Pac [OC10]. Pac-Man [OC10]. Packet [JNK11]. Page [Dij08, Dij63q, Dij64h, Dijxx-62]. pages [Bar75, Dij64b]. pagina [Dijxx-62]. pagina-administratie [Dijxx-62]. paginaadministratie [Dij64h]. painting [Dij95h]. Pair [DCF^{+}02, Dij95-32]. Pairing [ZJM12]. pairs [Dij89y, MT84]. palindrome [Udd90]. Palo [Dij87y]. Pamphlet [Dij78u, Dij77w]. Panel [Dij76-70, Dij82-105]. Paoli [Dij77-39]. paper [Dij77z, Dij77x, Dij77y, Dij78u, Dij83c, Dij84n, Dij87t, Dij88o, Dijxxv, Jef78, SL88, Dij92u]. paperback [Mil89]. Papers [Dij63b, Lap96, Lew21]. parabel [Dijxx-66]. Parable [Dij82-96, Dij76-46]. parade [Dij94y]. Paradigm [DCF^{+}02, Dij82-30]. Paradox [Dij85-52]. Parallel [ND00, AdBO90, Dij77-28, JAE^{+}12, Nep00, SPPR92]. Parallelism [DS72a, DS72b, DS82b]. parallelizable [Gou90]. Parallelization [CMMS98]. parallellism [Dij73b]. Parametric [KM83]. parentheses [Dij80o]. Pareto [FC14]. Parijs [Dij67g]. Paris [Dij67g]. Parnas [Kil79]. Part [Dij87b, Dij87c, Dij90c]. Particle [KLK08]. partition [Dij99l]. Partitioned [MCX13]. Partitioning [Dij80-27, Dij86-31, Dij86-32, MSS^{+}06, Dij84j]. partitions [Dij81p]. Pascal [BD79, Wir01b]. Passes [Coc02]. Path [BB99b, BB99a, CMMS98, ELCdCG14, FJ09, GLFM07, HCSW09, Joh73, KLK08, ND00, TS01, AAEG11, CF04, Cha10, CH10, Dij98a, Du10, FC14, HTVW08, HKS93, HSWW05, HSW08, JAE^{+}12, JHLC14, KGZ06, Koh76, Kun80, LSJT12, Mis01, MT84, Muk12, NA13, NMM78, Nos85, NS00, PRV09, SNR15, SS10, SxXq11, Trä95, WWZ05, WYY11, xXL10, XLH^{+}07, Yap83, ZLY13]. path-planning [WYY11]. Paths [Der85, MP91, Smi11, FNS03, Pet71]. Patrick [Dij94w]. pattern [Dij73g, Dij76-47, Dijxx-67]. pattern-matching [Dijxx-67]. Patterns [DS82a, GHJV01, SD82, DS77, DS82d, Dij95s]. Paying [Dij77v]. PC [Coc02]. PEARL [Dij82-39]. pebble [Rei08]. people [Dij91b, Dij99e]. Perfect [Dij82p]. Performance [GLFM07, AHHK93, AHH^{+}95, CSZ10, JAE^{+}12, PS14]. performance-driven [AHHK93, AHH^{+}95]. performing [BS01]. perhaps [Dij95-27]. period [Dij83a, Dij83t]. Perlis [DLP78, Kil79, Van78]. Permutation [Dij76-54]. Perry [Dij92t]. persistence [Dij82k, Dij82l]. person [Dij81d]. Personal [Bac80, Dij72o, Dij82-97, Dij82-69, Tro80, Dij76-48, Dij82-35, Har99, Rem90]. Perspective [Dij72o, Dij82-69, May80, Rem90]. Persuasive [Dij06, IdKM06, IdKM06]. Peter [Dij77b, Dij91k]. Peterson [Dij81e]. Petri [Kos73, Pop95]. Petri-nets [Pop95]. phenomenon [Dij86-27]. Philadelphia [Dij78-45]. Philips [Dij01i]. Philosophers [BBM07a]. Physical [Wil80b]. Physicist [Mar02]. picture [Dij93g]. pictures [Dij99h]. piece [Dij77o]. Pierre [Dij76-76, Dij82-113, Dij82-53]. Pierre-de-Chartreuse [Dij76-76, Dij82-113, Dij82-53]. pigeon [Dij91-31, Dij86-40]. pigeon-hole [Dij91-31, Dij86-40]. Pilot [Wil80b]. Ping [Dij94u, Dij94b, Dij95a]. Ping-pong [Dij94u, Dij94b, Dij95a]. Pioneer [Gol02]. Pioneering [Ano02a, Goo80]. Pioneers [BD01b, BD02, Wei88, BD01b]. Pittsburgh [Dij77-39]. plaatsbepaling [Dij72l]. Plan [LW11, Dij62c, YL10]. Planar [MP91]. plane [Dij89y, Dij95u, Dij95h, Dij99a, Pos90]. Planning [KLK08, AAEG11, Du10, ELCdCG14, FS10, FC14, LSJT12, PG13, RPGS14, WYY11, YJ10, YL11, zFmLcWcJ10]. plans [DvG83b]. plateau [Dij85t]. platforms [PG13]. played [Rei08]. players [Dij78-34]. plight [Dij80c, Dij81c]. Plotting [Dijxxa]. plus [Bov86, Pet71]. Poging [Dij72l]. Point [Dij54a, Dij65c, DvG84b, DvG84c, DvG84d, DvG84e, DvG85b, Mis01, NF95, dB76]. point-to-point [NF95]. pointless [Dij91q]. Points [Dij95u, Pos90]. Poland [Dij76-75, Dij82-112, Dij82-62, Win78]. Policy [Dij68b, Dij80u, Dij67a, Dijxxa]. Political [Dij78u, Dij77w]. polygons [Dij78g, Dij88i]. pong [Dij94b, Dij94u, Dij95a]. ponsband [Dijxx-59, Dijxx-81]. ponsbandorganisatie [Dijxxv]. ponskaarten [Dijxxf]. Portland [Dij80-45]. Portrait [Apt02]. posed [Dij79s, Dij85y]. Position [Dij77z, Dij87t, Dij88o, Dij77x, Dij77y, Dij83c, Jef78, SL88]. possibilities [Dij79-36, Dij79-37]. possibly [Dij89w, Dij89x, Dij90n]. Post [Eds82, Dij75-46, Dij80t, Dij82-36, DvG81b]. post-propadeuse [Dij80t]. Post-Scriptum [Eds82, Dij75-46, Dij82-36]. postcondition [DS90k]. postfix [Dij94g]. postscript [Dij80-28]. potentially [DFR74]. Potter [Dij91u]. pour [Bov86, Pet71]. power [PRM11]. Powers [Dij79-28, Dij82-102, Dij77-35]. Practice [CK02a]. pragmatic [Dij78-33]. pre [Dij94g]. pre- [Dij94g]. Precision [Dij02a]. precondition [BBM07a, BBM07c]. preconditions [DS82h, Luk94]. predecessor [Dijxx-79]. Predicate [Dij83p, Dij85-30, DS90g, BES95, BN98, DS82c, Dij82-32, Dij82s, DvG83a, Dij85u, DS90b, DS90j, Dij92n, Dij82y, Dij89-38, Bör94a, Bör94b]. predicates [DvG82a, Dij82s, Dij84j, Dij86-31, Dij86-32, DS90c]. Preface [Ano80j, Dij76-49, Dij89a]. Prehistory [Zem80]. preliminary [Dij75k, DLM^{+}75c, Dijxx-68]. Prelis [Dij78u]. preorders [Dij91-33]. presence [Yap83]. presentation [Are90, DvG81a, DvG81b, Dij93h]. Press [Bar75]. Pretense [Dijxx-69]. Pretentie [Dijxx-69]. prevent [Dijxxu]. Price [Bar75, Mil89]. Prim [AHHK93, AHH^{+}95, Nep96, Nep00]. Primality [Dij57a, Dij57b]. prime [Dij76-56, Dij88h, Dij93m, Dij93n, Dij96-33]. primer [Dij62e]. Primes [Bul15, Dij95v]. primitives [Dij73e, Dij73p, Hor90, Vv72, Kos73]. principle [Dij70a, Dij91-31, Dij94b, Dij94u, Dij86-40, Dij90l, Dij96z]. Principles [Ano13, Dij06, Tur90a, Dijxxe]. printer [Dijxxa]. priority [MT84]. Private [Dij82-82, Dij75-32]. Prize [AMM^{+}12, Alp18, Ano13, Mos10, Sch18]. Probabilistic [HSM97, GGAVGGRC15]. probably [DFxx]. Problem [DMS83, Dij65h, Dij68h, Dij82a, Dij83u, Dij94v, Dij96-37, Dij98d, Dij01j, Dij02i, Dij21b, Dijxx-71, Lam74, MMP87, MP91, Sha81, Swi11, BBM07a, Bau90, Dij67b, Dij68g, Dij73r, Dij76-40, Dij76-50, Dij76-51, Dij76-53, Dij76-52, Dij76-47, Dij76-54, Dij76-56, Dij76-55, Dij77-34, Dij77-43, Dij78-34, Dij78-35, Dij79s, Dij80m, Dij80a, Dij80-29, Dij82d, Dij84l, Dij85-31, Dij85y, Dij87p, Dij88t, Dij89-34, Dij89-39, Dij89-27, Dij89-40, Dij90j, Dij92c, Dij94w, Dij94-29, Dij95o, Dij96-34, Dij99a, Dij01a, Dij01p, Dijxx-70, Dijxx-67, Ein90, EM72, HKS93, Knu66a, Knu66b, Koh76, Lam19, MHMO10, MT84, Muk12, NA13, NF95, Pop95, Pri90, SNR15, vT90, vdH17]. Problem/Solution [Dijxx-71]. Problems [Nep96, Dij59c, Dij76-33, Dij96-42, Dij96-40, PJ89]. Procedure [Bri02, Day11b]. procedures [DS83c, vdH17]. Proceedings [Bro96, Dij77b, NRB76, Sam76, Maz76, Win78, HW06, IEE79b, IdKM06, Sam76, BMW93, HP72]. procesbeschrijvingen [Dijxx-58]. Process [Dij96-37, Leh80, Rua90, ZXZ^{+}09, BBM07b, Dij64i, Dijxx-58, MP90, Ter93]. Processes [DLP79, Dij71a, Dij72f, Dij82-71, Dij96j, Dij02b, Dij02h, Hoa78b, Hoa78a, Hoa85, Dij62g, Dij68e, Dij68j, Dij68k, Dij77d, Dijxxx, Dijxx-35, Dij65a, Dij66d, Dij67e, Dij68r, Dij69j, Dijxx-85, Dijxx-86]. Processing [Bri01, Dij63r, KM83, BS01, Dij64j]. processor [Nep00]. Processors [CMS92, ND00]. Procurement [Dij68c]. product [Dij83o, Dij97i]. production [Du10]. productivity [Dij80x]. Prof [DS76a]. prof. [Dij86y]. Professional [Dij95k]. professor [Dij70d, Dijxx-99, Dij76-31, Dij76-64, Gaw64]. Program [Bör94a, Bur80, CCD86, Dij68h, Dij78o, Dij78-36, Dij79a, Dij79x, Dij82-98, Dij82-45, Dij82-116, DS90g, Dij21b, Fag01b, Jac01, Wir01a, Are90, BMW93, Bör94b, BGM90, Bro96, CT90, Dij55c, Dij62b, Dij63n, Dij64c, Dij67b, Dij68g, Dij68o, Dij70a, Dij77i, Dij77j, Dij77k, Dij77-47, Dij78a, Dij78b, Dij81e, Dij82-35, Dij84b, Dij93c, Dij93h, Dijxx-61, Dijxxl, Dijxxq, Dijxxr, Knu90, LH90, MHMO10, ST99, BB79, Dij89-38]. programma’s [Dij73c, Dijxxw]. programmed [Dijxxx]. programmeermethodologie [Dijxx-90]. Programmer [Dij68a, Dij72h, Dij79d, Dij80-30, Ano02a, Dij55a, Dij55b, Dij72g, Dij76-57, Dij87j, DS72c]. Programmeren [Dij82v, Dij60b, Dij70b, DF84, Dijxx-36, Dijxx-54, DDvW57, Dij83w]. Programmering [Dij56c, Dij56d, Dij56e, Dij57c, Dij57d, Dij63m, Dij69f, Dijxx-56, Dijxx-89]. programmeur [Dij55a, Dij55b]. Programmierens [DF85]. Programmierung [Str72]. Programming [Bac80, Bri02, CK02a, Che02b, DDH72, DDvW57, Dij60d, Dij60e, Dij62f, Dij63c, Dij65f, Dij65g, Dij65h, Dij69a, Dij69e, Dij69b, Dij72a, Dij72m, Dij73j, Dij73s, Dij74m, Dij76-58, Dij77-27, Dij79u, Dij79b, Dij79v, Dij82o, Dij82v, Dij83u, Dij83w, Dij96-35, Dij01j, Dij01o, Dij02d, Dij02i, DS06, Dijxxy, Dijxx-56, Dijxx-73, Dijxx-77, Dijxx-83, DCF^{+}02, ESB80, HW06, Hoa01a, HB94, Knu74, KP80, Knu92, Lam74, MB97, Mor90b, Pay14, Sha81, Wil80a, Wir01b, BD79, CW75, Che02a, Dij60b, Dij61d, Dij63m, Dij66a, Dij69g, Dij69i, Dij70b, Dij70f, Dij70h, Dij71g, Dij73q, Dij73a, Dij74p, Dij75k, Dij75-42, Dij75-60, Dij76m, Dij76-60, Dij76-61, Dij76a, Dij76-45, Dij77s, Dij78-28, Dij79c, Dij80c, Dij81c, DF84, DF85, DF88a, Dijxx-36, Dijxx-54, Dijxx-72, Dijxx-78]. programming [Dijxx-79, Dijxx-89, Dijxx-90, EM72, Knu66a, Knu66b, Lam19, ŁMB02, NB93, Sni06, Str72, VU05, Bar74, Dij56c, Dij56d, Dij56e, Dij57c, Dij57d, Dij69f, Dij70e, Dij75-31, Dijxx-68, Bar75, Flo77, Llo74, Mil89]. programmirovanie [DDK75, DDH75]. Programs [CMS92, DLP79, Dij75y, Dij76z, Dij76-29, Dij82-77, Dij82r, Dij89p, Dij96v, KM83, van19, AdBO90, BS99, Ber99, Dij56a, Dij73c, Dij74e, Dij75v, Dij75x, Dij75w, Dij76y, Dij77-28, DvG81a, Dij81k, Dij82f, Dij85-36, DS90i, Dij90r, Dijxxw, Dijxx-50, Dijxx-93, Dijxx-52, Gou90, Har81, MP90, Mau78, RR64, SWS^{+}13, Dij84o]. Project [AVL02, Dijxx-69, Dijxx-72, SS10, Tro80]. promiscuous [Dij93f]. Proof [Dij78o, Dij79a, Dij82-71, DS82i, Dij86f, Hoa01b, Lee06, Dij73f, DS76b, Dij77d, Dij77i, Dij77j, Dij77k, Dij78a, Dij78b, Dij80-31, Dij80-36, Dij81e, Dij81p, DvG81b, Dij83b, Dij83m, Dij85c, Dij85v, Dij85-32, Dij86y, Dij87s, Dij88f, Dij89c, Dij89h, Dij89-33, DR90a, Dij90d, DR90b, DS90d, Dij91g, Dij91k, Dij91u, Dij91w, Dij92q, Dij92p, Dij92t, Dij93b, Dij94i, Dij94k, Dij94j, Dij95b, Dij95j, Dij95l, Dij98j, DM00, DM01, Dij01f, DFxx, Knu90, Mis90, RD92]. Proofs [DLP79, Dij82i, Dij89p, BS99, Dij75g, Dij81k, Dij89j, Dij91p, Dij92d, Dij92l, Dij92w, Dij93k, Sch90a, Zem90]. propadeuse [Dij80t]. properly [Dij76-44]. properties [DS90j, Urb01]. property [Pos90]. Proposal [Dijxx-74, Dij75-38, DvG83b, Dij64j]. proposed [Dij71b, Dij71c]. propositional [DD98]. propositions [Dij81q]. prose [Dij98a]. prove [Dij88p]. Proved [Dij82-99, Dij77-30, Dij94-27]. Proven [Dij82u]. Proving [AdBO90, DvG81c, Dij86-33, Dij90p, Dij91v, Dij97j, Kes88, NB93]. Pruning [Dij97k]. Pseudo [Dij02a]. Pseudo-Spectral [Dij02a]. PSO [PG13]. PSST [Dij91-27]. Psychology [Dij06, Dij81s]. Ptolemaeus [Dij86-34]. Public [SWW99, SWW00]. pudding [Dij85-32]. punch [Dijxxf]. punched [Dijxx-59, Dijxx-81]. punctual [Dij86-31, Dij94d]. purpose [Dijxx-69]. Pythagoras [Dij86-29, Dij09, Zem90]. Pythagorean [Dij94x].

QoS [Smi11]. quadrangular [Dij86-28]. Qualification [MB97]. quality [Dij89-35]. quantification [Dij80v, DS82g, Dij84i, Dij86-38, Dij90o, Dij94n, Dij94t, MZS02]. quasi [Bij90]. quasi-boolean [Bij90]. queries [HSW08]. quest [Tur90b]. question [Dij66d, Dij67e, Dij68r, Dij69j, Dijxx-85, Dijxx-86]. questionnaire [Dij75-31]. questions [Dij74q, Dij84q, Dij00a]. queue [MT84].

R [Bar75, Bar74, Dij70d, Dij74a, Dij75-61, Dij76v, Dij77l, Dij77q, Dij82-76, Dij82x, Dij82-89, Dij82s, Dij82-66, Dij85y, DR90b, Dijxxz, Hoa82, Llo74]. R. [DLP78, Dij76u, DvG81b, Dij90j, Dij94w]. Raad [Dij76b, Dij76c]. rabbit [Dij95w, Dij02c]. Rabin [Dij92c]. race [Dij80-49]. Radical [DCF^{+}02]. Railroad [SWW99, SWW00]. Rajeev [Dij96l]. Rallying [Day11a, Day11b, Day12b]. Ramification [MB97]. random [PS14]. range [Dij96g]. Rao [DR90b]. rapport [Dij63n]. raster [Wir90]. rational [Dij99a]. Rationals [CW00]. Raw [Dijxx-75]. Ray [Uch14]. RCA [Raj80]. Rd1 [Dij54a]. Rd2 [Dij54a]. Re [Dij77-28, Dij95v]. reaching [Dij76-62]. reactie [Dij63g, Dij63d, Dij63e, Dij63h, Dij79i]. reaction [Dij63g, Dij63d, Dij63e, Dij63h, Dij79n]. reactive [Bro90b]. reading [Sch90a]. Real [Dij75-35, SZHT10, Boa90, BGM90, Mau78, Van78, Dij88p]. Real-Time [Dij75-35, Boa90]. realistic [GGAVGGRC15]. reality [Dij86-36]. Realization [KLMZ11]. Really [Dij89-32, Dij88m, Dij91b]. rearranging [Dijxxb]. Reasoning [ŁMB02, Dij96n, Dij96p, LCL^{+}94]. Recensie [Dij70g]. Rechenmaschine [BS01]. recherche [Bov86]. recipients [Dij75-29]. Recollections [Svo80, Tro80, Dij01g, Dij05]. Reconfigurable [TS01]. reconsidered [RD68]. Reconsidering [Nag18]. reconstruction [Dij79l]. Record [DS82b, DS72a, DS72b, Dij84b, Dij85y, Dij85k, Dij85l, Dij85m, Dij91i, Dij95h, Dij98c, Dijxxz]. Recording [Dij80-32, Dij76-59, Dij95-27]. Recounting [CW00]. recovery [Dij92u]. rectangle [Dij88t]. rectangles [Dij92f]. rectangular [Dij94h]. recursion [Dij74d, Dij82-34]. Recursive [Day11b, Dij60d, Dij60e, Dij80x, DS83c]. RED [Dij78-30, Dij78e]. Rediscovered [Bri90]. redone [Dij80-48]. Reduce [Fag01b]. Reducing [Dij83q, Dij83r, Dij83s]. reduction [Dij81f, PRM11]. referee [Dij85-44, Dij92u]. referees [Dij75-30]. Refinement [Wir01a, Mor90b, vG90]. refinements [Dijxxo]. Reflections [Wel80, Dij82-37]. reflexivity [Dij86p]. regarding [Van78]. Region [MP91]. regions [Dij77-33]. regrettable [Dij93o]. regularity [Dij84m, vGDxx]. reis [Dij70i, Dij71h, Dij72r]. Reisverslag [Dij71e, Dij72n, Dij73k, Dijxx-76, Dijxx-77]. Rekenautomaat [Dij68l]. Rekenautomaten [Dij68c]. rekenmachine [Dijxxb]. rekenmachines [DDvW57]. Rekentechniek [Dij54a]. related [CSZ10]. relation [Bac90, Dij96-41, Dij00d, Dij00e, GB96, Urb01]. Relational [Cod01, Dij86-35, HB94, Ber99, Dij90q, Dij91q, Dij91-32, Dij92d, Dij92o, Dij92v]. relations [Dij87r, Dij88l, vdW90]. Relationship [Che01]. Relaxation [Dij82a, Dij73r, Dij82d]. relay [UOZ11]. Reliability [PRC12, Bro90b, Dij77y, Dijxx-50, Jef78, PRD14, Dij73c, Dij77x]. Rem [Dij76-59, Dij77t, Dij78v, Dij82-93]. remark [Dij87m, Str72]. remarkable [Bro90a]. Remarks [DvG82h, Moo02, Zus80]. remembered [She03b]. Reminiscences [Hou80, Har99]. Remote [Bri02]. removal [Dij95w]. renaming [Dij82-33]. renewed [Bro90a]. reordering [PRM11]. Repaired [HCSW09]. Repaying [Dij80-33, Dij82-38]. repetition [Dij85-33]. repetitions [DS90h]. rephrased [Dij99o]. Replacement [Dij82-83, Dij66c, Dij74h, Dij86z, DS90e]. Reply [Dij89i, Dij80f, Dij89t, Gaw64, Van78]. Report [BR70, DHPS59, DDG^{+}70, Dij82-87, Dij82c, Dij82-57, Dij82b, Dij82-109, Dij82-106, Dij82-110, Dij82-108, Dij82-112, Dij82-114, Dij82-113, Dij82-107, Dij82-111, NBB^{+}60, NR69, WTD^{+}69, vWMPK69, Dij62e, Dij63n, Dij65d, Dij65j, Dij71e, Dij72n, Dij72q, Dij73k, Dij73u, Dij73v, Dij73t, Dij73s, Dij74u, Dij74x, Dij74w, Dij74t, Dij74v, Dij75-53, Dij75-52, Dij76-37, Dij78-44, Dij79-33, Dij79-30, Dij79-29, Dij79-31, Dij79-32, Dij80-44, Dij80-47, Dij80-45, Dij80-46, Dij81-28, Dij81z, Dij81-27, Dij81-30, Dij81-31, Dij81-29, Dij81-32, Dij81y, Dij81-33, Dij82-54, Dij82-51, Dij82-49, Dij82-55, Dij82-52, Dij82-50, Dij82-56, Dij82-53, Dij83x, Dij83-27, Dij83y, Dij83-28, Dij83z, Dij84u, Dij84t, Dij84w, Dij84y, Dij84z, Dij84s, Dij84x, Dij84v, Dij85-48, Dij85-47, Dij85-50, Dij85-49, Dij85-45, Dij85-46, Dij86-44, Dij86-45, Dij87y]. report [Dij88r, Dij88s, Dijxx-43, Dijxx-76, Dij65i, Dij67g, Dij69k, Dij70i, Dij71h, Dij72r, Dij73x, Dij92u, Dijxx-97, Dijxx-96]. reports [Dij01l]. representational [Dij73h]. Representations [Hoa01b]. Reprint [Dij83v, Dij83u]. Requirements [Dijxx-78, BW00, Dij75j]. Research [Dij82-103, DvG83b, Dij86-46, Raj80, Tom80, WM14, Dij69f, Dij86t, Dij77-36, Du10]. réseau [Bov86]. resented [Dij74c, Dij75l]. residue [Dij87v]. respect [Dij67a, Dij72p]. respectable [Ano02a]. Respond [PSvE^{+}89]. Response [DLP78, PRM11]. restoration [SVSR04]. Restriction [DvG86c, DvG84b, DvG84c, DvG84d, DvG84e, DvG85b]. result [Dij95w]. results [DS84a, HHM94]. Retraction [Du10, zFmLcWcJ10]. Retrieval [WLXX09]. retrospect [Dij76-30]. Reuse [GHJV01]. Reverse [Coc02]. reversed [Rei08]. Review [Bar75, Bar74, Bör94a, Bör94b, Dij75-47, Dij77b, Flo77, Llo74, Mil89, Dij63r, Dij72c, Dij78h, Dij78i, Dij78-37, Dij79f, Dij82-39, Dij84o, Dij84n, Dij70g, Dij71f]. revised [VU05]. Revisited [DF89, BH91, Dij73e, Dij73p, Dij76s, Dij76-33, Dij81i, Dij82m, DS82g, Dij83p, Dij84-28, Dij87n, DF88b, Dij91d, Sni06]. Revolution [Pay14]. Revolutions [Bur80]. Rham [Urb01]. Ribbon [Dij85-31]. Richard [Dij99h, Dij94k]. Ridge [Hou80]. Ring [CMS92, BDB93]. river [Dij96k, Dij00i]. road [Dij82-48]. robot [AAEG11, LSJT12, WYY11]. Robust [Dij82-77, Dij75v, GB96, YJ10]. robustness [Bro90b, Dij75a, KF08]. Role [Dij82-92, Dij74l, Dij76-60, Dij77h, Dij78m, Dij98k]. rolled [Dij01l]. Rome [BR70, Dijxx-96]. ronde [Dij76c]. root [DS81, Dij99j]. Ross [Dij75-45]. Rotate [Hab90]. Rotating [Dij85-34]. Rough [Dij64i]. roulette [Dij89w, Dij89x, Dij90n]. round [Dij76c]. Route [LSMS12, Anw03, FS10, zFmLcWcJ10, KYTY09, LCL^{+}94, RPGS14, WM14]. Routed [GLFM07]. Routes [WLXX09]. routine [Dij91d, Dijxxb]. routines [Dijxx-64]. Routing [GLFM07, KKNI14, yLsLtZ09, yLsLtZ10, TS01, AHHK93, AHH^{+}95, EKP96, KF08, MTZ12, PS14, PRV09, YL10]. Royal [Dij82s, Dij82-48]. rule [Dij86g]. Ruler [DS06]. Rules [Dij82-103, DS06, Dij77-36]. Rush [SEEM01]. Russell [Dij85-52]. Rutger [Dij80-31]. Rutishauser [Bau80]. Ruwe [Dij64i].

S [Bör94a, Bör94b, Dij86h, Dij88f, Dij89h, Dij90i, Dij90d, Dij90j, Dij94k, Dijxx-29]. S. [DS76b, DS82i]. Saddleback [Dij85-35]. safety [Rem90]. Safra [Dij87u]. sake [Dij89o]. Salute [FvGGM90]. same [Dij94s, Dij96-32]. Samenvatting [Dij64j]. Sans [Dijxx-79]. Santa [Dij79-31]. SASL [Dij81h, Dij82f, Dij82-67]. satisfying [DvG83a]. save [Dij85d]. scalar [Dij94t]. scale [Dij97f, GL98]. scan [PRM11]. scanner [Dij66e]. scenes [Mos10]. schaats [Dij96m]. Scheduling [LWQ09, WLXX09, AVL02]. Schemes [CCD86, BT79]. schets [Dij64i]. scheve [Dij96m]. Schieder [Dij95e]. Schneider [Rem90]. Scholten [Bör94b, BN98, Bör94a, Dij86h, Dij89c, Dij92t, Dij92z, Dijxx-29, Rei08]. Scholten/Dijkstra [Rei08]. School [Dij71e, Dij73t, Dij75-55, Dij82b, Dij82-54, PGB00, BS76, Dij00j, VU05, Dij95-29, BB79]. Sci [Dij82o]. Science [BR70, Dij67a, Dij79r, Dij86-36, Dij89-32, Dijxxa, Dijxxm, Dijxx-96, Lap96, Lew21, NR69, Tat23, Dij71d, Dij72e, Dij76-88, Dij79q, Dij80t, Dij82-28, Dij82-29, Dij84k, Dij84r, Dij85z, Dij85d, Dij85-28, Dij86o, Dij87-27, Dij88m, Dij88j, Dij88k, Dij88u, Dij91a, Dij93e, Dij01c, Dijxx-32, LSJT12, Maz76, Win78, Dij78t, Dij87d, Dij90k, Dij95-31, Dij99b, Dijxxm, Dijxx-60]. Sciences [Ano12]. Scientific [Dij77-36, Dij82-92, Dij82-103, Dij74l, Dij76-58, Dij77-27, Dij78-33, Dij86-27, Dij91x, Dij68b]. Scientist [Dij75n, Dij82-72, Dij75m, Dij88d, Dij89g, Dij89-44, Dijxx-61]. Scientists [May80, SL98]. scope [Dij76r]. Scotland [Dij81-32]. Scriptum [Eds82, Dij75-46, Dij82-36]. sd&m [BD01b]. sdb [Dij80t, Dij80t]. SEAC [Slu80]. Search [Bri90, Dij76-32, DF89, Dij99c, FNS03, KP09, Cha10, CH10, Dij82-27, DF88b, Dij97k, KP10, NS00, Dij85m, Dij85-35]. second [BES95, BMW93, Dij76c, Dij84m, Dij94-29, Dijxx-96, Dij63d, Dij75-59, Dij76-81, Dij76-80, Dij77-42, Dijxx-95]. second-order [BES95]. Section [Dijxxc]. security [ZLY13]. See [Dij96x, Dij96-43, Dij83m, Dij86-31, Dij92n, Dij92z, Dij96b, Dij96-39, Dij89-37]. seem [DvG81d]. Seemingly [Dij89-40]. seems [DvG84h, DvG84i, Dij95-31]. Seen [Bir80]. Segers [Dij94-27]. Segment [Dij64k, DvG83a]. Segmentation [BDBR01]. segments [Kal90]. seinpalen [Dijxx-63]. Selected [Dij72o, Dij82-69]. Selection [CC12, LX10, LSMS12, SWX10, Dij89-36, KYTY09, Dijxxb]. Self [Dij73m, Dij73l, Dij73o, Dij73n, Dij74n, Dij74o, Dij82e, Dij82-40, Dij86f, DH01, CSZ10, Dij85c, Kes88]. Self-Stabilization [Dij82-40, Dij86f, Dij73m, Dij73l, Dij73n, Dij82e, Dij85c, Kes88]. Self-Stabilizing [DH01, Dij73o, Dij74n, Dij74o, CSZ10]. semantic [Dij76-61, TZDM10]. Semantics [Bij90, DS83c, Dij85-36, DS90g, DS90h, DS90i, Dij90r, MB97, Dij73i, Dij73k, Dij76h, Dijxxj, DLMB99, JH93, Kot90, ŁMB02, Luk94, dB76, Dij89-38, Dij73k, Bör94a, Bör94b]. semaphore [Dij79-34, Dij80-42, Dij82-64, Vv72, Kos73]. Semaphores [Bri02, Dijxx-63]. semester [Dijxxi]. Seminar [Dij82u, Dij82c, Dij73u, Dijxx-77]. SEN [Kru03]. Sensor [MTZ12]. Sep [Dij76-77, Dij78-47, Dij80-46, Dij81-32, Dij82-52, Dij82-50, Dij82-56, Dij82-63]. separately [BDR85]. Sept [Dij73v, Dij77-40, Dij82c, Dij82-114, Dij82-111, Dij82-61, Dij73u]. September [Dij74t, Dij75-57, Dij75-56, Dij76-75, Dij79-31, Dij79-32, Dij82-110, Dij82-112, Dij82-60, Dij82-62, Maz76, Win78, Dij72q]. Sequel [Dij82-84, Dij67c, Dij67d, Dij73e, Dij75-48, Dij76-34, Dij77-29, Dij78m, Dij78s, Dij79-39, Dij80p, Dij80-34, DS82f, Dij82-41, DS83d, DS84c, Dij84p, Dij85-37, Dij86-37, Dij91h, Dij91t, Dij92l, Dij96-36, Dijxxr, Dijxx-80]. sequence [Dij83t, Udd90, Urb01]. sequences [Dij75-40, Dij79-35, Dij80y, Dij82-32, Dij82-47, Dij91v, Dij93p, Dijxx-75, vDO01]. Sequencing [Dij73p, Dij78-38, Dijxx-84, Dij73e]. Sequential [Dij65a, Dij66d, Dij67e, Dij68r, Dij69j, Dij71a, Dij72f, Dij96j, Dij02b, Dij02h, Dijxx-85, Dijxx-86, Hoa78b, Hoa78a, Hoa85, Dij68e, Dij68j, Dij68k, Dij76-83, Dijxx-35, Dijxx-59, Dijxx-81, SB01]. sequentialiteit [Dijxx-58]. sequentiality [Dijxx-58]. Sequentiele [Dijxx-81, Dijxx-59]. Sequentielle [SB01]. serial [Dij62b, JAE^{+}12]. Serializable [Gou90]. Series [Flo77, Dij75-29]. Service [SWX10, SZHT10, JSW04]. Service-Correlation [SWX10]. Services [LX10]. Set [HB94]. Sets [Dij81t, Dij80y, Dij84-28, Dij00c]. Seventh [Dij75-67, Dij75-68, Dij76-93, Dij77-49]. Shaped [Mar02, Dij80z]. Shared [Cod01, BW00]. sharing [Dij72i, Dij72j]. sharp [Dij91-28]. Sheffer [Dijxx-29]. Shift [DCF^{+}02]. Shiloach [Dij98c]. Shiri [Ano13]. Shmuel [Dij87u]. short [Dij65d, Dij71g, Dij79y, Dij80-35, Dij80-36, Dij82-41, Dij84p, Dij86-38, Dijxx-79, Dij56b]. Shortest [BB99b, BB99a, CMMS98, Der85, FJ09, GLFM07, HCSW09, Joh73, Kal90, MP91, ND00, TS01, CF04, CH10, Dij76-55, Dij92b, Dij96h, Dij98a, HKS93, HSWW05, HSW08, JAE^{+}12, JHLC14, KGZ06, Koh76, Kun80, Lee06, Mis01, MT84, Muk12, NA13, NMM78, Nos85, NS00, Pet71, PRV09, SNR15, SxXq11, Trä95, WWZ05, xXL10, XLH^{+}07, Yap83, Dij60c]. shortest-path [HKS93, HSWW05, HSW08, Koh76, WWZ05]. Should [Dij79z, Dij82-68, Dij93g]. Shuffle [Dij82p]. shunting [Dij98i]. siblings [SZHT10]. side [Dij80i, Dij86r]. sided [Dij75-40]. Sides [Dij82-91, Dij77r]. Sieve [Leh80]. Simple [Dij79p, DS82a, DvG86c, ND00, SD82, BGM90, Dij73q, Dij73a, Dij74p, Dij76-35, Dij76-62, DS77, Dij77p, DS82d, Dij82-67, DvG84b, DvG84c, DvG84d, DvG84e, DvG85b, Dij85-39, Dij85-40, Dij85-38, Dij89-33, Dij92q, Dij93c, Dij94j, Dij95x, Dij98i, Dij98j, Knu90, Mis90]. Simpler [Sch90a]. simplest [Dij96-28]. Simplicial [YL11]. simplicity [PGB00]. simplification [Dij78x]. Simplifying [Dij91w]. simulate [GGAVGGRC15]. Simulation [GLFM07, UNTH14, BDB93, Cha10, CH10]. simulator [BGM90]. single [Trä95]. single-source [Trä95]. Sintzoff [Dij77-28]. Situ [DvG82f, Dij82-43, Dij81b, Dij82, Dij81u, Dij81v, Dij82o, DvG82g, DvG82d, DvG82e, Dij82-42]. Situation [JNK11]. Sixth [Dij75-65, Dij75-66, Dij76-92, Dij76-91, Dij77-48, JSW04]. size [Dij67f, Dij86u]. Sizeable [Dij76z, Dij82r, Dij76y]. skate [Dij96m]. skeletonization [VPM^{+}10]. sketch [Dij62c]. skew [DS81]. skewed [Dij96m]. Small [Bir90, Dij82-71, BDS82, Dij76-27, Dij76-63, Dij77d, Dijxx-28]. smallest [Dij76-56]. Smalltalk [Coc02]. smart [GGAVGGRC15]. Smithsonian [Tro80]. smoothing [Dij75-40]. Smoothsort [BH91, Dij82, Dij81u, Dij81v, Dij82o, Dij82-42, Dij82-43]. snapshot [Dij83e, Dij85i, Dij83f]. snapshots [Dij01i]. Snepscheut [Dij94k, Dij94m, Dij94l, Dij95o]. Snoopy [Nel90]. Snoopy-caching [Nel90]. sobering [Dij75d]. Social [DLP79, KKNI14, Dij81a]. Social-Graph-assisted [KKNI14]. Society [Dij98k, Dij82s]. Software [Boe01, BD01b, BD02, BR70, Day12a, Dij75-47, Dij76-70, Dij79z, Dij82u, Dij82-105, Dij00a, Dijxxm, Dijxx-97, Dijxx-96, Fag01a, Har95, IEE79a, IEE79b, NR69, NRB76, Par18, Ran02, Dij77y, Dij77-46, Dij82k, Dij82l, Dij82-115, Jef78, JHLC14, Neu90, RPGS14, Van78, Dij77b, Dij77x, Dijxxm, Dijxx-96]. Solution [Dij65h, Dij82a, Dij83u, Dij96-37, Dij01j, Dij02i, Dij21b, Dijxx-71, Lam74, Bau90, Dij73r, Dij79w, Dij80a, Dij82d, Dij86-39, Dij88t, Dij94w, Dij94s, Dij96-32, Lam19, PJ89]. Solutions [Dij82-89, Dij82-93, Dij77q, Dij77t, Dij77-43, DS83a, DS83b, DS84b, Dij85j, Dij86j, Dij87e, DS90c, Dij91s, Dij92m, SNR15]. Solve [MB97]. solved [Dij78-35, Dij80-29, Dij84l]. Solving [Dij02a, MHMO10, Muk12, Pop95]. Some [Bac80, BBM07c, DMS83, Dij60a, Dij62f, Dij64a, Dij74q, Dij78-39, Dij79-27, Dij80-37, Dij84q, DvG84f, Dij85-39, Dij85-40, DS90j, Dij92w, Dijxx-82, Dijxx-83, Nel92, Nep96, Tro80, Urb01, Zus80, Dij76-27, Dij77-32, Dij82-46, Dij89f, Dij95s, Dijxx-28, Hor90]. Somewhat [Dij82-99, Dij76-64, Dij77-30, Dij78v, Dij80-41, Dij80-40, Dij80-38, Dij80-39, Dij81w, Dij87q, Dij89-41, Dij91y, Dij92x, Dij96-38]. SOO [Dij80t, Dij80t]. SOO/jc/sdb [Dij80t, Dij80t]. Sorry [San05, Dij94y]. Sorted [Swi11]. Sorting [DvG82f, Dij82-43, Dij77-31, Dij81b, Dij82, Dij81u, Dij81v, Dij82o, DvG82g, DvG82d, DvG82e, Dij82-42, Dij85q, Dij85-39, Dij85-40, Dij88q]. Source [BD66, CC12, GLFM07, EKP96, Trä95]. Soviet [Tat23]. Space [yLsLtZ10, BDS82, Dij79h, Dij95c, yLsLtZ09]. spanning [Dij60a, LvdS90]. sparse [XLH^{+}07]. Speaking [Dijxx-92]. special [AG01, Dij77-31]. Specialised [NB93]. specialization [Dij72e, Dijxx-56]. Specification [DeM01, Bir90]. specifications [ST99, Tur90b]. specifying [Dij92u]. Spectral [Dij02a]. Speech [Dij74r, Dij82-70, Dij82-44, Dij74b, Dij96i, Mos10, Dijxx-90]. speed [BDn^{+}10, HSWW05]. speed-up [BDn^{+}10, HSWW05]. Speedup [NAGK09, MSS^{+}06]. spell [Dij00k, Dij01n]. Sphere [MJC07]. SPI [DCF^{+}02]. Spite [Dij74o, Dij82-40, Dij73l, Dij74n, Dij82e]. split [Dij79-34, Dij82-64]. sponsored [BR70, NR69]. Spring [Dij75s, Dij76e, Dij76l, Dij76k, Dij76q, Dij76-36, Dij76-81, Dij76-67, Dij76-79, Dij76-85, Dij76-87, Dij76-92, Dij76-93, Dij77a, Dij77e, Dij77f, Dij77-42, Dij77-44, Dij77-45, Dij77-48, Dij77-49, Dij96y, Dij97h]. Springer [Bör94b, Dij71f, Dij77b]. Springer-Verlag [Dij77b]. square [Dij97i]. squared [Dij95h]. squares [Dij76-89, Dij88e, Dij93m, Dij93n, Dij96-33]. St [Dij76-76, Dij82-113, Dij82-53]. Stabilization [Dij82-40, Dij86f, Dij73m, Dij73l, Dij73n, Dij82e, Dij85c, Kes88]. Stabilizing [DH01, CSZ10, Dij73o, Dij74n, Dij74o]. stack [Dij63p, Dijxx-65]. Stacker [WLXX09]. standaard [Dij56a]. standaard-typprogramma [Dij56a]. standaardroutines [Dijxx-64]. standard [Dijxx-64, PS14, Dij95k]. Standards [Hus80, Slu80]. Stanford [Dij87y]. Stanford/Palo [Dij87y]. Stanley [DvG81b]. stapeladministratie [Dij63p, Dijxx-65]. Start [Tom80, Dij82-68]. Starts [Dij82-100, Dij78-40]. starvation [Ban07, Dij77-43]. starvation-free [Dij77-43]. State [LSMS12, Dij73m, Dij73n, Dij78v, Dij79h, Lam22]. Statement [Dij96u, Dij01e, Dij02g, Dij21a, Dij68d, Dij68t, Dij68s, Gal73, RD68]. Statements [Knu74]. States [Dij76-65]. Static [Bai86, GLFM07]. Stationary [Dij77-32, Dij79-27]. status [Dij91-31]. stay [Dijxx-99]. STD [Dij80e]. Steinhaus [Dij96-42, Dij01a]. step [LCZW14]. stepping [Tur90b]. Stepwise [Dij68o, Dij82-45, Wir01a]. Stern [Urb01]. still [Dij93r]. Storage [WLXX09, Dijxx-51]. Storage/Retrieval [WLXX09]. Store [Dij82-104, Dij74s, Dij76-68, Dij78n, Dij76-69, Dijxx-87]. Stored [Bur80]. Stored-Program [Bur80]. stories [Dij85t]. Story [Dij82-100, Dij78-40]. straight [Dij85-36, DS90i, Dij90r]. straight-line [Dij85-36, DS90i, Dij90r]. straightly [Rei08]. strange [Dij86-40]. strategies [Dij72d]. Strategy [HZW09]. streamlining [Dij85-41, Dij85-42]. streams [Dij72i, Dij72j]. strengths [Dij94z]. string [BD89, Dij81f, Dij86-30]. striptease [Hel90]. Stroke [Dijxx-29]. Strong [Dij82q, Dij82-94, Are90, Dij73d, Dij76w, Dij77-33, Dij78-32, Dijxx-27]. strongest [DS90k]. strongly [Dij77-28]. Structure [BMWC01, DeM01, Dij66a, Dij69h, Dij83v, Dij01k, Dij02j, Dij21c, Par01a, van19, Dij68p, Dij68q, Dij79h, Dij80-32, Dij81k, EBR09, Kot90, Dij75h]. Structured [Bar74, Che02b, DDH72, Dij69a, Dij69e, Dij69i, Dij69b, Dij70h, Dij72a, Dij76a, Dij82u, Dij01o, Dij02d, Knu74, BT79, CW75, Che02a, Dij70f, Dijxx-94, Dij70e, Bar75, Llo74]. Structures [Par18, WG01, Dij85-27, Dij87b, Dij87c, DS90a, Dij90c, DS90m, Leh77, SAID13]. structuring [Dijxxx, Dijxx-53]. Structuurplan [Dij75h]. Strukturnoe [DDH75, DDK75]. student [Dij91n, Dij93t, Dijxx-92]. studenten [Dij75c, Dij75p, Dij75s, Dij75-63, Dij75-66, Dij75-68, Dij76e, Dij76d, Dij76l, Dij76k, Dij76j, Dij76q, Dij76n, Dij76-36, Dij76-81, Dij76-67, Dij76-79, Dij76-80, Dij76-85, Dij76-87, Dij76-84, Dij76-86, Dij76-92, Dij76-93, Dij76-91, Dij77a, Dij77e, Dij77f, Dij77-42, Dij77-44, Dij77-45, Dij77-48, Dij77-49, Dijxx-95, Dijxx-98]. students [Dij75c, Dij75p, Dij75s, Dij75-63, Dij75-66, Dij75-68, Dij76e, Dij76d, Dij76l, Dij76k, Dij76j, Dij76q, Dij76n, Dij76-36, Dij76-81, Dij76-67, Dij76-79, Dij76-80, Dij76-85, Dij76-87, Dij76-84, Dij76-86, Dij76-92, Dij76-93, Dij76-91, Dij77a, Dij77e, Dij77f, Dij77-42, Dij77-44, Dij77-45, Dij77-48, Dij77-49, Dij79y, Dij86c, Dij86d, Dij86e, Dij00a, Dijxx-92, Dijxx-95, Dijxx-98]. studied [Urb01]. studievereniging [Dij93t]. Study [Bro96, SWW99, BS99, Dij87v, MZS02, SWW00, Big80]. stukje [Dij77o]. stupid [Dij81x]. stupidity [Dij80g]. Style [Pad85, Dij90k, Dij91a, Dij95t]. Sub [DHPS59, Dijxxc]. Sub-Committee [DHPS59]. Sub-Section [Dijxxc]. subalgebras [Tse97]. Subclass [DN01]. subcommissie [Dij68c]. subcommittee [Dij68c]. Subdivision [MP91, Dij76-88]. Subgoal [Dij82-90, Dij76-43]. subgraphs [Dij86u]. subject [Ano02a]. subjectively [KYTY09]. submarines [Dij89r]. submitted [Dij78z, Dij78c, Dij78-29, Dij78d, Dij78-30, Dij78e, Dij78-31, Dij78f]. Subroutines [D^{+}66a, Dij54a]. subscripted [Dij74k]. subsequence [Dij76-52, Dij77-34]. subsequences [Pri90]. subspanning [Dij76-55]. Substitution [Dij62g, Dij82-33, Dij86-32, Dij86z, DS90e]. subtrees [Dij60a]. subtypes [Dij93g, Dij94t]. Successful [Dij82-103, Dij77-36]. suggested [Dijxxz]. sum [Dij76-89, Dij88e, Dij93m, Dij93n, Dij96-33]. Summary [Dij79d, Dij82-97, Dij76-48, Dij82-35, Dij82-46, Dij85-43, Dij90q, Dij92y, Dij64j]. Summer [Dij71e, Dij73t, Dij75-55, Dij82b, Dij82-54, BS76, BB79, Dij71e]. summing [Dij78k, Dij78-41]. summing-up [Dij78-41]. superficial [Dij76-66]. superfluity [Dij80-42]. supplement [Dij77-34, Dij86-41, Dij87w, Dij95y]. Supplementary [Ran80b]. Supply [SVSR04, Du10]. support [Bis85, Ter93]. supported [SNR15]. Suppose [Dij86-42]. suprema [Dij00c]. surfaces [HTVW08]. survey [Dijxxh, FCTB08]. Susanna [Dij90i]. Susono [HW06]. SWAC [Hus80]. swapping [Dij96-43]. Swarm [KLK08]. sweet [Dij84a]. swim [Dij89r]. Swinkels [Dijxx-100]. Swiss [Dij78-34]. Sybrand [Dij84l]. Sylvester [Dij88d, Dij89g, Dij96-39]. symbol [KP10]. symbolic [VPM^{+}10]. symbols [Dij95-30]. symcharf [Dijxx-41]. symchart [Dijxx-41]. symmetric [Dij80-35, Dij99p]. symmetry [Dij88q, Dij91u, Dij86-34]. Symposium [Dij82j, HW06, Maz76, Win78, Dij65d, Dij75i, Dijxxe]. Synchronisatie [Dijxx-84]. Synchronization [Rua90, Hor90, Dijxx-84]. Synge [Dij92x]. Synthesis [Dij82-101, Dij02k, Dij75-49, JM80]. synthesized [Dij92k]. System [DeM01, Dij66a, Dij75-35, Dij83v, Dij01k, Dij02j, Dij21c, Dijxxe, XWZH12, van19, BBM07b, BBM07c, Dij65e, D^{+}66c, D^{+}66b, D^{+}66a, Dij66f, Dij66e, Dij68p, Dij68q, Dij69h, Dij01g, Dij05, Dijxx-74, Dijxx-94, NB93, NBK95, BD66, Dij63r]. Systems [Bri01, Dij74o, Dij76p, Dij79f, Dij82-116, MJC07, Par01b, Ban07, Ban09, Boa90, Bro90b, Dij73b, Dij73o, Dij74n, Dij76o, Dij77-47, Dij78-50, Dij91b, Dij92u, HP72]. systolic [LH90].

T.H.E. [Dij65b]. Table [DvW55, Dij66f]. Tables [D^{+}66c, TZDM10]. tail [Dij93i, Dij85n]. takes [Che16]. tale [Dij89-42]. talk [Dij79y]. tape [Dijxxv, Dijxx-59, Dijxx-81]. target [Dij86-47]. Targets [BMST03, MS01]. Taris [vDO01]. Tartu [VU05]. Taschenbücher [Dij71f]. task [Dij69g, Dij88q]. te [Dij65j, Dij65i, Dijxx-76, Dijxx-96]. teachability [Dij80-43]. Teaching [BB99b, BB99a, DH96, Dij89-32, Dijxx-77, Dij69f, Dij75-42, Dij76-45, Dij88m]. technical [Dij77c]. Technique [Dij54a]. Techniques [BR70, Dij76z, Dij82u, Dij82r, NRB76, BDn^{+}10, Dij76y, HP72, HSWW05]. Technology [Dij06, IdKM06]. Teil [Dij71f, Dij72c]. teleprinter [Dij63m]. Tell [Dij82-79, Dij75-27, Dij82t]. Tentamen [Dij69j, Dijxx-85]. Tentamenopgave [Dij66d, Dij67e, Dij68r, Dijxx-86]. tentative [Dij93p]. Tenth [Dij75-50, Dij76-67]. Ter [Dij72p, Dij83w, Dij67a, Dijxxu]. terminating [Dij76-44]. Termination [DS78a, DS78b, DS79, DS80b, Dij82-94, DFvG83, AdBO90, BD89, Dij76-90, Dij78-32, DS82e, Dij82m, DFvG82, DvG84g, DFvG85, Dij87u, Dij95z, Hes95]. terms [Dij94o]. Ternary [Dij79-27, Dij77-32]. Terracost [HTVW08]. terrains [HTVW08]. terrible [Dij93q]. test [PRM11]. Testing [LW11]. Texas [Dij79-30, Dij80-45, Dij93t, TLR^{+}91, Dij91-27]. text [WTD^{+}69]. th [Dij99j, Dij77c]. Their [BD01b, Dij82-83, SL98, Dij72m, Dij74h, Dij76-65, Dij79-36, Dij79-37, Dij80-32, Dij87r, Dij88l, DS90c, Dijxx-73]. Theme [Hoa82, Dij75-61, Dij82-66, SS04a, SS04b]. Theorem [Dij76-32, Dij79-28, DS82i, Dij82-102, DS76b, Dij77p, Dij77-35, Dij80w, Dij81p, DvG81b, Dij82-32, Dij82-47, DvG84g, Dij85-38, Dij86-43, Dij86-29, Dij88d, Dij89c, Dij89g, Dij90p, Dij91z, Dij92b, Dij92q, Dij92t, Dij92z, Dij94x, Dij94-27, Dij95x, Dij95-27, Dij96h, Dij96-28, Dij96-39, Dij97d, Dij97e, Dij97l, Dij99l, DM00, DM01, Dij09, DFxx, Her90, Hoo90, NB93, Nos85, RD92, dB76, Dij81r, Dij83b, Dij83m, Dij86k, Dij86-33, Dij89-33, DR90a, DR90b, Dij92p, Dij98j]. theorem-proving [NB93]. Theorems [DLP79, Dij60a, Dij76-82, Dij78y, Dij79-35, Dij80-36, Dij85-40]. theoretic [BT79]. Theory [CK02a, CRT78, Dij76-48, Dij80-48, Dij82-46, Dij94-28, Dij98f, Dijxx-42, Hoa90, Dij65i, Dij82-97]. There [Dij93r, Dij82-48]. thereof [Dijxxo]. things [Cre02, Dij74c, Dij75l]. think [Dij85e]. Thinking [Dij79d, Dij75-42, Dij76-45, Dij80-43]. third [CSZ10, Dij63e, Dij75o, Dij75p, Dij76k, Dij76j, Dij77e]. Thirteenth [Dij75q, Dij76l]. Thought [Dij82-92, Ham80, Dij74l]. thousand [Dij99h]. Threads [NAGK09]. threats [Dij84r]. Three [DvG82f, Dij82-103, Dij96-40, Dij01l, Dij73n, Dij76-50, DvG82g, DvG82d, DvG82e, Dij89y, Dij91l, Dij92-29, Dij77-36]. three-coloured [Dij89y]. three-state [Dij73n]. Throughput [KP09, KP10]. Thrown [Dij76-39, Dij75-36, Dij75-37, Dij76-38]. TH’s [Dij77c]. Thursday [Dij93q]. ti [Dijxx-76]. Tide [Dij97m]. Tiende [Dij75-50, Dij76-67]. tijdschrift [Dij70g]. tiling [Dij95o]. Time [Dij75-35, Dij82-104, Gri90, SZHT10, Boa90, BGM90, Dij74s, Dij76-68, Dij76-69, Dijxx-51, Dijxx-87, Lee06]. time-varying [Lee06]. Time-Wise [Dij82-104, Dij74s, Dij76-68, Dij76-69, Dijxx-87]. Timebase [RPGS14]. timeless [Tur90b]. titre [Dijxx-79]. toekomst [Dijxx-60]. Toekomstverwachting [Dijxx-89]. toespraak [Dij75b, Dij75c, Dij75q, Dij75o, Dij75p, Dij75r, Dij75t, Dij75s, Dij75-34, Dij75-59, Dij75-50, Dij75-58, Dij75-62, Dij75-64, Dij75-63, Dij75-65, Dij75-67, Dij75-66, Dij75-68, Dij76e, Dij76d, Dij76l, Dij76k, Dij76j, Dij76q, Dij76n, Dij76-36, Dij76-81, Dij76-67, Dij76-79, Dij76-80, Dij76-85, Dij76-87, Dij76-84, Dij76-86, Dij76-92, Dij76-93, Dij76-91, Dij77a, Dij77e, Dij77f, Dij77-42, Dij77-44, Dij77-45, Dij77-48, Dij77-49, Dijxx-92, Dijxx-95, Dijxx-98, Dij75-51, Dijxx-90, Dijxx-91]. Tokyo [Dij76-77, Dij82-114, Dij82-63]. told [Dij92c]. tomography [Dij79l, Dij79o, DvG82b, DvG82c, Dij85o]. Tomorrow [Dij80-43]. Tompkins [Rei83b, Rei83a]. Tony [Dij89-28, Dij99e]. Too [Dij91-27]. tool [Dij68l]. Tools [Dij82u, Dijxx-78, Ter93]. topologies [FNS03]. Topology [JNK11, LvdS90]. Topology-independent [LvdS90]. tot [Dij64j, Dij72l, Dij72n, Dij75c, Dij75p, Dij75s, Dij75-51, Dij75-63, Dij75-66, Dij75-68, Dij76e, Dij76d, Dij76l, Dij76k, Dij76j, Dij76q, Dij76n, Dij76-36, Dij76-81, Dij76-67, Dij76-79, Dij76-80, Dij76-85, Dij76-87, Dij76-84, Dij76-86, Dij76-92, Dij76-93, Dij76-91, Dij77a, Dij77e, Dij77f, Dij77-42, Dij77-44, Dij77-45, Dij77-48, Dij77-49, Dij83w, Dijxx-61, Dijxx-54, Dijxx-91, Dijxx-92, Dijxx-95, Dijxx-98]. Total [Dij92-27, Dij89d, Har81]. Total-order [Dij92-27]. tough [Dij99n]. tour [Dij92j]. Tracing [Uch14]. tradeoffs [AHH^{+}95]. trading [Dijxx-51]. traffic [Dij83q, Dij83r, Dij83s, GGAVGGRC15]. trailer [PG13]. training [Dij80t]. trajectory [AAEG11]. Transactional [NAGK09, ANGK09]. Transactions [DS82b, DS72a, DS72b]. transform [FCTB08, Dij82p]. Transformation [Dij82-116, Dij77-47, Dij78-50, Dij82-35]. transformational [Dij92u]. Transformations [Bul15]. transformer [BES95]. transformers [DS82c, Dij82y, Dij82s, Dij85-30, DS90b, DS90j, Dij92n]. transitive [Dij74d, DvG90b, Dij91q, Dij96-41, Dij00d, Dij00e, Leh77, vGD92]. Transitivity [Dij93s, Dij86p, Dij88c, Dij92-28, Dij98i]. translation [Dij61c, Dij64i, RR64, SB01]. Translator [Dij63b, Dij61c, Dij61a, Dij63f, Dij63i, Dij63n, Dij66f]. translators [Dij78w]. transmitted [Dij87p, Dij89-27, Dij89-40]. Transport [SWW99, Dij63q, SWW00]. Traub [Dij77q, Dij82-89]. traveling [ZTZY12]. Travelogue [Dijxx-77]. travesty [Dij80h]. Treatment [Dij79p, Dij76-27, Dij76-35, Dij95-28, Dijxx-28]. Tree [HB94, AHH^{+}95, Dij76-55, Dij92b, Dij97k, Dijxx-94, HKS93, VPM^{+}10]. tree-structured [Dijxx-94]. trees [Dij60c, Dij80-27, Dij80-32, DS81, Dij96h, LvdS90]. trends [Bra90]. trial [Dij84m, vGDxx]. triangle [Dij97a, Dij99n]. Triangles [Dij91-28, Dij88n, Dij94h, Dij00h]. Tribute [CK02b, She03a, TLR^{+}91]. trickle [Dij85-29]. trifle [Dij78-43]. Triggered [Dij91-29, Dij95-28, Dij00j]. Trip [Dij73u, Dij73v, Dij73t, Dij73s, Dij74u, Dij74x, Dij74w, Dij74t, Dij74v, Dij75-53, Dij75-52, Dij78-44, Dij79-33, Dij79-30, Dij79-29, Dij79-31, Dij79-32, Dij80-44, Dij80-47, Dij80-45, Dij80-46, Dij81-28, Dij81z, Dij81-27, Dij81-30, Dij81-31, Dij81-29, Dij81-32, Dij81y, Dij81-33, Dij82-87, Dij82c, Dij82-57, Dij82b, Dij82-109, Dij82-106, Dij82-110, Dij82-108, Dij82-112, Dij82-114, Dij82-113, Dij82-107, Dij82-111, Dij82-54, Dij82-51, Dij82-49, Dij82-55, Dij82-52, Dij82-50, Dij82-56, Dij82-53, Dij83x, Dij83-27, Dij83y, Dij83-28, Dij83z, Dij84u, Dij84t, Dij84w, Dij84y, Dij84z, Dij84s, Dij84x, Dij84v, Dij85-48, Dij85-47, Dij85-50, Dij85-49, Dij85-45, Dij85-46, Dij86-44, Dij86-45, Dij87y, Dij88r, Dij88s, Dij70i, Dij71h, Dij72r, Dij73w, Dij73x, Dij74y, Dij76-37, Dij95-29, Dij01l, Dij71e, Dij72n, Dij72q, Dij73k]. Trip [Dijxx-76]. triples [Dij94x]. Tripreport [Dij75-54, Dij75-55, Dij75-57, Dij75-56, Dij76-73, Dij76-78, Dij76-72, Dij76-75, Dij76-77, Dij76-71, Dij76-74, Dij76-76, Dij77-39, Dij77-38, Dij77-37, Dij77-40, Dij77-41, Dij78-49, Dij78-46, Dij78-47, Dij78-48, Dij78-45, Dij82-60, Dij82-59, Dij82-62, Dij82-63, Dij82-58, Dij82-61]. trommelpagina’s [Dij64b]. trommelpaginatransporten [Dij63q]. Truths [Dij82-79, Dij75-27, Dij82t]. try [Dij99i]. Tsunami [UNTH14]. Tuesday [Dij78-42, Dij85b, Dij80-27, Dij80y, Dij82s]. Tulsa [Dij85-49]. Turing [Ash87, Gol02, Day12a, Dij78-37, Wil80b]. Turner [Dij80f, Dij80-38, Dij92y]. Turns [Coc02, KYTY09]. Turski [Dij80-41]. tussen [Dijxx-41]. tutorial [Dij79-34, Dij82-64]. tutoring [NB93, NBK95]. Twaalfde [Dij75-58, Dij76-79]. twee [Dij75-51]. Tweede [Dij75-59, Dij76-81, Dij76-80, Dij77-42, Dijxx-95, Dij76c, Dijxx-96]. Twelfth [Dij75-58, Dij76-79]. Twentieth [MHR80]. Twenty [Dij87z, Ash87, Dij88u]. Twenty-eight [Dij87z]. twins [Dij93j]. Two [Bur80, CF04, Dij63b, Dij75-60, Dij76-82, Dij77-43, Dij79-35, Dij82-91, Dij82-93, Dij82-65, Dij82-104, Dij91-30, Dij96-42, HZW09, MCX13, Zem90, BBM07b, Dij59c, Dij74s, Dij75-51, Dij76-68, Dij76-89, Dij77r, Dij77t, Dij79l, Dij80l, Dij80y, Dij81d, Dij82-27, Dij76-69, Dij88e, Dij89-42, Dij94s, Dij96-32, Dij97i, DvdS06, Dijxx-87, HKS93, Trä95, Yap83]. two-dimensional [Dij79l, Dij82-27]. Two-Function [HZW09]. Two-Level [Dij82-104, Dij74s, Dij76-68, Dij76-69, Dijxx-87]. Two-levels-greedy [CF04]. two-person [Dij81d]. two-process [BBM07b]. two-tree [HKS93]. Tyne [Dij77-40, Dij80-46, Dij82-111, Dij82-52, Dij82-61]. Type [DS06, Dij56a]. typed [ST99]. Types [WG01, Dij80y]. typewriter [Dij85s]. Typing [DS04]. typprogramma [Dij56a].

U.K [Dij76-78, Dij78-44]. U.S.A [Dij75-52, Dij76-78, Dij78-44, Dij80-47, Dij82-106, Dij82-49]. U.S.A. [Bir80, Dij74y]. Übersetzer [Dij61a]. uit [Dij60c, vWBD56]. uitg [Dij70b]. uitreiking [Dijxx-92]. uitvoer [Dij54b]. UK [BMW93]. Ulam [Dij95v]. Ullman [Dij76-40]. Ulrich [Dij88t, Dij99o]. ultra [DFR74]. ultra-high [DFR74]. unary [Dij94g]. unavoidable [Dij92-31, Dij01m]. uncle [Dij02f]. uncomfortable [KYTY09]. unconstrained [DvG84g]. undelivered [Mos10]. understand [Dij86-48]. understanding [Dijxx-52]. undeserved [Dij91-31]. undirected [Dij77p]. unexpected [Dij93s]. unfinished [Dij85n, Dij95u]. unfortunate [Dij82-30, Dij95-30]. Unibags [Dij81t]. unification [Dij92-29]. Unified [Che01]. unify [Dij62b]. unimportant [Dij89r]. union [Dij78-38]. unique [Dij95r]. unit [BDS82]. United [Mil89, Dij87d]. units [Dij75-43]. universal [DS82g, Dij86-38, Dij90o]. University [Dij79-33, Dij93t, Dijxx-77, TLR^{+}91, Dij77c, Boo80, Dij92u, Lav80]. Unlikely [Dij82-116, Dij77-47, Dij78-50]. unordered [Dij95-32]. unspecified [Dij73b]. Unstructured [CCD86]. Unsupportive [DH01]. Untitled [Dij89-44]. updates [DLMB99]. Updating [Dij76-83]. upon [Dij74s, Dij76-68, Dij77-40, Dij80-46, Dij82-111, Dij82-52, Dij82-61, Dij82-104, Dij76-69, Dijxx-87]. Upsequence [DMS83]. urban [Cha10]. Uri [Dij89-41]. USA [Dij74w, Dij76-73, Dij77-41, Dij81-33, Dij82-55, Dij83-28, Dij83z, Dij84y, Dij84x, Dij72n]. Use [BMST03, Dij82-104, MS01, BGM90, Dij74s, Dij76-68, Dij76-69, Dij85u, Dij92h, Dij95-30, Dijxx-87, RR64]. Used [Par01b, Dij96-39, zFmLcWcJ10]. useful [Dij79h, DvG84f, Dijxx-53]. useless [DFxx]. User [Dij84-27, Dij81s]. User-friendly [Dij84-27]. Users [Dij82-95, Dij77u]. Using [Dij80-37, KLK08, LSMS12, MB97, SS10, Sha81, Smi11, ANGK09, BBM07c, Dij78-39, DLMB99, EBR09, HHM94, ŁMB02, Nep00, PRM11, PPS^{+}13, PS14, SPPR92, SVSR04]. USSR [Dij76-75, Dij82-112, Dij82-62, ESB80]. UT [Ano12, Dij93t]. Utrecht [Dij82-56]. UTS [Rua90].

V-implementation [Dij77-33]. Valley [Dij78-48]. value [Dij91-30]. valued [Dijxx-29]. values [Dij95i]. Variability [Dij69c, Dij76i, Dij69d, Dijxxk]. Variable [Bai86, Bis85, Dij67f, Dij74k]. variables [Dij76f, Dij76r, Dijxxg]. variant [Dij76-63, Dij80q, Kes88]. Variations [Dij75-61, Dij82-66, Hoa82]. varying [Lee06]. vascular [VPM^{+}10]. VC [ZXZ^{+}09]. vector [Dijxx-34, PRM11]. vectorgeheugen [Dijxx-34]. vehicle [EKP96, zFmLcWcJ10, PPS^{+}13]. Veltkamp [Dij68f]. Venture [Dij86-46]. Verarbeitung [BS01]. Verfahren [BS01]. Verfahrens [BS01, Str72]. verfijningen [Dijxxo]. verification [MP90]. verify [BGM90]. Verifying [SWS^{+}13, Swi11]. Verkavelde [Dij79-36, Dij79-37]. verkenning [Dijxxw]. Verlag [Dij71f, Dij77b]. verlos [Dijxx-30]. Version [Nep96, DLM^{+}75c, Dij87u]. Verslag [Dij65j, Dij65i, Dij67g, Dij69k, Dij70i, Dij71h, Dij72q, Dij72r, Dij73x, Dijxx-97, Dijxx-96, Dij65d, Dijxx-43]. versus [Dij74d, Dij78-33, Dij88g, Dijxxl, Neu90]. vertaalproces [Dij64i]. vertaler [Dij63n]. vertices [Bar98, vT90]. VERTROUWELIJK [Dij62d, Dij68n]. vervanging [Dij66c]. Vervolg [Dij63k, Dijxx-46]. Very [Dij80-48, Dij82-100, Dij78-40, Dij82-67, Dij89-45, Dij93c, Dij94j, Dij94-28, Dij95s, Dij96-40, Dij68n, Dij65d]. via [Dij97j]. Viejo [Dij79-31]. vier [Dij71b, Dij71c]. Vierde [Dij75-62, Dij75-63, Dij76-85, Dij76-84, Dij77-44].

View [Che01, CM88, Kil79, Par78]. viewed [Mis01]. Views [Coc02, Dij75-60]. Vijfde [Dij75-64, Dij76-87, Dij76-86, Dij77-45, Dijxx-98]. villages [Dij76-53, Dijxx-70]. vindicated [Dij90m]. Visit [Dij82-57, Bro90a, Dij65d, Dij65i, Dij67g, Dij69k, Dij72n, Dij75-53, Dijxx-97, Dij65j, Dij72b]. visits [Dijxx-77]. visualization [Dij91x]. Visualizing [EBR09]. Visuals [Dij86-46]. viz [Dij87v]. Vizing [DR90a, DR90b, RD92]. VLSI [PRV09]. Voermans [Dij91f]. volatile [Dij79t]. Volume [Dij94v]. voor [DDvW57, Dij55a, Dij55b, Dij56b, Dij56c, Dij56d, Dij56e, Dij57c, Dij57d, Dij64j, Dij65b, Dij68c, Dij70g, Dij80t, Dij93t, Dijxx-55, Dijxxv]. voorgestelde [Dij71b, Dij71c]. voorjaar [Dij75s, Dij76e, Dij76l, Dij76k, Dij76q, Dij76-36, Dij76-81, Dij76-67, Dij76-79, Dij76-85, Dij76-87, Dij76-92, Dij76-93, Dij77a, Dij77e, Dij77f, Dij77-42, Dij77-44, Dij77-45, Dij77-48, Dij77-49]. voorkoming [Dijxxu]. Voorstel [Dij64j]. Voorwoord [Dij93t, Dijxx-56]. vote [Dij89v, Dij91l]. Vries [Dij96-34]. vs [Kil79].

W [AMM^{+}12, Alp18, Ano06b, Bar75, Bar74, BB03, Bör94a, Bör94b, BFG^{+}02, CK02b, Che02a, Che02b, Che16, Cre02, Dij71e, Dij71h, Dij72b, Dij72n, Dij72r, Dij73k, Dij73t, Dij73s, Dij74u, Dij74x, Dij74w, Dij74t, Dij74v, Dij75-53, Dij75-52, Dij75-54, Dij75-55, Dij75-57, Dij75-56, Dij76u, Dij76-37, Dij76-73, Dij76-78, Dij76-72, Dij76-75, Dij76-77, Dij76-71, Dij76-74, Dij76-76, Dij77-39, Dij77-38, Dij77-37, Dij77-40, Dij77-41, Dij78-44, Dij78-49, Dij78-46, Dij78-47, Dij78-48, Dij78-45, Dij79-33, Dij79-30, Dij79-29, Dij79-31, Dij79-32, Dij80-44, Dij80-47, Dij80-45, Dij80-46, Dij81-28, Dij81z, Dij81-27, Dij81-30, Dij81-31, Dij81-29, Dij81-32, Dij81y, Dij81-33, Dij82-87, Dij82-57, Dij82b, Dij82-109, Dij82-106, Dij82-110, Dij82-108, Dij82-112, Dij82-114, Dij82-113, Dij82-107, Dij82-111, Dij82-54, Dij82-51, Dij82-49, Dij82-55, Dij82-52, Dij82-50]. W [Dij82-56, Dij82-60, Dij82-59, Dij82-62, Dij82-63, Dij82-53, Dij82-58, Dij82-61, Dij83x, Dij83-27, Dij83y, Dij83-28, Dij83z, Dij84u, Dij84t, Dij84w, Dij84y, Dij84z, Dij84s, Dij84x, Dij84v, Dij85-48, Dij85-47, Dij85-50, Dij85-49, Dij85-45, Dij85-46, Dij86-44, Dij86-45, Dij87y, Dij88r, Dij88s, Dijxx-76, FvGGM90, Flo77, Llo74, Mil89, Mis10, Moo02, SL88, Sch18, She03a, TLR^{+}91]. W. [Ano06a, Dij79j, Dij79s, Dij79w, Dij86-30]. W.G.2.1 [Dijxx-43]. W.G.2.3 [Dij73s, Dij74x, Dij74v, Dij75-57, Dij82-109, Dij82-51, Dijxx-76]. W.O. [Dij75h]. Waarom [Dij76-88, Dijxx-99]. walk [Mis01]. Walker [Dij91h]. want [Dijxx-99]. war [Dij93r, Tat23]. Warning [Dij82-85, Dij82-88, Dij75-33, Dij75-39]. Warshall [FURD13, DK13, HM12]. Warwick [Dij92u, Dijxx-76]. Washington [Dij73x]. Wasserman [Dij77b]. wat [Dij77o]. wave [ZTZY12]. Waves [Dij97m]. Way [Dij82-99, LX10, BD89, Dij77-30, Dij89e, Dij93m, Dij93n, Dij96-33]. ways [Dij91-30]. Weak [Dij82-94, BES95, Dij78-32]. Weakest [DS82h, BBM07a, BBM07c, Dij95q, Luk94]. web [Dij04]. Webster [Dij77u, Dij82-95]. Wegner [DS72c]. weight [Dij71d, Dij92-31, RPGS14]. Weighted [BMST03, MS01, MP91, Bar98, Dij00f, Dij00g]. Welcome [Dij81-34, Dij81-35]. Weledelgestrenge [Dijxx-100, Dijxx-39]. Weledelgetrenge [Dij83j]. Welfare [Dij88b]. Well [CCD86, Dij84-28, DvG90a, DvG90b, Dij91-32, IdKM06, vGD92, Dij80z]. Well-Being [IdKM06]. Well-formed [CCD86]. Well-founded [Dij84-28]. Well-foundedness [DvG90a, DvG90b, Dij91-32, vGD92]. well-shaped [Dij80z]. wellfounded [Dij96-41, Dij00d, Dij00e]. were [Cre02]. Wesley [Mil89]. Westboro [Dij85-50]. Westboro-Boston [Dij85-50]. Western [Hus80]. Wetenschapsbeleid [Dij67a, Dij68b, Dijxxa]. WF [Dij96-36]. WG2.3 [Dij70c]. Where [Dij85-52]. Whether [Nag18]. which [PJ89]. WHIRLWIND [Eve80]. Whitaker [Dij78w]. white [EBR09, Dij99k]. Who [Dij86-47, Dij92-30, Mar02, Ano02a]. whole [Kot90]. whose [Dij99k, Knu90]. wife [Dij96-43]. wife-swapping [Dij96-43]. wil [Dijxx-99]. William [Dij78w]. Wiltink [Dij86h]. Wim [Dij95-28]. Winner [Gol02, Sch18]. winters [Dij89-42]. Wireless [JNK11, KKNI14, MTZ12]. Wirth [PGB00]. Wisdom [Gol02]. Wise [Dij82-104, Dij74s, Dij76-68, Dij76-69, Dijxx-87, KGZ06]. wiskunde [Dij76-88, vWBD56, Dijxxc]. Wiskundefreaks [Dij86a]. wiskundig [Dijxx-36]. wiskundige [Dij72p, Dijxx-38]. Wit [Gol02]. within [Dij77c, DvdS06]. Without [DvG86c, Dij02a, DvG84b, DvG84c, DvG84d, DvG84e, DvG85b]. witness [Dij96-31]. Władisław [Dij80-41]. WLOG [Dij95-32]. WO [Dij75h]. Wokingham [Mil89]. Woodenman [Dij75j]. Word [Dij81-34, Dij81-35]. words [D^{+}66b]. Work [AH22, Dij82-116, Goo80, Wil80b, Dij77-47, Dij78-50]. Workflow [LWQ09]. Working [Dij73s]. Works [Mal19]. Workshop [Dij65i, Dij77b]. worth [Dij95-27, Dij99h]. Woude [Dij88f, Dij89h, Dij90d]. Would [Ham80]. wp [BES95, Mor90a]. write [Dij84f]. Writing [Dij76-89, Dij89-35, Sch90a]. Writings [Dij72o, Dij82-69]. Written [Dij78-51]. WTM [PRM11]. Wybe [Ano02b, Ano02c, Ano12, Apt02, AH22, Hoa03, MR03, vL04].

X [Dij61a, Dij61a, Dij79g]. X1 [Dij63b, Dij63f, Kru03, Dij61c]. X8 [Dij65e, Dij63k, Dij65b, Dij65c, Dij66c, Dijxxt, Dijxx-45, Dijxx-46, Dijxxv].

year [Dij85-43]. yearbook [Dij93t]. Years [DM97, Hoa22, Ash87, DvG83b, Dij87z, Dij88u, Dij89z, Dij96-29, Dij96-30]. YELLOW [Dij78-31, Dij78f]. yield [Dij94d]. YoP [Dij87f, Dij88a, Dij90g, Dij90h]. YoP-institute [Dij87f, Dij90g, Dij90h]. York [Dij65j, Dij77b, Dij65j]. York/Heidelberg/Berlin [Dij77b]. Yossi [Dij98c]. young [Dij89-44].

Z. [Dij94-29]. Z8 [Dij66b]. Z8-commissie [Dij66b]. Z8-commission [Dij66b]. Zagier [Dij93b]. zake [Dij67a, Dij72p]. Zakopane [Win78]. zero [BD89, Dij82-68]. zero-termination [BD89]. Zesde [Dij75-65, Dij75-66, Dij76-92, Dij76-91, Dij77-48]. Zevende [Dij75-67, Dij75-68, Dij76-93, Dij77-49]. Zhendong [Dij94w]. Zhiming [Dij92u]. zich [Dij76-88]. Zohar [Dij76-31]. Zonneveld [Kru03]. Zuckerman [Dij01p]. zur [BS01, Str72]. Zürich [Dij78-49, Dij82-57, Dij84z, Dij88s, Dij75-53]. Zuse [Bau80]. Zwanzig [Dij88u]. Zweiter [Dij72c]. zwemmen [Dij89r].

References

Asadi:2011:NGO

[AAEG11] S. Asadi, V. Azimirad, A. Eslami, and A. Ghanbari. A novel global optimal path planning and trajectory method based on adaptive Dijkstra-immune approach for mobile robot. In 2011 IEEE/ASME International Conference on Advanced Intelligent Mechatronics (AIM), pages 1093–1098. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2011. ISBN 1-4577-0838-8. ISSN 2159-

6247. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=6027073.

Apt:1990:PTP

[AdBO90] Krzysztof R. Apt, Frank S. de Boer, and Ernst-Rüdiger Olderog. Proving termination of parallel programs. In Feijen et al. [FvGGM90], chapter 0, pages 0–6. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN QA76 .B326

1990. URL http://www.zentralblatt-math.org/zmath/

en/search/?an=0718.68004. Contains important treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

Almstrum:2001:ESI

[AG01] Vicki L. Almstrum and David Gries. From the Editors of this special issue. Information Processing Letters, 77 (2–4):49–51, February 28, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://

www.elsevier.nl/gej-ng/10/23/20/68/29/24/abstract.
html;   http://www.elsevier.nl/gej-ng/10/23/20/68/

29/24/article.pdf. Papers from the Symposium “In Pursuit of Simplicity” in honor of Edsger W. Dijkstra held at the University of Texas, Austin, TX, May 12–13, 2000, Inform. Process. Lett. 77 (2001), no. 2–4.

Apt:2022:EWD

[AH22] Krzysztof R. Apt and Tony Hoare, editors. Edsger Wybe Dijkstra: His Life, Work, and Legacy, volume 45 of ACM books. ACM Press, New York, NY 10036, USA, 2022. ISBN 1-4503-9771-9 (paperback), 1-4503-9772-7 (epub), 1-4503-9773-5 (hardcover), 1-4503-9774-3 (e-book). ISSN 2374-6777. xxiv + 550 pp. LCCN DS119 .A685 2022.

Alpert:1995:PDT

[AHH^{+}95] C. J. Alpert, T. C. Hu, J. H. Huang, A. B. Kahng, and D. Karger. Prim–Dijkstra tradeoffs for improved performance-driven routing tree design. IEEE Transactions on Computer-Aided Design of Integrated Circuits and Systems, 14(7):890–896, 1995. CODEN ITCSDI. ISSN 0278-0070 (print), 1937-4151 (electronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?arnumber=391737.

Alpert:1993:DCP

[AHHK93] C. J. Alpert, T. C. Hu, J. H. Huang, and A. B. Kahng. A direct combination of the Prim and Dijkstra constructions for improved performance-driven global routing. In 1993 IEEE International Symposium on Circuits and Systems, 1993, ISCAS ’93, volume 3, pages 1869–1872. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-7803-

1281-3. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=394112;   http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=693037.

Alpern:2018:EWD

[Alp18] Bowen Alpern. Edsger W. Dijkstra: The man behind the prize. SIGACT News (ACM Special Interest Group on Automata and Computability Theory), 49(4):64–65, December 2018. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Aguilera:2012:AEW

[AMM^{+}12] Marcos K. Aguilera, Dahlia Malkhi, Keith Marzullo, Alessandro Panconesi, Andrzej Pelc, and Roger Wattenhofer. Announcing the 2012 Edsger W. Dijkstra Prize in Distributed Computing. ACM SIGARCH Computer Architecture News, 40(4):1–2, September 2012. CODEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Anastopoulos:2009:EEA

[ANGK09] N. Anastopoulos, K. Nikas, G. Goumas, and N. Koziris. Early experiences on accelerating Dijkstra’s algorithm using transactional memory. In IEEE, editor, IPDPS 2009 Rome: program, 23rd IEEE International Parallel and Distributed Processing Symposium: May 25–29, 2009 Rome, Italy, pages 1–8. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2009. ISBN 1-4244-3751-2 (print), 1-4244-3750-4. ISSN 1530-2075. URL

http://ieeexplore.ieee.org/servlet/opac?punumber=
5136864;   http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=5161103.

Anonymous:1980:A

[Ano80a] Anonymous. Acknowledgments. In Metropolis et al. [MHR80], page xviii. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5 .I63 1976. URL http://www.

sciencedirect.com/science/article/pii/B9780124916500500071.

Original versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Anonymous:1980:Ca

[Ano80b] Anonymous. Contributors. In Metropolis et al. [MHR80], page ii. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN

QA75.5 .I63 1976. URL http://www.sciencedirect.com/
science/article/pii/B9780124916500500010. Original

versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Anonymous:1980:Cb

[Ano80c] Anonymous. Copyright. In Metropolis et al. [MHR80], page iv. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN

QA75.5 .I63 1976. URL http://www.sciencedirect.com/
science/article/pii/B9780124916500500034. Original

versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Anonymous:1980:D

[Ano80d] Anonymous. Dedication. In Metropolis et al. [MHR80], pages v–x. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book).

LCCN QA75.5 .I63 1976. URL http://www.sciencedirect.
com/science/article/pii/B9780124916500500046. Origi-

nal versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Anonymous:1980:FM

[Ano80e] Anonymous. Front matter. In Metropolis et al. [MHR80], pages iii–ii. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book).

LCCN QA75.5 .I63 1976. URL http://www.sciencedirect.
com/science/article/pii/B9780124916500500022. Origi-

nal versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Anonymous:1980:GI

[Ano80f] Anonymous. General index. In Metropolis et al. [MHR80], pages 661–683. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book).

LCCN QA75.5 .I63 1976. URL http://www.sciencedirect.
com/science/article/pii/B9780124916500500472. Origi-

nal versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Anonymous:1980:ICF

[Ano80g] Anonymous. Index of claimed firsts and inventions. In Metropolis et al. [MHR80], pages 684–685. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5 .I63 1976.

URL http://www.sciencedirect.com/science/article/

pii/B9780124916500500484. Original versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Anonymous:1980:IN

[Ano80h] Anonymous. Index of names. In Metropolis et al. [MHR80], pages 686–693. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book).

LCCN QA75.5 .I63 1976. URL http://www.sciencedirect.
com/science/article/pii/B9780124916500500496. Origi-

nal versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Anonymous:1980:LC

[Ano80i] Anonymous. List of contributors. In Metropolis et al. [MHR80], pages xi–xiii. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5 .I63 1976. URL http://www.

sciencedirect.com/science/article/pii/B9780124916500500058.

Original versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Anonymous:1980:P

[Ano80j] Anonymous. Preface. In Metropolis et al. [MHR80], pages xv–xvii. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book).

LCCN QA75.5 .I63 1976. URL http://www.sciencedirect.
com/science/article/pii/B978012491650050006X. Origi-

nal versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Anonymous:2002:EDP

[Ano02a] Anonymous. Edsger Dijkstra: Pioneering computer programmer who made his subject intellectually respectable. The Guardian, ??(??):??, August 19, 2002. ISSN 0261-3077 (print), 1756-3224 (electronic). URL http://www.

theguardian.com/news/2002/aug/19/guardianobituaries.
obituaries1.

Anonymous:2002:EWD

[Ano02b] Anonymous. Edsger Wybe Dijkstra: 11 May 1930 – 6 August 2002. Web site., August 2002. URL http://www.cs.utexas.

edu/users/EWD/obituary.html.

Anonymous:2002:OOJ

[Ano02c] Anonymous. Obituary: Ole-Johan Dahl, 1931–2002; Edsger Wybe Dijkstra, 1930–2002; Kristen Nygaard, 1926–2002. ACM Transactions on Software Engineering and Methodology, 11(4):385, 2002. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic).

Anonymous:2006:KTE

[Ano06a] Anonymous. Dijkstra, E. W. In De kleine TU/e encyclopedie 1956–2006, pages 99–102. ????, ????, 2006. URL http://

www.win.tue.nl/lotgevallen/em/KleineTUEEncyclopedie19562006_
EWDijkstra.pdf.

Anonymous:2006:EWD

[Ano06b] Anonymous. Edsger W. Dijkstra Archive: Obituaries and memorials. Web site., December 3, 2006. URL http://www.

cs.utexas.edu/users/EWD/obituaries/obits.html.

Anonymous:2012:UCS

[Ano12] Anonymous. UT Computer Sciences Department’s obituary: Edsger Wybe Dijkstra: 11 May 1930–6 August 2002. Edsger W. Dijkstra Archive document., March 6, 2012.

Anonymous:2013:DPD

[Ano13] Anonymous. 2013 Dijkstra Prize in Distributed Computing to Nati Linial: 2013 Principles of Distributed Computing Doctoral Dissertation Award to Shiri Chechik and Danupon Nanongkai. SIGACT News (ACM Special Interest Group on Automata and Computability Theory), 44(4):81–82, December 2013. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Anwar:2003:IKB

[Anw03] M. A. Anwar. Integrating knowledge-base and Dijkstra’s algorithm for finding best alternate route dynamically. In 7th International Multi Topic Conference, 2003. INMIC 2003., pages 428–433. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2003. ISBN

0-7803-8183-1. URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=1416765.

Apt:2002:EWD

[Apt02] Krzysztof R. Apt. Edsger Wybe Dijkstra (1930–2002): A portrait of a genius. Formal Aspects of Computing, 14(2):92–98, December 2002. CODEN FACME5. ISSN 0934-5043 (print),

1433-299X (electronic). URL http://link.springer.com/
article/10.1007/s001650200029;   http://tinman.cs.
gsu.edu/~raj/8710/f03/dijkstra.pdf.

Aretz:1990:MSC

[Are90] F. E. J. Kruseman Aretz. Maximal strong components: an exercise in program presentation. In Feijen et al. [FvGGM90], chapter 29, pages 251–261. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN QA76 .B326 1990. URL

http://www.zentralblatt-math.org/zmath/en/search/?

an=0718.68004. Contains important treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

Ashenhurst:1987:ATA

[Ash87] Robert L. Ashenhurst, editor. ACM Turing Award Lectures: the first twenty years, 1966–1985. ACM Press anthology series. ACM Press and Addison-Wesley, New York, NY 10036, USA and Reading, MA, USA, 1987. ISBN 0-201-07794-9. xviii + 483 pp. LCCN QA76.24 .A33 1987.

Adelson-Velsky:2002:PSG

[AVL02] George M. Adelson-Velsky and Eugene Levner. Project scheduling in AND–OR graphs: a generalization of Dijkstra’s algorithm. Mathematics of Operations Research, 27 (3):504–517, 2002. CODEN MOREDQ. ISSN 0364-765x (print), 1526-5471 (electronic). URL http://www.jstor.

org/stable/3690449.

Backus:1980:PAS

[Bac80] John Backus. Programming in America in the 1950s — some personal impressions. In Metropolis et al. [MHR80], pages 125–135. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book).

LCCN QA75.5 .I63 1976. URL http://www.sciencedirect.
com/science/article/pii/B9780124916500500174. Origi-

nal versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Backhouse:1990:RF

[Bac90] Roland C. Backhouse. On a relation on functions. In Feijen et al. [FvGGM90], chapter 1, pages 7–18. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN QA76 .B326

1990. URL http://www.zentralblatt-math.org/zmath/

en/search/?an=0718.68004. Contains important treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

Bailes:1986:SCV

[Bai86] Paul A. Bailes. Static checking of variable handling in Dijkstra’s Guarded Commands Language. Computer Languages, 11(3-4):123–142, ???? 1986. CODEN COLADA. ISSN 0096-0551 (print), 1873-6742 (electronic).

Bandyopadhyay:2007:MFS

[Ban07] Anup Kumar Bandyopadhyay. Modeling fairness and starvation in concurrent systems. ACM SIGSOFT Software Engineering Notes, 32(6):2:1–2:??, November 2007. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Bandyopadhyay:2009:FCC

[Ban09] Anup Kumar Bandyopadhyay. Fairness and conspiracy concepts in concurrent systems. ACM SIGSOFT Software Engineering Notes, 34(2):1–8, March 2009. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Barth:1974:BRJ

[Bar74] W. Barth. Book review: O. J. Dahl, W. Dijkstra und C. A. R. Hoare, Structured Programming. Computing: Archiv für informatik und numerik, 12(3):285, September 1974. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic).

Barron:1975:BRB

[Bar75] D. W. Barron. Book review: Structured Programming, O.-J. Dahl, E. W. Dijkstra and C. A. R. Hoare, Academic Press, London, 1972. No. of pages: 220. Price: £4.20. Software — Practice and Experience, 5(3):314–315, July 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barbehenn:1998:NCD

[Bar98] Michael Barbehenn. A note on the complexity of Dijkstra’s algorithm for graphs with weighted vertices. IEEE Transactions on Computers, 47(2):263, February 1998. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

URL http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=
&arnumber=663776.

Bauer:1980:BZR

[Bau80] Friedrich L. Bauer. Between Zuse and Rutishauser — the early development of digital computing in Central Europe. In Metropolis et al. [MHR80], pages 505–524. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5

.I63 1976. URL http://www.sciencedirect.com/science/
article/pii/B978012491650050037X. Original versions of

these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Bauer:1990:ESN

[Bau90] F. L. Bauer. Efficient solution of a non-monotonic inverse problem. In Feijen et al. [FvGGM90], chapter 2, pages 19–26. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN QA76 .B326

1990. URL http://www.zentralblatt-math.org/zmath/

en/search/?an=0718.68004. Contains important treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

Bauer:1979:PCI

[BB79] Friedrich Ludwig Bauer and M. Broy, editors. Program construction: international summer school, volume 69 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CODEN LNCSD9. ISBN 0-387-09251-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

QA76.6 .P75117. URL http://link.springer-ny.com/
link/service/series/0558/tocs/t0069.htm;   http:/
/www.springerlink.com/content/978-0-387-09251-5;
http://www.springerlink.com/openurl.asp?genre=issue&
issn=0302-9743&volume=69. “The international summer

school took place from July 26 to August 6, 1978, in Marktoberdorf … and was sponsored by the NATO Scientific Affairs Division under the 1978 Advanced Study Institutes programme.”.

Boskovic:1999:TND

[BB99a] Ronald Boskovic and Paul Belcher. 83.03 A teaching note on Dijkstra’s shortest path algorithm. Mathematical Gazette, 83(496):94–97, March 1999. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL http://www.

jstor.org/stable/3618691.

Boskovic:1999:NTN

[BB99b] Ronald Boskovic and Paul Belcher. Notes: 83.03 A teaching note on Dijkstra’s shortest path algorithm. Mathematical Gazette, 83(497):251–254, July 1999. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL http:/

/www.jstor.org/stable/3619048.

Bauer:2003:EWD

[BB03] Friedrich L. Bauer and Manfred Broy. Edsger W. Dijkstra — Acta Informatica and Marktoberdorf. Acta Informatica, 39 (3):141–142, March 2003. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Banerjee:2007:ADW

[BBM07a] Jayasri Banerjee, Anup Kumar Bandyopadhyay, and Ajit Kumar Mandal. Application of Dijkstra’s weakest precondition calculus to Dining Philosophers problem. ACM SIGSOFT Software Engineering Notes, 32(4):2:1–2:??, July 2007. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Banerjee:2007:OET

[BBM07b] Jayasri Banerjee, Anup Kumar Bandyopadhyay, and Ajit Kumar Mandal. Ordering of events in two-process concurrent system. ACM SIGSOFT Software Engineering Notes, 32(4): 1:1–1:??, July 2007. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Banerjee:2007:SID

[BBM07c] Jayasri Banerjee, Anup Kumar Bandyopadhyay, and Ajit Kumar Mandal. Some investigations on deadlock freedom issues of a cyclically connected system using Dijkstra’s weakest precondition calculus. ACM SIGPLAN Notices, 42(5):10–15, May 2007. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bron:1966:SCO

[BD66] Coen Bron and Edsger W. Dijkstra. Source code of THE Operating System. Web archive, 1966. URL http:/

/archive.computerhistory.org/resources/text/Knuth_
Don_X4100/PDF_index/k-1-pdf/k-1-C1063.6-source-THE-
os.pdf.

Bron:1979:DPI

[BD79] C. Bron and E. J. Dijkstra. A discipline for the programming of interactive I/O in Pascal. ACM SIGPLAN Notices, 14(12): 59–61, December 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bron:1989:BWC

[BD89] C. Bron and E. J. Dijkstra. A better way to combine efficient string length encoding and zero-termination. ACM SIGPLAN Notices, 24(6):11–19, June 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Broy:2001:LE

[BD01a] Manfred Broy and Ernst Denert. Letters to the Editor. In Pioneers and Their Contributions to Software Engineering: sd&m Conference on Software Pioneers, Bonn, June 28/29, 2001, Original Historic Contributions [BD01b], page 295. ISBN 3-540-42290-0 (print), 3-642-48354-2 (e-book). LCCN

QA76.758. URL http://link.springer.com/chapter/10.
1007/978-3-642-48354-7_11.

Broy:2001:PTC

[BD01b] Manfred Broy and Ernst Denert, editors. Pioneers and Their Contributions to Software Engineering: sd&m Conference on Software Pioneers, Bonn, June 28/29, 2001, Original Historic Contributions. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-42290-0 (print), 3-642-48354-2 (e-book). LCCN QA76.758.

Broy:2002:SPC

[BD02] Manfred Broy and Ernst Denert, editors. Software Pioneers: Contributions to Software Engineering. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. ISBN 3-540-43081-4, 3-642-48354-2 (e-book), 3-642-59412-3 (e-book), 3-642-59413-1, 3-642-63970-4 (print). 728 pp. LCCN QA76.2.A2 S62

2002. URL http://aleph.unisg.ch/hsgscan/hm00048094.
pdf;   http://link.springer.com/10.1007/978-3-642-
59412-0;   http://www.loc.gov/catdir/enhancements/
fy0817/2002070484-d.html;   http://www.springerlink.
com/content/978-3-642-48354-7;   z3950.loc.gov:7090/
Voyager.

Bekker:1993:MDS

[BDB93] H. Bekker, E. J. Dijkstra, and H. J. C. Berendsen. Molecular dynamics simulation on an i860 based ring architecture. Supercomputer, 10(2):4–10, March 1993. CODEN SPCOEL. ISSN 0168-7875.

Bovenkamp:2001:CMA

[BDBR01] E. G. P. Bovenkamp, J. Dijkstra, J. G. Bosch, and J. H. C. Reiber. Collaborative multi-agent IVUS image segmentation. Lecture Notes in Computer Science, 2208:1185–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer-ny.com/
link/service/series/0558/bibs/2208/22081185.htm;
http://link.springer-ny.com/link/service/series/0558/
papers/2208/22081185.pdf.

Bauer:2010:CHG

[BDn^{+}10] Reinhard Bauer, Daniel Delling, Peter Sanders n, Dennis Schieferdecker, Dominik Schultes, and Dorothea Wagner. Combining hierarchical and goal-directed speed-up techniques for Dijkstra’s algorithm. ACM Journal of Experimental Algorithmics, 15(1):2.3:1–2.3:31, March 2010. CODEN ???? ISSN 1084-6654.

Bron:1985:NCI

[BDR85] C. Bron, E. J. Dijkstra, and T. J. Rossingh. A note on the checking of interfaces between separately compiled modules. ACM SIGPLAN Notices, 20(8):60–63, August 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bron:1982:MMU

[BDS82] C. Bron, E. J. Dijkstra, and S. D. Swierstra. A memory management unit for the optimal exploitation of a small address space. Information Processing Letters, 15(1):20–22, August 19, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Berghammer:1999:CRC

[Ber99] Rudolf Berghammer. Combining relational calculus and the Dijkstra–Gries method for deriving relational programs. Information Sciences, 119(3-4):155–171, 1999. CODEN ISI-JBC. ISSN 0020-0255 (print), 1872-6291 (electronic).

Berghammer:1995:FDP

[BES95] Rudolf Berghammer, Birgit Elbl, and Ulf Schmerl. Formalizing Dijkstra’s predicate transformer wp in weak second-order logic. Theoretical Computer Science, 146(1–2):185–197, July 24, 1995. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-

2294 (electronic). URL http://www.elsevier.com/cgi-
bin/cas/tree/store/tcs/cas_sub/browse/browse.cgi?
year=1995&volume=146&issue=1-2&aid=1818.

Boyer:2002:MEW

[BFG^{+}02] R. S. Boyer, W. Feijen, D. Gries, C. A. R. Hoare, J. Misra, J. Moore, and H. Richards. In memoriam: Edsger W. Dijkstra 1930–2002. Communications of the Association for Computing Machinery, 45(10):21–22, October 2002. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Boyer:1990:UFS

[BGM90] Robert S. Boyer, Milton W. Green, and J. Strother Moore. The use of a formal simulator to verify a simple real time control program. In Feijen et al. [FvGGM90], chapter 6, pages 54–66. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN

QA76 .B326 1990. URL http://www.zentralblatt-math.
org/zmath/en/search/?an=0718.68004. Contains impor-

tant treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

Bron:1991:SR

[BH91] Coenraad Bron and Wim H. Hesselink. Smoothsort revisited. Information Processing Letters, 39(5):269–276, September 13, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bigelow:1980:CDI

[Big80] Julian Bigelow. Computer development at the Institute for Advanced Study. In Metropolis et al. [MHR80], pages 291– 310. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN

QA75.5 .I63 1976. URL http://www.sciencedirect.com/
science/article/pii/B9780124916500500216. Original

versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Bijlsma:1990:SQB

[Bij90] A. Bijlsma. Semantics of quasi-boolean expressions. In Feijen et al. [FvGGM90], chapter 3, pages 27–35. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN QA76 .B326

1990. URL http://www.zentralblatt-math.org/zmath/

en/search/?an=0718.68004. Contains important treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

Birkhoff:1980:CDS

[Bir80] Garrett Birkhoff. Computing developments 1935–1955, as seen from Cambridge, U.S.A. In Metropolis et al. [MHR80], pages 21–30. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book).

LCCN QA75.5 .I63 1976. URL http://www.sciencedirect.
com/science/article/pii/B9780124916500500113. Origi-

nal versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Bird:1990:SSE

[Bir90] Richard S. Bird. Small specification exercises. In Feijen et al. [FvGGM90], chapter 4, pages 36–43. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN QA76 .B326

1990. URL http://www.zentralblatt-math.org/zmath/

en/search/?an=0718.68004. Contains important treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

Biswas:1985:ASV

[Bis85] P. Biswas. Architectural support for variable addressing in Ada — a design approach. International Journal of Computer and Information Sciences, 14(1):51–72, February 1985. CODEN IJCIAH. ISSN 0091-7036.

Bayer:2001:OML

[BM01] Prof. Dr. R. Bayer and Dr. E. M. McCreight. Organization and maintenance of large ordered indexes. In Broy and Denert [BD01b], pages 41–59. ISBN 3-540-42290-0 (print), 3-642-48354-2 (e-book). LCCN QA76.758. URL http://link.

springer.com/chapter/10.1007/978-3-642-48354-7_3/.

Bast:2003:HDA

[BMST03] Holger Bast, Kurt Mehlhorn, Guido Schäfer, and Hisao Tamaki. A heuristic for Dijkstra’s algorithm with many targets and its use in weighted matching algorithms. Algorithmica, 36(1):75–88, February 2003. CODEN AL-GOEJ. ISSN 0178-4617 (print), 1432-0541 (electronic).

URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0178-4617&volume=36&issue=1&spage=75.

Bird:1993:MPC

[BMW93] R. S. Bird, Carroll Morgan, and Jim Woodcock, editors. Mathematics of program construction: second international conference, Oxford, UK, June 29–July 3, 1992: proceedings, volume 669 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. CODEN LNCSD9. ISBN 3-540-56625-2 (Berlin), 0-387-56625-2 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .M367

1993. DM72.00. URL http://link.springer-ny.com/
link/service/series/0558/tocs/t0669.htm;   http:/
/www.springerlink.com/content/978-3-540-56625-0;
http://www.springerlink.com/openurl.asp?genre=issue&
issn=0302-9743&volume=669. Papers . . . presented at the

Second International Conference on the Mathematics of Program Construction, held at St Catherine’s College, Oxford, during the week of 29 June–3 July, 1992.

Brooks:2001:FS

[BMWC01] Fred Brooks, G. H. Mealy, B. I. Witt, and W. A. Clark. The functional structure of OS/360. In Broy and Denert [BD01b], pages 151–203. ISBN 3-540-42290-0 (print), 3-642-48354-2 (e-book). LCCN QA76.758. URL http://link.springer.

com/chapter/10.1007/978-3-642-48354-7_6/.

Broy:1994:AFC

[BN94] Manfred Broy and Greg Nelson. Adding fair choice to Dijkstra’s calculus. ACM Transactions on Programming Languages and Systems, 16(3):924–938, May 1994. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL http://www.acm.org/pubs/toc/Abstracts/
0164-0925/177727.html.

Bijlsma:1998:DSP

[BN98] Lex Bijlsma and Rob Nederpelt. Dijkstra–Scholten predicate calculus: concepts and misconceptions. Acta Informatica, 35(12):1007–1036, December 1998. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:/

/link.springer-ny.com/link/service/journals/00236/
bibs/8035012/80351007.htm;   http://link.springer-
ny.com/link/service/journals/00236/papers/8035012/
80351007.pdf.

Boasson:1990:ART

[Boa90] Maarten Boasson. Architecture of real-time systems. In Feijen et al. [FvGGM90], chapter 5, pages 44–53. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN QA76 .B326

1990. URL http://www.zentralblatt-math.org/zmath/

en/search/?an=0718.68004. Contains important treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

Boehm:2001:SEE

[Boe01] Barry W. Boehm. Software engineering economics. In Broy and Denert [BD01b], pages 99–150. ISBN 3-540-42290-0 (print), 3-642-48354-2 (e-book). LCCN QA76.758.

URL http://link.springer.com/chapter/10.1007/978-
3-642-48354-7_5/.

BohorquezV:2008:ILA

[Boh08] Jaime A. Bohórquez V. Intuitionistic logic according to Dijkstra’s calculus of equational deduction. Notre Dame Journal of Formal Logic, 49(4):361–384, 2008. CODEN NDJFAM. ISSN 0029-4527 (print), 1939-0726 (electronic).

Booth:1980:CUL

[Boo80] Andrew D. Booth. Computers in the University of London, 1945–1962. In Metropolis et al. [MHR80], pages 551–561. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5

.I63 1976. URL http://www.sciencedirect.com/science/
article/pii/B978012491650050040X. Original versions of

these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Borger:1994:BRBa

[Bör94a] Egon Börger. Book review: Predicate Calculus and Program Semantics by Edsger W. Dijkstra, Carel S. Scholten. Journal of Symbolic Logic, 52(2):673–678, June 1994. CODEN JSYLA6. ISSN 0022-4812 (print), 1943-5886 (electronic).

URL http://www.jstor.org/stable/2275420.

Borger:1994:BRBb

[Bör94b] Egon Börger. Book review: Predicate calculus and program semantics: E. W. Dijkstra and C. S. Scholten, (Springer, Berlin, 1989). Science of Computer Programming, 23(1):91–101, October 1994. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0167642394900027.

Bovet:1986:AMD

[Bov86] Jean Bovet. Une amélioration de la méthode de Dijkstra pour la recherche d’un plus court chemin dans un réseau. Discrete Applied Mathematics, 13(1):93–96, 1986. CODEN DAMADU. ISSN 0166-218X (print), 1872-6771 (electronic).

Buxton:1970:SET

[BR70] J. N. Buxton and B. Randell, editors. Software Engineering Techniques: Report on a Conference sponsored by the NATO Science Committee, Rome, Italy, 27th to 31st October 1969. NATO Scientific Affairs Division, Brussels, Belgium,

April 1970. URL http://homepages.cs.ncl.ac.uk/brian.
randell/NATO/nato1969.PDF. Reprinted in [NRB76]. PDF

file reconstructed by Robert M. McLure (details on page 2).

Braben:1990:EFT

[Bra90] Donald W. Braben. Exploring the future: trends and discontinuities. In Feijen et al. [FvGGM90], chapter 7, pages 67–74. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN QA76 .B326

1990. URL http://www.zentralblatt-math.org/zmath/

en/search/?an=0718.68004. Contains important treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

BrinchHansen:1990:LSR

[Bri90] Per Brinch Hansen. The linear search rediscovered. Structured Programming, 11:53–55, 1990. CODEN STPGEM.

ISSN 0935-1183. URL http://brinch-hansen.net/papers/

1990a.pdf. See [DF89].

BrinchHansen:2001:COS

[Bri01] Per Brinch Hansen. Classic Operating Systems: from Batch Processing to Distributed Systems. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 0-387-95113-X. x + 597 pp. LCCN QA76.76.O63 B7425 2001. URL http:

//www.loc.gov/catdir/enhancements/fy0816/00045036-
d.html;   http://www.loc.gov/catdir/enhancements/
fy0816/00045036-t.html;   https://link.springer.com/
book/10.1007/978-1-4757-3510-9.

BrinchHansen:2002:OCP

[Bri02] Per Brinch Hansen, editor. The Origin of Concurrent Programming: From Semaphores to Remote Procedure Calls. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. ISBN 0-387-95401-5, 1-4419-2986-X, 1-4757-3472-7, 1-4757-3472-7. x + 534 pp. LCCN

QA76.642 .O75 2002. URL http://link.springer.com/10.
1007/978-1-4757-3472-0.

Bron:1990:RVB

[Bro90a] Coen Bron. On a renewed visit to the banker and remarkable analogy. In Feijen et al. [FvGGM90], chapter 8, pages 76–82. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN QA76 .B326

1990. URL http://www.zentralblatt-math.org/zmath/

en/search/?an=0718.68004. Contains important treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

Broy:1990:BBM

[Bro90b] Manfred Broy. On bounded buffers: modularity, robustness, and reliability in reactive systems. In Feijen et al. [FvGGM90], chapter 9, pages 83–93. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN QA76 .B326

1990. URL http://www.zentralblatt-math.org/zmath/

en/search/?an=0718.68004. Contains important treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

Broy:1996:DPD

[Bro96] M. Broy, editor. Deductive program design: Proceedings of the NATO Advanced Study Institute on Deductive Program Design, held in Marktoberdorf, Germany, July 26–August 7, 1994, volume 152 of NATO ASI series. Series F, Computer and systems sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-60947-4 (hardcover). LCCN QA76.9.D5 D38 1996.

URL http://www.loc.gov/catdir/enhancements/fy0812/
96010788-d.html.

Bauer:1976:LHI

[BS76] Friedrich Ludwig Bauer and Klaus Samelson, editors. Language hierarchies and interfaces: international summer school [held July 23 to August 2, 1975 in Marktoberdorf], volume 46 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. ISBN 0-387-07994-7 (New York), 3-540-07994-7 (Berlin), 3-540-37972-X (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

QA76.6 .L335.  URL http://link.springer-ny.com/
link/service/series/0558/tocs/t0046.htm;   http:/
/www.springerlink.com/content/978-0-387-07994-3;
http://www.springerlink.com/openurl.asp?genre=book&
isbn=978-3-540-07994-1;   http://www.springerlink.
com/openurl.asp?genre=issue&issn=0302-9743&volume=
46.

Benl:1999:FCP

[BS99] Holger Benl and Helmut Schwichtenberg. Formal correctness proofs of functional programs: Dijkstra’s algorithm, a case study. In Computational logic (Marktoberdorf, 1997), volume 165 of NATO Adv. Sci. Inst. Ser. F Comput. Systems Sci., pages 113–126. Springer, Berlin, 1999.

Bauer:2001:VAV

[BS01] Dr. Friedrich Ludwig Bauer and Dr. Klaus Samelson. Verfahren zur automatischen Verarbeitung von kodierten Daten und Rechenmaschine zur Ausübung des Verfahrens. (German) [A method for automatic processing of coded data and calculating machine for performing the method]. In Broy and Denert [BD01b], pages 29–40. ISBN 3-540-42290-0 (print), 3-642-48354-2 (e-book). LCCN QA76.758. URL http://link.

springer.com/chapter/10.1007/978-3-642-48354-7_2/.

Bernstein:1990:AFD

[BSE^{+}90] Danielle R. Bernstein, Joshua Stern, William Ehrich, Michael R. Dunlavey, Leonard Zettel, Bruce B. Herbert, Mark Nelson, W. F. Smyth, Bruce W. Walker, Robert M. Gordon, Stephen N. Cole, and David B. Lightstone. ACM Forum: Debatable. Communications of the Association for Computing Machinery, 33(3):264–271, March 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Dij89-32, WWSD90].

Bloom:1979:AGT

[BT79] S. L. Bloom and R. Tendell. Algebraic and graph theoretic characterizations of structured flowchart schemes. Theoretical Computer Science, 9(2):265–286, October 1979. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bullynck:2015:CPT

[Bul15] Maarten Bullynck. Computing primes (1929–1949): Transformations in the early days of digital computing. IEEE Annals of the History of Computing, 37(3):44–54, July/September 2015. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547

(electronic). URL http://www.computer.org/csdl/mags/
an/2015/03/man2015030044-abs.html.

Burks:1980:ESP

[Bur80] Arthur W. Burks. From ENIAC to the stored-program computer: Two revolutions in computers. In Metropolis et al. [MHR80], pages 311–344. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5 .I63 1976.

URL http://www.sciencedirect.com/science/article/

pii/B9780124916500500228. Original versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Brzezinski:2000:CRD

[BW00] J. Brzezinski and D. Wawrzyniak. Consistency requirements of distributed shared memory for Dijkstra’s mutual exclusion algorithm. In Proceedings. 20th International Conference on Distributed Computing Systems, 2000, pages 618– 625. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2000. ISBN 0-7695-0601-1. ISSN 1063-6927. URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?arnumber=840978.

Cuan:2012:AID

[CC12] Ying Cuan and Xiaoni Chen. Application of improved Dijkstra algorithm in selection of gas source node in gas network. In 2012 International Conference on Industrial Control and Electronics Engineering (ICICEE), pages 1558–1560. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2012. ISBN 1-4673-1450-

1. URL http://ieeexplore.ieee.org/stamp/stamp.jsp?
arnumber=6322700.

Cantone:1986:WFC

[CCD86] G. Cantone, A. Cimitile, and U. De Carlini. Well-formed conversion of unstructured One-in/one-out schemes for complexity measurement and program maintenance. The Computer Journal, 29(4):322–329, August 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/29/4/322.full.
pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_04/tiff/322.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_29/Issue_04/
tiff/323.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_04/tiff/324.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_29/Issue_04/
tiff/325.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_04/tiff/326.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_29/Issue_04/
tiff/327.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_29/Issue_04/tiff/328.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_29/Issue_04/
tiff/329.tif.

Cantone:2004:TLG

[CF04] Domenico Cantone and Simone Faro. Two-levels-greedy: a generalization of Dijkstra’s shortest path algorithm. In Workshop on Graphs and Combinatorial Optimization, volume 17 of Electron. Notes Discrete Math., pages 81–86 (electronic). Elsevier, Amsterdam, 2004.

Chao:2010:DDS

[CH10] Yin Chao and Wang Hongxia. Developed Dijkstra shortest path search algorithm and simulation. In Bin Wang and Jinkuan Wang, editors, ICCDA 2010: 2010 International Conference on Computer Design and Applications: Qinhuangdao, Hebei, China: June 25–27, 2010, volume 1, pages V1–116–V1–119. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2010. ISBN 1-4244-7164-8, 1-4244-7163-X. LCCN

TK7885.A1. URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5541129.

Chao:2010:DDA

[Cha10] Yin Chao. A developed Dijkstra algorithm and simulation of urban path search. In 2010 5th International Conference on Computer Science and Education (ICCSE), pages 1164–1167. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2010. ISBN 1-4244-6002-

6. URL http://ieeexplore.ieee.org/stamp/stamp.jsp?
arnumber=5593700.

Chen:2001:ERM

[Che01] Peter Pin-Shan Chen. The entity relationship model — toward a unified view of data. In Broy and Denert [BD01b], pages 205–234. ISBN 3-540-42290-0 (print), 3-642-48354-2 (e-book). LCCN QA76.758. URL http://link.springer.

com/chapter/10.1007/978-3-642-48354-7_7/.

Chen:2002:GLS

[Che02a] P. P. Chen. From goto-less to structured programming: the legacy of Edger W. Dijkstra. IEEE Software, 19(5):21, 2002. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=1032847.

Chen:2002:MGL

[Che02b] Peter P. Chen. Memorial: From goto-less to structured programming: The legacy of Edsger W. Dijkstra. IEEE Software, 19(5):21, September/October 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http:/

/dlib.computer.org/so/books/so2002/pdf/s5021.pdf.

Chen:2016:GAT

[Che16] Jiahao Chen. “This guy’s arrogance takes your breath away”: Letters between John W Backus and Edsger W Dijkstra, 1979. Web site., May 29, 2016. URL https:

//medium.com/@acidflask/this-guys-arrogance-takes-
your-breath-away-5b903624ca5f. See Dijkstra’s response

[Dij78-37] to Backus’ Turing Award Lecture.

Chu:1980:CDA

[Chu80] J. C. Chu. Computer development at Argonne National Laboratory. In Metropolis et al. [MHR80], pages 345–346. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5

.I63 1976. URL http://www.sciencedirect.com/science/
article/pii/B978012491650050023X. Original versions of

these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Campbell-Kelly:2002:EDD

[CK02a] Martin Campbell-Kelly. Edsger Dijkstra, doyen of the theory and practice of computer programming. The Independent, August 22, 2002.

Campbell-Kelly:2002:TEW

[CK02b] Martin Campbell-Kelly. A tribute to Edsger W. Dijkstra. The British Computer Society, Advanced Programming Specialist Group, Computer Conservation Society, 2002.

Chandy:1988:AVF

[CM88] K. Mani Chandy and J. Misra. Another view on “fairness”. ACM SIGSOFT Software Engineering Notes, 13(3):20, July 1988. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). See [Dij88o, SL88].

Crauser:1998:PDS

[CMMS98] A. Crauser, K. Mehlhorn, U. Meyer, and P. Sanders. A parallelization of Dijkstra’s shortest path algorithm. Lecture Notes in Computer Science, 1450:722–731, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer-ny.com/link/
service/series/0558/bibs/1450/14500722.htm;   http:
//link.springer-ny.com/link/service/series/0558/papers/
1450/14500722.pdf.

Clark:1992:CDF

[CMS92] W. Edwin Clark, Gregory L. McColm, and W. Richard Stark. On the complexity of deadlock-free programs on a ring of processors. Journal of Parallel and Distributed Computing, 16(1):67–71, September 1992. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Cochran:2002:NVS

[Coc02] Shannon Cochran. News and views: Smalltalk turns 30; the billionth PC; reverse challenge concludes; Edsger Dijkstra passes away. Dr. Dobb’s Journal of Software Tools, 27 (10):14, October 2002. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/.

Codd:2001:RMD

[Cod01] E. F. Codd. A relational model of data for large shared data banks. In Broy and Denert [BD01b], pages 61–98. ISBN 3-540-42290-0 (print), 3-642-48354-2 (e-book). LCCN

QA76.758. URL http://link.springer.com/chapter/10.
1007/978-3-642-48354-7_4/.

Creak:2002:HTW

[Cre02] Alan Creak. How things were: Edsger W. Dijkstra. ACM SIGPLAN Notices, 37(12):14–16, December 2002. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Casadei:1978:ATG

[CRT78] Giorgio Casadei, Roberto Righi, and Antonio G. B. Teolis. An algebraic theory for the “guarded commands” of Dijkstra. Rend. Sem. Mat. Univ. Padova, 59:269–283 (1979),

1978. ISSN 0041-8994. URL http://www.numdam.org/item?
id=RSMUP_1978__59__269_0.

Chernoy:2010:PDT

[CSZ10] Viacheslav Chernoy, Mordechai Shalom, and Shmuel Zaks. On the performance of Dijkstra’s third self-stabilizing algorithm for mutual exclusion and related algorithms. Distributed Computing, 23(1):43–60, September 2010. CO-DEN DICOEB. ISSN 0178-2770 (print), 1432-0452 (elec-

tronic). URL https://link.springer.com/article/10.
1007/s00446-010-0104-6.

Chandy:1990:EPC

[CT90] K. Mani Chandy and Stephen Taylor. Examples in program composition. In Feijen et al. [FvGGM90], chapter 10, pages 94–101. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN

QA76 .B326 1990. URL http://www.zentralblatt-math.
org/zmath/en/search/?an=0718.68004. Contains impor-

tant treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

Charmonman:1975:SPF

[CW75] S. Charmonman and J. L. Wagener. On structured programming in FORTRAN. ACM SIGNUM Newsletter, 10(1):21–23, January 1975. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Calkin:2000:RR

[CW00] Neil Calkin and Herbert S. Wilf. Recounting the rationals. American Mathematical Monthly, 107(4):360–363, April 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2589182.

Dijkstra:1966:OSAc

[D^{+}66a] Edsger W. Dijkstra et al. THE operating system — arithmetic and system subroutines. Web archive, 1966.

URL http://archive.computerhistory.org/resources/
text/Knuth_Don_X4100/PDF_index/k-9-pdf/k-9-c1063-
3-THEos-Arith-System-Subroutines.pdf.

Dijkstra:1966:OSAb

[D^{+}66b] Edsger W. Dijkstra et al. THE operating system — list of operator message words. Web archive, 1966.

URL http://archive.computerhistory.org/resources/
text/Knuth_Don_X4100/PDF_index/k-9-pdf/k-9-c1063-
2-THEos-message-words.pdf.

Dijkstra:1966:OSAa

[D^{+}66c] Edsger W. Dijkstra et al. THE operating system — system tables. Web archive, 1966. URL http://

archive.computerhistory.org/resources/text/Knuth_
Don_X4100/PDF_index/k-9-pdf/k-9-c1063-1-THEos-SystemTables.
pdf.

Daylight:2011:DRCa

[Day11a] Edgar G. Daylight. Dijkstra’s rallying cry for generalization. Web site, February 22, 2011. URL http:/

/www.dijkstrascry.com/;   http://www.dijkstrascry.
com/node/41.

Daylight:2011:DRCb

[Day11b] Edgar G. Daylight. Dijkstra’s rallying cry for generalization: the advent of the recursive procedure, late 1950s–early 1960s. The Computer Journal, 54(11):1756–1772, November 2011. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL https://citeseerx.ist.psu.edu/viewdoc/
download?doi=10.1.1.366.3916&rep=rep1&type=pdf.

Daylight:2012:DSE

[Day12a] Edgar G. Daylight. The Dawn of Software Engineering: from Turing to Dijkstra. Lonely Scholar, Heverlee, Belgium, 2012. ISBN 94-91386-02-6. vi + 239 pp. LCCN QA76.17 .D38 2012. Edited by Kurt De Grave.

Daylight:2012:DRC

[Day12b] Edgar G. Daylight. Dijkstra’s rallying cry for generalization. In The Dawn of Software Engineering: from Turing to Dijkstra [Day12a], pages 43–78. ISBN 94-91386-02-6. LCCN QA76.17 .D38 2012. Edited by Kurt De Grave.

deBakker:1976:FPS

[dB76] J. W. de Bakker. Fixed point semantics and Dijkstra’s fundamental invariance theorem. Mathematisch Centrum, Amsterdam, Amsterdam, The Netherlands, 1976. ii + 21 pp. Second printing, Afdeling Informatica, No. IW 29/76. [Department of Computer Science, No. IW 29/76].

Durbin:2002:LAS

[DCF^{+}02] John P. Durbin, Reidar Conradi, Alfonso Fuggetta, Paul Freedman, Ferdinand Wagner, and Peter Wolstenholme. Letters: Not another SPI framework; pair programming à la Dijkstra; a radical paradigm shift? IEEE Software, 19(6):8–9, November/December 2002. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL http://dlib.

computer.org/so/books/so2002/pdf/s6008.pdf.

Dijkstra:1998:ACP

[DD98] Edsger W. Dijkstra and Rutger M. Dijkstra. The algebraic core of a propositional logic. Circulated privately.,

July 1998. URL http://www.cs.utexas.edu/users/EWD/
ewd12xx/EWD1276.PDF.

Dijkstra:1970:NIM

[DDG^{+}70] Edsger W. Dijkstra, ?. Duncan, Jan V. Garwick, C. A. R. Hoare, Brian Randell, G. Seegmueller, Władisław M. Turski, and Michael Woodger. News item — minority report. AL-GOL Bulletin (Amsterdam: Mathematisch Centrum), 31:7, March 1970. CODEN ALGOBG. ISSN 0084-6198. Item AB31.1.1.

Dahl:1972:SP

[DDH72] Ole-Johan Dahl, Edsger W. Dijkstra, and C. A. R. (Charles Antony Richard) Hoare. Structured Programming, volume 8 of APIC studies in data processing. Academic Press, New York, NY, USA, third edition, 1972. ISBN 0-12-200550-3 (hardcover), 0-12-200556-2 (paperback). viii + 220 pp. LCCN QA76.6 .D33 1972.

Dahl:1975:CSP

[DDH75] Ole-Johan Dahl, Edsger Dijkstra, and C. A. R. Hoare, editors. Strukturnoe programmirovanie. Matematicheskoe Obespechenie ÈVM. [Monographs in Mathematical Computer Software]. Izdat. “Mir”, Moscow, USSR, 1975. ISBN 2-01-507541-0. 247 pp. Translated from the English by S. D. Zeleneckiĭ, V. V. Martynjuk and L. V. Uhov, Edited by È. Z. Ljubimskiĭ and V. V. Martynjuk.

Dal:1975:SP

[DDK75] U. Dal, E. Deĭkstra, and K. Khoor. Strukturnoe programmirovanie. Izdat. “Mir”, Moscow, USSR, 1975. 247 pp. Translated from the English by S. D. Zeleneckiĭ, V. V. Martynjuk and L. V. Uhov, Edited by È. Z. Ljubimskiĭ and V. V. Martynjuk, Matematicheskoe Obespechenie EVM. [Monographs in Mathematical Computer Software].

Dekker:1957:CPV

[DDvW57] T. J. Dekker, Edsger W. Dijkstra, and A. van Wijngaarden. Cursus: Programmeren voor automatische rekenmachines. (Dutch) [Course: Programming automatic calculators]. Technical Report 9, Mathematisch Centrum, Amster-

dam, 1957. URL http://www.cs.utexas.edu/users/EWD/
MCReps/CR1957-009.PDF.

DeMarco:2001:SAS

[DeM01] Tom DeMarco. Structure analysis and system specification. In Broy and Denert [BD01b], pages 255–288. ISBN 3-540-42290-0 (print), 3-642-48354-2 (e-book). LCCN

QA76.758. URL http://link.springer.com/chapter/10.
1007/978-3-642-48354-7_9/.

Derigs:1985:EDL

[Der85] Ulrich Derigs. An efficient Dijkstra-like labeling method for computing shortest odd/even paths. Information Processing Letters, 21(5):253–258, November 18, 1985. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dijkstra:1984:MVP

[DF84] E. W. Dijkstra and W. H. J. Feijen. Een methode van programmeren. (Dutch) [A method of programming]. Academic Service, Den Haag, The Netherlands, 1984. Also available in German [DF85] and English [DF88a].

Dijkstra:1985:MPG

[DF85] E. W. Dijkstra and W. H. J. Feijen. Methodik des Programmierens. (German) [A method of programming]. Addison-Wesley, Reading, MA, USA, 1985. ISBN 3-925118-18-7. 239 pp. German translation by ?? of the Dutch original [DF84].

Dijkstra:1988:MP

[DF88a] Edsger W. Dijkstra and W. H. Feijen. A method of programming. Addison-Wesley, Reading, MA, USA, 1988. ISBN 0-201-17536-3. viii + 188 pp. LCCN ???? URL

http://www.zentralblatt-math.org/zmath/en/search/?

an=0681.68001. English translation by Joke Sterringa of the Dutch original [DF85].

Dijkstra:1988:LSR

[DF88b] Edsger W. Dijkstra and W. H. J. Feijen. The linear search revisited. Published as [DF89]., 1988. URL http://www.cs.

utexas.edu/users/EWD/ewd10xx/EWD1029.PDF.

Dijkstra:1989:LSR

[DF89] Edsger W. Dijkstra and W. H. J. Feijen. The linear search revisited. Structured Programming, 10(1):5–9, 1989. CODEN STPGEM. ISSN 0935-1183.

Dijkstra:20xx:BPP

[DFxx] Edsger W. Dijkstra and W. H. J. Feijen. A beautiful proof of a probably useless theorem. Circulated pri-

vately., 20xx. URL http://www.cs.utexas.edu/users/
EWD/ewd04xx/EWD415.PDF.

Dijkstra:1974:AET

[DFR74] Edsger W. Dijkstra, W. H. J. Feijen, and M. Rem. Associons: an effort towards accommodating potentially ultra-high concurrency. Circulated privately., July 1974. URL http://www.

cs.utexas.edu/users/EWD/ewd04xx/EWD435.PDF.

Dijkstra:20xx:AC

[DFRxx] Edsger W. Dijkstra, W. H. J. Feijen, and M. Rem. Associons continued. Circulated privately., 20xx. URL http://www.

cs.utexas.edu/users/EWD/ewd04xx/EWD439.PDF.

Dijkstra:1982:DTDb

[DFvG82] Edsger W. Dijkstra, W. H. J. Feijen, and A. J. M. van Gasteren. Derivation of a termination detection algorithm for distributed computations. Published as

[DFvG83]., 1982. URL http://www.cs.utexas.edu/users/
EWD/ewd08xx/EWD840.PDF.

Dijkstra:1983:DTDa

[DFvG83] Edsger W. Dijkstra, W. H. J. Feijen, and A. J. M. van Gasteren. Derivation of a termination detection algorithm for distributed computations. Information Processing Letters, 16(5):217–219, June 10, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dijkstra:1985:DTD

[DFvG85] Edsger W. Dijkstra, W. H. J. Feijen, and A. J. M. van Gasteren. Derivation of a termination detection algorithm for distributed computations. In Control flow and data flow: concepts of distributed programming (Marktoberdorf, 1984), volume 14 of NATO Adv. Sci. Inst. Ser. F Comput. Systems Sci., pages 507–512. Springer, Berlin, 1985.

Dean:1996:TLF

[DH96] C. Neville Dean and Michael Gerard Hinchey. Teaching and Learning Formal Methods. Academic Press international series in formal methods. Academic Press, New York, NY, USA, 1996. ISBN 0-12-349040-5. xvi + 285 pp. LCCN

QA76.76D47 T42. URL http://www.gbv.de/dms/bowker/

toc/9780123490407.pdf. Foreword by Edsger W. Dijkstra.

Dolev:2001:DSS

[DH01] Shlomi Dolev and Ted Herman. Dijkstra’s self-stabilizing algorithm in unsupportive environments. Lecture Notes in Computer Science, 2194:67–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/series/0558/bibs/
2194/21940067.htm;   http://link.springer-ny.com/
link/service/series/0558/papers/2194/21940067.pdf.

Dijkstra:1959:ASC

[DHPS59] Edsger W. Dijkstra, W. Heise, Alan J. Perlis, and K. Samelson. ALGOL sub-committee report — extensions. Communications of the Association for Computing Machinery, 2(9):24– ??, ???? 1959. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dijkstra:1953:FBV

[Dij53] Edsger W. Dijkstra. Functionele beschrijving van de ARRA. (Dutch) [Functional description of the ARRA]. Technical Report 12, Mathematisch Centrum, Amster-

dam, 1953. URL http://www.cs.utexas.edu/users/EWD/
MCReps/MR12.PDF.

Dijkstra:1954:DKR

[Dij54a] Edsger W. Dijkstra. “Drijvende-komma”-rekentechniek: (ARRA-subroutines Rd1 en Rd2). (Dutch) [“Floating-point calculation technique: (ARRA subroutines Rd1 and Rd2)]. Technical Report 16, Mathematisch Centrum, Amsterdam,

September 1954. 27 pp. URL http://www.cs.utexas.edu/
users/EWD/MCReps/MR16.PDF.

Dijkstra:1954:UVA

[Dij54b] Edsger W. Dijkstra. In- en uitvoer van de ARRA. (Dutch) [Export and import of the ARRA]. Technical Report 14, Mathematisch Centrum, Amsterdam, 1954. URL http://

www.cs.utexas.edu/users/EWD/MCReps/MR14.PDF.

Dijkstra:1955:HVPa

[Dij55a] Edsger W. Dijkstra. Handboek voor de programmeur (FERTA), 1. (Dutch) [Manual for the programmer (FERTA) 1]. Technical Report 17, Mathematisch Centrum, Amster-

dam, 1955. URL http://www.cs.utexas.edu/users/EWD/
MCReps/MR17.PDF.

Dijkstra:1955:HVPb

[Dij55b] Edsger W. Dijkstra. Handboek voor de programmeur (FERTA), 2. (Dutch) [Manual for the programmer (FERTA) 2]. Technical Report 20, Mathematisch Centrum, Amster-

dam, 1955. URL http://www.cs.utexas.edu/users/EWD/
MCReps/MR20.PDF.

Dijkstra:1955:CVA

[Dij55c] Edsger W. Dijkstra. Het communicatieprogramma van de ARRA. (Dutch) [The communication program of the ARRA]. Technical Report 21, Mathematisch Centrum, Am-

sterdam, 1955. URL http://www.cs.utexas.edu/users/
EWD/MCReps/MR21.PDF.

Dijkstra:1956:STV

[Dij56a] Edsger W. Dijkstra. Het standaard-typprogramma van ARMAC. (Dutch) [The default type programs of AR-MAC]. Technical Report 24, Mathematisch Centrum, Am-

sterdam, 1956. URL http://www.cs.utexas.edu/users/
EWD/MCReps/MR24.PDF.

Dijkstra:1956:KBV

[Dij56b] Edsger W. Dijkstra. Korte beschrijving van de opdrachtencode etc. voor ARMAC. (Dutch) [Short description of the command code, etc. for ARMAC]. Technical Report 23, Mathematisch Centrum, Amsterdam, 1956. URL http://

www.cs.utexas.edu/users/EWD/MCReps/MR23.PDF.

Dijkstra:1956:PVAa

[Dij56c] Edsger W. Dijkstra. Programmering voor de ARMAC, 1: Algemeen. (Dutch) [Programming for the ARMAC 1: General]. Technical Report 25, Mathematisch Centrum, Amster-

dam, 1956. URL http://www.cs.utexas.edu/users/EWD/
MCReps/MR25.PDF.

Dijkstra:1956:PVAb

[Dij56d] Edsger W. Dijkstra. Programmering voor de ARMAC, 2: De inhoud der geblokkeerde kanalen. (Dutch) [Programming for the ARMAC 2: The contents of blocked channels]. Technical Report 26, Mathematisch Centrum, Amster-

dam, 1956. URL http://www.cs.utexas.edu/users/EWD/
MCReps/MR26.PDF.

Dijkstra:1956:PVAc

[Dij56e] Edsger W. Dijkstra. Programmering voor de ARMAC, 3. (Dutch) [Programming for the ARMAC 3]. Technical Report 27, Mathematisch Centrum, Amsterdam, 1956. URL

http://www.cs.utexas.edu/users/EWD/MCReps/MR27.PDF.

Dijkstra:1957:MIPa

[Dij57a] E. W. Dijkstra. A method to investigate primality. Mathematical Tables and Other Aids to Computation, 11(59):195–196, July 1957. CODEN MTTCAS. ISSN 0891-6837. URL http:

//www.jstor.org/stable/2002081. In Technical Notes and

Short Papers.

Dijkstra:1957:MIPb

[Dij57b] Edsger W. Dijkstra. A method to investigate primality. Technical Report 2, Mathematisch Centrum, Amsterdam, Jan-

uary 1957. URL http://www.cs.utexas.edu/users/EWD/
MCReps/ZW1957-002.PDF.

Dijkstra:1957:PVAa

[Dij57c] Edsger W. Dijkstra. Programmering voor de ARMAC, 1a: Algemeen. (Dutch) [Programming for the ARMAC 1a: General]. Technical Report 25a, Mathematisch Centrum, Amster-

dam, 1957. URL http://www.cs.utexas.edu/users/EWD/
MCReps/MR25a.PDF.

Dijkstra:1957:PVAb

[Dij57d] Edsger W. Dijkstra. Programmering voor de ARMAC, 5. (Dutch) [Programming for the ARMAC, 5]. Technical Report 29, Mathematisch Centrum, Amsterdam, 1957. URL

http://www.cs.utexas.edu/users/EWD/MCReps/MR29.PDF.

Dijkstra:1959:CACa

[Dij59a] Edsger W. Dijkstra. Communication with an Automatic Computer. PhD thesis, University of Amsterdam, Amsterdam, The Netherlands, 1959. v + 166 + 3 pp.

URL http://www.cs.utexas.edu/users/EWD/PhDthesis/
PhDthesis.PDF.

Dijkstra:1959:CACb

[Dij59b] Edsger W. Dijkstra. Communication with an automatic computer. Communications of the Association for Computing Machinery, 2(9):24, September 1959. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dijkstra:1959:NTP

[Dij59c] Edsger W. Dijkstra. A note on two problems in connexion with graphs. Numerische Mathematik, 1(1):269–271, December 1959. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://gdzdoc.sub.

uni-goettingen.de/sub/digbib/loader?did=D196313;
http://link.springer.com/article/10.1007/BF01386390;
 http://www.springerlink.com/openurl.asp?genre=article&
issn=0029-599X&volume=1&issue=1&spage=269.

Dijkstra:1960:STS

[Dij60a] E. W. Dijkstra. Some theorems on spanning subtrees of a graph. Indagationes Mathematicae (Proceedings), 63: 196–199, 1960. ISSN 1385-7258 (print), 1878-5972 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1385725860500266.

Dijkstra:1960:CPA

[Dij60b] Edsger W. Dijkstra. Cursus programmeren in ALGOL 60 (Dutch) [Course in programming in ALGOL 60]. Technical report, Mathematisch Centrum, Amsterdam, Amsterdam, The Netherlands, 1960.

Dijkstra:1960:KBU

[Dij60c] Edsger W. Dijkstra. Kortste bomen uit een graph. (Dutch) [Shortest trees from a graph]. Technical Report 1, Mathematisch Centrum, Amsterdam, January 1960. URL http://

www.cs.utexas.edu/users/EWD/MCReps/ZW1960-001.PDF.

Dijkstra:1960:RPa

[Dij60d] Edsger W. Dijkstra. Recursive programming. In Saul Rosen, editor, Programming Systems and Languages, chapter 3C, pages 221–227. McGraw-Hill, New York, NY, USA, 1960.

Dijkstra:1960:RPb

[Dij60e] Edsger W. Dijkstra. Recursive programming. Numerische Mathematik, 2(1):312–318, December 1960. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (elec-

tronic). URL http://gdzdoc.sub.uni-goettingen.de/
sub/digbib/loader?did=D196346. See [vdH17].

Dijkstra:1961:AUX

[Dij61a] E. W. Dijkstra. Ein ALGOL-60-Übersetzer für die X 1. II. (German) [an ALGOL-60 translator for the X 1. II]. Math.- Tech.-Wirtschaft, 8:115–119, 1961.

Dijkstra:1961:LED

[Dij61b] E. W. Dijkstra. Letter to the Editor: Defense of ALGOL 60. Communications of the Association for Computing Machinery, 4(11):502–503, November 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dijkstra:1961:ATA

[Dij61c] Edsger W. Dijkstra. Algol 60 translation: an Algol 60 translator for the x1 and Making a translator for Algol 60. Technical Report 35, Mathematisch Centrum, Amster-

dam, 1961. URL http://www.cs.utexas.edu/users/EWD/
MCReps/MR35.PDF.

Dijkstra:1961:DMI

[Dij61d] Edsger W. Dijkstra. On the design of machine independent programming languages. Technical Report 34, Mathematisch Centrum, Amsterdam, October 1961. URL http://www.cs.

utexas.edu/users/EWD/MCReps/MR34.PDF.

Dijkstra:1962:OEA

[Dij62a] E. W. Dijkstra. Operating experience with ALGOL 60. The Computer Journal, 5(2):125–127, August 1962. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/
content/5/2/125.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_05/Issue_02/050125.
sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_02/tiff/125.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_05/Issue_02/
tiff/126.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_05/Issue_02/tiff/127.tif.

Dijkstra:1962:AUC

[Dij62b] Edsger W. Dijkstra. An attempt to unify the constituent concepts of serial program execution. Technical Report 46, Mathematisch Centrum, Amsterdam, January 1962. URL http:/

/www.cs.utexas.edu/users/EWD/MCReps/MR46.PDF.

Dijkstra:1962:OOO

[Dij62c] Edsger W. Dijkstra. Ontwerpschets ontwikkelingsplan. 24 oktober 1962. (Dutch) [Design sketch development plan. 24 October 1962]. Circulated privately., October 1962. URL http:

//www.cs.utexas.edu/users/EWD/OtherDocs/NN033.PDF.

Dijkstra:1962:HLV

[Dij62d] Edsger W. Dijkstra. Over de “hot line” (VERTROUWELIJK). (Dutch) [On the “hot line” (CONFIDENTIAL)]. Circulated privately., February 1962. URL http://www.cs.utexas.

edu/users/EWD/OtherDocs/NN188.PDF.

Dijkstra:1962:PAR

[Dij62e] Edsger W. Dijkstra. A primer of ALGOL 60: report on the algorithmic language ALGOL 60, volume 2 of APIC studies in data processing. Academic Press, New York, NY, USA, 1962. ISBN 0-12-216250-1. ISSN 0067-2483. xi + 114 pp. LCCN QA76.73.A24.

Dijkstra:1962:SMA

[Dij62f] Edsger W. Dijkstra. Some meditations on advanced programming. In IFIP Congress, pages 535–538. ????, ????, 1962.

Dijkstra:1962:SP

[Dij62g] Edsger W. Dijkstra. Substitution processes. Circulated pri-

vately., January 1962. URL http://www.cs.utexas.edu/
users/EWD/ewd00xx/EWD28.PDF.

Dijkstra:1963:A

[Dij63a] E. W. Dijkstra. ,,Algol 60“. Statistica Neerlandica, 17(2):187– 188, June 1963. CODEN ???? ISSN 0039-0402 (print), 1467-

9574 (electronic). URL https://onlinelibrary.wiley.
com/doi/epdf/10.1111/j.1467-9574.1963.tb00317.x.

Dijkstra:1963:ATP

[Dij63b] E. W. Dijkstra. Appendix — two papers on an AL-GOL translator for the X1. Annual Review in Automatic Programming, 3:327–??, 1963. CODEN ARVAAM. ISSN 0066-4138 (print), 1878-545X (electronic). URL

http://www.sciencedirect.com/science/article/pii/
B9780080097633500175;   http://www.sciencedirect.
com/science/article/pii/S0066413863800142.

Dijkstra:1963:DMI

[Dij63c] E. W. Dijkstra. On the design of machine independent programming languages. Annual Review in Automatic Programming, 3(0):27–42, 1963. CODEN ARVAAM. ISSN 0066-4138 (print), 1878-545X (electronic). URL

http://www.sciencedirect.com/science/article/pii/
B9780080097633500072;   http://www.sciencedirect.
com/science/article/pii/S0066413863800038. Originally

issued as Report M.R.34 of the Mathematics Centre, Amsterdam.

Dijkstra:1963:ERVb

[Dij63d] Edsger W. Dijkstra. 2de Excerpt reactie van fabrikanten. (Dutch) [Second excerpt reaction from manufacturers]. Circulated privately., January 1963. URL http://www.cs.

utexas.edu/users/EWD/OtherDocs/NN037.PDF.

Dijkstra:1963:ERVc

[Dij63e] Edsger W. Dijkstra. 3de Excerpt reactie’s van fabrikanten. (Dutch) [Third excerpt reaction from manufacturers]. Circulated privately., February 1963. URL http://www.cs.

utexas.edu/users/EWD/OtherDocs/NN041.PDF.

Dijkstra:1963:ATX

[Dij63f] Edsger W. Dijkstra. An ALGOL-60 translator for the X1. Annual Review in Automatic Programming, 3(??):329–345, 1963. CODEN ARVAAM. ISSN 0066-4138 (print), 1878-545X (electronic).

Dijkstra:1963:ERVa

[Dij63g] Edsger W. Dijkstra. Excerpt reactie’s van fabrikanten. (Dutch) [Excerpt reaction from manufacturers]. Circulated privately., January 1963. URL http://www.cs.utexas.edu/

users/EWD/OtherDocs/NN036.PDF.

Dijkstra:1963:ERVd

[Dij63h] Edsger W. Dijkstra. Excerpt reactie’s van fabrikanten. no. 4. (Dutch) [Excerpt reaction from manufacturers, no. 4]. Circulated privately., March 1963. URL http://www.cs.utexas.

edu/users/EWD/OtherDocs/NN042.PDF.

Dijkstra:1963:MTA

[Dij63i] Edsger W. Dijkstra. Making a translator for ALGOL-60. Annual Review in Automatic Programming, 3(??):347–356, 1963. CODEN ARVAAM. ISSN 0066-4138 (print), 1878-545X (electronic).

Dijkstra:1963:MVK

[Dij63j] Edsger W. Dijkstra. Motivering van de keuze. (Dutch) [Justification for the choice]. Circulated privately., June 1963.

URL http://www.cs.utexas.edu/users/EWD/OtherDocs/
NN062.PDF.

Dijkstra:1963:MXV

[Dij63k] Edsger W. Dijkstra. Multiprogrammering en de X8 (Vervolg van EWD54). (Dutch) [Multiprogramming and X8 (continuation of EWD54)]. Circulated privately., June 1963. URL

http://www.cs.utexas.edu/users/EWD/ewd00xx/EWD57.
PDF.

Dijkstra:1963:MDF

[Dij63l] Edsger W. Dijkstra. My dear friends. Cover letter for submission of [Dij63r] to Commun. ACM; circulated privately., Jan-

uary 1963. URL http://www.cs.utexas.edu/users/EWD/
OtherDocs/NN038.PDF.

Dijkstra:1963:NAP

[Dij63m] Edsger W. Dijkstra. Notitie over de aansluiting en de programmering van de teleprinter. (Dutch) [Note on connecting and programming the teleprinter]. Circulated privately.,

July 1963. URL http://www.cs.utexas.edu/users/EWD/
ewd00xx/EWD62.PDF.

Dijkstra:1963:OGD

[Dij63n] Edsger W. Dijkstra. Objectprogramma, gegenereerd door de M.C.-vertaler: (intern rapport). (Dutch) [Object program generated by the M.C.-translator (internal report)]. Technical Report 55, Mathematisch Centrum, Amsterdam, Jan-

uary 1963. URL http://www.cs.utexas.edu/users/EWD/
MCReps/MR55.PDF.

Dijkstra:1963:JDJ

[Dij63o] Edsger W. Dijkstra. Over Jansen. (Dutch) [On Jansen]. Circulated privately., September 1963. URL http://www.cs.

utexas.edu/users/EWD/ewd00xx/EWD65.PDF.

Dijkstra:1963:SDS

[Dij63p] Edsger W. Dijkstra. Over stapeladministratie. (Dutch) [On stack management]. Circulated privately., Novem-

ber 1963. URL http://www.cs.utexas.edu/users/EWD/
ewd00xx/EWD69.PDF.

Dijkstra:1963:TDD

[Dij63q] Edsger W. Dijkstra. Over trommelpaginatransporten. (Dutch) [On drum page transport]. Circulated privately.,

December 1963. URL http://www.cs.utexas.edu/users/
EWD/ewd00xx/EWD72.PDF.

Dijkstra:1963:RID

[Dij63r] Edsger W. Dijkstra. A review of the IBM 1620 Data Processing System. Circulated privately., January 1963. URL http:

//www.cs.utexas.edu/users/EWD/ewd00xx/EWD37.PDF.

Dijkstra:1964:SCA

[Dij64a] E. W. Dijkstra. Some comments on the aims of MIRFAC. Communications of the Association for Computing Machinery, 7(3):190, March 1964. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dijkstra:1964:BTD

[Dij64b] Edsger W. Dijkstra. Bezettingsadministratie der trommelpagina’s. (Dutch) [Occupation administration of drum pages]. Circulated privately., May 1964. URL http://www.cs.

utexas.edu/users/EWD/ewd00xx/EWD86.PDF.

Dijkstra:1964:DOP

[Dij64c] Edsger W. Dijkstra. Description of the object program, II. Circulated privately., September 1964. URL http://www.cs.

utexas.edu/users/EWD/ewd01xx/EWD102.PDF.

Dijkstra:1964:ECA

[Dij64d] Edsger W. Dijkstra. Embedding complex arithmetic. Circulated privately., August 1964. URL http://www.cs.utexas.

edu/users/EWD/ewd00xx/EWD96.PDF.

Dijkstra:1964:EC

[Dij64e] Edsger W. Dijkstra. Error checking. Circulated privately.,

December 1964. URL http://www.cs.utexas.edu/users/
EWD/ewd01xx/EWD111.PDF.

Dijkstra:1964:F

[Dij64f] Edsger W. Dijkstra. Foreword. In ALGOL 60 implementation; the translation and use of ALGOL 60 programs on a computer [RR64], page ?? LCCN QA76.5 .R35 1964. With a foreword by E. W. Dijkstra. Published for the Automatic Programming Information Centre, Brighton College of Technology, England. Russian translation in 1967.

Dijkstra:1964:FL

[Dij64g] Edsger W. Dijkstra. Over formal locations. Circulated

privately., June 1964. URL http://www.cs.utexas.edu/
users/EWD/ewd00xx/EWD89.PDF.

Dijkstra:1964:PDP

[Dij64h] Edsger W. Dijkstra. Over paginaadministratie. (Dutch) [On page management]. Circulated privately., May 1964. URL

http://www.cs.utexas.edu/users/EWD/ewd00xx/EWD84.
PDF.

Dijkstra:1964:RSV

[Dij64i] Edsger W. Dijkstra. Ruwe schets vertaalproces. (Dutch) [Rough translation process]. Circulated privately., Decem-

ber 1964. URL http://www.cs.utexas.edu/users/EWD/
ewd01xx/EWD112.PDF.

Dijkstra:1964:SVO

[Dij64j] Edsger W. Dijkstra. Samenvatting van oordeel over het “Voorstel tot aanschaffing van een digitale informatieverwerkende machine voor de afdeling der elektrotechniek”. (Dutch) [Summary judgment on the “Proposal to acquire a digital information processing machine for the Department of Electrical Engineering”]. Circulated privately.,

June 1964. URL http://www.cs.utexas.edu/users/EWD/
ewd00xx/EWD91.PDF.

Dijkstra:1964:SC

[Dij64k] Edsger W. Dijkstra. Segment control. Circulated privately.,

December 1964. URL http://www.cs.utexas.edu/users/
EWD/ewd01xx/EWD113.PDF.

Dijkstra:1965:CSP

[Dij65a] E. W. Dijkstra. Cooperating Sequential Processes. Technical report, Technische Universiteit Eindhoven, Eindhoven, The Netherlands, September 1965.

Dijkstra:1965:CVX

[Dij65b] Edsger W. Dijkstra. Communicatiebuffering voor de EL-X8 – T.H.E.. (Dutch) [Communication buffering for EL-X8 — T.H.E.]. Circulated privately., February 1965. URL http://

www.cs.utexas.edu/users/EWD/ewd01xx/EWD118.PDF.

Dijkstra:1965:GBV

[Dij65c] Edsger W. Dijkstra. Globale beschrijving van de drijvende arithmetiek van de EL X8. (Dutch) [General description of the floating-point arithmetic of the X8 EL]. Circulated privately.,

December 1965. URL http://www.cs.utexas.edu/users/
EWD/ewd01xx/EWD145.PDF.

Dijkstra:1965:HKV

[Dij65d] Edsger W. Dijkstra. Heel kort verslag bezoek aan het symposium over Multi Access Computers, 2–4 november 1965 in het NPL. (Dutch) [Very short report of a visit to the symposium Multi Access Computers, 2 to 4 November 1965 in the NPL]. Circulated privately., November 1965. URL http://

www.cs.utexas.edu/users/EWD/ewd01xx/EWD137.PDF.

Dijkstra:1965:MSX

[Dij65e] Edsger W. Dijkstra. The multiprogramming system for the EL X8 THE. Circulated privately., June 1965. URL http:/

/www.cs.utexas.edu/users/EWD/ewd01xx/EWD126.PDF.

Dijkstra:1965:PCHa

[Dij65f] Edsger W. Dijkstra. Programming considered as a human activity. Published as [Dij65g]., 1965. URL http://www.cs.

utexas.edu/users/EWD/ewd01xx/EWD117.PDF.

Dijkstra:1965:PCHb

[Dij65g] Edsger W. Dijkstra. Programming considered as a human activity. In Proceedings of the IFIP Congress, volume 65, pages 213–217. ????, ????, 1965.

Dijkstra:1965:SPCa

[Dij65h] Edsger W. Dijkstra. Solution of a problem in concurrent programming control. Communications of the Association for Computing Machinery, 8(9):569, September 1965. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dijkstra:1965:VVBb

[Dij65i] Edsger W. Dijkstra. Verslag van bezoek aan de Arbeitstagung für Automatentheorie te Hanover. (Dutch) [Report of visit to the Workshop on Automata Theory in Hanover]. Circulated

privately., 1965. URL http://www.cs.utexas.edu/users/
EWD/ewd01xx/EWD131.PDF.

Dijkstra:1965:VVBa

[Dij65j] Edsger W. Dijkstra. Verslag van het bezoek aan het IFIPcongres te New York, 24–29 mei 1965. (Dutch) [Visit report to the IFIP-congress in New York, 24–29 May 1965]. Circulated privately., June 1965. URL http://www.cs.utexas.

edu/users/EWD/OtherDocs/NN130.PDF.

Dijkstra:1966:SMP

[Dij66a] E. W. Dijkstra. The structure of the THE multi-programming system. Communications of the Association for Computing Machinery, 9(5):??, May 1966. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dijkstra:1966:ALV

[Dij66b] Edsger W. Dijkstra. Aan de leden van de Z8-commissie [Over Michael]. (Dutch) [To the members of the Z8-commission [on Michael]]. Circulated privately., May 1966. URL http://

www.cs.utexas.edu/users/EWD/OtherDocs/NN158.PDF.

Dijkstra:1966:DCA

[Dij66c] Edsger W. Dijkstra. Documentatie over de communicatieapparatuur aan de EL X8 (vervanging van EWD140). (Dutch) [Documentation of the communication to the EL X8 (replacement of EWD140)]. Circulated privately., Jan-

uary 1966. URL http://www.cs.utexas.edu/users/EWD/
ewd01xx/EWD149.PDF.

Dijkstra:1966:TCS

[Dij66d] Edsger W. Dijkstra. Tentamenopgave “Cooperating Sequential Processes” (april 1966). (Dutch) [Exam question “Cooperating Sequential Processes” (April 1966)]. Circulated

privately., April 1966. URL http://www.cs.utexas.edu/
users/EWD/ewd01xx/EWD158.PDF.

Dijkstra:1966:OSAe

[Dij66e] Edsger W. Dijkstra. THE operating system — AL-GOL lex scanner. Web archive, 1966. URL http:/

/archive.computerhistory.org/resources/text/Knuth_
Don_X4100/PDF_index/KnuthDigitalArchive-Index.html#
C1063-5.

Dijkstra:1966:OSAd

[Dij66f] Edsger W. Dijkstra. THE operating system — table of the ALGOL translator. Web archive, 1966.

URL http://archive.computerhistory.org/resources/
text/Knuth_Don_X4100/PDF_index/k-9-pdf/k-9-c1063-
4-THEOS-Translator-Table.pdf.

Dijkstra:1967:ACW

[Dij67a] Edsger W. Dijkstra. Aan de Commissie Wetenschapsbeleid ter zake van Automatiseringswiskunde aan de THE. (Dutch) [The Committee on Science Policy in respect of automation of mathematics in the THE]. Circulated privately., Febru-

ary 1967. URL http://www.cs.utexas.edu/users/EWD/
ewd01xx/EWD191.PDF.

Dijkstra:1967:CAP

[Dij67b] Edsger W. Dijkstra. A constructive approach to the problem of program correctness. Circulated privately., Au-

gust 1967. URL http://www.cs.utexas.edu/users/EWD/
ewd02xx/EWD209.PDF.

Dijkstra:1967:SEa

[Dij67c] Edsger W. Dijkstra. A sequel to EWD200. Circulated

privately., June 1967. URL http://www.cs.utexas.edu/
users/EWD/ewd02xx/EWD201.PDF.

Dijkstra:1967:SEb

[Dij67d] Edsger W. Dijkstra. A sequel to EWD201. Circulated

privately., July 1967. URL http://www.cs.utexas.edu/
users/EWD/ewd02xx/EWD206.PDF.

Dijkstra:1967:TCS

[Dij67e] Edsger W. Dijkstra. Tentamenopgave “Cooperating Sequential Processes” (21 juni 1967). (Dutch) [Exam question “Cooperating Sequential Processes” (21 June 1967)]. Circulated

privately., June 1967. URL http://www.cs.utexas.edu/
users/EWD/OtherDocs/NN215.PDF.

Dijkstra:1967:VSM

[Dij67f] Edsger W. Dijkstra. The variable size machine. Circulated

privately., June 1967. URL http://www.cs.utexas.edu/
users/EWD/ewd02xx/EWD200.PDF.

Dijkstra:1967:VVB

[Dij67g] Edsger W. Dijkstra. Verslag van het bezoek aan Grenoble en Parijs (6–11 december 1967). (Dutch) [Report of the visit to Grenoble and Paris (6 to 11 December 1967)]. Circulated

privately., 1967. URL http://www.cs.utexas.edu/users/
EWD/ewd02xx/EWD222.PDF.

Dijkstra:1968:HP

[Dij68a] E. W. Dijkstra. The humble programmer. In Proceedings of the International Conference on Software Engineering at Garmisch, West Germany. ????, ????, 1968.

Dijkstra:1968:ALV

[Dij68b] Edsger W. Dijkstra. Aan de leden van de Commissie Wetenschapsbeleid. (Dutch) [To the members of the Scientific Policy Committee]. Circulated privately., March 1968. URL http:

//www.cs.utexas.edu/users/EWD/ewd02xx/EWD229.PDF.

Dijkstra:1968:BVS

[Dij68c] Edsger W. Dijkstra. Bijdrage voor de subcommissie Aanschaffingsbeleid Rekenautomaten. (Dutch) [Contribution to the subcommittee Procurement of Computers]. Circulated privately., September 1968. URL http://www.cs.utexas.

edu/users/EWD/ewd02xx/EWD243.PDF.

Dijkstra:1968:CAG

[Dij68d] Edsger W. Dijkstra. A case against the GO TO statement. Letter to the Editor: “Go To Statement Considered Harmful”; published as [Dij68t]., 1968. URL http://www.cs.

utexas.edu/users/EWD/ewd02xx/EWD215.PDF.

Dijkstra:1968:COS

[Dij68e] Edsger W. Dijkstra. Co-operating sequential processes. In F. Genuys, editor, Programming Languages, pages 43–112. Academic Press, New York, NY, USA, 1968.

Dijkstra:1968:COC

[Dij68f] Edsger W. Dijkstra. Commentaar op het Concept van Benders en Veltkamp. (Dutch) [Commentary on the concept of Benders and Veltkamp]. Circulated privately., Decem-

ber 1968. URL http://www.cs.utexas.edu/users/EWD/
OtherDocs/NN247.PDF.

Dijkstra:1968:CAPa

[Dij68g] Edsger W. Dijkstra. A constructive approach to the problem of program correctness. Published as [Dij68h]., Febru-

ary 1968. URL http://www.cs.utexas.edu/users/EWD/
ewd02xx/EWD209A.PDF.

Dijkstra:1968:CAPb

[Dij68h] Edsger W. Dijkstra. A constructive approach to the problem of program correctness. BIT (Nordisk tidskrift for informationsbehandling), 8(3):174–186, September 1968. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-

9125 (electronic). URL http://www.springerlink.com/
openurl.asp?genre=article&issn=0006-3835&volume=8&
issue=3&spage=174.

Dijkstra:1968:CED

[Dij68i] Edsger W. Dijkstra. Contractie en expansie. (Dutch) [Contraction and expansion]. Circulated privately., June 1968.

URL http://www.cs.utexas.edu/users/EWD/ewd02xx/EWD235.
PDF.

Dijkstra:1968:CSPa

[Dij68j] Edsger W. Dijkstra. Cooperating sequential processes. Published as [Dij68k]., 1968. URL http://www.cs.utexas.edu/

users/EWD/ewd01xx/EWD123.PDF.

Dijkstra:1968:CSPb

[Dij68k] Edsger W. Dijkstra. Cooperating sequential processes. In F. Genuys, editor, Programming Languages: NATO Advanced Study Institute, pages 43–112. Academic Press, 1968.

Dijkstra:1968:RGD

[Dij68l] Edsger W. Dijkstra. De Rekenautomaat als Gebruiksvoorwerp. (Dutch) [The computer as tool]. Circulated privately.,

April 1968. URL http://www.cs.utexas.edu/users/EWD/
ewd02xx/EWD231.PDF.

Dijkstra:1968:EDD

[Dij68m] Edsger W. Dijkstra. Een educatief dilemma. (Dutch) [An educational dilemma]. Circulated privately., Septem-

ber 1968. URL http://www.cs.utexas.edu/users/EWD/
ewd02xx/EWD244.PDF.

Dijkstra:1968:HV

[Dij68n] Edsger W. Dijkstra. HEEL VERTROUWELIJK. (Dutch) [VERY CONFIDENTIAL]. Circulated privately., Jan-

uary 1968. URL http://www.cs.utexas.edu/users/EWD/
ewd02xx/EWD224.PDF.

Dijkstra:1968:SPC

[Dij68o] Edsger W. Dijkstra. Stepwise program construction. Published as [Dij82-45]., February 1968. URL http://www.cs.

utexas.edu/users/EWD/ewd02xx/EWD227.PDF.

Dijkstra:1968:SMSa

[Dij68p] Edsger W. Dijkstra. The structure of the “THE”- multiprogramming system. Published as [Dij68q]., 1968. URL

http://www.cs.utexas.edu/users/EWD/ewd01xx/EWD196.
PDF.

Dijkstra:1968:SMSb

[Dij68q] Edsger W. Dijkstra. The structure of the “THE”- multiprogramming system. Communications of the Association for Computing Machinery, 11(5):341–346, May 1968.

CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dijkstra:1968:TCO

[Dij68r] Edsger W. Dijkstra. Tentamenopgave “Co-operating Sequential Processes” [22 mei 1968]. (Dutch) [Exam question “Cooperating Sequential Processes” [22 May 1968]]. Circulated

privately., May 1968. URL http://www.cs.utexas.edu/
users/EWD/OtherDocs/NN231.PDF.

Dijkstra:1968:LEG

[Dij68s] Edsger Wybe Dijkstra. Letter to the Editor: Go to statement considered harmful. Communications of the Association for Computing Machinery, 11(3):147–148, March 1968. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). This letter in support of structured programming, and in favor of eliminating control-flow disruption caused by go to statements, inspired scores of others, published mainly in SIGPLAN Notices up to the mid-1980s. The best-known is [Knu74].

Dijkstra:1968:GSC

[Dij68t] Edsger Wybe Dijkstra. Go to statement considered harmful. Communications of the Association for Computing Machinery, 11(3):147–148, March 1968. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). This letter to the Editor inspired scores of others, published mainly in SIGPLAN Notices up to the mid-1980s. The best-known is [Knu74].

Dijkstra:1969:NSPa

[Dij69a] E. W. Dijkstra. Notes on structured programming. Technical report, Teshnische Hogeschool, Eindhoven, The Netherlands, August 1969.

Dijkstra:1969:SPb

[Dij69b] E. W. Dijkstra. Structured programming. In Buxton and Randell [BR70], pages 84–88. URL http://homepages.cs.

ncl.ac.uk/brian.randell/NATO/nato1969.PDF. See also

one-paragraph comment by Dijkstra on page 21 on the importance of precise terminology.

Dijkstra:1968:CCH

[Dij69c] Edsger W. Dijkstra. Complexity controlled by hierarchical ordering of function and variability. In Naur and Randell [NR69], pages 181–185. LCCN QA 264 W675 1968.

Dijkstra:1969:CCH

[Dij69d] Edsger W. Dijkstra. Complexity controlled by hierarchical ordering of function and variability. In Software Engineering. NATO Science Committee, 1969. URL http://homepages.

cs.ncl.ac.uk/brian.randell/NATO/nato1968.PDF.

Dijkstra:1969:NSPb

[Dij69e] Edsger W. Dijkstra. Notes on structured programming. In Structured Programming. Academic Press, New York, NY, USA, 1969.

Dijkstra:1969:OVO

[Dij69f] Edsger W. Dijkstra. Organisatie van onderzoek en onderwijs van Fundamentele Programmering. (Dutch) [Organization of research and teaching of Basic Programming]. Circulated pri-

vately., November 1969. URL http://www.cs.utexas.edu/
users/EWD/ewd02xx/EWD278.PDF.

Dijkstra:1969:PTC

[Dij69g] Edsger W. Dijkstra. The programming task considered as an intellectual challenge. Circulated privately., Decem-

ber 1969. URL http://www.cs.utexas.edu/users/EWD/
ewd02xx/EWD273.PDF.

Dijkstra:1969:SEO

[Dij69h] Edsger W. Dijkstra. Structure of an extendable operating system. Circulated privately., November 1969. URL http:/

/www.cs.utexas.edu/users/EWD/ewd02xx/EWD275.PDF.

Dijkstra:1969:SPa

[Dij69i] Edsger W. Dijkstra. Structured programming. Circulated

privately., August 1969. URL http://www.cs.utexas.edu/
users/EWD/ewd02xx/EWD268.PDF.

Dijkstra:1969:TCO

[Dij69j] Edsger W. Dijkstra. Tentamen Co-operating Sequential Processes (jan. 1969). (Dutch) [Exam question Co-operating Sequential Processes (January 1969)]. Circulated privately.,

January 1969. URL http://www.cs.utexas.edu/users/
EWD/ewd02xx/EWD253.PDF.

Dijkstra:1969:VBA

[Dij69k] Edsger W. Dijkstra. Verslag bezoek aan MIT. (Dutch) [Report of a visit to MIT]. Circulated privately., April 1969. URL

http://www.cs.utexas.edu/users/EWD/ewd02xx/EWD257.
PDF.

Dijkstra:1970:CCG

[Dij70a] Edsger W. Dijkstra. Concern for correctness as a guiding principle for program construction. Circulated privately.,

July 1970. URL http://www.cs.utexas.edu/users/EWD/
ewd02xx/EWD288.PDF.

Dijkstra:1970:CPA

[Dij70b] Edsger W. Dijkstra. Cursus programmeren in Algol 60: 11e uitg. (Dutch) [Course in programming in Algol 60: 11th edition]. Technical report, Mathematisch Centrum, Amsterdam, Amsterdam, The Netherlands, 1970. URL http://www.cs.

utexas.edu/users/EWD/MCReps/CR1970-013.PDF.

Dijkstra:1970:DWM

[Dij70c] Edsger W. Dijkstra. Dear WG2.3 Member. Circulated pri-

vately., February 1970. URL http://www.cs.utexas.edu/
users/EWD/OtherDocs/NN282.PDF.

Dijkstra:1970:LPC

[Dij70d] Edsger W. Dijkstra. Letter to professor C. A. R. Hoare (31 August 1970). Circulated privately., August 1970. URL http:

//www.cs.utexas.edu/users/EWD/ewd02xx/EWD292.PDF.

Dijkstra:1970:NSPa

[Dij70e] Edsger W. Dijkstra. Notes on Structured Programming. Circulated privately., April 1970. URL http://www.cs.utexas.

edu/users/EWD/ewd02xx/EWD249.PDF.

Dijkstra:1970:NSPb

[Dij70f] Edsger W. Dijkstra. Notes on structured programming. Technical Report 70-WSK-03, Technological University Eindhoven, April 1970.

Dijkstra:1970:RVT

[Dij70g] Edsger W. Dijkstra. Recensie voor het tijdschrift “Informatie”. (Dutch) [Review for the journal “Information”]. Circulated privately., June 1970. URL http://www.cs.utexas.

edu/users/EWD/ewd02xx/EWD288a.PDF.

Dijkstra:1970:SP

[Dij70h] Edsger W. Dijkstra. Structured programming. In Software Engineering Techniques. NATO Science Committee, Au-

gust 1970. URL http://homepages.cs.ncl.ac.uk/brian.
randell/NATO/nato1969.PDF.

Dijkstra:1970:VVM

[Dij70i] Edsger W. Dijkstra. Verslag van mijn reis naar California. (Dutch) [Report of my trip to California]. Circulated

privately., April 1970. URL http://www.cs.utexas.edu/
users/EWD/ewd02xx/EWD286.PDF.

Dijkstra:1971:HOS

[Dij71a] Edsger W. Dijkstra. Hierarchical ordering of sequential processes. Acta Informatica, 1(2):115–138, October 1971. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). Reprinted in Operating Systems Techniques, C. A. R. Hoare and R. H. Perrot, Eds., Academic Press, 1972, pp. 72– 93. This paper introduces the classical synchronization problem of Dining Philosophers.

Dijkstra:1971:VCVa

[Dij71b] Edsger W. Dijkstra. (over een voorgestelde configuratie van een P1400 en vier P880’s). (Dutch) [(On a proposed configuration of a P1400 and four P880’s)]. Circulated

privately., 1971. URL http://www.cs.utexas.edu/users/
EWD/ewd03xx/EWD307.PDF.

Dijkstra:1971:VCVb

[Dij71c] Edsger W. Dijkstra. (over een voorgestelde configuratie van een P1400 en vier P880’s). (Dutch) [(On a proposed configuration of a P1400 and four P880’s)]. Circulated privately.,

April 1971. URL http://www.cs.utexas.edu/users/EWD/
ewd03xx/EWD309.PDF.

Dijkstra:1971:GVI

[Dij71d] Edsger W. Dijkstra. Over het gewicht van een informaticaopleiding. (Dutch) [About the weight of a computer science course]. Circulated privately., March 1971. URL http://

www.cs.utexas.edu/users/EWD/ewd03xx/EWD306.PDF.

Dijkstra:1971:RVE

[Dij71e] Edsger W. Dijkstra. Reisverslag van Edsger W. Dijkstra aan Summer School Marktoberdorf, juli 1971. (Dutch) [Trip report of Edsger W. Dijkstra to the Marktoberdorf Summer School, July 1971]. Circulated privately., Au-

gust 1971. URL http://www.cs.utexas.edu/users/EWD/
ewd03xx/EWD316A.PDF.

Dijkstra:1971:RBG

[Dij71f] Edsger W. Dijkstra. [Review of] Bauer, Goos: “Informatik, Erster Teil” Heidelberger Taschenbücher, Band 80, Springer Verlag, 1971. Circulated privately., May 1971. URL http:/

/www.cs.utexas.edu/users/EWD/ewd03xx/EWD311.PDF.

Dijkstra:1971:SIA

[Dij71g] Edsger W. Dijkstra. A short introduction to the art of programming. Circulated privately., August 1971. URL http:/

/www.cs.utexas.edu/users/EWD/ewd03xx/EWD316.PDF.

Dijkstra:1971:VVR

[Dij71h] Edsger W. Dijkstra. Verslag van reis van E. W. Dijkstra naar Noord Amerika. (Dutch) [Report of a trip by E. W. Dijkstra to North America]. Circulated privately.,

June 1971. URL http://www.cs.utexas.edu/users/EWD/
ewd03xx/EWD312.PDF.

Dijkstra:1972:NSP

[Dij72a] E. W. Dijkstra. Notes on structured programming. In Structured Programming [DDH72], pages 1–82. ISBN 0-12-200550-3 (hardcover), 0-12-200556-2 (paperback). LCCN QA76.6 .D33 1972.

Dijkstra:1972:BVW

[Dij72b] Edsger W. Dijkstra. Bezoek van E. W. Dijkstra aan l’Alpe d’Huez, 8–15 december 1972. (Dutch) [Visit by E. W. Dijkstra to l’Alpe d’Huez, 8–15 December 1972]. Circulated

privately., 1972. URL http://www.cs.utexas.edu/users/
EWD/ewd03xx/EWD362.PDF.

Dijkstra:1972:BRB

[Dij72c] Edsger W. Dijkstra. [Book review of: Bauer–Goos, Informatik, Zweiter Teil]. Circulated privately., March 1972. URL

http://www.cs.utexas.edu/users/EWD/ewd03xx/EWD330.
PDF.

Dijkstra:1972:CAS

[Dij72d] Edsger W. Dijkstra. A class of allocation strategies inducing bounded delays only. Circulated privately., 1972. URL http:

//www.cs.utexas.edu/users/EWD/ewd03xx/EWD319.PDF.

Dijkstra:1972:EEA

[Dij72e] Edsger W. Dijkstra. Elementen ener afstudeerrichting in de informatica. (Dutch) [Elements of energy specialization in computer science]. Circulated privately., Octo-

ber 1972. URL http://www.cs.utexas.edu/users/EWD/
ewd03xx/EWD353.PDF.

Dijkstra:1972:HOS

[Dij72f] Edsger W. Dijkstra. Hierarchical ordering of sequential processes. In Hoare and Perrott [HP72], pages 72–93. ISBN 0-12-350650-6. LCCN QA76.6 .I57 1971.

Dijkstra:1972:HPa

[Dij72g] Edsger W. Dijkstra. The humble programmer. Turing Award lecture; published as [Dij72h]., 1972. URL http://www.cs.

utexas.edu/users/EWD/ewd03xx/EWD340.PDF.

Dijkstra:1972:HPb

[Dij72h] Edsger W. Dijkstra. The humble programmer. Communications of the Association for Computing Machinery, 15(10): 859–866, October 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Turing Award lecture.

Dijkstra:1972:ISSa

[Dij72i] Edsger W. Dijkstra. Information streams sharing a finite buffer. Published as [Dij72j]., 1972. URL http://www.cs.

utexas.edu/users/EWD/ewd03xx/EWD329.PDF.

Dijkstra:1972:ISSb

[Dij72j] Edsger W. Dijkstra. Information streams sharing a finite buffer. Information Processing Letters, 1(5):179–180, October ??, 1972. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dijkstra:1972:OID

[Dij72k] Edsger W. Dijkstra. Ontwikkelingsplan Informatica. (Dutch) [Development of informatics]. Circulated privately., Septem-

ber 1972. URL http://www.cs.utexas.edu/users/EWD/
ewd03xx/EWD349.PDF.

Dijkstra:1972:PTP

[Dij72l] Edsger W. Dijkstra. Poging tot plaatsbepaling van de Informatica. (Dutch) [Attempt at locating informatics]. Circulated privately., April 1972. URL http://www.cs.utexas.

edu/users/EWD/ewd03xx/EWD325.PDF.

Dijkstra:1972:PMT

[Dij72m] Edsger W. Dijkstra. Programming methodologies, their objectives and their nature. In Structured Programming [DDH72], pages 203–216. ISBN 0-12-200550-3 (hardcover), 0-12-200556-2 (paperback). LCCN QA76.6 .D33 1972.

Dijkstra:1972:RBB

[Dij72n] Edsger W. Dijkstra. Reisverslag betreffende het bezoek aan de USA van 14 tot 30 mei 1972 door E. W. Dijkstra. (Dutch) [Trip report on the visit to the USA from 14 to 30 May 1972 by E. W. Dijkstra ]. Circulated privately., June 1972. URL http:

//www.cs.utexas.edu/users/EWD/ewd03xx/EWD343.PDF.

Dijkstra:1972:SWC

[Dij72o] Edsger W. Dijkstra. Selected writings on computing: a personal perspective. EWD, 477:60–66, 1972.

Dijkstra:1972:TZV

[Dij72p] Edsger W. Dijkstra. Ter zake van wiskundige modelvorming. (Dutch) [In respect of mathematical modeling]. Circulated privately., June 1972. URL http://www.cs.utexas.

edu/users/EWD/ewd03xx/EWD345.PDF.

Dijkstra:1972:VBE

[Dij72q] Edsger W. Dijkstra. Verslag bezoek Engeland 30 augustus– 9 september 1972. (Dutch) [Trip report to England 30 August–9 September 1972]. Circulated privately., Septem-

ber 1972. URL http://www.cs.utexas.edu/users/EWD/
ewd03xx/EWD351.PDF.

Dijkstra:1972:VVR

[Dij72r] Edsger W. Dijkstra. Verslag van de reis van E. W. Dijkstra naar Boston, 12–18 augustus 1972. (Dutch) [Report of the trip by E. W. Dijkstra to Boston, 12–18 August 1972]. Circulated

privately., August 1972. URL http://www.cs.utexas.edu/
users/EWD/ewd03xx/EWD348.PDF.

Dijkstra:1973:SABb

[Dij73a] E. W. Dijkstra. A simple axiomatic basis for programming language constructs. Technical report, Technical University of Eindhoven, Eindhoven, The Netherlands, 1973.

Dijkstra:1973:AMS

[Dij73b] Edsger W. Dijkstra. The analysis of multiprogrammed systems of unspecified degree of parallellism. Circulated

privately., May 1973. URL http://www.cs.utexas.edu/
users/EWD/ewd03xx/EWD374.PDF.

Dijkstra:1973:BVP

[Dij73c] Edsger W. Dijkstra. Betrouwbaarheid van programma’s. (Dutch) [Reliability of programs]. Circulated privately.,

July 1973. URL http://www.cs.utexas.edu/users/EWD/
ewd03xx/EWD384.PDF.

Dijkstra:1973:FMS

[Dij73d] Edsger W. Dijkstra. Finding the maximum strong components in a directed graph. Published as [Dij82q].,

May 1973. URL http://www.cs.utexas.edu/users/EWD/
ewd03xx/EWD376.PDF.

Dijkstra:1973:ISE

[Dij73e] Edsger W. Dijkstra. An immediate sequel to EWD398: “Sequencing primitives revisited”. Circulated privately., Novem-

ber 1973. URL http://www.cs.utexas.edu/users/EWD/
ewd03xx/EWD399.PDF.

Dijkstra:1973:NAE

[Dij73f] Edsger W. Dijkstra. A non algebraic example of a constructive correctness proof. Circulated privately., May 1973. URL

http://www.cs.utexas.edu/users/EWD/ewd03xx/EWD375.
PDF.

Dijkstra:1973:CPB

[Dij73g] Edsger W. Dijkstra. On a connection pattern between 2**N elements. Circulated privately., June 1973. URL http://

www.cs.utexas.edu/users/EWD/ewd03xx/EWD379.PDF.

Dijkstra:1973:RA

[Dij73h] Edsger W. Dijkstra. On representational abstraction. Circulated privately., November 1973. URL http://www.cs.

utexas.edu/users/EWD/ewd03xx/EWD393.PDF.

Dijkstra:1973:ADS

[Dij73i] Edsger W. Dijkstra. On the axiomatic definition of semantics. Circulated privately., April 1973. URL http://www.cs.

utexas.edu/users/EWD/ewd03xx/EWD367.PDF.

Dijkstra:1973:PDM

[Dij73j] Edsger W. Dijkstra. Programming as a discipline of mathematical nature. Published as [Dij74m]., May 1973. URL

http://www.cs.utexas.edu/users/EWD/ewd03xx/EWD361.
PDF.

Dijkstra:1973:RWD

[Dij73k] Edsger W. Dijkstra. Reisverslag E. W. Dijkstra: Open house on semantics, Aarhus, 7–18 jan. 1973. (Dutch) [Trip report of E. W. Dijkstra: Open house on semantics, Aarhus, 7–18 January 1973]. Circulated privately., 1973. URL http://

www.cs.utexas.edu/users/EWD/ewd03xx/EWD363.PDF.

Dijkstra:1973:SSSa

[Dij73l] Edsger W. Dijkstra. Self-stabilization in spite of distributed control. Published as [Dij82-40]., October 1973. URL http:

//www.cs.utexas.edu/users/EWD/ewd03xx/EWD391.PDF.

Dijkstra:1973:SSF

[Dij73m] Edsger W. Dijkstra. Self-stabilization with four-state machines. Circulated privately., October 1973. URL http://

www.cs.utexas.edu/users/EWD/ewd03xx/EWD392.PDF.

Dijkstra:1973:SST

[Dij73n] Edsger W. Dijkstra. Self-stabilization with three-state machines. Circulated privately., November 1973. URL http://

www.cs.utexas.edu/users/EWD/ewd03xx/EWD396.PDF.

Dijkstra:1973:SSSb

[Dij73o] Edsger W. Dijkstra. Self-stabilizing systems with distributed control. Circulated privately., November 1973. URL http:/

/www.cs.utexas.edu/users/EWD/ewd03xx/EWD397.PDF.

Dijkstra:1973:SPR

[Dij73p] Edsger W. Dijkstra. Sequencing primitives revisited. Circulated privately., November 1973. URL http://www.cs.

utexas.edu/users/EWD/ewd03xx/EWD398.PDF.

Dijkstra:1973:SABa

[Dij73q] Edsger W. Dijkstra. A simple axiomatic basis for programming language constructs. Circulated privately.,

May 1973. URL http://www.cs.utexas.edu/users/EWD/
ewd03xx/EWD372.PDF.

Dijkstra:1973:SCR

[Dij73r] Edsger W. Dijkstra. The solution to a cyclic relaxation problem. Published as [Dij82a]., August 1973. URL http://www.

cs.utexas.edu/users/EWD/ewd03xx/EWD386.PDF.

Dijkstra:1973:TRWb

[Dij73s] Edsger W. Dijkstra. Trip report E. W. Dijkstra IFIP Working Group W.G.2.3 “On Programming Methodology”, 21st– 26th October 1973, Blanchland, England. Circulated pri-

vately., October 1973. URL http://www.cs.utexas.edu/
users/EWD/ewd03xx/EWD395.PDF.

Dijkstra:1973:TRWa

[Dij73t] Edsger W. Dijkstra. Trip report E. W. Dijkstra Summer School Munich, July 25 to August 4, 1973. Published as

[?]., August 1973. URL http://www.cs.utexas.edu/users/
EWD/ewd03xx/EWD385.PDF.

Dijkstra:1973:TRI

[Dij73u] Edsger W. Dijkstra. Trip report IBM Seminar “Communication and Computers”, Newcastle, Sept. 1973. Published as [Dij82c]., September 1973. URL http://www.cs.utexas.

edu/users/EWD/ewd03xx/EWD387.PDF.

Dijkstra:1973:TRU

[Dij73v] Edsger W. Dijkstra. Trip report I.U.C.C. Colloquium, Canterbury, 18th–21st sept. 1973. Circulated privately., Septem-

ber 1973. URL http://www.cs.utexas.edu/users/EWD/
ewd03xx/EWD389.PDF.

Dijkstra:1973:TFD

[Dij73w] Edsger W. Dijkstra. A trip to France: 13th–20th December 1973. Circulated privately., December 1973. URL http://

www.cs.utexas.edu/users/EWD/ewd04xx/EWD404.PDF.

Dijkstra:1973:VVB

[Dij73x] Edsger W. Dijkstra. Verslag van bezoek aan Washington, maart 1973. (Dutch) [Report of a trip to Washington, March 1973]. Circulated privately., March 1973. URL http://www.

cs.utexas.edu/users/EWD/ewd03xx/EWD368.PDF.

Dijkstra:1974:LCR

[Dij74a] Edsger W. Dijkstra. [A letter to C. A. R. Hoare, 13 July 1974]. Circulated privately., July 1974. URL http://www.

cs.utexas.edu/users/EWD/ewd04xx/EWD432.PDF.

Dijkstra:1974:ASA

[Dij74b] Edsger W. Dijkstra. Acceptance speech for the AFIPS Harry Goode Memorial Award 1974. Published as [Dij82-70]., 1974.

URL http://www.cs.utexas.edu/users/EWD/ewd04xx/EWD407.
PDF.

Dijkstra:1974:CCA

[Dij74c] Edsger W. Dijkstra. Correctness concerns and, among other things, why they are resented. invited paper, to be presented at the International Conference on Reliable Software, Los Angeles, 21–23 April 1975; circulated privately, Novem-

ber 1974. URL http://www.cs.utexas.edu/users/EWD/
ewd04xx/EWD450.PDF.

Dijkstra:1974:DRV

[Dij74d] Edsger W. Dijkstra. Determinism and recursion versus nondeterminism and the transitive closure. Circulated privately.,

October 1974. URL http://www.cs.utexas.edu/users/
EWD/ewd04xx/EWD456.PDF.

Dijkstra:1974:GCN

[Dij74e] Edsger W. Dijkstra. Guarded commands, non-determinacy and a calculus for the derivation of programs. See [Dij75y]; circulated privately, June 1974. URL http://www.cs.

utexas.edu/users/EWD/ewd04xx/EWD418.PDF.

Dijkstra:1974:IMI

[Dij74f] Edsger W. Dijkstra. Inside “Mathematics Inc.”. Published as

[Dij82w]., 1974. URL http://www.cs.utexas.edu/users/
EWD/ewd04xx/EWD442.PDF.

Dijkstra:1974:LDH

[Dij74g] Edsger W. Dijkstra. [Letter to Dr. H. Bekic, 8 October 1974]. Circulated privately., October 1974. URL http://www.cs.

utexas.edu/users/EWD/ewd04xx/EWD454.PDF.

Dijkstra:1974:MRA

[Dij74h] Edsger W. Dijkstra. Monotonic replacement algorithms and their implementation. Published as [Dij82-83]., Decem-

ber 1974. URL http://www.cs.utexas.edu/users/EWD/
ewd04xx/EWD465.PDF.

Dijkstra:1974:MAM

[Dij74i] Edsger W. Dijkstra. A multidisciplinary approach to mathematics. Published as [Dij82-31]., August 1974. URL http:/

/www.cs.utexas.edu/users/EWD/ewd04xx/EWD443.PDF.

Dijkstra:1974:NEB

[Dij74j] Edsger W. Dijkstra. A new elephant built from mosquitos humming in harmony. Published as [Dij82-86]., Novem-

ber 1974. URL http://www.cs.utexas.edu/users/EWD/
ewd04xx/EWD464.PDF.

Dijkstra:1974:ASV

[Dij74k] Edsger W. Dijkstra. On the abolishment of the subscripted variable. Circulated privately., March 1974. URL http://

www.cs.utexas.edu/users/EWD/ewd04xx/EWD417.PDF.

Dijkstra:1974:RST

[Dij74l] Edsger W. Dijkstra. On the role of scientific thought. Published as [Dij82-92]. Reprinted in [Dij82-69, pages 60–66]., Au-

gust 1974. URL http://www.cs.utexas.edu/users/EWD/
ewd04xx/EWD447.PDF.

Dijkstra:1974:PDMa

[Dij74m] Edsger W. Dijkstra. Programming as a discipline of mathematical nature. American Mathematical Monthly, 81(6):608– 612, June/July 1974. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.

org/stable/2319209.

Dijkstra:1974:SSSa

[Dij74n] Edsger W. Dijkstra. Self-stabilizing systems in spite of distributed control. Published as [Dij74o]., June 1974. URL

http://www.cs.utexas.edu/users/EWD/ewd04xx/EWD426.
PDF.

Dijkstra:1974:SSSb

[Dij74o] Edsger W. Dijkstra. Self-stabilizing systems in spite of distributed control. Communications of the Association for Computing Machinery, 17(11):643–644, November 1974. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Winner of the 2002 ACM Influential Paper Award.

Dijkstra:1974:SAB

[Dij74p] Edsger W. Dijkstra. A simple axiomatic basis for programming language constructs. Nederl. Akad. Wetensch. Proc. Ser. A 77=Indag. Math., 36:1–15, 1974.

Dijkstra:1974:SQ

[Dij74q] Edsger W. Dijkstra. Some questions. Circulated privately.,

November 1974. URL http://www.cs.utexas.edu/users/
EWD/ewd04xx/EWD463.PDF.

Dijkstra:1974:SOA

[Dij74r] Edsger W. Dijkstra. Speech at the occasion of an anniversary. Published as [Dij82-44]., June 1974. URL http://www.cs.

utexas.edu/users/EWD/ewd04xx/EWD427.PDF.

Dijkstra:1974:TWH

[Dij74s] Edsger W. Dijkstra. A time-wise hierarchy imposed upon the use of a two-level store. Published as [Dij82-104]., Decem-

ber 1974. URL http://www.cs.utexas.edu/users/EWD/
ewd04xx/EWD462.PDF.

Dijkstra:1974:TRWc

[Dij74t] Edsger W. Dijkstra. Trip report E. W. Dijkstra, Edinburgh and Newcastle, 1–6 September 1974. Circulated privately.,

September 1974. URL http://www.cs.utexas.edu/users/
EWD/ewd04xx/EWD448.PDF.

Dijkstra:1974:TRW

[Dij74u] Edsger W. Dijkstra. Trip report E. W. Dijkstra, Luxembourg, 7–12 April 1974. Circulated privately., April 1974. URL

http://www.cs.utexas.edu/users/EWD/ewd04xx/EWD420.
PDF.

Dijkstra:1974:TRWd

[Dij74v] Edsger W. Dijkstra. Trip report E. W. Dijkstra, Meeting IFIP W.G.2.3., Munich, 8 –14 December 1974. Published as [Dij82-51]., December 1974. URL http://www.cs.utexas.

edu/users/EWD/ewd04xx/EWD466.PDF.

Dijkstra:1974:TRWb

[Dij74w] Edsger W. Dijkstra. Trip report E. W. Dijkstra, USA and Canada, 5th–25th May 1974. Circulated privately.,

June 1974. URL http://www.cs.utexas.edu/users/EWD/
ewd04xx/EWD425.PDF.

Dijkstra:1974:TRWa

[Dij74x] Edsger W. Dijkstra. Trip report E. W. Dijkstra, W.G.2.3 Meeting “Boldern” 28th April–3rd May 1974. Circulated

privately., June 1974. URL http://www.cs.utexas.edu/
users/EWD/ewd04xx/EWD423.PDF.

Dijkstra:1974:TUJ

[Dij74y] Edsger W. Dijkstra. A trip to the U.S.A., 5th–25th January 1974. Circulated privately., January 1974. URL http://www.

cs.utexas.edu/users/EWD/ewd04xx/EWD406.PDF.

Dijkstra:1975:ARL

[Dij75a] Edsger W. Dijkstra. About robustness and the like. Circulated privately., January 1975. URL http://www.cs.

utexas.edu/users/EWD/ewd04xx/EWD452.PDF.

Dijkstra:1975:ATN

[Dij75b] Edsger W. Dijkstra. Achtste toespraak, najaar 1975. (Dutch) [Eighth lecture, Autumn 1975]. Circulated privately., Novem-

ber 1975. URL http://www.cs.utexas.edu/users/EWD/
ewd05xx/EWD529.PDF.

Dijkstra:1975:ATT

[Dij75c] Edsger W. Dijkstra. Achtste toespraak tot mijn studenten. (Dutch) [Eighth lecture to my students]. Circulated privately.,

May 1975. URL http://www.cs.utexas.edu/users/EWD/
ewd04xx/EWD495.PDF.

Dijkstra:1975:AMS

[Dij75d] Edsger W. Dijkstra. After many a sobering experience. Circulated privately., June 1975. URL http://www.cs.utexas.

edu/users/EWD/ewd05xx/EWD500.PDF.

Dijkstra:1975:AJM

[Dij75e] Edsger W. Dijkstra. An answer to Jack Mazola. Circulated privately., November 1975. URL http://www.cs.utexas.

edu/users/EWD/ewd05xx/EWD535.PDF.

Dijkstra:1975:CMD

[Dij75f] Edsger W. Dijkstra. A challenge to memory designers? Circulated privately., June 1975. URL http://www.cs.utexas.

edu/users/EWD/ewd04xx/EWD497.PDF.

Dijkstra:1975:CBP

[Dij75g] Edsger W. Dijkstra. A collection of beautiful proofs. Published as [Dij82i]., November 1975. URL http://www.cs.

utexas.edu/users/EWD/ewd05xx/EWD538.PDF.

Dijkstra:1975:COS

[Dij75h] Edsger W. Dijkstra. Commentaar op “Structuurplan Informatica (W.O.)” van de ARSI. (Dutch) [Comments on “Structure Informatics (WO)” of the ARSI]. Circulated privately.,

February 1975. URL http://www.cs.utexas.edu/users/
EWD/ewd04xx/EWD479.PDF.

Dijkstra:1975:CSc

[Dij75i] Edsger W. Dijkstra. Comments at a symposium. Published as [Dij82j]., September 1975. URL http://www.cs.utexas.

edu/users/EWD/ewd05xx/EWD512.PDF.

Dijkstra:1975:CWH

[Dij75j] Edsger W. Dijkstra. Comments on “Woodenman” HOL Requirements for the DoD. Circulated privately., Novem-

ber 1975. URL http://www.cs.utexas.edu/users/EWD/
ewd05xx/EWD526.PDF.

Dijkstra:1975:CPP

[Dij75k] Edsger W. Dijkstra. Concurrent programming: a preliminary investigation. Circulated privately., February 1975. URL

http://www.cs.utexas.edu/users/EWD/ewd04xx/EWD476.
PDF.

Dijkstra:1975:CCA

[Dij75l] Edsger W. Dijkstra. Correctness concerns and, among other things, why they are resented. ACM SIGPLAN Notices, 10 (6):546–550, June 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dijkstra:1975:CSa

[Dij75m] Edsger W. Dijkstra. “Craftsman or scientist?”. Published as [Dij82-72]., February 1975. URL http://www.cs.utexas.

edu/users/EWD/ewd04xx/EWD480.PDF.

Dijkstra:1975:CSb

[Dij75n] Edsger W. Dijkstra. Craftsman or scientist? In Proceedings ACM Pacific 75, pages 217–223. ACM Press, New York, NY 10036, USA, 1975.

Dijkstra:1975:DTN

[Dij75o] Edsger W. Dijkstra. Derde toespraak, najaar 1975. (Dutch) [Third lecture, Autumn 1975]. Circulated privately., Septem-

ber 1975. URL http://www.cs.utexas.edu/users/EWD/
ewd05xx/EWD516.PDF.

Dijkstra:1975:DTT

[Dij75p] Edsger W. Dijkstra. Derde toespraak tot mijn studenten. (Dutch) [Third lecture to my students]. Circulated privately.,

March 1975. URL http://www.cs.utexas.edu/users/EWD/
ewd04xx/EWD481.PDF.

Dijkstra:1975:DLT

[Dij75q] Edsger W. Dijkstra. Dertiende en laatste toespraak, najaar 1975. (Dutch) [Thirteenth and last lecture, Autumn 1975]. Circulated privately., December 1975. URL http://www.cs.

utexas.edu/users/EWD/ewd05xx/EWD541.PDF.

Dijkstra:1975:ETNa

[Dij75r] Edsger W. Dijkstra. Eerste toespraak, najaar 1975. (Dutch) [First lecture, Autumn 1975]. Circulated pri-

vately., 1975. URL http://www.cs.utexas.edu/users/
EWD/ewd05xx/EWD509.PDF.

Dijkstra:1975:ETT

[Dij75s] Edsger W. Dijkstra. Eerste toespraak tot mijn studenten, voorjaar 1976. (Dutch) [First lecture to my students, Spring 1976]. Circulated privately., December 1975. URL http://

www.cs.utexas.edu/users/EWD/ewd05xx/EWD543.PDF.

Dijkstra:1975:ETNb

[Dij75t] Edsger W. Dijkstra. Elfde toespraak, najaar 1975. (Dutch) [Eleventh lecture, Autumn 1975]. Circulated privately.,

November 1975. URL http://www.cs.utexas.edu/users/
EWD/ewd05xx/EWD534.PDF.

Dijkstra:1975:EEE

[Dij75u] Edsger W. Dijkstra. Erratum and embellishments of EWD503. Published as [Dij82-75]., July 1975. URL http:/

/www.cs.utexas.edu/users/EWD/ewd05xx/EWD504.PDF.

Dijkstra:1975:EMP

[Dij75v] Edsger W. Dijkstra. Exercises in making programs robust. Published as [Dij82-77]., March 1975. URL http://www.cs.

utexas.edu/users/EWD/ewd04xx/EWD482.PDF.

Dijkstra:1975:GCNc

[Dij75w] Edsger W. Dijkstra. Guarded commands, non-determinacy and a calculus for the derivation of programs. ACM SIG-PLAN Notices, 10(6):2, 2.1–2.13, June 1975. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dijkstra:1975:GCNa

[Dij75x] Edsger W. Dijkstra. Guarded commands, non-determinacy and formal derivation of programs. Published as [Dij75y].,

January 1975. URL http://www.cs.utexas.edu/users/
EWD/ewd04xx/EWD472.PDF.

Dijkstra:1975:GCNd

[Dij75y] Edsger W. Dijkstra. Guarded commands, nondeterminacy and formal derivation of programs. Communications of the Association for Computing Machinery, 18(8):453–457, August 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dijkstra:1975:HC

[Dij75z] Edsger W. Dijkstra. HOMO COGITANS. Circulated pri-

vately., November 1975. URL http://www.cs.utexas.edu/
users/EWD/ewd05xx/EWD533.PDF.

Dijkstra:1975:HDW

[Dij75-27] Edsger W. Dijkstra. How do we tell truths that might hurt? Published as [Dij82-79]., June 1975. URL http://www.cs.

utexas.edu/users/EWD/ewd04xx/EWD498.PDF.

Dijkstra:1975:LMO

[Dij75-28] Edsger W. Dijkstra. A letter to my old friend Jonathan. Published as [Dij82z]., February 1975. URL http://www.

cs.utexas.edu/users/EWD/ewd04xx/EWD475.PDF.

Dijkstra:1975:LBR

[Dij75-29] Edsger W. Dijkstra. Letter to the Burroughs recipients of the EWD-series. Circulated privately., March 1975. URL http:

//www.cs.utexas.edu/users/EWD/ewd04xx/EWD487.PDF.

Dijkstra:1975:LRE

[Dij75-30] Edsger W. Dijkstra. Letter to the referees of EWD418. Circulated privately., January 1975. URL http://www.cs.

utexas.edu/users/EWD/ewd04xx/EWD470.PDF.

Dijkstra:1975:MQD

[Dij75-31] Edsger W. Dijkstra. Marketing questionnaire “A Discipline of Programming”. Circulated privately., March 1975. URL

http://www.cs.utexas.edu/users/EWD/ewd04xx/EWD485.
PDF.

Dijkstra:1975:MIP

[Dij75-32] Edsger W. Dijkstra. Mathematics inc., a private letter from its chairman. Published as [Dij82-82]., Decem-

ber 1975. URL http://www.cs.utexas.edu/users/EWD/
ewd05xx/EWD539.PDF.

Dijkstra:1975:MHW

[Dij75-33] Edsger W. Dijkstra. More on Hauck’s warning. Published as [Dij82-85]., November 1975. URL http://www.cs.utexas.

edu/users/EWD/ewd05xx/EWD528.PDF.

Dijkstra:1975:NTN

[Dij75-34] Edsger W. Dijkstra. Negende toespraak, najaar 1975. (Dutch) [Ninth lecture, Autumn 1975]. Circulated privately., Novem-

ber 1975. URL http://www.cs.utexas.edu/users/EWD/
ewd05xx/EWD530.PDF.

Dijkstra:1975:NRT

[Dij75-35] Edsger W. Dijkstra. Notes on a real-time garbage collection system. From a conversation with D. E. Knuth (private collection of D. E. Knuth), 1975.

Dijkstra:1975:GTDa

[Dij75-36] Edsger W. Dijkstra. On a gauntlet thrown by David Gries. Circulated privately., August 1975. URL http://www.cs.

utexas.edu/users/EWD/ewd05xx/EWD502.PDF.

Dijkstra:1975:GTDb

[Dij75-37] Edsger W. Dijkstra. On a gauntlet thrown by David Gries. Circulated privately., August 1975. URL http://www.cs.

utexas.edu/users/EWD/ewd05xx/EWD507.PDF.

Dijkstra:1975:LPD

[Dij75-38] Edsger W. Dijkstra. On a language proposal for the Department of Defense. Circulated privately., Septem-

ber 1975. URL http://www.cs.utexas.edu/users/EWD/
ewd05xx/EWD514.PDF.

Dijkstra:1975:WH

[Dij75-39] Edsger W. Dijkstra. On a warning from E. A. Hauck. Published as [Dij82-88]., October 1975. URL http://www.cs.

utexas.edu/users/EWD/ewd05xx/EWD525.PDF.

Dijkstra:1975:OSS

[Dij75-40] Edsger W. Dijkstra. On one-sided smoothing of event sequences. Circulated privately., February 1975. URL http:/

/www.cs.utexas.edu/users/EWD/ewd04xx/EWD478.PDF.

Dijkstra:1975:FGCa

[Dij75-41] Edsger W. Dijkstra. On-the-fly garbage collection: an exercise in multiprocessing. Circulated privately. Published in

[DLM^{+}78]., April 1975. URL http://www.cs.utexas.edu/
users/EWD/ewd04xx/EWD492.PDF.

Dijkstra:1975:TPTa

[Dij75-42] Edsger W. Dijkstra. On the teaching of programming, i.e. on the teaching of thinking. Circulated privately., Febru-

ary 1975. URL http://www.cs.utexas.edu/users/EWD/
ewd04xx/EWD473.PDF.

Dijkstra:1975:UC

[Dij75-43] Edsger W. Dijkstra. On units of consistency. Circulated pri-

vately., November 1975. URL http://www.cs.utexas.edu/
users/EWD/ewd05xx/EWD527.PDF.

Dijkstra:1975:OLB

[Dij75-44] Edsger W. Dijkstra. An open letter to L. Bass. Circulated privately., November 1975. URL http://www.cs.utexas.

edu/users/EWD/ewd05xx/EWD532.PDF.

Dijkstra:1975:OLR

[Dij75-45] Edsger W. Dijkstra. An open letter to Ross Honsberger. Circulated privately., December 1975. URL http://www.cs.

utexas.edu/users/EWD/ewd05xx/EWD544.PDF.

Dijkstra:1975:PSE

[Dij75-46] Edsger W. Dijkstra. A post-scriptum to EWD501. Published as [Dij82-36]., July 1975. URL http://www.cs.utexas.edu/

users/EWD/ewd05xx/EWD503.PDF.

Dijkstra:1975:RFS

[Dij75-47] Edsger W. Dijkstra. Review of “On the Feasibility of Software Certification”. Circulated privately., October 1975. URL

http://www.cs.utexas.edu/users/EWD/ewd05xx/EWD524.
PDF.

Dijkstra:1975:SEb

[Dij75-48] Edsger W. Dijkstra. A sequel to EWD535. Circulated pri-

vately., November 1975. URL http://www.cs.utexas.edu/
users/EWD/ewd05xx/EWD536.PDF.

Dijkstra:1975:SEa

[Dij75-49] Edsger W. Dijkstra. A synthesis emerging? Published as [Dij82-101]., August 1975. URL http://www.cs.utexas.

edu/users/EWD/ewd05xx/EWD508.PDF.

Dijkstra:1975:TTNb

[Dij75-50] Edsger W. Dijkstra. Tiende toespraak, najaar 1975. (Dutch) [Tenth lecture, Autumn 1975]. Circulated privately., Novem-

ber 1975. URL http://www.cs.utexas.edu/users/EWD/
ewd05xx/EWD531.PDF.

Dijkstra:1975:TTT

[Dij75-51] Edsger W. Dijkstra. [Toespraak tot twee afstudeerders]. (Dutch) [Addressing two graduations]. Circulated pri-

vately., 1975. URL http://www.cs.utexas.edu/users/
EWD/ewd05xx/EWD542.PDF.

Dijkstra:1975:TRW

[Dij75-52] Edsger W. Dijkstra. Trip report E. W. Dijkstra 16th April/7th May, U.S.A. and Canada. Published as [Dij82-49].,

May 1975. URL http://www.cs.utexas.edu/users/EWD/
ewd04xx/EWD494.PDF.

Dijkstra:1975:TRV

[Dij75-53] Edsger W. Dijkstra. Trip report visit ETH Zürich, 3–4 February 1975 by E. W. Dijkstra. Published as [Dij82-57].,

February 1975. URL http://www.cs.utexas.edu/users/
EWD/ewd04xx/EWD474.PDF.

Dijkstra:1975:TWDa

[Dij75-54] Edsger W. Dijkstra. Tripreport E. W. Dijkstra, London, 9–10 March 1975. Circulated privately., March 1975. URL http:

//www.cs.utexas.edu/users/EWD/ewd04xx/EWD484.PDF.

Dijkstra:1975:TWDb

[Dij75-55] Edsger W. Dijkstra. Tripreport E. W. Dijkstra: NATO Summer School Marktoberdorf 1975. Circulated privately., Au-

gust 1975. URL http://www.cs.utexas.edu/users/EWD/
ewd05xx/EWD506.PDF.

Dijkstra:1975:TWDd

[Dij75-56] Edsger W. Dijkstra. Tripreport E. W. Dijkstra Newcastle, 8–12 September 1975. Published as [Dij82-60]., Septem-

ber 1975. URL http://www.cs.utexas.edu/users/EWD/
ewd05xx/EWD513.PDF.

Dijkstra:1975:TWDc

[Dij75-57] Edsger W. Dijkstra. Tripreport E. W. Dijkstra W.G.2.3, Baden, 1–5 September 1975. Circulated privately., Septem-

ber 1975. URL http://www.cs.utexas.edu/users/EWD/
ewd05xx/EWD511.PDF.

Dijkstra:1975:TTNc

[Dij75-58] Edsger W. Dijkstra. Twaalfde toespraak, najaar 1975. (Dutch) [Twelfth lecture, Autumn 1975]. Circulated pri-

vately., November 1975. URL http://www.cs.utexas.edu/
users/EWD/ewd05xx/EWD537.PDF.

Dijkstra:1975:TTNa

[Dij75-59] Edsger W. Dijkstra. Tweede toespraak, najaar 1975. (Dutch) [Second lecture, Autumn 1975]. Circulated privately.,

September 1975. URL http://www.cs.utexas.edu/users/
EWD/ewd05xx/EWD515.PDF.

Dijkstra:1975:TVP

[Dij75-60] Edsger W. Dijkstra. Two views of programming. Circulated

privately., 1975. URL http://www.cs.utexas.edu/users/
EWD/ewd05xx/EWD540.PDF.

Dijkstra:1975:VTO

[Dij75-61] Edsger W. Dijkstra. Variations on a theme: an open letter to C. A. R. Hoare. Published as [Dij82-66]., July 1975. URL

http://www.cs.utexas.edu/users/EWD/ewd05xx/EWD501.
PDF.

Dijkstra:1975:VTNa

[Dij75-62] Edsger W. Dijkstra. Vierde toespraak, najaar 1975. (Dutch) [Fourth lecture, Autumn 1975]. Circulated privately., Septem-

ber 1975. URL http://www.cs.utexas.edu/users/EWD/
ewd05xx/EWD518.PDF.

Dijkstra:1975:VTT

[Dij75-63] Edsger W. Dijkstra. Vierde toespraak tot mijn studenten. (Dutch) [Fourth lecture to my students]. Circulated

privately., March 1975. URL http://www.cs.utexas.edu/
users/EWD/ewd04xx/EWD483.PDF.

Dijkstra:1975:VTNb

[Dij75-64] Edsger W. Dijkstra. Vijfde toespraak, najaar 1975. (Dutch) [Fifth lecture, Autumn 1975]. Circulated privately., Octo-

ber 1975. URL http://www.cs.utexas.edu/users/EWD/
ewd05xx/EWD521.PDF.

Dijkstra:1975:ZTNa

[Dij75-65] Edsger W. Dijkstra. Zesde toespraak, najaar 1975. (Dutch) [Sixth lecture, Autumn 1975]. Circulated privately., Oc-

tober 1975. URL http://www.cs.utexas.edu/users/EWD/
ewd05xx/EWD522.PDF.

Dijkstra:1975:ZTTa

[Dij75-66] Edsger W. Dijkstra. Zesde toespraak tot mijn studenten. (Dutch) [Sixth lecture to my students]. Circulated privately.,

March 1975. URL http://www.cs.utexas.edu/users/EWD/
ewd04xx/EWD489.PDF.

Dijkstra:1975:ZTNb

[Dij75-67] Edsger W. Dijkstra. Zevende toespraak, najaar 1975. (Dutch) [Seventh lecture, Autumn 1975]. Circulated privately., Oc-

tober 1975. URL http://www.cs.utexas.edu/users/EWD/
ewd05xx/EWD523.PDF.

Dijkstra:1975:ZTTb

[Dij75-68] Edsger W. Dijkstra. Zevende toespraak tot mijn studenten. (Dutch) [Seventh lecture to my students]. Circulated

privately., April 1975. URL http://www.cs.utexas.edu/
users/EWD/ewd04xx/EWD494a.PDF.

Dijkstra:1976:SP

[Dij76a] E. W. Dijkstra. Structured programming. In Naur et al. [NRB76], pages 222–226. ISBN 0-88405-334-2. LCCN QA76.6 S618 1968. Republication of [Dij69b].

Dijkstra:1976:ARVa

[Dij76b] Edsger W. Dijkstra. Aan de Raad van Advies. (Dutch) [The Advice Council]. Circulated privately., March 1976. URL

http://www.cs.utexas.edu/users/EWD/ewd05xx/EWD555.
PDF.

Dijkstra:1976:ARVb

[Dij76c] Edsger W. Dijkstra. Aan de Raad van Advies, in tweede ronde. (Dutch) [The Advice Council, second round]. Circulated privately., April 1976. URL http://www.cs.utexas.

edu/users/EWD/ewd05xx/EWD564a.PDF.

Dijkstra:1976:ATTb

[Dij76d] Edsger W. Dijkstra. Achtste toespraak tot mijn studenten, najaar 1976. (Dutch) [Eighth lecture to my students, Autumn 1976]. Circulated privately., 1976. URL http://www.cs.

utexas.edu/users/EWD/ewd05xx/EWD599.PDF.

Dijkstra:1976:ATTa

[Dij76e] Edsger W. Dijkstra. Achtste toespraak tot mijn studenten, voorjaar 1976. (Dutch) [Eighth lecture to my students, Spring 1976]. Circulated privately., 1976. URL http://www.cs.

utexas.edu/users/EWD/ewd05xx/EWD558.PDF.

Dijkstra:1976:AV

[Dij76f] Edsger W. Dijkstra. Array variables, chapter 11. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1976.

Dijkstra:1976:BMB

[Dij76g] Edsger W. Dijkstra. A bug in my book! Circulated

privately., 1976. URL http://www.cs.utexas.edu/users/
EWD/ewd05xx/EWD598.PDF.

Dijkstra:1976:CS

[Dij76h] Edsger W. Dijkstra. The characterization of semantics, chapter 3. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1976.

Dijkstra:1976:CCH

[Dij76i] Edsger W. Dijkstra. Complexity controlled by hierarchical ordering of function and variability. In Naur et al. [NRB76], pages 114–116. ISBN 0-88405-334-2. LCCN QA76.6 S618 1968. Conferences sponsored by the NATO Science Committee, held at Garmisch, Germany, 7–11 October 1968, and at Rome, Italy, 27–31 October 1969.

Dijkstra:1976:DTTb

[Dij76j] Edsger W. Dijkstra. Derde toespraak tot mijn studenten, najaar 1976. (Dutch) [Third lecture to my students, Autumn 1976]. Circulated privately., 1976. URL http://www.cs.

utexas.edu/users/EWD/ewd05xx/EWD587.PDF.

Dijkstra:1976:DTTa

[Dij76k] Edsger W. Dijkstra. Derde toespraak tot mijn studenten, voorjaar 1976. (Dutch) [Third lecture to my students, Spring 1976]. Circulated privately., 1976. URL http://www.cs.

utexas.edu/users/EWD/ewd05xx/EWD549.PDF.

Dijkstra:1976:DLT

[Dij76l] Edsger W. Dijkstra. Dertiende en laatste toespraak tot mijn studenten, voorjaar 1976. (Dutch) [Thirteenth and last lecture to my students, Spring 1976]. Circulated pri-

vately., 1976. URL http://www.cs.utexas.edu/users/
EWD/ewd05xx/EWD569.PDF.

Dijkstra:1976:DP

[Dij76m] Edsger W. Dijkstra. A discipline of programming. Prentice-Hall Series in Automatic Computation. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1976. ISBN 0-13-215871-X. xvii + 217 pp. LCCN QA76.6 .D54. With a foreword by C. A. R. Hoare.

Dijkstra:1976:ETTb

[Dij76n] Edsger W. Dijkstra. Eerste toespraak tot mijn studenten, najaar 1976. (Dutch) [First lecture to my students, Autumn 1976]. Circulated privately., 1976. URL http://www.cs.

utexas.edu/users/EWD/ewd05xx/EWD583.PDF.

Dijkstra:1976:EALa

[Dij76o] Edsger W. Dijkstra. The effective arrangement of logical systems. Circulated privately., May 1976. URL http://www.

cs.utexas.edu/users/EWD/ewd05xx/EWD562.PDF.

Dijkstra:1976:EALb

[Dij76p] Edsger W. Dijkstra. The effective arrangement of logical systems. In Mazurkiewicz [Maz76], pages 40–51. CODEN LNCSD9. ISBN 0-387-07854-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.5.S4

S989 1976. URL http://link.springer-ny.com/link/
service/series/0558/tocs/t0045.htm;   http://www.
springerlink.com/content/978-0-387-07854-0;   http:
//www.springerlink.com/openurl.asp?genre=issue&issn=

0302-9743&volume=45. “The Symposium is being organized by the Computation Centre of the Polish Academy of Sciences in cooperation with the University of Gdansk.”.

Dijkstra:1976:ETTa

[Dij76q] Edsger W. Dijkstra. Elfde toespraak tot mijn studenten, voorjaar 1976. (Dutch) [Eleventh lecture to my students, Spring 1976]. Circulated privately., 1976. URL http://www.cs.

utexas.edu/users/EWD/ewd05xx/EWD565.PDF.

Dijkstra:1976:ENS

[Dij76r] Edsger W. Dijkstra. An essay on the notion: “the scope of variables”, chapter 10. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1976.

Dijkstra:1976:EARa

[Dij76s] Edsger W. Dijkstra. Euclid’s Algorithm revisited, chapter 7. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1976.

Dijkstra:1976:EA

[Dij76t] Edsger W. Dijkstra. Executional abstraction, chapter 0. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1976.

Dijkstra:1976:EARb

[Dij76u] Edsger W. Dijkstra. An exercise attributed to R. W. Hamming, chapter 17. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1976.

Dijkstra:1976:EDR

[Dij76v] Edsger W. Dijkstra. An exercise for Dr. R. M. Burstall. Published as [Dij82-76]., May 1976. URL http://www.cs.

utexas.edu/users/EWD/ewd05xx/EWD570.PDF.

Dijkstra:1976:FMS

[Dij76w] Edsger W. Dijkstra. Finding the maximal strong components in a directed graph, chapter 25. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1976.

Dijkstra:1976:FIC

[Dij76x] Edsger W. Dijkstra. A first investigation of the crossflow computer. Circulated privately., October 1976. URL http:/

/www.cs.utexas.edu/users/EWD/ewd05xx/EWD590.PDF.

Dijkstra:1976:FTSa

[Dij76y] Edsger W. Dijkstra. Formal techniques and sizeable programs. Published as [Dij82r]., May 1976. URL http://www.

cs.utexas.edu/users/EWD/ewd05xx/EWD563.PDF.

Dijkstra:1976:FTSb

[Dij76z] Edsger W. Dijkstra. Formal techniques and sizeable programs. In Samelson [Sam76], pages 225–235. ISBN 3-540-07804-5.

Dijkstra:1976:FTSc

[Dij76-27] Edsger W. Dijkstra. The formal treatment of some small examples, chapter 8. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1976.

Dijkstra:1976:GI

[Dij76-28] Edsger W. Dijkstra. A great improvement. Published as

[Dij82-78]., July 1976. URL http://www.cs.utexas.edu/
users/EWD/ewd05xx/EWD573.PDF.

Dijkstra:1976:GCN

[Dij76-29] Edsger W. Dijkstra. Guarded commands, non-determinacy and a calculus for the derivation of programs. In Bauer and Samelson [BS76], pages 111–124. ISBN 0-387-07994-7 (New York), 3-540-07994-7 (Berlin), 3-540-37972-X (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

QA76.6 .L335.  URL http://link.springer-ny.com/
link/service/series/0558/tocs/t0046.htm;   http:/
/www.springerlink.com/content/978-0-387-07994-3;
http://www.springerlink.com/openurl.asp?genre=book&
isbn=978-3-540-07994-1;   http://www.springerlink.
com/openurl.asp?genre=issue&issn=0302-9743&volume=
46.

Dijkstra:1976:R

[Dij76-30] Edsger W. Dijkstra. In retrospect, chapter 27. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1976.

Dijkstra:1976:LPZ

[Dij76-31] Edsger W. Dijkstra. A letter to Professor Zohar Manna, 26 July 1976. Circulated privately., July 1976. URL http://

www.cs.utexas.edu/users/EWD/ewd05xx/EWD574.PDF.

Dijkstra:1976:LST

[Dij76-32] Edsger W. Dijkstra. The Linear Search Theorem, chapter 12. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1976.

Dijkstra:1976:MPR

[Dij76-33] Edsger W. Dijkstra. Merging problems revisited, chapter 16. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1976.

Dijkstra:1976:MAF

[Dij76-34] Edsger W. Dijkstra. More about the function “fusc” (a sequel to EWD570). Published as [Dij82-84]., Au-

gust 1976. URL http://www.cs.utexas.edu/users/EWD/
ewd05xx/EWD578.PDF.

Dijkstra:1976:MFT

[Dij76-35] Edsger W. Dijkstra. A more formal treatment of a less simple example. Circulated privately., March 1976. URL http://

www.cs.utexas.edu/users/EWD/ewd05xx/EWD550.PDF.

Dijkstra:1976:NTT

[Dij76-36] Edsger W. Dijkstra. Negende toespraak tot mijn studenten, voorjaar 1976. (Dutch) [Ninth lecture to my students, Spring 1976]. Circulated privately., 1976. URL http://www.cs.

utexas.edu/users/EWD/ewd05xx/EWD559.PDF.

Dijkstra:1976:NTR

[Dij76-37] Edsger W. Dijkstra. A “non trip report” from E. W. Dijkstra. Published as [Dij82-87]., April 1976. URL http://www.cs.

utexas.edu/users/EWD/ewd05xx/EWD561.PDF.

Dijkstra:1976:GTDa

[Dij76-38] Edsger W. Dijkstra. On a gauntlet thrown by David Gries. Published as [Dij76-39]., 1976. URL http://www.

cs.utexas.edu/users/EWD/ewd05xx/EWD553.PDF.

Dijkstra:1976:GTDb

[Dij76-39] Edsger W. Dijkstra. On a gauntlet thrown by David Gries. Acta Informatica, 6(4):357–359, ???? 1976. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Dijkstra:1976:PAH

[Dij76-40] Edsger W. Dijkstra. “On a problem from Aho, Hopcroft, and Ullman”. Circulated privately., 1976. URL http://www.cs.

utexas.edu/users/EWD/ewd05xx/EWD596.PDF.

Dijkstra:1976:MI

[Dij76-41] Edsger W. Dijkstra. On manuals and implementations, chapter 26. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1976.

Dijkstra:1976:NDB

[Dij76-42] Edsger W. Dijkstra. On non-determinacy being bounded, chapter 9. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1976.

Dijkstra:1976:SI

[Dij76-43] Edsger W. Dijkstra. On subgoal induction. Published as

[Dij82-90]., 1976. URL http://www.cs.utexas.edu/users/
EWD/ewd05xx/EWD576.PDF.

Dijkstra:1976:DPT

[Dij76-44] Edsger W. Dijkstra. On the design of properly terminating constructs, chapter 6. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1976.

Dijkstra:1976:TPT

[Dij76-45] Edsger W. Dijkstra. On the teaching of programming, i.e. on the teaching of thinking. In Bauer and Samelson [BS76], pages 1–10. ISBN 0-387-07994-7 (New York), 3-540-07994-7 (Berlin), 3-540-37972-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .L335. URL http:

//link.springer.com/chapter/10.1007/3-540-07994-7_
46.

Dijkstra:1976:Pb

[Dij76-46] Edsger W. Dijkstra. A parable. Published as [Dij82-96]., 1976.

URL http://www.cs.utexas.edu/users/EWD/ewd05xx/EWD594.
PDF.

Dijkstra:1976:PMP

[Dij76-47] Edsger W. Dijkstra. The pattern matching problem, chapter 18. Prentice-Hall, Upper Saddle River, NJ 07458, USA,

Dijkstra:1976:PSG

[Dij76-48] Edsger W. Dijkstra. A personal summary of the Gries–Owicki theory. Published as [Dij82-97]., March 1976. URL http://

www.cs.utexas.edu/users/EWD/ewd05xx/EWD554.PDF.

Dijkstra:1976:Pa

[Dij76-49] Edsger W. Dijkstra. Preface. In A discipline of programming [Dij76m], page ?? ISBN 0-13-215871-X. LCCN QA76.6 .D54. With a foreword by C. A. R. Hoare.

Dijkstra:1976:PCH

[Dij76-50] Edsger W. Dijkstra. The problem of the convex hull in three dimensions, chapter 24. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1976.

Dijkstra:1976:PDN

[Dij76-51] Edsger W. Dijkstra. The problem of The Dutch National Flag, chapter 14. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1976.

Dijkstra:1976:PML

[Dij76-52] Edsger W. Dijkstra. The problem of the maximum length of an ascending subsequence. Circulated privately., 1976. URL

http://www.cs.utexas.edu/users/EWD/ewd05xx/EWD591.
PDF.

Dijkstra:1976:PMI

[Dij76-53] Edsger W. Dijkstra. The problem of the most isolated villages, chapter 21. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1976.

Dijkstra:1976:PNP

[Dij76-54] Edsger W. Dijkstra. The problem of The Next Permutation, chapter 13. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1976.

Dijkstra:1976:PSS

[Dij76-55] Edsger W. Dijkstra. The problem of the shortest subspanning tree, chapter 22. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1976.

Dijkstra:1976:PSP

[Dij76-56] Edsger W. Dijkstra. The problem of the smallest prime factor of a large number, chapter 20. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1976.

Dijkstra:1976:PEM

[Dij76-57] Edsger W. Dijkstra. A programmer’s early memories. Published as [Dij80-30]., May 1976. URL http://www.cs.

utexas.edu/users/EWD/ewd05xx/EWD568.PDF.

Dijkstra:1976:PCS

[Dij76-58] Edsger W. Dijkstra. Programming: from craft to scientific discipline. Circulated privately., 1976. URL http://www.cs.

utexas.edu/users/EWD/ewd05xx/EWD566.PDF.

Dijkstra:1976:RAR

[Dij76-59] Edsger W. Dijkstra. Rem’s algorithm for the recording of equivalence classes, chapter 23. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1976.

Dijkstra:1976:RPL

[Dij76-60] Edsger W. Dijkstra. The role of programming languages, chapter 1. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1976.

Dijkstra:1976:SCP

[Dij76-61] Edsger W. Dijkstra. The semantic characterization of a programming language, chapter 4. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1976.

Dijkstra:1976:SCF

[Dij76-62] Edsger W. Dijkstra. A simple consideration with far-reaching consequences (DRAFT). Circulated privately., 1976. URL

http://www.cs.utexas.edu/users/EWD/ewd05xx/EWD571.
PDF.

Dijkstra:1976:SNA

[Dij76-63] Edsger W. Dijkstra. A small note on the additive composition of variant functions. Circulated privately., 1976. URL http:

//www.cs.utexas.edu/users/EWD/ewd05xx/EWD592.PDF.

Dijkstra:1976:SOL

[Dij76-64] Edsger W. Dijkstra. A somewhat open letter to Professor John McCarthy. Circulated privately., August 1976. URL

http://www.cs.utexas.edu/users/EWD/ewd05xx/EWD581.
PDF.

Dijkstra:1976:STC

[Dij76-65] Edsger W. Dijkstra. States and their characterization, chapter 2. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1976.

Dijkstra:1976:SB

[Dij76-66] Edsger W. Dijkstra. A superficial book. Circulated privately.,

April 1976. URL http://www.cs.utexas.edu/users/EWD/
ewd05xx/EWD564.PDF.

Dijkstra:1976:TTTb

[Dij76-67] Edsger W. Dijkstra. Tiende toespraak tot mijn studenten, voorjaar 1976. (Dutch) [Tenth lecture to my students, Spring 1976]. Circulated privately., 1976. URL http://www.cs.

utexas.edu/users/EWD/ewd05xx/EWD560.PDF.

Dijkstra:1976:TWH

[Dij76-68] Edsger W. Dijkstra. A time-wise hierarchy imposed upon the use of a two-level store. In Bauer and Samelson [BS76], pages 345–357. ISBN 0-387-07994-7 (New York), 3-540-07994-7 (Berlin), 3-540-37972-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

QA76.6 .L335.  URL http://link.springer-ny.com/
link/service/series/0558/tocs/t0046.htm;   http:/
/www.springerlink.com/content/978-0-387-07994-3;
http://www.springerlink.com/openurl.asp?genre=book&
isbn=978-3-540-07994-1;   http://www.springerlink.
com/openurl.asp?genre=issue&issn=0302-9743&volume=
46.

Dijkstra:1982:TWHb

[Dij76-69] Edsger W. Dijkstra. A time-wise hierarchy imposed upon the use of a two-level store. In Bauer and Samelson [BS76], pages 345–357. ISBN 0-387-07994-7 (New York), 3-540-07994-7 (Berlin), 3-540-37972-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

QA76.6 .L335.  URL http://link.springer-ny.com/
link/service/series/0558/tocs/t0046.htm;   http:/
/www.springerlink.com/content/978-0-387-07994-3;
http://www.springerlink.com/openurl.asp?genre=book&
isbn=978-3-540-07994-1;   http://www.springerlink.
com/openurl.asp?genre=issue&issn=0302-9743&volume=
46.

Dijkstra:1976:HDM

[Dij76-70] Edsger W. Dijkstra. To H. D. Mills, Chairman Software Methodology Panel. Published as [Dij82-105].,

July 1976. URL http://www.cs.utexas.edu/users/EWD/
ewd05xx/EWD575.PDF.

Dijkstra:1976:TWDf

[Dij76-71] Edsger W. Dijkstra. Tripreport E. W. Dijkstra, Copenhagen 10–12 Nov. 1976. Circulated privately., 1976. URL http://

www.cs.utexas.edu/users/EWD/ewd05xx/EWD596a.PDF.

Dijkstra:1976:TWDc

[Dij76-72] Edsger W. Dijkstra. Tripreport E. W. Dijkstra, ECIconference 9–12 August 1976, Amsterdam. Published as

[Dij82-59]., 1976. URL http://www.cs.utexas.edu/users/
EWD/ewd05xx/EWD577.PDF.

Dijkstra:1976:TWDa

[Dij76-73] Edsger W. Dijkstra. Tripreport E. W. Dijkstra, England and USA, 3–26 January 1976. Circulated privately., Febru-

ary 1976. URL http://www.cs.utexas.edu/users/EWD/
ewd05xx/EWD547.PDF.

Dijkstra:1976:TWDg

[Dij76-74] Edsger W. Dijkstra. Tripreport E. W. Dijkstra, Munich 24–26 November 1976. Circulated privately., 1976. URL http://

www.cs.utexas.edu/users/EWD/ewd06xx/EWD601.PDF.

Dijkstra:1976:TWDd

[Dij76-75] Edsger W. Dijkstra. Tripreport E. W. Dijkstra, Poland and USSR, 4–25 September 1976. Published as [Dij82-62]., Oc-

tober 1976. URL http://www.cs.utexas.edu/users/EWD/
ewd05xx/EWD584.PDF.

Dijkstra:1976:TWDh

[Dij76-76] Edsger W. Dijkstra. Tripreport E. W. Dijkstra, St. Pierre-de-Chartreuse, 12–19 Dec. 1976. Published as

[Dij82-53]., 1976. URL http://www.cs.utexas.edu/users/
EWD/ewd06xx/EWD603.PDF.

Dijkstra:1976:TWDe

[Dij76-77] Edsger W. Dijkstra. Tripreport E. W. Dijkstra, Tokyo, 28 Sep.–3 Oct. 1976. Published as [Dij82-63]., Octo-

ber 1976. URL http://www.cs.utexas.edu/users/EWD/
ewd05xx/EWD585.PDF.

Dijkstra:1976:TWDb

[Dij76-78] Edsger W. Dijkstra. Tripreport E. W. Dijkstra, U.S.A. and U.K., 8 June–10 July 1976. Circulated privately.,

July 1976. URL http://www.cs.utexas.edu/users/EWD/
ewd05xx/EWD572.PDF.

Dijkstra:1976:TTTc

[Dij76-79] Edsger W. Dijkstra. Twaalfde toespraak tot mijn studenten, voorjaar 1976. (Dutch) [Twelfth lecture to my students, Spring 1976]. Circulated privately., 1976. URL http://www.

cs.utexas.edu/users/EWD/ewd05xx/EWD567.PDF.

Dijkstra:1976:TTTd

[Dij76-80] Edsger W. Dijkstra. Tweede toespraak tot mijn studenten, najaar 1976. (Dutch) [Second lecture to my students, Autumn 1976]. Circulated privately., 1976. URL http://www.cs.

utexas.edu/users/EWD/ewd05xx/EWD586.PDF.

Dijkstra:1976:TTTa

[Dij76-81] Edsger W. Dijkstra. Tweede toespraak tot mijn studenten, voorjaar 1976. (Dutch) [Second lecture to my students, Spring 1976]. Circulated privately., 1976. URL http://www.cs.

utexas.edu/users/EWD/ewd05xx/EWD548.PDF.

Dijkstra:1976:TT

[Dij76-82] Edsger W. Dijkstra. Two theorems, chapter 5. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1976.

Dijkstra:1976:USF

[Dij76-83] Edsger W. Dijkstra. Updating a sequential file, chapter 15. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1976.

Dijkstra:1976:VTTc

[Dij76-84] Edsger W. Dijkstra. Vierde toespraak tot mijn studenten, najaar 1976. (Dutch) [Fourth lecture to my students, Autumn 1976]. Circulated privately., 1976. URL http://www.cs.

utexas.edu/users/EWD/ewd05xx/EWD588.PDF.

Dijkstra:1976:VTTa

[Dij76-85] Edsger W. Dijkstra. Vierde toespraak tot mijn studenten, voorjaar 1976. (Dutch) [Fourth lecture to my students, Spring 1976]. Circulated privately., 1976. URL http://www.cs.

utexas.edu/users/EWD/ewd05xx/EWD551.PDF.

Dijkstra:1976:VTTd

[Dij76-86] Edsger W. Dijkstra. Vijfde toespraak tot mijn studenten, najaar 1976. (Dutch) [Fifth lecture to my students, Autumn 1976]. Circulated privately., 1976. URL http://www.cs.

utexas.edu/users/EWD/ewd05xx/EWD589.PDF.

Dijkstra:1976:VTTb

[Dij76-87] Edsger W. Dijkstra. Vijfde toespraak tot mijn studenten, voorjaar 1976. (Dutch) [Fifth lecture to my students, Spring 1976]. Circulated privately., 1976. URL http://www.cs.

utexas.edu/users/EWD/ewd05xx/EWD552.PDF.

Dijkstra:1976:WOW

[Dij76-88] Edsger W. Dijkstra. Waarom de onderafdeling der wiskunde zich met de informatica moet bezighouden. (Dutch) [Why the subdivision of mathematics must deal with computer science]. Circulated privately., August 1976. URL http://www.cs.

utexas.edu/users/EWD/ewd05xx/EWD580.PDF.

Dijkstra:1976:WNS

[Dij76-89] Edsger W. Dijkstra. Writing a number as the sum of two squares, chapter 19. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1976.

Dijkstra:1976:YAN

[Dij76-90] Edsger W. Dijkstra. Yet another note about termination. Circulated privately., 1976. URL http://www.cs.utexas.

edu/users/EWD/ewd05xx/EWD597.PDF.

Dijkstra:1976:ZTTc

[Dij76-91] Edsger W. Dijkstra. Zesde toespraak tot mijn studenten, najaar 1976. (Dutch) [Sixth lecture to my students, Autumn 1976]. Circulated privately., 1976. URL http://www.cs.

utexas.edu/users/EWD/ewd05xx/EWD593.PDF.

Dijkstra:1976:ZTTa

[Dij76-92] Edsger W. Dijkstra. Zesde toespraak tot mijn studenten, voorjaar 1976. (Dutch) [Sixth lecture to my students, Spring 1976]. Circulated privately., 1976. URL http://www.cs.

utexas.edu/users/EWD/ewd05xx/EWD556.PDF.

Dijkstra:1976:ZTTb

[Dij76-93] Edsger W. Dijkstra. Zevende toespraak tot mijn studenten, voorjaar 1976. (Dutch) [Seventh lecture to my students, Spring 1976]. Circulated privately., 1976. URL http://www.

cs.utexas.edu/users/EWD/ewd05xx/EWD557.PDF.

Dijkstra:1977:ATT

[Dij77a] Edsger W. Dijkstra. Achtste toespraak tot mijn studenten, voorjaar 1977. (Dutch) [Eighth lecture to my students, Spring 1977]. Circulated privately., 1977. URL http://www.cs.

utexas.edu/users/EWD/ewd06xx/EWD621.PDF.

Dijkstra:1977:BRW

[Dij77b] Edsger W. Dijkstra. [Book review:] Wasserman, Antony I. en Peter Freeman (Eds.). Software Engineering Education. Needs and Objectives. Proceedings of an Interface Workshop. New York/Heidelberg/Berlin. Springer-Verlag, 1976. Circulated

privately., 1977. URL http://www.cs.utexas.edu/users/
EWD/ewd06xx/EWD626.PDF.

Dijkstra:1977:COO

[Dij77c] Edsger W. Dijkstra. Commentaar op een ontwikkeling binnen de TH’s. (Dutch) [Comments on development within the TH [technical university]’s]. Circulated privately., 1977. URL

http://www.cs.utexas.edu/users/EWD/ewd06xx/EWD647.
PDF.

Dijkstra:1977:CPC

[Dij77d]  Edsger W. Dijkstra.  A correctness proof for com-

municating processes: a small exercise. Published as

[Dij82-71]., 1977. URL http://www.cs.utexas.edu/users/
EWD/ewd06xx/EWD607.PDF.

Dijkstra:1977:DTT

[Dij77e] Edsger W. Dijkstra. Derde toespraak tot mijn studenten, voorjaar 1977. (Dutch) [Third lecture to my students, Spring 1977]. Circulated privately., 1977. URL http://www.cs.

utexas.edu/users/EWD/ewd06xx/EWD612.PDF.

Dijkstra:1977:ETT

[Dij77f] Edsger W. Dijkstra. Eerste toespraak tot mijn studenten, voorjaar 1977. (Dutch) [First lecture to my students, Spring 1977]. Circulated privately., 1977. URL http://www.cs.

utexas.edu/users/EWD/ewd06xx/EWD606.PDF.

Dijkstra:1977:EID

[Dij77g] Edsger W. Dijkstra. An elephant inspired by the Dutch National Flag. Published as [Dij82-73]., 1977. URL http://

www.cs.utexas.edu/users/EWD/ewd06xx/EWD608.PDF.

Dijkstra:1977:ENR

[Dij77h] Edsger W. Dijkstra. Essays on the nature and role of mathematical elegance (1). Circulated privately., 1977. URL http:

//www.cs.utexas.edu/users/EWD/ewd06xx/EWD619.PDF.

Dijkstra:1977:FCPa

[Dij77i] Edsger W. Dijkstra. Finding the correctness proof of a concurrent program. Circulated privately., 1977. URL http://

www.cs.utexas.edu/users/EWD/ewd06xx/EWD640.PDF.

Dijkstra:1977:FCPb

[Dij77j] Edsger W. Dijkstra. Finding the correctness proof of a concurrent program. Published as [Dij77k]., 1977. URL http:/

/www.cs.utexas.edu/users/EWD/ewd06xx/EWD640a.PDF.

Dijkstra:1977:FCPc

[Dij77k] Edsger W. Dijkstra. Finding the correctness proof of a concurrent program. Indagationes Mathematicae (Proceedings), 81(2):207–215, 1977. ISSN 1385-7258 (print), 1878-5972 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/1385725878900380.

Dijkstra:1977:IMR

[Dij77l] Edsger W. Dijkstra. The introduction of MAES(R). Published as [Dij82x]., 1977. URL http://www.cs.utexas.edu/

users/EWD/ewd06xx/EWD639.PDF.

Dijkstra:1977:LJA

[Dij77m] Edsger W. Dijkstra. Letter to J. Arsac. Circulated pri-

vately., October 1977. URL http://www.cs.utexas.edu/
users/EWD/ewd06xx/EWD644.PDF.

Dijkstra:1977:MBB

[Dij77n] Edsger W. Dijkstra. The mathematics behind the Banker’s Algorithm. Published as [Dij82-81]., 1977. URL http://www.

cs.utexas.edu/users/EWD/ewd06xx/EWD623.PDF.

Dijkstra:1977:NAV

[Dij77o] Edsger W. Dijkstra. Naar aanleiding van een eigenlijk wat beledigend stukje. (Dutch) [Following a basically offensive piece]. Circulated privately., 1977. URL http://www.cs.

utexas.edu/users/EWD/ewd06xx/EWD628.PDF.

Dijkstra:1977:STA

[Dij77p] Edsger W. Dijkstra. A not so simple theorem about undirected graphs. Circulated privately., 1977. URL http://

www.cs.utexas.edu/users/EWD/ewd06xx/EWD646.PDF.

Dijkstra:1977:MSM

[Dij77q] Edsger W. Dijkstra. On making solutions more and more fine-grained (in gratitude dedicated to C. A. R. Hoare, D. E. Knuth, and J. F. Traub.). Published as [Dij82-89].,

May 1977. URL http://www.cs.utexas.edu/users/EWD/
ewd06xx/EWD622.PDF.

Dijkstra:1977:FAO

[Dij77r] Edsger W. Dijkstra. On the fact that the Atlantic Ocean has two sides. Published as [Dij82-91]., 1977. URL http://www.

cs.utexas.edu/users/EWD/ewd06xx/EWD611.PDF.

Dijkstra:1977:IBM

[Dij77s] Edsger W. Dijkstra. On the interplay between mathematics and programming. Circulated privately., 1977. URL http:/

/www.cs.utexas.edu/users/EWD/ewd06xx/EWD641.PDF.

Dijkstra:1977:TBS

[Dij77t] Edsger W. Dijkstra. On two beautiful solutions designed by Martin Rem. Published as [Dij82-93]., 1977. URL http://

www.cs.utexas.edu/users/EWD/ewd06xx/EWD629.PDF.

Dijkstra:1977:WUB

[Dij77u] Edsger W. Dijkstra. On Webster, users, bugs and Aristotle. Published as [Dij82-95]., 1977. URL http://www.cs.

utexas.edu/users/EWD/ewd06xx/EWD618.PDF.

Dijkstra:1977:PLC

[Dij77v] Edsger W. Dijkstra. Paying logical conscience-money to the fair demon. Circulated privately., January 1977. URL http:

//www.cs.utexas.edu/users/EWD/ewd06xx/EWD604.PDF.

Dijkstra:1977:PPM

[Dij77w] Edsger W. Dijkstra. A political pamphlet from the Middle Ages. Negative commentary on [DLP79]. Circulated

privately., 1977. URL http://www.cs.utexas.edu/users/
EWD/ewd06xx/EWD638.PDF.

Dijkstra:1977:PPSa

[Dij77x] Edsger W. Dijkstra. A position paper on Software Reliability. Circulated privately., 1977. URL http://www.cs.utexas.

edu/users/EWD/ewd06xx/EWD627.PDF.

Dijkstra:1977:PPSb

[Dij77y] Edsger W. Dijkstra. A position paper on software reliability. ACM SIGSOFT Software Engineering Notes, 2(5):3–5, October 1977. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dijkstra:1977:PPI

[Dij77z] Edsger W. Dijkstra. Position paper on the impact of microprocessors (IFIP77). Circulated privately., 1977. URL http:

//www.cs.utexas.edu/users/EWD/ewd06xx/EWD634.PDF.

Dijkstra:1977:PCS

[Dij77-27] Edsger W. Dijkstra. Programming: from craft to scientific discipline. In International Computing Symposium 1977 (Liège, 1977), pages 23–30. North-Holland, Amsterdam,

Dijkstra:1977:RFD

[Dij77-28] Edsger W. Dijkstra. Re: “Formal derivation of strongly correct parallel programs” by Axel van Lamsweerde and M. Sintzoff. Circulated privately., 1977. URL http://www.cs.

utexas.edu/users/EWD/ewd06xx/EWD631.PDF.

Dijkstra:1977:SE

[Dij77-29] Edsger W. Dijkstra. A sequel to EWD592. Circulated pri-

vately., January 1977. URL http://www.cs.utexas.edu/
users/EWD/ewd06xx/EWD600.PDF.

Dijkstra:1977:SOL

[Dij77-30] Edsger W. Dijkstra. A somewhat open letter to EAA or: why I proved the boundedness of the non-determinacy in the way I did. Published as [Dij82-99]., 1977. URL http://www.cs.

utexas.edu/users/EWD/ewd06xx/EWD614.PDF.

Dijkstra:1977:SSE

[Dij77-31] Edsger W. Dijkstra. A special sorting elephant. Circulated

privately., 1977. URL http://www.cs.utexas.edu/users/
EWD/ewd06xx/EWD642.PDF.

Dijkstra:1977:SBS

[Dij77-32] Edsger W. Dijkstra. Stationary behaviour of some ternary networks. Circulated privately., 1977. URL http://www.cs.

utexas.edu/users/EWD/ewd06xx/EWD624.PDF.

Dijkstra:1977:SPV

[Dij77-33] Edsger W. Dijkstra. A strong P/V-implementation of conditional critical regions. Circulated privately., 1977. URL http:

//www.cs.utexas.edu/users/EWD/ewd06xx/EWD651.PDF.

Dijkstra:1977:SEP

[Dij77-34] Edsger W. Dijkstra. A supplement to EWD591 “The problem of the maximum length of an ascending subsequence”. Circulated privately., 1977. URL http://www.cs.utexas.

edu/users/EWD/ewd06xx/EWD649.PDF.

Dijkstra:1977:TAO

[Dij77-35] Edsger W. Dijkstra. A theorem about odd powers of odd integers. Published as [Dij82-102]., 1977. URL http://www.

cs.utexas.edu/users/EWD/ewd06xx/EWD650.PDF.

Dijkstra:1977:TGR

[Dij77-36] Edsger W. Dijkstra. The Three Golden Rules for Successful Scientific Research. Published as [Dij82-103]., 1977. URL

http://www.cs.utexas.edu/users/EWD/ewd06xx/EWD637.
PDF.

Dijkstra:1977:TWDc

[Dij77-37] Edsger W. Dijkstra. Tripreport E. W. Dijkstra, 29 July–21 August 1977. Circulated privately., 1977. URL http://www.

cs.utexas.edu/users/EWD/ewd06xx/EWD633.PDF.

Dijkstra:1977:TWDb

[Dij77-38] Edsger W. Dijkstra. Tripreport E. W. Dijkstra, Australia, 16 February 1977–21 March 1977. Published as [Dij82-58].,

March 1977. URL http://www.cs.utexas.edu/users/EWD/
ewd06xx/EWD613.PDF.

Dijkstra:1977:TWDa

[Dij77-39] Edsger W. Dijkstra. Tripreport E. W. Dijkstra, Columbus — Pittsburgh — Paoli, 12–22 January 1977. Circulated

privately., 1977. URL http://www.cs.utexas.edu/users/
EWD/ewd06xx/EWD605.PDF.

Dijkstra:1977:TWDd

[Dij77-40] Edsger W. Dijkstra. Tripreport E. W. Dijkstra, Newcastleupon-Tyne, 5–10 Sept. 1977. Published as [Dij82-61]., 1977.

URL http://www.cs.utexas.edu/users/EWD/ewd06xx/EWD635.
PDF.

Dijkstra:1977:TWDe

[Dij77-41] Edsger W. Dijkstra. Tripreport E. W. Dijkstra, USA 19–30 October 1977. Circulated privately., November 1977. URL

http://www.cs.utexas.edu/users/EWD/ewd06xx/EWD645.
PDF.

Dijkstra:1977:TTT

[Dij77-42] Edsger W. Dijkstra. Tweede toespraak tot mijn studenten, voorjaar 1977. (Dutch) [Second lecture to my students, Spring 1977]. Circulated privately., 1977. URL http://www.cs.

utexas.edu/users/EWD/ewd06xx/EWD609.PDF.

Dijkstra:1977:TSF

[Dij77-43] Edsger W. Dijkstra. Two starvation-free solutions of a general exclusion problem. Circulated privately., 1977. URL http:/

/www.cs.utexas.edu/users/EWD/ewd06xx/EWD625.PDF.

Dijkstra:1977:VTTa

[Dij77-44] Edsger W. Dijkstra. Vierde toespraak tot mijn studenten, voorjaar 1977. (Dutch) [Fourth lecture to my students, Spring 1977]. Circulated privately., 1977. URL http://www.cs.

utexas.edu/users/EWD/ewd06xx/EWD615.PDF.

Dijkstra:1977:VTTb

[Dij77-45] Edsger W. Dijkstra. Vijfde toespraak tot mijn studenten, voorjaar 1977. (Dutch) [Fifth lecture to my students, Spring 1977]. Circulated privately., 1977. URL http://www.cs.

utexas.edu/users/EWD/ewd06xx/EWD616.PDF.

Dijkstra:1977:WSE

[Dij77-46] Edsger W. Dijkstra. “Why is software so expensive?” An explanation to the hardware designer. Published

as [Dij82-115]., 1977. URL http://www.cs.utexas.edu/
users/EWD/ewd06xx/EWD648.PDF.

Dijkstra:1977:WNP

[Dij77-47] Edsger W. Dijkstra. Why naive program transformation systems are unlikely to work. Published as [Dij82-116]., 1977.

URL http://www.cs.utexas.edu/users/EWD/ewd06xx/EWD636.
PDF.

Dijkstra:1977:ZTTa

[Dij77-48] Edsger W. Dijkstra. Zesde toespraak tot mijn studenten, voorjaar 1977. (Dutch) [Sixth lecture to my students, Spring 1977]. Circulated privately., 1977. URL http://www.cs.

utexas.edu/users/EWD/ewd06xx/EWD617.PDF.

Dijkstra:1977:ZTTb

[Dij77-49] Edsger W. Dijkstra. Zevende toespraak tot mijn studenten, voorjaar 1977. (Dutch) [Seventh lecture to my students, Spring 1977]. Circulated privately., 1977. URL http://www.

cs.utexas.edu/users/EWD/ewd06xx/EWD620.PDF.

Dijkstra:1978:FCPc

[Dij78a] E. W. Dijkstra. Finding the correctness proof of a concurrent program. Nederl. Akad. Wetensch. Indag. Math., 40(2):207– 215, 1978. CODEN IMTHBJ. ISSN 0019-3577, 0023-3358.

Dijkstra:1978:FCPd

[Dij78b] E. W. Dijkstra. Finding the correctness proof of a concurrent program. Indagationes Mathematicae (Proceedings), 81 (1):207–215, 1978. ISSN 1385-7258 (print), 1878-5972 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S1385725878800110.

Dijkstra:1978:BLSb

[Dij78c] E. W. Dijkstra. On the BLUE language submitted to the DoD. ACM SIGPLAN Notices, 13(10):10–15, October 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dijkstra:1978:GLSb

[Dij78d] E. W. Dijkstra. On the GREEN language submitted to the DoD. ACM SIGPLAN Notices, 13(10):16–21, October 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dijkstra:1978:RLSb

[Dij78e] E. W. Dijkstra. On the RED language submitted to the DoD. ACM SIGPLAN Notices, 13(10):27–32, October 1978. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dijkstra:1978:YLSb

[Dij78f] E. W. Dijkstra. On the YELLOW language submitted to the DoD. ACM SIGPLAN Notices, 13(10):22–26, October 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dijkstra:1978:APD

[Dij78g] Edsger W. Dijkstra. About polygons in Detroit. Circulated privately., December 1978. URL http://www.cs.utexas.

edu/users/EWD/ewd06xx/EWD693.PDF.

Dijkstra:1978:BRa

[Dij78h] Edsger W. Dijkstra. A book review. Circulated pri-

vately., 1978. URL http://www.cs.utexas.edu/users/
EWD/ewd06xx/EWD669.PDF.

Dijkstra:1978:BRb

[Dij78i] Edsger W. Dijkstra. A book review. Circulated privately.,

September 1978. URL http://www.cs.utexas.edu/users/
EWD/ewd06xx/EWD680.PDF.

Dijkstra:1978:CE

[Dij78j] Edsger W. Dijkstra. A correction on EWD651. Circulated

privately., 1978. URL http://www.cs.utexas.edu/users/
EWD/ewd06xx/EWD652.PDF.

Dijkstra:1978:DS

[Dij78k] Edsger W. Dijkstra. DoD-I: the summing up. ACM SIG-PLAN Notices, 13(7):21–26, July 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dijkstra:1978:EBN

[Dij78l] Edsger W. Dijkstra. The equivalence of bounded nondeterminacy and continuity. Published as [Dij82n]., 1978. URL http:

//www.cs.utexas.edu/users/EWD/ewd06xx/EWD675.PDF.

Dijkstra:1978:ENR

[Dij78m] Edsger W. Dijkstra. Essays on the nature and role of mathematical elegance (3): On notation (a sequel to EWD619). Circulated privately., 1978. URL http://www.cs.utexas.

edu/users/EWD/ewd06xx/EWD655.PDF.

Dijkstra:1978:ECL

[Dij78n] Edsger W. Dijkstra. Exploiting contiguity in a linear store. Circulated privately., 1978. URL http://www.cs.utexas.

edu/users/EWD/ewd06xx/EWD653.PDF.

Dijkstra:1978:FCPb

[Dij78o] Edsger W. Dijkstra. Finding the correctness proof of a concurrent program. In Winkowski [Win78], pages 31– 38. CODEN LNCSD9. ISBN 0-387-08917-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6

.S9194 1978. URL http://link.springer-ny.com/link/
service/series/0558/tocs/t0064.htm;   http://www.
springerlink.com/content/978-0-387-08917-1;   http:
//www.springerlink.com/openurl.asp?genre=issue&issn=
0302-9743&volume=64.

Dijkstra:1978:HFa

[Dij78p] Edsger W. Dijkstra. In honour of Fibonacci. Circulated

privately., 1978. URL http://www.cs.utexas.edu/users/
EWD/ewd06xx/EWD654.PDF.

Dijkstra:1978:III

[Dij78q] Edsger W. Dijkstra. An introduction to implementation issues. Circulated privately., 1978. URL http://www.cs.

utexas.edu/users/EWD/ewd06xx/EWD656.PDF.

Dijkstra:1978:MEC

[Dij78r] Edsger W. Dijkstra. Mathematics in an easy chair. Circulated privately., December 1978. URL http://www.cs.utexas.

edu/users/EWD/ewd06xx/EWD695.PDF.

Dijkstra:1978:MJM

[Dij78s] Edsger W. Dijkstra. More on A. J. Martin’s design (a sequel to EWD668). Circulated privately., 1978. URL http://www.

cs.utexas.edu/users/EWD/ewd06xx/EWD672.PDF.

Dijkstra:1978:NCS

[Dij78t] Edsger W. Dijkstra. The nature of Computer Science (first draft). Circulated privately., 1978. URL http://www.cs.

utexas.edu/users/EWD/ewd06xx/EWD682.PDF.

Dijkstra:1978:PPM

[Dij78u] Edsger W. Dijkstra. On a political pamphlet from the Middle Ages (commentary on a paper by DeMillo, Lipton, and Prelis). ACM SIGSOFT Software Engineering Notes, 3(2): 14–16, April 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). Negative commentary on [DLP79]. See responses [DLP78, Jef78, Van78, Mau78].

Dijkstra:1978:ISA

[Dij78v] Edsger W. Dijkstra. On improving the state of the art (a somewhat open letter to dr. Martin Rem). Circulated pri-

vately., October 1978. URL http://www.cs.utexas.edu/
users/EWD/ewd06xx/EWD691.PDF.

Dijkstra:1978:LCE

[Dij78w] Edsger W. Dijkstra. On language constraints enforceable by translators (an open letter to Lt. Col. William A. Whitaker). Circulated privately., March 1978. URL http://www.cs.

utexas.edu/users/EWD/ewd06xx/EWD658.PDF.

Dijkstra:1978:LNS

[Dij78x] Edsger W. Dijkstra. On leaves and nodes: a simplification of EWD653. Circulated privately., 1978. URL http://www.cs.

utexas.edu/users/EWD/ewd06xx/EWD657.PDF.

Dijkstra:1978:OCT

[Dij78y] Edsger W. Dijkstra. On one of Cayley’s theorems. Circulated

privately., August 1978. URL http://www.cs.utexas.edu/
users/EWD/ewd06xx/EWD677.PDF.

Dijkstra:1978:BLSa

[Dij78z] Edsger W. Dijkstra. On the BLUE Language submitted to DoD. Circulated privately., 1978. URL http://www.cs.

utexas.edu/users/EWD/ewd06xx/EWD659.PDF.

Dijkstra:1978:CDA

[Dij78-27] Edsger W. Dijkstra. On the correctness of a design by Alain J. Martin. Circulated privately., 1978. URL http://www.cs.

utexas.edu/users/EWD/ewd06xx/EWD668.PDF.

Dijkstra:1978:FNLa

[Dij78-28] Edsger W. Dijkstra. On the foolishness of “natural language programming”. Circulated privately., 1978. URL http://

www.cs.utexas.edu/users/EWD/ewd06xx/EWD667.PDF.

Dijkstra:1978:GLSa

[Dij78-29] Edsger W. Dijkstra. On the GREEN Language submitted to the DoD. Circulated privately., 1978. URL http://www.cs.

utexas.edu/users/EWD/ewd06xx/EWD660.PDF.

Dijkstra:1978:RLSa

[Dij78-30] Edsger W. Dijkstra. On the RED Language submitted to the DoD. Circulated privately., 1978. URL http://www.cs.

utexas.edu/users/EWD/ewd06xx/EWD661.PDF.

Dijkstra:1978:YLSa

[Dij78-31] Edsger W. Dijkstra. On the YELLOW Language submitted to the DoD. Circulated privately., 1978. URL http://www.

cs.utexas.edu/users/EWD/ewd06xx/EWD662.PDF.

Dijkstra:1978:WST

[Dij78-32] Edsger W. Dijkstra. On weak and strong termination. Published as [Dij82-94]., 1978. URL http://www.cs.utexas.

edu/users/EWD/ewd06xx/EWD673.PDF.

Dijkstra:1978:PEV

[Dij78-33] Edsger W. Dijkstra. The pragmatic engineer versus the scientific designer. Circulated privately., November 1978. URL

http://www.cs.utexas.edu/users/EWD/ewd06xx/EWD690.
PDF.

Dijkstra:1978:PSF

[Dij78-34] Edsger W. Dijkstra. The problem of the Swiss football players. Circulated privately., 1978. URL http://www.cs.

utexas.edu/users/EWD/ewd06xx/EWD685.PDF.

Dijkstra:1978:PSM

[Dij78-35] Edsger W. Dijkstra. A problem solved in my head. Circulated

privately., 1978. URL http://www.cs.utexas.edu/users/
EWD/ewd06xx/EWD666.PDF.

Dijkstra:1978:PIa

[Dij78-36] Edsger W. Dijkstra. Program inversion. Published as

[Dij82-98]., 1978. URL http://www.cs.utexas.edu/users/
EWD/ewd06xx/EWD671.PDF.

Dijkstra:1978:RTA

[Dij78-37] Edsger W. Dijkstra. A review of the 1977 Turing Award Lecture by John Backus. Circulated privately., 1978. URL http:

//www.cs.utexas.edu/users/EWD/ewd06xx/EWD692.PDF.

Dijkstra:1978:SDU

[Dij78-38] Edsger W. Dijkstra. Sequencing and the discriminated union. Circulated privately., 1978. URL http://www.cs.utexas.

edu/users/EWD/ewd06xx/EWD670.PDF.

Dijkstra:1978:SBA

[Dij78-39] Edsger W. Dijkstra. Some beautiful arguments using mathematical induction. Published as [Dij80-37]., Decem-

ber 1978. URL http://www.cs.utexas.edu/users/EWD/
ewd06xx/EWD697.PDF.

Dijkstra:1978:SSV

[Dij78-40] Edsger W. Dijkstra. A story that starts with a very good computer. Published as [Dij82-100]., 1978. URL http://

www.cs.utexas.edu/users/EWD/ewd06xx/EWD678.PDF.

Dijkstra:1978:S

[Dij78-41] Edsger W. Dijkstra. The summing-up. Circulated privately.,

April 1978. URL http://www.cs.utexas.edu/users/EWD/
ewd06xx/EWD663.PDF.

Dijkstra:1978:NMT

[Dij78-42] Edsger W. Dijkstra. To a new member of The Tuesday Afternoon Club. Circulated privately., October 1978. URL http:

//www.cs.utexas.edu/users/EWD/ewd06xx/EWD683.PDF.

Dijkstra:1978:T

[Dij78-43] Edsger W. Dijkstra. A trifle. Circulated privately., Oc-

tober 1978. URL http://www.cs.utexas.edu/users/EWD/
ewd06xx/EWD688.PDF.

Dijkstra:1978:TRW

[Dij78-44] Edsger W. Dijkstra. Trip report E. W. Dijkstra, U.K. — Bahamas — U.S.A., 11–30 April 1978. Circulated privately.,

May 1978. URL http://www.cs.utexas.edu/users/EWD/
ewd06xx/EWD665.PDF.

Dijkstra:1978:TWDe

[Dij78-45] Edsger W. Dijkstra. Tripreport E. W. Dijkstra, Detroit, Austin, Philadelphia, 25 Nov.–9 Dec. 1978. Circulated pri-

vately., December 1978. URL http://www.cs.utexas.edu/
users/EWD/ewd06xx/EWD694.PDF.

Dijkstra:1978:TWDb

[Dij78-46] Edsger W. Dijkstra. Tripreport E. W. Dijkstra, Marktoberdorf 24 July–6 August 1978. Circulated privately., Au-

gust 1978. URL http://www.cs.utexas.edu/users/EWD/
ewd06xx/EWD676.PDF.

Dijkstra:1978:TWDc

[Dij78-47] Edsger W. Dijkstra. Tripreport E. W. Dijkstra, Newcastle— London—Lancaster, 4–13 Sep. 1978. Circulated privately.,

September 1978. URL http://www.cs.utexas.edu/users/
EWD/ewd06xx/EWD679.PDF.

Dijkstra:1978:TWDd

[Dij78-48] Edsger W. Dijkstra. Tripreport E. W. Dijkstra, Valley Forge, 28 Oct. –2 Nov. 1978. Circulated privately., Novem-

ber 1978. URL http://www.cs.utexas.edu/users/EWD/
ewd06xx/EWD689.PDF.

Dijkstra:1978:TWDa

[Dij78-49] Edsger W. Dijkstra. Tripreport E. W. Dijkstra, Zürich, 2–7 July 1978. Circulated privately., July 1978. URL http://

www.cs.utexas.edu/users/EWD/ewd06xx/EWD674.PDF.

Dijkstra:1978:WNT

[Dij78-50] Edsger W. Dijkstra. Why naive transformation systems are unlikely to work. Working paper, Burroughs, Nuenen, The Netherlands, 1978.

Dijkstra:1978:WA

[Dij78-51] Edsger W. Dijkstra. Written in anger. Circulated privately.,

December 1978. URL http://www.cs.utexas.edu/users/
EWD/ewd06xx/EWD696.PDF.

Dijkstra:1979:FCP

[Dij79a] E. W. Dijkstra. Finding the correctness proof of a concurrent program. In Bauer and Broy [BB79], pages 24–34. CO-DEN LNCSD9. ISBN 0-387-09251-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .P75117. URL http:/

/link.springer.com/chapter/10.1007/BFb0014652. “The

international summer school took place from July 26 to August 6, 1978, in Marktoberdorf … and was sponsored by the NATO Scientific Affairs Division under the 1978 Advanced Study Institutes programme.”.

Dijkstra:1979:IBM

[Dij79b] E. W. Dijkstra. On the interplay between mathematics and programming. In Bauer and Broy [BB79], pages 35–46. CODEN LNCSD9. ISBN 0-387-09251-X.

ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

QA76.6 .P75117. URL http://link.springer-ny.com/
link/service/series/0558/tocs/t0069.htm;   http:/
/www.springerlink.com/content/978-0-387-09251-5;
http://www.springerlink.com/openurl.asp?genre=issue&
issn=0302-9743&volume=69. “The international summer

school took place from July 26 to August 6, 1978, in Marktoberdorf … and was sponsored by the NATO Scientific Affairs Division under the 1978 Advanced Study Institutes programme.”.

Dijkstra:1979:IBMa

[Dij79c] E. W. Dijkstra. On the interplay between mathematics and programming. In Proceedings, Bicentennial Congress Wiskundig Genootschap (Vrije Univ., Amsterdam, 1978), Part I, volume 100 of Math. Centre Tracts, pages 177–188. Math. Centrum, Amsterdam, 1979.

Dijkstra:1979:TPS

[Dij79d] E. W. Dijkstra. The thinking programmer: Summary. In Bauer and Broy [BB79], page 1. CODEN LNCSD9. ISBN 0-387-09251-X. ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN QA76.6 .P75117. URL http://link.springer-
ny.com/link/service/series/0558/tocs/t0069.htm;
http://www.springerlink.com/content/978-0-387-09251-
5;   http://www.springerlink.com/openurl.asp?genre=
issue&issn=0302-9743&volume=69. “The international

summer school took place from July 26 to August 6, 1978, in Marktoberdorf … and was sponsored by the NATO Scientific Affairs Division under the 1978 Advanced Study Institutes programme.”.

Dijkstra:1979:ACE

[Dij79e] Edsger W. Dijkstra. Assembly conventions for the EDSAC. Circulated privately., October 1979. URL http://www.cs.

utexas.edu/users/EWD/ewd07xx/EWD718.PDF.

Dijkstra:1979:BRB

[Dij79f] Edsger W. Dijkstra. A book review for the IBM Systems Journal. Circulated privately., 1979. URL http://www.cs.

utexas.edu/users/EWD/ewd07xx/EWD722.PDF.

Dijkstra:1979:DMX

[Dij79g] Edsger W. Dijkstra. Dear Mr. X of Company Y. Circulated

privately., 1979. URL http://www.cs.utexas.edu/users/
EWD/ewd07xx/EWD707.PDF.

Dijkstra:1979:DSS

[Dij79h] Edsger W. Dijkstra. The design of a state space with a useful structure (I). Circulated privately., November 1979. URL

http://www.cs.utexas.edu/users/EWD/ewd07xx/EWD721.
PDF.

Dijkstra:1979:RDC

[Dij79i] Edsger W. Dijkstra. Een reactie. (Dutch) [A comment]. Circulated privately., January 1979. URL http://www.cs.

utexas.edu/users/EWD/ewd06xx/EWD699.PDF.

Dijkstra:1979:EED

[Dij79j] Edsger W. Dijkstra. An examination exercise, designed by W. H. J. Feijen. Circulated privately., February 1979. URL http:

//www.cs.utexas.edu/users/EWD/ewd07xx/EWD700.PDF.

Dijkstra:1979:EE

[Dij79k] Edsger W. Dijkstra. An exercise in exposition. Circulated privately., October 1979. URL http://www.cs.utexas.edu/

users/EWD/ewd07xx/EWD717.PDF.

Dijkstra:1979:IRT

[Dij79l] Edsger W. Dijkstra. Image reconstruction in two-dimensional tomography. Circulated privately., March 1979. URL http:

//www.cs.utexas.edu/users/EWD/ewd07xx/EWD706.PDF.

Dijkstra:1979:HF

[Dij79m] Edsger W. Dijkstra. In honour of Fibonacci. In Bauer and Broy [BB79], pages 49–50. CODEN LNCSD9. ISBN 0-387-09251-X. ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN QA76.6 .P75117. URL http://link.springer.com/
chapter/10.1007/BFb0014655. “The international summer

school took place from July 26 to August 6, 1978, in Marktoberdorf … and was sponsored by the NATO Scientific Affairs Division under the 1978 Advanced Study Institutes programme.”.

Dijkstra:1979:REC

[Dij79n] Edsger W. Dijkstra. In reaction to Ernest Chang’s “Deadlock Detection”. Circulated privately., February 1979. URL http:

//www.cs.utexas.edu/users/EWD/ewd07xx/EWD702.PDF.

Dijkstra:1979:MIC

[Dij79o] Edsger W. Dijkstra. A machine for image construction in tomography. Circulated privately., February 1979. URL http:

//www.cs.utexas.edu/users/EWD/ewd07xx/EWD704.PDF.

Dijkstra:1979:MFT

[Dij79p] Edsger W. Dijkstra. A more formal treatment of a less simple example. In Bauer and Broy [BB79], pages 1– 20. CODEN LNCSD9. ISBN 0-387-09251-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6

.P75117.  URL http://link.springer-ny.com/link/
service/series/0558/tocs/t0069.htm;   http://www.
springerlink.com/content/978-0-387-09251-5;   http:
//www.springerlink.com/openurl.asp?genre=issue&issn=

0302-9743&volume=69. “The international summer school took place from July 26 to August 6, 1978, in Marktoberdorf … and was sponsored by the NATO Scientific Affairs Division under the 1978 Advanced Study Institutes programme.”.

Dijkstra:1979:MHCa

[Dij79q] Edsger W. Dijkstra. My hopes of computing science. Published as [Dij79r]., April 1979. URL http://www.cs.utexas.

edu/users/EWD/ewd07xx/EWD709.PDF.

Dijkstra:1979:MHCb

[Dij79r] Edsger W. Dijkstra. My hopes of computing science (EWD709). In IEEE [IEE79b], pages 442–448.

Dijkstra:1979:PPW

[Dij79s] Edsger W. Dijkstra. On a problem posed by W. H. J. Feijen. Circulated privately., 1979. URL http://www.cs.utexas.

edu/users/EWD/ewd07xx/EWD713.PDF.

Dijkstra:1979:DVI

[Dij79t] Edsger W. Dijkstra. On not duplicating volatile information. Circulated privately., October 1979. URL http://www.cs.

utexas.edu/users/EWD/ewd07xx/EWD719.PDF.

Dijkstra:1979:FNL

[Dij79u] Edsger W. Dijkstra. On the foolishness of “Natural language programming”. In Bauer and Broy [BB79], pages 51–53. CO-DEN LNCSD9. ISBN 0-387-09251-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .P75117. URL http:/

/link.springer.com/chapter/10.1007/BFb0014656. “The

international summer school took place from July 26 to August 6, 1978, in Marktoberdorf … and was sponsored by the NATO Scientific Affairs Division under the 1978 Advanced Study Institutes programme.”.

Dijkstra:1979:IBMb

[Dij79v] Edsger W. Dijkstra. On the interplay between mathematics and programming. In Bauer and Broy [BB79], pages 35–46. CODEN LNCSD9. ISBN 0-387-09251-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

QA76.6 .P75117. URL http://link.springer-ny.com/
link/service/series/0558/tocs/t0069.htm;   http:/
/www.springerlink.com/content/978-0-387-09251-5;
http://www.springerlink.com/openurl.asp?genre=issue&
issn=0302-9743&volume=69. “The international summer

school took place from July 26 to August 6, 1978, in Marktoberdorf … and was sponsored by the NATO Scientific Affairs Division under the 1978 Advanced Study Institutes programme.”.

Dijkstra:1979:WHJ

[Dij79w] Edsger W. Dijkstra. On W. H. J. Feijen’s solution for the lexicographic minimum of a circular list. Circulated privately.,

December 1979. URL http://www.cs.utexas.edu/users/
EWD/ewd07xx/EWD723.PDF.

Dijkstra:1979:PI

[Dij79x] Edsger W. Dijkstra. Program inversion. In Bauer and Broy [BB79], pages 54–57. CODEN LNCSD9. ISBN 0-387-09251-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

QA76.6 .P75117. URL http://link.springer-ny.com/
link/service/series/0558/tocs/t0069.htm;   http:/
/www.springerlink.com/content/978-0-387-09251-5;
http://www.springerlink.com/openurl.asp?genre=issue&
issn=0302-9743&volume=69. “The international summer

school took place from July 26 to August 6, 1978, in Marktoberdorf … and was sponsored by the NATO Scientific Affairs Division under the 1978 Advanced Study Institutes programme.”.

Dijkstra:1979:STM

[Dij79y] Edsger W. Dijkstra. A short talk to my students about money. Circulated privately., October 1979. URL http://www.cs.

utexas.edu/users/EWD/ewd07xx/EWD716.PDF.

Dijkstra:1979:SEI

[Dij79z] Edsger W. Dijkstra. Software engineering: As it should be. In IEEE [IEE79a], pages 442–448.

Dijkstra:1979:SBS

[Dij79-27] Edsger W. Dijkstra. Stationary behaviour of some ternary networks. In Bauer and Broy [BB79], pages 21–23. CO-DEN LNCSD9. ISBN 0-387-09251-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .P75117. URL http:/

/link.springer.com/chapter/10.1007/BFb0014651. “The

international summer school took place from July 26 to August 6, 1978, in Marktoberdorf … and was sponsored by the NATO Scientific Affairs Division under the 1978 Advanced Study Institutes programme.”.

Dijkstra:1979:TAO

[Dij79-28] Edsger W. Dijkstra. A theorem about odd powers of odd integers. In Bauer and Broy [BB79], pages 47– 48. CODEN LNCSD9. ISBN 0-387-09251-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6

.P75117.  URL http://link.springer-ny.com/link/
service/series/0558/tocs/t0069.htm;   http://www.
springerlink.com/content/978-0-387-09251-5;   http:
//www.springerlink.com/openurl.asp?genre=issue&issn=

0302-9743&volume=69. “The international summer school took place from July 26 to August 6, 1978, in Marktoberdorf … and was sponsored by the NATO Scientific Affairs Division under the 1978 Advanced Study Institutes programme.”.

Dijkstra:1979:TRWc

[Dij79-29] Edsger W. Dijkstra. Trip report E. W. Dijkstra, Antwerp, 24–29 June 1979. Circulated privately., 1979. URL http://

www.cs.utexas.edu/users/EWD/ewd07xx/EWD712.PDF.

Dijkstra:1979:TRWb

[Dij79-30] Edsger W. Dijkstra. Trip report E. W. Dijkstra, Ithaca, Albany, Austin (Texas), 26/5–10/6 1979. Circulated privately.,

June 1979. URL http://www.cs.utexas.edu/users/EWD/
ewd07xx/EWD711.PDF.

Dijkstra:1979:TRWd

[Dij79-31] Edsger W. Dijkstra. Trip report E. W. Dijkstra, Mission Viejo, Santa Cruz, Austin, 29 July–8 September 1979. Circulated privately., September 1979. URL http://www.cs.

utexas.edu/users/EWD/ewd07xx/EWD714.PDF.

Dijkstra:1979:TRWe

[Dij79-32] Edsger W. Dijkstra. Trip report E. W. Dijkstra, Munich — London, 16–29 September 1979. Circulated privately., Oc-

tober 1979. URL http://www.cs.utexas.edu/users/EWD/
ewd07xx/EWD715.PDF.

Dijkstra:1979:TRWa

[Dij79-33] Edsger W. Dijkstra. Trip report E. W. Dijkstra, Oxford University, 10–15 Jan. 1979. Circulated privately., Jan-

uary 1979. URL http://www.cs.utexas.edu/users/EWD/
ewd06xx/EWD698.PDF.

Dijkstra:1979:TSBb

[Dij79-34] Edsger W. Dijkstra. A tutorial on the split binary semaphore. Circulated privately. Available via URL

http://www/cs/utexas.edu/users/EWD/index07xx.html,
March 1979. URL http://www.cs.utexas.edu/users/EWD/
ewd07xx/EWD703.PDF.

Dijkstra:1979:TTW

[Dij79-35] Edsger W. Dijkstra. Two theorems on (what I have called) continuously mixed sequences. Circulated privately., Febru-

ary 1979. URL http://www.cs.utexas.edu/users/EWD/
ewd07xx/EWD701.PDF.

Dijkstra:1979:VBHa

[Dij79-36] Edsger W. Dijkstra. Verkavelde berekeningen, hun mogelijkheden en moeilijkheden. (Dutch) [Allotted calculations, their possibilities and difficulties]. Published as [Dij79-37].,

March 1979. URL http://www.cs.utexas.edu/users/EWD/
ewd07xx/EWD705.PDF.

Dijkstra:1979:VBHb

[Dij79-37] Edsger W. Dijkstra. Verkavelde berekeningen, hun mogelijkheden en moeilijkheden. (Dutch) [Allotted calculations, their possibilities and difficulties]. Proc. Koninklijke Nederlandse Akademie van Wetenschappen 88, 3:21–23, 1979.

Dijkstra:1979:WMMa

[Dij79-38] Edsger W. Dijkstra. When messages may crawl. Circulated

privately., March 1979. URL http://www.cs.utexas.edu/
users/EWD/ewd07xx/EWD708.PDF.

Dijkstra:1979:WMMb

[Dij79-39] Edsger W. Dijkstra. When messages may crawl, II (A sequel to EWD708). Circulated privately., 1979. URL http://www.

cs.utexas.edu/users/EWD/ewd07xx/EWD710.PDF.

Dijkstra:1979:WCM

[Dij79-40] Edsger W. Dijkstra. Why correctness must be a mathematical concern. Circulated privately., November 1979. URL http:/

/www.cs.utexas.edu/users/EWD/ewd07xx/EWD720.PDF.

Dijkstra:1980:JMS

[Dij80a] Edsger W. Dijkstra. A. J. Martin’s solution of the Hungarian problem. Circulated privately., December 1980. URL http:

//www.cs.utexas.edu/users/EWD/ewd07xx/EWD767.PDF.

Dijkstra:1980:ACG

[Dij80b] Edsger W. Dijkstra. About 2-coloured 6-graphs. Circulated privately., December 1980. URL http://www.cs.utexas.

edu/users/EWD/ewd07xx/EWD771.PDF.

Dijkstra:1980:APP

[Dij80c] Edsger W. Dijkstra. American programming’s plight. Circulated privately., September 1980. URL http://www.cs.

utexas.edu/users/EWD/ewd07xx/EWD750.PDF.

Dijkstra:1980:C

[Dij80d] Edsger W. Dijkstra. A chutspa. Circulated privately.,

March 1980. URL http://www.cs.utexas.edu/users/EWD/
ewd07xx/EWD724E.PDF.

Dijkstra:1980:CMS

[Dij80e] Edsger W. Dijkstra. Comments on MIL-STD-1862, 28 May 1980. Circulated privately., September 1980. URL http://

www.cs.utexas.edu/users/EWD/ewd07xx/EWD747.PDF.

Dijkstra:1980:DTR

[Dij80f] Edsger W. Dijkstra. D. A. Turner’s reply. Circulated pri-

vately., December 1980. URL http://www.cs.utexas.edu/
users/EWD/ewd07xx/EWD770.PDF.

Dijkstra:1980:ES

[Dij80g] Edsger W. Dijkstra. An educational stupidity. Circulated privately., December 1980. URL http://www.cs.utexas.

edu/users/EWD/ewd07xx/EWD766.PDF.

Dijkstra:1980:GDT

[Dij80h] Edsger W. Dijkstra. Een gotspe. (Dutch) [A travesty]. Circulated privately., March 1980. URL http://www.cs.utexas.

edu/users/EWD/ewd07xx/EWD724.PDF.

Dijkstra:1980:KDS

[Dij80i] Edsger W. Dijkstra. Een kanttekening. (Dutch) [A side note]. Circulated privately., November 1980. URL http://www.cs.

utexas.edu/users/EWD/ewd07xx/EWD761.PDF.

Dijkstra:1980:EE

[Dij80j] Edsger W. Dijkstra. An error in EWD744. Circulated pri-

vately., October 1980. URL http://www.cs.utexas.edu/
users/EWD/ewd07xx/EWD754.PDF.

Dijkstra:1980:EME

[Dij80k] Edsger W. Dijkstra. An experiment in mathematical exposition. Circulated privately., February 1980. URL http://

www.cs.utexas.edu/users/EWD/ewd07xx/EWD731.PDF.

Dijkstra:1980:HDI

[Dij80l] Edsger W. Dijkstra. How Dutch Informatics fell between two chairs. Circulated privately., January 1980. URL http://

www.cs.utexas.edu/users/EWD/ewd07xx/EWD726.PDF.

Dijkstra:1980:HP

[Dij80m] Edsger W. Dijkstra. A Hungarian problem. Circulated pri-

vately., December 1980. URL http://www.cs.utexas.edu/
users/EWD/ewd07xx/EWD765.PDF.

Dijkstra:1980:IE

[Dij80n] Edsger W. Dijkstra. An intriguing example. Circulated pri-

vately., November 1980. URL http://www.cs.utexas.edu/
users/EWD/ewd07xx/EWD758.PDF.

Dijkstra:1980:MOP

[Dij80o] Edsger W. Dijkstra. Mainly on the omission of parentheses. Circulated privately., September 1980. URL http://www.cs.

utexas.edu/users/EWD/ewd07xx/EWD751.PDF.

Dijkstra:1980:MSE

[Dij80p] Edsger W. Dijkstra. A methodological sequel to EWD771. Circulated privately., December 1980. URL http://www.cs.

utexas.edu/users/EWD/ewd07xx/EWD772.PDF.

Dijkstra:1980:MVC

[Dij80q] Edsger W. Dijkstra. A mild variant of Combinatory Logic. Circulated privately., May 1980. URL http://www.cs.

utexas.edu/users/EWD/ewd07xx/EWD735.PDF.

Dijkstra:1980:MEE

[Dij80r] Edsger W. Dijkstra. A misguided educational effort. Circulated privately., October 1980. URL http://www.cs.

utexas.edu/users/EWD/ewd07xx/EWD757.PDF.

Dijkstra:1980:MMF

[Dij80s] Edsger W. Dijkstra. More mathematical folklore. Circulated privately., May 1980. URL http://www.cs.utexas.

edu/users/EWD/ewd07xx/EWD738.PDF.

Dijkstra:1980:NAV

[Dij80t] Edsger W. Dijkstra. Naar aanleiding van “Experimentele post-propadeuse opleiding voor informatica-ingenieur aan de THE” (Eindhoven februari 1980 SOO/jc/sdb). (Dutch) [Following “Experimental post-propadeuse training for computer science engineering at THE” (Eindhoven February 1980 SOO/jc/sdb)]. Circulated privately., March 1980. URL http:

//www.cs.utexas.edu/users/EWD/ewd07xx/EWD733.PDF.

Dijkstra:1980:NPM

[Dij80u] Edsger W. Dijkstra. A new policy for Mathematics Inc.? Circulated privately., 1980. URL http://www.cs.utexas.

edu/users/EWD/ewd07xx/EWD743.PDF.

Dijkstra:1980:NAQ

[Dij80v] Edsger W. Dijkstra. A notational alternative for quantification. Circulated privately., April 1980. URL http://www.

cs.utexas.edu/users/EWD/ewd07xx/EWD737.PDF.

Dijkstra:1980:TLM

[Dij80w] Edsger W. Dijkstra. On a theorem by Lambek and Moser. Circulated privately., October 1980. URL http://www.cs.

utexas.edu/users/EWD/ewd07xx/EWD753.PDF.

Dijkstra:1980:PRD

[Dij80x] Edsger W. Dijkstra. On the productivity of recursive definitions. Circulated privately., September 1980. URL http://

www.cs.utexas.edu/users/EWD/ewd07xx/EWD749.PDF.

Dijkstra:1980:TTI

[Dij80y] Edsger W. Dijkstra. On two types of infinite sets of infinite sequences (by the Tuesday Afternoon Club). Circulated pri-

vately., February 1980. URL http://www.cs.utexas.edu/
users/EWD/ewd07xx/EWD730.PDF.

Dijkstra:1980:WSM

[Dij80z] Edsger W. Dijkstra. On well-shaped mathematical arguments. Circulated privately., February 1980. URL http:/

/www.cs.utexas.edu/users/EWD/ewd07xx/EWD729.PDF.

Dijkstra:1980:PEC

[Dij80-27] Edsger W. Dijkstra. Partitioning the edges of the complete graphs into trees or cycles (by the Tuesday Afternoon Club). Circulated privately., May 1980. URL http://www.

cs.utexas.edu/users/EWD/ewd07xx/EWD741.PDF.

Dijkstra:1980:PE

[Dij80-28] Edsger W. Dijkstra. A postscript to EWD755. Circulated privately., October 1980. URL http://www.cs.utexas.edu/

users/EWD/ewd07xx/EWD756.PDF.

Dijkstra:1980:PSN

[Dij80-29] Edsger W. Dijkstra. A problem solved by Netty van Gasteren and me. Circulated privately., February 1980. URL http://

www.cs.utexas.edu/users/EWD/ewd07xx/EWD728.PDF.

Dijkstra:1980:PEM

[Dij80-30] Edsger W. Dijkstra. A programmer’s early memories. In Metropolis et al. [MHR80], pages 563–573. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5 .I63 1976.

URL http://www.sciencedirect.com/science/article/

pii/B9780124916500500411. Original versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Dijkstra:1980:PRM

[Dij80-31] Edsger W. Dijkstra. A proof by Rutger M. Dijkstra and me. Circulated privately., November 1980. URL http://www.cs.

utexas.edu/users/EWD/ewd07xx/EWD763.PDF.

Dijkstra:1980:RST

[Dij80-32] Edsger W. Dijkstra. Recording the structure of trees in their leaves. Circulated privately., April 1980. URL http://www.

cs.utexas.edu/users/EWD/ewd07xx/EWD736.PDF.

Dijkstra:1980:ROD

[Dij80-33] Edsger W. Dijkstra. Repaying our debts. Circulated pri-

vately., December 1980. URL http://www.cs.utexas.edu/
users/EWD/ewd07xx/EWD764.PDF.

Dijkstra:1980:SE

[Dij80-34] Edsger W. Dijkstra. A sequel to EWD740. Circulated

privately., May 1980. URL http://www.cs.utexas.edu/
users/EWD/ewd07xx/EWD742.PDF.

Dijkstra:1980:SNS

[Dij80-35] Edsger W. Dijkstra. A short note on symmetric distributed arbitration. Circulated privately., July 1980. URL http://

www.cs.utexas.edu/users/EWD/ewd07xx/EWD744.PDF.

Dijkstra:1980:SPO

[Dij80-36] Edsger W. Dijkstra. A short proof of one of Fermat’s theorems. Circulated privately., May 1980. URL http://www.

cs.utexas.edu/users/EWD/ewd07xx/EWD740.PDF.

Dijkstra:1980:SBAa

[Dij80-37] Edsger W. Dijkstra. Some beautiful arguments using mathematical induction. Acta Informatica, 13(1):1–8, January 1980. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Dijkstra:1980:SOLc

[Dij80-38] Edsger W. Dijkstra. A somewhat open letter to D. A. Turner. Circulated privately., November 1980. URL http://www.cs.

utexas.edu/users/EWD/ewd07xx/EWD759.PDF.

Dijkstra:1980:SOLd

[Dij80-39] Edsger W. Dijkstra. A somewhat open letter to F. Kroeger. Circulated privately., November 1980. URL http://www.cs.

utexas.edu/users/EWD/ewd07xx/EWD760.PDF.

Dijkstra:1980:SOLb

[Dij80-40] Edsger W. Dijkstra. A somewhat open letter to the Editor-in-Chief of Acta Informatica. Circulated privately., 1980. URL

http://www.cs.utexas.edu/users/EWD/ewd07xx/EWD739.
PDF.

Dijkstra:1980:SOLa

[Dij80-41] Edsger W. Dijkstra. A somewhat open letter to Władisław M. Turski. Circulated privately., September 1980. URL http:/

/www.cs.utexas.edu/users/EWD/ewd07xx/EWD748.PDF.

Dijkstra:1980:SGS

[Dij80-42] Edsger W. Dijkstra. The superfluity of the general semaphore. Circulated privately., April 1980. URL http://www.cs.

utexas.edu/users/EWD/ewd07xx/EWD734.PDF.

Dijkstra:1980:TMT

[Dij80-43] Edsger W. Dijkstra. The teachability of mathematical thinking (Draft contribution to ‘ ‘Mathematics Tomorrow”). Circulated privately., March 1980. URL http://www.cs.utexas.

edu/users/EWD/ewd07xx/EWD732.PDF.

Dijkstra:1980:TRWa

[Dij80-44] Edsger W. Dijkstra. Trip report E. W. Dijkstra, Liege, Nov.– Dec.1979. Circulated privately., January 1980. URL http:/

/www.cs.utexas.edu/users/EWD/ewd07xx/EWD725.PDF.

Dijkstra:1980:TRWc

[Dij80-45] Edsger W. Dijkstra. Trip report E. W. Dijkstra, Los Angeles, Austin (Texas), and Portland, 6–28 August 1980. Circulated

privately., August 1980. URL http://www.cs.utexas.edu/
users/EWD/ewd07xx/EWD745.PDF.

Dijkstra:1980:TRWd

[Dij80-46] Edsger W. Dijkstra. Trip report E. W. Dijkstra, Newcastleupon-Tyne, 1–5 Sep. 1980. Circulated privately., Septem-

ber 1980. URL http://www.cs.utexas.edu/users/EWD/
ewd07xx/EWD746.PDF.

Dijkstra:1980:TRWb

[Dij80-47] Edsger W. Dijkstra. Trip report E. W. Dijkstra, U.S.A. 12 Jan.–2 Feb. 1980. Circulated privately., February 1980. URL

http://www.cs.utexas.edu/users/EWD/ewd07xx/EWD727.
PDF.

Dijkstra:1980:VEN

[Dij80-48] Edsger W. Dijkstra. Very elementary number theory redone. Circulated privately., October 1980. URL http://www.cs.

utexas.edu/users/EWD/ewd07xx/EWD755.PDF.

Dijkstra:1980:WML

[Dij80-49] Edsger W. Dijkstra. We mathematicians are losing the race. Circulated privately., November 1980. URL http://www.cs.

utexas.edu/users/EWD/ewd07xx/EWD762.PDF.

Dijkstra:1981:ASD

[Dij81a] Edsger W. Dijkstra. The administration as social disease. Circulated privately., March 1981. URL http://www.cs.

utexas.edu/users/EWD/ewd07xx/EWD783.PDF.

Dijkstra:1981:AHS

[Dij81b] Edsger W. Dijkstra. An alternative to heapsort for sorting in situ. Circulated privately., June 1981. URL http://www.cs.

utexas.edu/users/EWD/ewd07xx/EWD795.PDF.

Dijkstra:1981:APP

[Dij81c] Edsger W. Dijkstra. American programming’s plight. ACM SIGSOFT Software Engineering Notes, 6(1):5, January 1981. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Dijkstra:1981:ATP

[Dij81d] Edsger W. Dijkstra. The analysis of a two-person game. Circulated privately., June 1981. URL http://www.cs.utexas.

edu/users/EWD/ewd07xx/EWD793.PDF.

Dijkstra:1981:APPa

[Dij81e] Edsger W. Dijkstra. An assertional proof of a program by G. L. Peterson. Circulated privately., February 1981. URL http:

//www.cs.utexas.edu/users/EWD/ewd07xx/EWD779.PDF.

Dijkstra:1981:CSR

[Dij81f] Edsger W. Dijkstra. Canonical string reduction. Circulated privately., December 1981. URL http://www.cs.utexas.

edu/users/EWD/ewd08xx/EWD808.PDF.

Dijkstra:1981:FNL

[Dij81g] Edsger W. Dijkstra. Fibonacci numbers and Leonardo numbers. Circulated privately., July 1981. URL http://www.cs.

utexas.edu/users/EWD/ewd07xx/EWD797.PDF.

Dijkstra:1981:HES

[Dij81h] Edsger W. Dijkstra. Hamming’s exercise in SASL. Circulated privately., June 1981. URL http://www.cs.utexas.

edu/users/EWD/ewd07xx/EWD792.PDF.

Dijkstra:1981:LMR

[Dij81i] Edsger W. Dijkstra. Lambek and Moser revisited. Circulated privately., February 1981. URL http://www.cs.utexas.

edu/users/EWD/ewd07xx/EWD776.PDF.

Dijkstra:1981:LN

[Dij81j] Edsger W. Dijkstra. Largely on nomenclature. Circulated

privately., March 1981. URL http://www.cs.utexas.edu/
users/EWD/ewd07xx/EWD768.PDF.

Dijkstra:1981:LNS

[Dij81k] Edsger W. Dijkstra. Lecture notes on the structure of programs and proofs. Circulated privately., Novem-

ber 1981. URL http://www.cs.utexas.edu/users/EWD/
ewd08xx/EWD803.PDF.

Dijkstra:1981:MIHa

[Dij81l] Edsger W. Dijkstra. A minor improvement of Heapsort. Circulated privately., May 1981. URL http://www.cs.utexas.

edu/users/EWD/ewd07xx/EWD789.PDF.

Dijkstra:1981:MIHb

[Dij81m] Edsger W. Dijkstra. A minor improvement of Heapsort. Circulated privately., May 1981. URL http://www.cs.utexas.

edu/users/EWD/ewd07xx/EWD789a.PDF.

Dijkstra:1981:MIHc

[Dij81n] Edsger W. Dijkstra. A minor improvement of Heapsort. Circulated privately., May 1981. URL http://www.cs.utexas.

edu/users/EWD/ewd07xx/EWD789b.PDF.

Dijkstra:1981:MMC

[Dij81o] Edsger W. Dijkstra. My mother’s contribution to Honsberger’s collection. Circulated privately., April 1981. URL

http://www.cs.utexas.edu/users/EWD/ewd07xx/EWD784.
PDF.

Dijkstra:1981:NPT

[Dij81p] Edsger W. Dijkstra. A (new?) proof of a theorem of Euler’s on partitions. Circulated privately., April 1981. URL http:

//www.cs.utexas.edu/users/EWD/ewd07xx/EWD787.PDF.

Dijkstra:1981:EP

[Dij81q] Edsger W. Dijkstra. On equality of propositions. Circulated privately., October 1981. URL http://www.cs.utexas.edu/

users/EWD/ewd08xx/EWD804.PDF.

Dijkstra:1981:KT

[Dij81r] Edsger W. Dijkstra. On Kleinrock’s Theorem. Circulated pri-

vately., February 1981. URL http://www.cs.utexas.edu/
users/EWD/ewd07xx/EWD775.PDF.

Dijkstra:1981:PU

[Dij81s] Edsger W. Dijkstra. The psychology of the user. Circulated privately., June 1981. URL http://www.cs.utexas.

edu/users/EWD/ewd07xx/EWD791.PDF.

Dijkstra:1981:SU

[Dij81t] Edsger W. Dijkstra. Sets are Unibags. Circulated privately.,

April 1981. URL http://www.cs.utexas.edu/users/EWD/
ewd07xx/EWD786a.PDF.

Dijkstra:1981:SASb

[Dij81u] Edsger W. Dijkstra. Smoothsort, an alternative for sorting in situ. Circulated privately., July 1981. URL http://www.cs.

utexas.edu/users/EWD/ewd07xx/EWD796.PDF.

Dijkstra:1981:SASc

[Dij81v] Edsger W. Dijkstra. Smoothsort, an alternative for sorting in situ. Circulated privately., August 1981. URL http://www.

cs.utexas.edu/users/EWD/ewd07xx/EWD796a.PDF.

Dijkstra:1981:SOL

[Dij81w] Edsger W. Dijkstra. A somewhat open letter to Nils J. Nilsson. Circulated privately., February 1981. URL http://www.

cs.utexas.edu/users/EWD/ewd07xx/EWD778.PDF.

Dijkstra:1981:SN

[Dij81x] Edsger W. Dijkstra. A stupid notation. Circulated privately.,

March 1981. URL http://www.cs.utexas.edu/users/EWD/
ewd07xx/EWD782.PDF.

Dijkstra:1981:TRWh

[Dij81y] Edsger W. Dijkstra. Trip report E. W. Dijkstra, Amsterdam, 26–29 Oct. 1981. Circulated privately., Octo-

ber 1981. URL http://www.cs.utexas.edu/users/EWD/
ewd08xx/EWD805.PDF.

Dijkstra:1981:TRWb

[Dij81z] Edsger W. Dijkstra. Trip report E. W. Dijkstra, California, 17–31 January 1981. Circulated privately., Febru-

ary 1981. URL http://www.cs.utexas.edu/users/EWD/
ewd07xx/EWD774.PDF.

Dijkstra:1981:TRWc

[Dij81-27] Edsger W. Dijkstra. Trip report E. W. Dijkstra, England, 23 April 1981. Circulated privately., April 1981. URL http://

www.cs.utexas.edu/users/EWD/ewd07xx/EWD786.PDF.

Dijkstra:1981:TRWa

[Dij81-28] Edsger W. Dijkstra. Trip report E. W. Dijkstra, Hansur-Lesse, 7–12 Jan. 1981. Circulated privately., Jan-

uary 1981. URL http://www.cs.utexas.edu/users/EWD/
ewd07xx/EWD773.PDF.

Dijkstra:1981:TRWf

[Dij81-29] Edsger W. Dijkstra. Trip report E. W. Dijkstra, Marktoberdorf, 26/7–10/8/81. Circulated privately., August 1981. URL

http://www.cs.utexas.edu/users/EWD/ewd07xx/EWD799.
PDF.

Dijkstra:1981:TRWd

[Dij81-30] Edsger W. Dijkstra. Trip report E. W. Dijkstra, München, 19 May 1981. Circulated privately., May 1981. URL http:/

/www.cs.utexas.edu/users/EWD/ewd07xx/EWD790.PDF.

Dijkstra:1981:TRWe

[Dij81-31] Edsger W. Dijkstra. Trip report E. W. Dijkstra, Newcastle, 19–25 July 1981. Circulated privately., July 1981. URL http:

//www.cs.utexas.edu/users/EWD/ewd07xx/EWD798.PDF.

Dijkstra:1981:TRWg

[Dij81-32] Edsger W. Dijkstra. Trip report E. W. Dijkstra, Scotland, Newcastle 31 Aug.–15 Sep. 1981. Circulated privately.,

September 1981. URL http://www.cs.utexas.edu/users/
EWD/ewd08xx/EWD802.PDF.

Dijkstra:1981:TRWi

[Dij81-33] Edsger W. Dijkstra. Trip report E. W. Dijkstra, USA, 1–27 Nov. 1981. Circulated privately., November 1981. URL http:

//www.cs.utexas.edu/users/EWD/ewd08xx/EWD806.PDF.

Dijkstra:1981:WW

[Dij81-34] Edsger W. Dijkstra. A word of welcome. Science of Computer Programming, 1(1–2):3–4, October 1981. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Dijkstra:1981:WWD

[Dij81-35] Edsger W. Dijkstra. A word of welcome (Draft). Circulated

privately., 1981. URL http://www.cs.utexas.edu/users/
EWD/ewd07xx/EWD777.PDF.

Dijkstra:1982:SCRa

[Dij82a] Dr. Edsger W. Dijkstra. The solution to a cyclic relaxation problem. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 34–35. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982.

URL http://link.springer.com/accesspage/chapter/
10.1007/978-1-4612-5695-3_5?coverImageUrl=/static%
2F0.9313%2Fsites%2Flink%2Fimages%2Fabstract_cover_

placeholder.png. Including a paper co-authored by C. S. Schölten.

Dijkstra:1982:TRWa

[Dij82b] Dr. Edsger W. Dijkstra. Trip report E. W. Dijkstra, Summer School Munich, July 25 to August 4, 1973. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 31–33. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982. URL http://

link.springer.com/chapter/10.1007/978-1-4612-5695-

3_4. Including a paper co-authored by C. S. Schölten.

Dijkstra:1982:TRIa

[Dij82c] Dr. Edsger W. Dijkstra. Trip report IBM Seminar “Communication and Computers”, Newcastle, Sept. 1973. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 36–40. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982. URL http://

link.springer.com/chapter/10.1007/978-1-4612-5695-

3_6. Including a paper co-authored by C. S. Schölten.

Dijkstra:1982:ESC

[Dij82d] E. W. Dijkstra. EWD386 the solution to a cyclic relaxation problem. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 34–35. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982. EWD386’s original date is 1973.

Dijkstra:1982:ESS

[Dij82e] E. W. Dijkstra. EWD391 self-stabilization in spite of distributed control. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 41–46. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982. EWD391’s original date is 1973.

Dijkstra:1982:ACF

[Dij82f] Edsger W. Dijkstra. About the correctness of a few SASL programs. Circulated privately., May 1982. URL http://

www.cs.utexas.edu/users/EWD/ewd08xx/EWD825.PDF.

Dijkstra:1982:AEA

[Dij82g] Edsger W. Dijkstra. An alternative ending for AvG16/ EWD809. Circulated privately., February 1982. URL http:

//www.cs.utexas.edu/users/EWD/ewd08xx/EWD811.PDF.

Dijkstra:1982:BLH

[Dij82h] Edsger W. Dijkstra. A bagatelle for the left hand. Circulated privately., January 1982. URL http://www.cs.utexas.edu/

users/EWD/ewd08xx/EWD800.PDF.

Dijkstra:1982:CBPa

[Dij82i] Edsger W. Dijkstra. A collection of beautiful proofs. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 174–183. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982.

URL http://link.springer.com/chapter/10.1007/978-

1-4612-5695-3_31. Including a paper co-authored by C. S. Schölten.

Dijkstra:1982:CSa

[Dij82j] Edsger W. Dijkstra. Comments at a symposium. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 161–164. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982. URL http://

link.springer.com/chapter/10.1007/978-1-4612-5695-

3_27. Including a paper co-authored by C. S. Schölten.

Dijkstra:1982:SCOa

[Dij82k] Edsger W. Dijkstra. De software crisis, ontstaan en hardnekkigheid. (Dutch) [The software crisis, emergence and persistence]. Circulated privately., July 1982. URL http://www.

cs.utexas.edu/users/EWD/ewd08xx/EWD829.PDF.

Dijkstra:1982:SCOb

[Dij82l] Edsger W. Dijkstra. De software crisis, ontstaan en hardnekkigheid. (Dutch) [The software crisis, emergence and persistence]. De computer krant, 3(20), October 1982.

Dijkstra:1982:DTDa

[Dij82m] Edsger W. Dijkstra. Distributed termination detection revisited. Circulated privately., June 1982. URL http://www.cs.

utexas.edu/users/EWD/ewd08xx/EWD828.PDF.

Dijkstra:1982:EBNb

[Dij82n] Edsger W. Dijkstra. The equivalence of bounded nondeterminacy and continuity. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 358–359. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982. Including a paper co-authored by C. S. Schölten.

Dijkstra:1982:ESA

[Dij82o] Edsger W. Dijkstra. Errata: “Smoothsort, an alternative for sorting in situ” [Sci. Comput. Programming 1 (1981/82), no. 3, 223–233; MR 83g:68090]. Science of Computer Programming, 2(1):85, October 1982. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). See [Dij82-42].

Dijkstra:1982:FFT

[Dij82p] Edsger W. Dijkstra. The Fast Fourier Transform and the Perfect Shuffle. Circulated privately., February 25, 1982. URL

http://www.cs.utexas.edu/users/EWD/ewd08xx/EWD807.
PDF.

Dijkstra:1982:FMSa

[Dij82q] Edsger W. Dijkstra. Finding the maximum strong components in a directed graph. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 22–30. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24

.D54 1982. URL http://link.springer.com/chapter/10.

1007/978-1-4612-5695-3_3. Including a paper co-authored by C. S. Schölten.

Dijkstra:1982:FTSa

[Dij82r] Edsger W. Dijkstra. Formal techniques and sizeable programs. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 205–214. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54

1982. URL http://link.springer.com/chapter/10.1007/

978-1-4612-5695-3_35. Including a paper co-authored by C. S. Schölten.

Dijkstra:1982:PTP

[Dij82s] Edsger W. Dijkstra. From predicate transformers to predicates (Dedicated by the Tuesday Afternoon Club to C. A. R. Hoare at the occasion of his being elected Fellow of the Royal Society). Circulated privately., April 1982. URL http://

www.cs.utexas.edu/users/EWD/ewd08xx/EWD821.PDF.

Dijkstra:1982:HDWb

[Dij82t] Edsger W. Dijkstra. How do we tell truths that might hurt? ACM SIGPLAN Notices, 17(5):13–15, May 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dijkstra:1982:DSP

[Dij82u] Edsger W. Dijkstra. An in-depth seminar on proven tools & techniques of structured methodology for effective software configuration management. Circulated privately., Au-

gust 1982. URL http://www.cs.utexas.edu/users/EWD/
ewd08xx/EWD833.PDF.

Dijkstra:1982:IP

[Dij82v] Edsger W. Dijkstra. Inleiding programmeren. (Dutch) [Introduction to programming]. Technical report, Technische Universiteit Eindhoven, Eindhoven, The Netherlands, Au-

gust 1982. URL http://www.cs.utexas.edu/users/EWD/
OtherDocs/NN832.PDF.

Dijkstra:1982:IMIa

[Dij82w] Edsger W. Dijkstra. Inside “Mathematics Inc.”. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 54–55. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982. URL http:/

/link.springer.com/accesspage/chapter/10.1007/978-
1-4612-5695-3_10?coverImageUrl=/static%2F0.9313%2Fsites%

2Flink%2Fimages%2Fabstract_cover_placeholder.png. In-

cluding a paper co-authored by C. S. Schölten.

Dijkstra:1982:IMR

[Dij82x] Edsger W. Dijkstra. The introduction of MAES(R). In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 331–333. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982. Including a paper co-authored by C. S. Schölten.

Dijkstra:1982:LNP

[Dij82y] Edsger W. Dijkstra. Lecture notes “Predicate transformers”(Draft). Circulated privately., November 1982. URL

http://www.cs.utexas.edu/users/EWD/ewd08xx/EWD835.
PDF.

Dijkstra:1982:LMOa

[Dij82z] Edsger W. Dijkstra. A letter to my old friend Jonathan. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 99–103. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982.

URL http://link.springer.com/chapter/10.1007/978-

1-4612-5695-3_18. Including a paper co-authored by C. S. Schölten.

Dijkstra:1982:LTD

[Dij82-27] Edsger W. Dijkstra. Linearization of a two-dimensional search. Circulated privately., May 1982. URL http://www.

cs.utexas.edu/users/EWD/ewd08xx/EWD824.PDF.

Dijkstra:1982:MICa

[Dij82-28] Edsger W. Dijkstra. Mathematical induction and computing science. Published as [Dij82-29]., April 4, 1982. URL http:

//www.cs.utexas.edu/users/EWD/ewd08xx/EWD819.PDF.

Dijkstra:1982:MICb

[Dij82-29] Edsger W. Dijkstra. Mathematical induction and computing science. Nieuw Archief voor Wiskunde (3), 30(2):117–123, 1982. CODEN NAWIA7. ISSN 0028-9825. URL https://

zbmath.org/?q=an:0492.68010.

Dijkstra:1982:MFU

[Dij82-30] Edsger W. Dijkstra. McCarthy’s g1-function: an unfortunate paradigm. Circulated privately., December 21, 1982. URL

http://www.cs.utexas.edu/users/EWD/ewd08xx/EWD845.
PDF.

Dijkstra:1982:MAMa

[Dij82-31] Edsger W. Dijkstra. A multidisciplinary approach to mathematics. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 56–59. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54

1982. URL http://link.springer.com/chapter/10.1007/

978-1-4612-5695-3_11. Including a paper co-authored by C. S. Schölten.

Dijkstra:1982:NTM

[Dij82-32] Edsger W. Dijkstra. A nice theorem on monotonic predicate sequences. Circulated privately., April 1982. URL http://

www.cs.utexas.edu/users/EWD/ewd08xx/EWD818.PDF.

Dijkstra:1982:NSR

[Dij82-33] Edsger W. Dijkstra. A note on substitution and renaming. Circulated privately., September 1982. URL http://www.cs.

utexas.edu/users/EWD/ewd08xx/EWD836.PDF.

Dijkstra:1982:IR

[Dij82-34] Edsger W. Dijkstra. On iteration and recursion. Circulated

privately., April 1982. URL http://www.cs.utexas.edu/
users/EWD/ewd08xx/EWD820.PDF.

Dijkstra:1982:PSJ

[Dij82-35] Edsger W. Dijkstra. A personal summary of Jonkers’s program transformation. Circulated privately., Febru-

ary 1982. URL http://www.cs.utexas.edu/users/EWD/
ewd08xx/EWD812.PDF.

Dijkstra:1982:PSE

[Dij82-36] Edsger W. Dijkstra. A post-scriptum to EWD501. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 141–144. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982. Including a paper co-authored by C. S. Schölten.

Dijkstra:1982:RCN

[Dij82-37] Edsger. W. Dijkstra. [Reflections on [?]]. Circulated

privately., 1982. URL http://www.cs.utexas.edu/users/
EWD/ewd08xx/EWD841a.PDF.

Dijkstra:1982:ROD

[Dij82-38] Edsger W. Dijkstra. Repaying our debts. In Theoretical foundations of programming methodology (Munich, 1981), volume 91 of NATO Adv. Study Inst. Ser. C: Math. Phys. Sci., pages 135–141. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1982.

Dijkstra:1982:RBP

[Dij82-39] Edsger W. Dijkstra. A review of a book on PEARL. Circulated privately., February 1982. URL http://www.cs.

utexas.edu/users/EWD/ewd08xx/EWD814.PDF.

Dijkstra:1982:SSSa

[Dij82-40] Edsger W. Dijkstra. Self-stabilization in spite of distributed control. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 41–46. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54

1982. URL http://link.springer.com/chapter/10.1007/

978-1-4612-5695-3_7. Including a paper co-authored by C. S. Schölten.

Dijkstra:1982:SSE

[Dij82-41] Edsger W. Dijkstra. A short sequel to EWD842. Circulated privately., October 1982. URL http://www.cs.utexas.edu/

users/EWD/ewd08xx/EWD843.PDF.

Dijkstra:1982:SASa

[Dij82-42] Edsger W. Dijkstra. Smoothsort, an alternative for sorting in situ. In Theoretical foundations of programming methodology (Munich, 1981), volume 91 of NATO Adv. Study Inst. Ser. C: Math. Phys. Sci., pages 3–17. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1982.

Dijkstra:1982:SASb

[Dij82-43] Edsger W. Dijkstra. Smoothsort, an alternative for sorting in situ. Science of Computer Programming, 1(3):223–233, May 1982. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). See errata [Dij82o].

Dijkstra:1982:SOAa

[Dij82-44] Edsger W. Dijkstra. Speech at the occasion of an anniversary. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 50–53. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54

1982. URL http://link.springer.com/chapter/10.1007/

978-1-4612-5695-3_9. Including a paper co-authored by C. S. Schölten.

Dijkstra:1982:SPCa

[Dij82-45] Edsger W. Dijkstra. Stepwise program construction. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 1–14. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982.

URL http://link.springer.com/chapter/10.1007/978-

1-4612-5695-3_1. Including a paper co-authored by C. S. Schölten.

Dijkstra:1982:SFS

[Dij82-46] Edsger W. Dijkstra. A summary of formulae (and of some theory). Circulated privately., May 1982. URL http://www.

cs.utexas.edu/users/EWD/ewd08xx/EWD822.PDF.

Dijkstra:1982:TAI

[Dij82-47] Edsger W. Dijkstra. A theorem about infinite sequences of numbers. Circulated privately., August 1982. URL http://

www.cs.utexas.edu/users/EWD/ewd08xx/EWD832.PDF.

Dijkstra:1982:TNR

[Dij82-48] Edsger W. Dijkstra. “There is no ‘royal road’ to geometry”. Circulated privately., October 1982. URL http://www.cs.

utexas.edu/users/EWD/ewd08xx/EWD841.PDF.

Dijkstra:1982:TRWm

[Dij82-49] Edsger W. Dijkstra. Trip report E. W. Dijkstra 16th April/7th May, U.S.A. and Canada. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 120– 128. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982. Including a paper co-authored by C. S. Schölten.

Dijkstra:1982:TRWp

[Dij82-50] Edsger W. Dijkstra. Trip report E. W. Dijkstra, Copenhagen, 10–16 Sep. 1982. Circulated privately., Septem-

ber 1982. URL http://www.cs.utexas.edu/users/EWD/
ewd08xx/EWD838.PDF.

Dijkstra:1982:TRWl

[Dij82-51] Edsger W. Dijkstra. Trip report E. W. Dijkstra, Meeting IFIP W.G.2.3., Munich, 8 –14 December 1974. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 89–94. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982. Including a paper co-authored by C. S. Schölten.

Dijkstra:1982:TRWo

[Dij82-52] Edsger W. Dijkstra. Trip report E. W. Dijkstra, Newcastleupon Tyne, 6–10 Sep. 1982. Circulated privately., Septem-

ber 1982. URL http://www.cs.utexas.edu/users/EWD/
ewd08xx/EWD837.PDF.

Dijkstra:1982:TWDe

[Dij82-53] Edsger W. Dijkstra. Trip report E. W. Dijkstra, St. Pierrede-Chartreuse, 12–19 Dec. 1976. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 253– 258. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982. Including a paper co-authored by C. S. Schölten.

Dijkstra:1982:TRWk

[Dij82-54] Edsger W. Dijkstra. Trip report E. W. Dijkstra Summer School Munich, July 25 to August 4, 1973. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 31–33. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982. Including a paper co-authored by C. S. Schölten.

Dijkstra:1982:TRWn

[Dij82-55] Edsger W. Dijkstra. Trip report E. W. Dijkstra, USA, 20 May 1982–14 June 1982. Circulated privately., June 1982. URL

http://www.cs.utexas.edu/users/EWD/ewd08xx/EWD826.
PDF.

Dijkstra:1982:TRWq

[Dij82-56] Edsger W. Dijkstra. Trip report E. W. Dijkstra, Utrecht 30 Sep.–1 Oct. 1982. Circulated privately., October 1982. URL

http://www.cs.utexas.edu/users/EWD/ewd08xx/EWD839.
PDF.

Dijkstra:1982:TRVa

[Dij82-57] Edsger W. Dijkstra. Trip report visit ETH Zürich, 3–4 February 1975 by E. W. Dijkstra. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 95– 98. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982. URL http://link.springer.

com/chapter/10.1007/978-1-4612-5695-3_17. Including a

paper co-authored by C. S. Schölten.

Dijkstra:1982:TWDf

[Dij82-58] Edsger W. Dijkstra. Tripreport E. W. Dijkstra, Australia, 16 February 1977–21 March 1977. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 277– 283. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982. Including a paper co-authored by C. S. Schölten.

Dijkstra:1982:TWDb

[Dij82-59] Edsger W. Dijkstra. Tripreport E. W. Dijkstra, ECIconference 9–12 August 1976, Amsterdam. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 225–229. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982. Including a paper coauthored by C. S. Schölten.

Dijkstra:1982:TWDa

[Dij82-60] Edsger W. Dijkstra. Tripreport E. W. Dijkstra Newcastle, 8–12 September 1975. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 165–168. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982. Including a paper co-authored by C. S. Schölten.

Dijkstra:1982:TWDg

[Dij82-61] Edsger W. Dijkstra. Tripreport E. W. Dijkstra, Newcastleupon-Tyne, 5–10 Sept. 1977. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 319–323. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982. Including a paper co-authored by C. S. Schölten.

Dijkstra:1982:TWDc

[Dij82-62] Edsger W. Dijkstra. Tripreport E. W. Dijkstra, Poland and USSR, 4–25 September 1976. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 235– 244. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982. Including a paper co-authored by C. S. Schölten.

Dijkstra:1982:TWDd

[Dij82-63] Edsger W. Dijkstra. Tripreport E. W. Dijkstra, Tokyo, 28 Sep.–3 Oct. 1976. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 245–250. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982. Including a paper co-authored by C. S. Schölten.

Dijkstra:1982:TSB

[Dij82-64] Edsger W. Dijkstra. A tutorial on the split binary semaphore. In Theoretical foundations of programming methodology (Munich, 1981), volume 91 of NATO Adv. Study Inst. Ser. C: Math. Phys. Sci., pages 555–564. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1982.

Dijkstra:1982:TCE

[Dij82-65] Edsger W. Dijkstra. Two cheers for equivalence. Circulated privately., October 1982. URL http://www.cs.utexas.edu/

users/EWD/ewd08xx/EWD842.PDF.

Dijkstra:1982:VTO

[Dij82-66] Edsger W. Dijkstra. Variations on a theme: an open letter to C. A. R. Hoare. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 132–140. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982. Including a paper co-authored by C. S. Schölten.

Dijkstra:1982:VSE

[Dij82-67] Edsger W. Dijkstra. A very simple exercise in SASL. Circulated privately., June 1982. URL http://www.cs.utexas.

edu/users/EWD/ewd08xx/EWD827.PDF.

Dijkstra:1982:WNS

[Dij82-68] Edsger W. Dijkstra. Why numbering should start at zero. Circulated privately., August 1982. URL http://www.cs.

utexas.edu/users/EWD/ewd08xx/EWD831.PDF.

Dijkstra:1982:SWC

[Dij82-69] Edsger Wybe Dijkstra. Selected Writings on Computing: a Personal Perspective. Texts and Monographs in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). xvii + 362 pp. LCCN QA76.24 .D54 1982. Including a paper co-authored by C. S. Schölten.

Dijkstra:1982:ASA

[Dij82-70] Prof. Dr. Edsger W. Dijkstra. Acceptance speech for the AFIPS Harry Goode Memorial Award 1974. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 47–49. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982. URL http://

link.springer.com/chapter/10.1007/978-1-4612-5695-

3_8. Including a paper co-authored by C. S. Schölten.

Dijkstra:1982:CPCa

[Dij82-71] Prof. Dr. Edsger W. Dijkstra. A correctness proof for communicating processes: A small exercise. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 259– 263. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982. URL http://link.springer.

com/chapter/10.1007/978-1-4612-5695-3_47. Including a

paper co-authored by C. S. Schölten.

Dijkstra:1982:CSb

[Dij82-72] Prof. Dr. Edsger W. Dijkstra. “Craftsman or scientist?”. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 104–109. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54

1982. URL http://link.springer.com/chapter/10.1007/

978-1-4612-5695-3_19. Including a paper co-authored by C. S. Schölten.

Dijkstra:1982:EIDa

[Dij82-73] Prof. Dr. Edsger W. Dijkstra. An elephant inspired by the Dutch National Flag. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 264–267. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN

QA76.24 .D54 1982. URL http://link.springer.com/
chapter/10.1007/978-1-4612-5695-3_48. Including a pa-

per co-authored by C. S. Schölten.

Dijkstra:1982:EBNa

[Dij82-74] Prof. Dr. Edsger W. Dijkstra. The equivalence of bounded nondeterminacy and continuity. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 358– 359. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982. URL http://link.springer.

com/accesspage/chapter/10.1007/978-1-4612-5695-3_
65?coverImageUrl=/static%2F0.9313%2Fsites%2Flink%
2Fimages%2Fabstract_cover_placeholder.png. Including

a paper co-authored by C. S. Schölten.

Dijkstra:1982:EEEa

[Dij82-75] Prof. Dr. Edsger W. Dijkstra. Erratum and embellishments of EWD503. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 145–146. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982. URL http://link.springer.

com/accesspage/chapter/10.1007/978-1-4612-5695-3_
25?coverImageUrl=/static%2F0.9313%2Fsites%2Flink%
2Fimages%2Fabstract_cover_placeholder.png. Including

a paper co-authored by C. S. Schölten.

Dijkstra:1982:EDRa

[Dij82-76] Prof. Dr. Edsger W. Dijkstra. An exercise for Dr. R. M. Burstall. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 215–216. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982.

URL http://link.springer.com/accesspage/chapter/
10.1007/978-1-4612-5695-3_36?coverImageUrl=/static%
2F0.9313%2Fsites%2Flink%2Fimages%2Fabstract_cover_

placeholder.png. Including a paper co-authored by C. S. Schölten.

Dijkstra:1982:EMPa

[Dij82-77] Prof. Dr. Edsger W. Dijkstra. Exercises in making programs robust. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 110–119. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54

1982. URL http://link.springer.com/chapter/10.1007/

978-1-4612-5695-3_20. Including a paper co-authored by C. S. Schölten.

Dijkstra:1982:GIa

[Dij82-78] Prof. Dr. Edsger W. Dijkstra. A great improvement. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 217–219. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982.

URL http://link.springer.com/chapter/10.1007/978-

1-4612-5695-3_37. Including a paper co-authored by C. S. Schölten.

Dijkstra:1982:HDWa

[Dij82-79] Prof. Dr. Edsger W. Dijkstra. How do we tell truths that might hurt? In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 129–131. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54

1982. URL http://link.springer.com/chapter/10.1007/

978-1-4612-5695-3_22. Including a paper co-authored by C. S. Schölten.

Dijkstra:1982:IM

[Dij82-80] Prof. Dr. Edsger W. Dijkstra. The introduction of MAES(R). In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 331–333. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982.

URL http://link.springer.com/chapter/10.1007/978-

1-4612-5695-3_59. Including a paper co-authored by C. S. Schölten.

Dijkstra:1982:MBBa

[Dij82-81] Prof. Dr. Edsger W. Dijkstra. The mathematics behind the Banker’s Algorithm. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 308–312. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN

QA76.24 .D54 1982. URL http://link.springer.com/
chapter/10.1007/978-1-4612-5695-3_54. Including a pa-

per co-authored by C. S. Schölten.

Dijkstra:1982:MIPa

[Dij82-82] Prof. Dr. Edsger W. Dijkstra. Mathematics Inc., a private letter from its Chairman. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 184–187. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN

QA76.24 .D54 1982. URL http://link.springer.com/

chapter/10.1007/978-1-4612-5695-3_32. Including a pa-

per co-authored by C. S. Schölten.

Dijkstra:1982:MRAa

[Dij82-83] Prof. Dr. Edsger W. Dijkstra. Monotonic replacement algorithms and their implementation. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 84– 88. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982. URL http://link.springer.

com/chapter/10.1007/978-1-4612-5695-3_15. Including a

paper co-authored by C. S. Schölten.

Dijkstra:1982:MAFa

[Dij82-84] Prof. Dr. Edsger W. Dijkstra. More about the function “fusc” (a sequel to EWD570). In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 230– 232. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982. URL http://link.springer.

com/chapter/10.1007/978-1-4612-5695-3_41. Including a

paper co-authored by C. S. Schölten.

Dijkstra:1982:MHWa

[Dij82-85] Prof. Dr. Edsger W. Dijkstra. More on Hauck’s warning. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 172–173. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982.

URL http://link.springer.com/accesspage/chapter/
10.1007/978-1-4612-5695-3_30?coverImageUrl=/static%
2F0.9313%2Fsites%2Flink%2Fimages%2Fabstract_cover_

placeholder.png. Including a paper co-authored by C. S. Schölten.

Dijkstra:1982:NEBa

[Dij82-86] Prof. Dr. Edsger W. Dijkstra. A new elephant built from mosquitoes humming in harmony. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 79– 83. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982. URL http://link.springer.

com/chapter/10.1007/978-1-4612-5695-3_14. Including a

paper co-authored by C. S. Schölten.

Dijkstra:1982:NTRa

[Dij82-87] Prof. Dr. Edsger W. Dijkstra. A “Non trip report” from E. W. Dijkstra. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 200–204. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54

1982. URL http://link.springer.com/chapter/10.1007/

978-1-4612-5695-3_34. Including a paper co-authored by C. S. Schölten.

Dijkstra:1982:WHa

[Dij82-88] Prof. Dr. Edsger W. Dijkstra. On a warning from E. A. Hauck. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 169–171. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54

1982. URL http://link.springer.com/chapter/10.1007/

978-1-4612-5695-3_29. Including a paper co-authored by C. S. Schölten.

Dijkstra:1982:MSMa

[Dij82-89] Prof. Dr. Edsger W. Dijkstra. On making solutions more and more fine-grained (In gratitude dedicated to C. A. R. Hoare, D. E. Knuth, and J. F. Traub). In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 292– 307. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982. URL http://link.springer.

com/chapter/10.1007/978-1-4612-5695-3_53. Including a

paper co-authored by C. S. Schölten.

Dijkstra:1982:SIa

[Dij82-90] Prof. Dr. Edsger W. Dijkstra. On subgoal induction. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 223–224. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982.

URL http://link.springer.com/accesspage/chapter/
10.1007/978-1-4612-5695-3_39?coverImageUrl=/static%
2F0.9313%2Fsites%2Flink%2Fimages%2Fabstract_cover_

placeholder.png. Including a paper co-authored by C. S. Schölten.

Dijkstra:1982:FAOa

[Dij82-91] Prof. Dr. Edsger W. Dijkstra. On the fact that the Atlantic Ocean has two sides. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 268–276. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN

QA76.24 .D54 1982. URL http://link.springer.com/
chapter/10.1007/978-1-4612-5695-3_49. Including a pa-

per co-authored by C. S. Schölten.

Dijkstra:1982:RSTa

[Dij82-92] Prof. Dr. Edsger W. Dijkstra. On the role of scientific thought. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 60–66. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54

1982. URL http://link.springer.com/chapter/10.1007/

978-1-4612-5695-3_12. Including a paper co-authored by C. S. Schölten.

Dijkstra:1982:TBSa

[Dij82-93] Prof. Dr. Edsger W. Dijkstra. On two beautiful solutions designed by Martin Rem. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 313–318. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN

QA76.24 .D54 1982. URL http://link.springer.com/
chapter/10.1007/978-1-4612-5695-3_55. Including a pa-

per co-authored by C. S. Schölten.

Dijkstra:1982:WSTa

[Dij82-94] Prof. Dr. Edsger W. Dijkstra. On weak and strong termination. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 355–357. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54

1982. URL http://link.springer.com/chapter/10.1007/

978-1-4612-5695-3_64. Including a paper co-authored by C. S. Schölten.

Dijkstra:1982:WUBa

[Dij82-95] Prof. Dr. Edsger W. Dijkstra. On Webster, users, bugs, and Aristotle. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 288–291. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54

1982. URL http://link.springer.com/chapter/10.1007/

978-1-4612-5695-3_52. Including a paper co-authored by C. S. Schölten.

Dijkstra:1982:Pa

[Dij82-96] Prof. Dr. Edsger W. Dijkstra. A parable. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 251– 252. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982. URL http://link.springer.

com/accesspage/chapter/10.1007/978-1-4612-5695-3_
45?coverImageUrl=/static%2F0.9313%2Fsites%2Flink%

2Fimages%2Fabstract_cover_placeholder.png. Including

a paper co-authored by C. S. Schölten.

Dijkstra:1982:PSGa

[Dij82-97] Prof. Dr. Edsger W. Dijkstra. A personal summary of the Gries–Owicki Theory. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 188–199. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN

QA76.24 .D54 1982. URL http://link.springer.com/
chapter/10.1007/978-1-4612-5695-3_33. Including a pa-

per co-authored by C. S. Schölten.

Dijkstra:1982:PIa

[Dij82-98] Prof. Dr. Edsger W. Dijkstra. Program inversion. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 351–354. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982. URL http://

link.springer.com/chapter/10.1007/978-1-4612-5695-

3_63. Including a paper co-authored by C. S. Schölten.

Dijkstra:1982:SOLa

[Dij82-99] Prof. Dr. Edsger W. Dijkstra. A somewhat open letter to EAA or: Why I proved the boundedness of the nondeterminacy in the way I did. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 284–287. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN

QA76.24 .D54 1982. URL http://link.springer.com/
chapter/10.1007/978-1-4612-5695-3_51. Including a pa-

per co-authored by C. S. Schölten.

Dijkstra:1982:SSVa

[Dij82-100] Prof. Dr. Edsger W. Dijkstra. A story that starts with a very good computer. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 360–362. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN

QA76.24 .D54 1982. URL http://link.springer.com/
chapter/10.1007/978-1-4612-5695-3_66. Including a pa-

per co-authored by C. S. Schölten.

Dijkstra:1982:SEa

[Dij82-101] Prof. Dr. Edsger W. Dijkstra. A synthesis emerging? In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 147–160. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982.

URL http://link.springer.com/chapter/10.1007/978-

1-4612-5695-3_26. Including a paper co-authored by C. S. Schölten.

Dijkstra:1982:TAOa

[Dij82-102] Prof. Dr. Edsger W. Dijkstra. A theorem about odd powers of odd integers. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 349–350. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982. URL http://link.springer.

com/accesspage/chapter/10.1007/978-1-4612-5695-3_
62?coverImageUrl=/static%2F0.9313%2Fsites%2Flink%
2Fimages%2Fabstract_cover_placeholder.png. Including

a paper co-authored by C. S. Schölten.

Dijkstra:1982:TGRa

[Dij82-103] Prof. Dr. Edsger W. Dijkstra. The three golden rules for successful scientific research. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 329–330. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982. URL http://link.springer.

com/accesspage/chapter/10.1007/978-1-4612-5695-3_
58?coverImageUrl=/static%2F0.9313%2Fsites%2Flink%
2Fimages%2Fabstract_cover_placeholder.png. Including

a paper co-authored by C. S. Schölten.

Dijkstra:1982:TWHa

[Dij82-104] Prof. Dr. Edsger W. Dijkstra. A time-wise hierarchy imposed upon the use of a two-level store. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 67– 78. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982. URL http://link.springer.

com/chapter/10.1007/978-1-4612-5695-3_13. Including a

paper co-authored by C. S. Schölten.

Dijkstra:1982:HDMa

[Dij82-105] Prof. Dr. Edsger W. Dijkstra. To H. D. Mills, Chairman Software Methodology Panel. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 220–222. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN

QA76.24 .D54 1982. URL http://link.springer.com/
chapter/10.1007/978-1-4612-5695-3_38. Including a pa-

per co-authored by C. S. Schölten.

Dijkstra:1982:TRWc

[Dij82-106] Prof. Dr. Edsger W. Dijkstra. Trip report E. W. Dijkstra 16th April / 7th May, 1975, U.S.A. and Canada. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 120–128. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982.

URL http://link.springer.com/chapter/10.1007/978-

1-4612-5695-3_21. Including a paper co-authored by C. S. Schölten.

Dijkstra:1982:TRWi

[Dij82-107] Prof. Dr. Edsger W. Dijkstra. Trip report E. W. Dijkstra, Australia, 16 February 1977–21 March 1977. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 277–283. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982. URL http://

link.springer.com/chapter/10.1007/978-1-4612-5695-

3_50. Including a paper co-authored by C. S. Schölten.

Dijkstra:1982:TRWe

[Dij82-108] Prof. Dr. Edsger W. Dijkstra. Trip report E. W. Dijkstra, ECI-Conference 9–12 August 1976, Amsterdam. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 225–229. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982.

URL http://link.springer.com/chapter/10.1007/978-

1-4612-5695-3_40. Including a paper co-authored by C. S. Schölten.

Dijkstra:1982:TRWb

[Dij82-109] Prof. Dr. Edsger W. Dijkstra. Trip report E. W. Dijkstra, Meeting IFIP W.G.2.3, Munich, 8—14 December 1974. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 89–94. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54

1982. URL http://link.springer.com/chapter/10.1007/

978-1-4612-5695-3_16. Including a paper co-authored by C. S. Schölten.

Dijkstra:1982:TRWd

[Dij82-110] Prof. Dr. Edsger W. Dijkstra. Trip report E. W. Dijkstra, Newcastle, 8–12 September 1975. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 165– 168. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982. URL http://link.springer.

com/chapter/10.1007/978-1-4612-5695-3_28. Including a

paper co-authored by C. S. Schölten.

Dijkstra:1982:TRWj

[Dij82-111] Prof. Dr. Edsger W. Dijkstra. Trip report E. W. Dijkstra, Newcastle-upon-Tyne, 5–10 Sept. 1977. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 319– 323. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982. URL http://link.springer.

com/chapter/10.1007/978-1-4612-5695-3_56. Including a

paper co-authored by C. S. Schölten.

Dijkstra:1982:TRWf

[Dij82-112] Prof. Dr. Edsger W. Dijkstra. Trip report E. W. Dijkstra, Poland and USSR, 4–25 September 1976. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 235– 244. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982. URL http://link.springer.

com/chapter/10.1007/978-1-4612-5695-3_43. Including a

paper co-authored by C. S. Schölten.

Dijkstra:1982:TRWh

[Dij82-113] Prof. Dr. Edsger W. Dijkstra. Trip report E. W. Dijkstra, St. Pierre-de-Chartreuse, 12–19 Dec. 1976. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 253– 258. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982. URL http://link.springer.

com/chapter/10.1007/978-1-4612-5695-3_46. Including a

paper co-authored by C. S. Schölten.

Dijkstra:1982:TRWg

[Dij82-114] Prof. Dr. Edsger W. Dijkstra. Trip report E. W. Dijkstra, Tokyo, 28 Sept–3 Oct. 1976. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 245– 250. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982. URL http://link.springer.

com/chapter/10.1007/978-1-4612-5695-3_44. Including a

paper co-authored by C. S. Schölten.

Dijkstra:1982:WSEa

[Dij82-115] Prof. Dr. Edsger W. Dijkstra. “Why is software so expensive?” An explanation to the hardware designer. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 338–348. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982.

URL http://link.springer.com/chapter/10.1007/978-

1-4612-5695-3_61. Including a paper co-authored by C. S. Schölten.

Dijkstra:1982:WNPa

[Dij82-116] Prof. Dr. Edsger W. Dijkstra. Why naive program transformation systems are unlikely to work. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 324– 328. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982. URL http://link.springer.

com/chapter/10.1007/978-1-4612-5695-3_57. Including a

paper co-authored by C. S. Schölten.

Dijkstra:1983:P

[Dij83a] Edsger W. Dijkstra. |x[n]| = x[n − 1] + x[n + 1] has period 9. Circulated privately., August 1983. URL http://www.cs.

utexas.edu/users/EWD/ewd08xx/EWD861a.PDF.

Dijkstra:1983:APB

[Dij83b] Edsger W. Dijkstra. An analytical proof of the Butterfly Theorem. Circulated privately., November 1983. URL http:

//www.cs.utexas.edu/users/EWD/ewd08xx/EWD866.PDF.

Dijkstra:1983:CGE

[Dij83c] Edsger W. Dijkstra. Computers and General Education: a position paper. Circulated privately., Decem-

ber 1983. URL http://www.cs.utexas.edu/users/EWD/
ewd08xx/EWD868.PDF.

Dijkstra:1983:CE

[Dij83d] Edsger W. Dijkstra. A correction of EWD851b. Circulated

privately., April 1983. URL http://www.cs.utexas.edu/
users/EWD/ewd08xx/EWD853.PDF.

Dijkstra:1983:DSC

[Dij83e] Edsger W. Dijkstra. The distributed snapshot of Chandy/ Lamport/Misra. Circulated privately., September 1983. URL

http://www.cs.utexas.edu/users/EWD/ewd08xx/EWD864.
PDF.

Dijkstra:1983:DSK

[Dij83f] Edsger W. Dijkstra. The Distributed Snapshot of K. M. Chandy and L. Lamport. Circulated privately., Novem-

ber 1983. URL http://www.cs.utexas.edu/users/EWD/
ewd08xx/EWD864a.PDF.

Dijkstra:1983:FMa

[Dij83g] Edsger W. Dijkstra. The fruits of misunderstanding. Published as [Dij83h]., May 1983. URL http://www.cs.utexas.

edu/users/EWD/ewd08xx/EWD854.PDF.

Dijkstra:1983:FMb

[Dij83h] Edsger W. Dijkstra. The fruits of misunderstanding. Elektronische Rechenanlagen, 25(6):10–13, 1983. CODEN ELRAA4. ISSN 0013-5720.

Dijkstra:1983:GOF

[Dij83i] Edsger W. Dijkstra. Generalizing an old formula. Circulated privately., June 1983. URL http://www.cs.utexas.

edu/users/EWD/ewd08xx/EWD857.PDF.

Dijkstra:1983:IEW

[Dij83j] Edsger W. Dijkstra. Ingenieur Ebergen, Weledelgetrenge Heer. (Dutch) [Engineer Ebergen, noble gentleman]. Circulated privately., 1983. URL http://www.cs.utexas.edu/

users/EWD/ewd08xx/EWD853a.PDF.

Dijkstra:1983:JHDa

[Dij83k] Edsger W. Dijkstra. Judging “HOS” from a distance. Circulated privately., April 1983. URL http://www.cs.utexas.

edu/users/EWD/ewd08xx/EWD852.PDF.

Dijkstra:1983:JHDb

[Dij83l] Edsger W. Dijkstra. Judging “HOS” from a distance. Circulated privately., April 1983. URL http://www.cs.utexas.

edu/users/EWD/ewd08xx/EWD852a.PDF.

Dijkstra:1983:MMP

[Dij83m] Edsger W. Dijkstra. My mother’s proof of the Butterfly Theorem (see EWD866). Circulated privately., Decem-

ber 1983. URL http://www.cs.utexas.edu/users/EWD/
ewd08xx/EWD870.PDF.

Dijkstra:1983:I

[Dij83n] Edsger W. Dijkstra. On IPW’s. Circulated privately., Novem-

ber 1983. URL http://www.cs.utexas.edu/users/EWD/
ewd08xx/EWD867.PDF.

Dijkstra:1983:MP

[Dij83o] Edsger W. Dijkstra. On maximizing a product. Circulated

privately., June 1983. URL http://www.cs.utexas.edu/
users/EWD/ewd08xx/EWD856.PDF.

Dijkstra:1983:PCR

[Dij83p] Edsger W. Dijkstra. Predicate calculus revisited. Circulated privately., October 1983. URL http://www.cs.utexas.edu/

users/EWD/ewd08xx/EWD863.PDF.

Dijkstra:1983:RCTa

[Dij83q] Edsger W. Dijkstra. Reducing control traffic in a distributed implementation of mutual exclusion. Circulated privately.,

March 1983. URL http://www.cs.utexas.edu/users/EWD/
ewd08xx/EWD851.PDF.

Dijkstra:1983:RCTb

[Dij83r] Edsger W. Dijkstra. Reducing control traffic in a distributed implementation of mutual exclusion. Circulated privately.,

March 1983. URL http://www.cs.utexas.edu/users/EWD/
ewd08xx/EWD851a.PDF.

Dijkstra:1983:RCTc

[Dij83s] Edsger W. Dijkstra. Reducing control traffic in a distributed implementation of mutual exclusion. Circulated privately.,

March 1983. URL http://www.cs.utexas.edu/users/EWD/
ewd08xx/EWD851b.PDF.

Dijkstra:1983:SP

[Dij83t] Edsger W. Dijkstra. A sequence with |x[n]| = x[n − 1] + x[n + 1] has period 9. Circulated privately., Septem-

ber 1983. URL http://www.cs.utexas.edu/users/EWD/
ewd08xx/EWD862.PDF.

Dijkstra:1983:SPC

[Dij83u] Edsger W. Dijkstra. Solution of a problem in concurrent programming control (reprint). Communications of the Association for Computing Machinery, 26(1):21–22, January 1983. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Reprint of [Dij65h].

Dijkstra:1983:SMS

[Dij83v] Edsger W. Dijkstra. The structure of the “THE”- multiprogramming system (reprint). Communications of the Association for Computing Machinery, 26(1):49–52, January 1983. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Reprint of [Dij68p].

Dijkstra:1983:TAV

[Dij83w] Edsger W. Dijkstra. Ter afsluiting van de “Inleiding tot de Kunst van het Programmeren”. (Dutch) [In conclusion to the “Introduction to Art of Programming”]. Circulated privately.,

December 1983. URL http://www.cs.utexas.edu/users/
EWD/ewd08xx/EWD869.PDF.

Dijkstra:1983:TRWa

[Dij83x] Edsger W. Dijkstra. Trip report E. W. Dijkstra, Australia, 19 Jan. 1983–12 Feb. 1983. Circulated privately., Febru-

ary 1983. URL http://www.cs.utexas.edu/users/EWD/
ewd08xx/EWD847.PDF.

Dijkstra:1983:TRWc

[Dij83y] Edsger W. Dijkstra. Trip report E. W. Dijkstra, Barcelona 5–7 June 1983. Circulated privately., June 1983. URL http:

//www.cs.utexas.edu/users/EWD/ewd08xx/EWD855.PDF.

Dijkstra:1983:TRWe

[Dij83z] Edsger W. Dijkstra. Trip report E. W. Dijkstra, Brasil & USA, 17 Oct.–6 Nov. 1983. Circulated privately., Novem-

ber 1983. URL http://www.cs.utexas.edu/users/EWD/
ewd08xx/EWD865.PDF.

Dijkstra:1983:TRWb

[Dij83-27] Edsger W. Dijkstra. Trip report E. W. Dijkstra, Helsinki, 1–4 March 1983. Circulated privately., March 1983. URL http:

//www.cs.utexas.edu/users/EWD/ewd08xx/EWD850.PDF.

Dijkstra:1983:TRWd

[Dij83-28] Edsger W. Dijkstra. Trip report E. W. Dijkstra, USA, 10 June–3 July 1983. Circulated privately., July 1983. URL

http://www.cs.utexas.edu/users/EWD/ewd08xx/EWD858.
PDF.

Dijkstra:1984:MLD

[Dij84a] Edsger W. Dijkstra. De microprocessor als lollie. (Dutch) [The microprocessor as lolly [hard candy/sweet]]. Circulated

privately., 1984. URL http://www.cs.utexas.edu/users/
EWD/ewd08xx/EWD876.PDF.

Dijkstra:1984:FPD

[Dij84b] Edsger W. Dijkstra. A formal program derivation for the record. Circulated privately., September 1984. URL http:/

/www.cs.utexas.edu/users/EWD/ewd08xx/EWD897.PDF.

Dijkstra:1984:INa

[Dij84c] Edsger W. Dijkstra. Invariance and nondeterminacy. manuscript version; published as [?, ?], February 1984. URL

http://www.cs.utexas.edu/users/EWD/ewd08xx/EWD871m.
PDF.

Dijkstra:1984:INb

[Dij84d] Edsger W. Dijkstra. Invariance and nondeterminacy. Typescript version; published as [Dij85r, Dij84e]., Febru-

ary 1984. URL http://www.cs.utexas.edu/users/EWD/
ewd08xx/EWD871t.PDF.

Dijkstra:1984:INc

[Dij84e] Edsger W. Dijkstra. Invariance and nondeterminacy. Philosophical Transactions of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 312(1522):491–499, October 1984. CODEN PTRMAD. ISSN 0080-4614. URL http:

//www.jstor.org/stable/37447. With discussion, Mathe-

matical logic and programming languages.

Dijkstra:1984:LEC

[Dij84f] Edsger W. Dijkstra. The little essay I could not write. Circulated privately., January 1984. URL http://www.cs.

utexas.edu/users/EWD/ewd08xx/EWD872.PDF.

Dijkstra:1984:Mb

[Dij84g] Edsger W. Dijkstra. Minsegsumtwodim. Circulated pri-

vately., 1984. URL http://www.cs.utexas.edu/users/
EWD/ewd09xx/EWD900.PDF.

Dijkstra:1984:Mc

[Dij84h] Edsger W. Dijkstra. Minsegsumtwodim. Circulated pri-

vately., December 1984. URL http://www.cs.utexas.edu/
users/EWD/ewd09xx/EWD900a.PDF.

Dijkstra:1984:MQ

[Dij84i] Edsger W. Dijkstra. Monotonicity and quantification. Circulated privately., January 1984. URL http://www.cs.

utexas.edu/users/EWD/ewd08xx/EWD875.PDF.

Dijkstra:1984:PP

[Dij84j] Edsger W. Dijkstra. On partitioning predicates. Circulated

privately., July 1984. URL http://www.cs.utexas.edu/
users/EWD/ewd08xx/EWD894.PDF.

Dijkstra:1984:NCS

[Dij84k] Edsger W. Dijkstra. On the nature of computing science. Circulated privately., August 1984. URL http://www.cs.

utexas.edu/users/EWD/ewd08xx/EWD896.PDF.

Dijkstra:1984:PSM

[Dij84l] Edsger W. Dijkstra. A problem solved by my nephew Sybrand L. Dijkstra. Circulated privately., December 1984. URL

http://www.cs.utexas.edu/users/EWD/ewd09xx/EWD904.
PDF.

Dijkstra:1984:RCS

[Dij84m] Edsger W. Dijkstra. The regularity calculus: a second trial. Circulated privately., 1984. URL http://www.cs.utexas.

edu/users/EWD/ewd09xx/EWD902.PDF.

Dijkstra:1984:RPA

[Dij84n] Edsger W. Dijkstra. A review of a paper in AI. Circulated privately., September 1984. URL http://www.cs.utexas.

edu/users/EWD/ewd08xx/EWD899.PDF.

Dijkstra:1984:REP

[Dij84o] Edsger W. Dijkstra. A review of “The Evolution of Programs”. Circulated privately., February 1984. URL http://

www.cs.utexas.edu/users/EWD/ewd08xx/EWD881.PDF.

Dijkstra:1984:SSE

[Dij84p] Edsger W. Dijkstra. A short sequel to EWD863. Circulated privately., December 1984. URL http://www.cs.utexas.

edu/users/EWD/ewd09xx/EWD903.PDF.

Dijkstra:1984:SQ

[Dij84q] Edsger W. Dijkstra. Some questions. Report ????, ????, ????, November 1984.

Dijkstra:1984:TCS

[Dij84r] Edsger W. Dijkstra. The threats to computing science. Circulated privately., November 1984. URL http://www.cs.

utexas.edu/users/EWD/ewd08xx/EWD898.PDF.

Dijkstra:1984:TRWf

[Dij84s] Edsger W. Dijkstra. Trip report E. W. Dijkstra, London, 12– 14 June 1984. Circulated privately., June 1984. URL http:

//www.cs.utexas.edu/users/EWD/ewd08xx/EWD892.PDF.

Dijkstra:1984:TRWb

[Dij84t] Edsger W. Dijkstra. Trip report E. W. Dijkstra, London, 14–17 February 1984. Circulated privately., Febru-

ary 1984. URL http://www.cs.utexas.edu/users/EWD/
ewd08xx/EWD879.PDF.

Dijkstra:1984:TRWa

[Dij84u] Edsger W. Dijkstra. Trip report E. W. Dijkstra, London and Colchester, 16–20 Jan. 1984. Circulated privately., Jan-

uary 1984. URL http://www.cs.utexas.edu/users/EWD/
ewd08xx/EWD873.PDF.

Dijkstra:1984:TRWh

[Dij84v] Edsger W. Dijkstra. Trip report E. W. Dijkstra, Marktoberdorf, 30 July–12 Aug. 1984. Circulated privately., Septem-

ber 1984. URL http://www.cs.utexas.edu/users/EWD/
ewd08xx/EWD895.PDF.

Dijkstra:1984:TRWc

[Dij84w] Edsger W. Dijkstra. Trip report E. W. Dijkstra, München, 12–14 April 1984. Circulated privately., April 1984. URL

http://www.cs.utexas.edu/users/EWD/ewd08xx/EWD885.
PDF.

Dijkstra:1984:TRWg

[Dij84x] Edsger W. Dijkstra. Trip report E. W. Dijkstra, USA, 17–26 June 1984. Circulated privately., June 1984. URL http://

www.cs.utexas.edu/users/EWD/ewd08xx/EWD893.PDF.

Dijkstra:1984:TRWd

[Dij84y] Edsger W. Dijkstra. Trip report E. W. Dijkstra, USA, 23 April–10 May 1984. Circulated privately., May 1984. URL

http://www.cs.utexas.edu/users/EWD/ewd08xx/EWD887.
PDF.

Dijkstra:1984:TRWe

[Dij84z] Edsger W. Dijkstra. Trip report E. W. Dijkstra, Zürich, 22–24 May 1984. Circulated privately., May 1984. URL http://

www.cs.utexas.edu/users/EWD/ewd08xx/EWD888.PDF.

Dijkstra:1984:UFM

[Dij84-27] Edsger W. Dijkstra. User-friendly Mathematics. Circulated

privately., May 1984. URL http://www.cs.utexas.edu/
users/EWD/ewd08xx/EWD889.PDF.

Dijkstra:1984:WFS

[Dij84-28] Edsger W. Dijkstra. Well-founded sets revisited. Circulated

privately., June 1984. URL http://www.cs.utexas.edu/
users/EWD/ewd08xx/EWD891.PDF.

Dijkstra:1985:AME

[Dij85a] Edsger W. Dijkstra. Another misguided effort. Circulated

privately., April 1985. URL http://www.cs.utexas.edu/
users/EWD/ewd09xx/EWD917.PDF.

Dijkstra:1985:AAT

[Dij85b] Edsger W. Dijkstra. The ATAC ( = Austin Tuesday Afternoon Club). Circulated privately., August 1985. URL http:

//www.cs.utexas.edu/users/EWD/ewd09xx/EWD927.PDF.

Dijkstra:1985:BPS

[Dij85c] Edsger W. Dijkstra. A belated proof of self-stabilization. Published as [Dij86f]., 1985. URL http://www.cs.utexas.edu/

users/EWD/ewd09xx/EWD922.PDF.

Dijkstra:1985:CCS

[Dij85d] Edsger W. Dijkstra. Can computing science save the computer industry? Circulated privately., May 1985. URL http:

//www.cs.utexas.edu/users/EWD/ewd09xx/EWD920.PDF.

Dijkstra:1985:CTA

[Dij85e] Edsger W. Dijkstra. A computer to think about. Circulated privately., July 1985. URL http://www.cs.utexas.

edu/users/EWD/ewd09xx/EWD926.PDF.

Dijkstra:1985:CEb

[Dij85f] Edsger W. Dijkstra. Concerning the equivalence. Circulated

privately., August 1985. URL http://www.cs.utexas.edu/
users/EWD/ewd09xx/EWD931.PDF.

Dijkstra:1985:CEE

[Dij85g] Edsger W. Dijkstra. Correction and extension of EWD932b. Circulated privately., September 1985. URL http://www.cs.

utexas.edu/users/EWD/ewd09xx/EWD937.PDF.

Dijkstra:1985:CEa

[Dij85h] Edsger W. Dijkstra. A correction of EWD914-15. Circulated privately., April 30, 1985. URL http://www.cs.utexas.

edu/users/EWD/ewd09xx/EWD921.PDF.

Dijkstra:1985:DSK

[Dij85i] Edsger W. Dijkstra. The distributed snapshot of K. M. Chandy and L. Lamport. In Control flow and data flow: concepts of distributed programming (Marktoberdorf, 1984), volume 14 of NATO Adv. Sci. Inst. Ser. F Comput. Systems Sci., pages 513–517. Springer, Berlin, 1985.

Dijkstra:1985:ESE

[Dij85j] Edsger W. Dijkstra. Extreme solutions of equations (Draft Ch.5). Became Chapter 8; see EWD 969 for a later version; published as [?]., March 1985. URL http://www.cs.utexas.

edu/users/EWD/ewd09xx/EWD912.PDF.

Dijkstra:1985:RBBa

[Dij85k] Edsger W. Dijkstra. For the record: Batcher’s Baffler. Circulated privately., September 1985. URL http://www.cs.

utexas.edu/users/EWD/ewd09xx/EWD935.PDF.

Dijkstra:1985:RBBb

[Dij85l] Edsger W. Dijkstra. For the record: Batcher’s Baffler. Circulated privately., September 1985. URL http://www.cs.

utexas.edu/users/EWD/ewd09xx/EWD935a.PDF.

Dijkstra:1985:RLS

[Dij85m] Edsger W. Dijkstra. For the record: the Linear Search. Circulated privately., January 1985. URL http://www.cs.

utexas.edu/users/EWD/ewd09xx/EWD930.PDF.

Dijkstra:1985:GFH

[Dij85n] Edsger W. Dijkstra. A generalization of the functions head and tail (unfinished). Circulated privately., 1985. URL http:

//www.cs.utexas.edu/users/EWD/ewd09xx/EWD929.PDF.

Dijkstra:1985:ICC

[Dij85o] Edsger W. Dijkstra. The image construction in computerized axial tomography (CAT). In Control flow and data flow: concepts of distributed programming (Marktoberdorf, 1984), volume 14 of NATO Adv. Sci. Inst. Ser. F Comput. Systems Sci., pages 503–506. Springer, Berlin, 1985.

Dijkstra:1985:IE

[Dij85p] Edsger W. Dijkstra. An improvement on EWD912. Circulated privately., April 1985. URL http://www.cs.utexas.

edu/users/EWD/ewd09xx/EWD916.PDF.

Dijkstra:1985:ISO

[Dij85q] Edsger W. Dijkstra. Incremental sorting once more. Circulated privately., September 1985. URL http://www.cs.

utexas.edu/users/EWD/ewd09xx/EWD939.PDF.

Dijkstra:1985:IN

[Dij85r] Edsger W. Dijkstra. Invariance and nondeterminacy. In Mathematical Logic and Programming Languages, pages 157– 164. Prentice-Hall, Upper Saddle River, NJ 07458, USA,

Dijkstra:1985:LTM

[Dij85s] Edsger W. Dijkstra. A letter to a typewriter manufacturer. Circulated privately., 1985. URL http://www.cs.utexas.

edu/users/EWD/ewd09xx/EWD947.PDF.

Dijkstra:1985:LPO

[Dij85t] Edsger W. Dijkstra. The longest plateau and other stories. Circulated privately., October 1985. URL http://www.cs.

utexas.edu/users/EWD/ewd09xx/EWD942.PDF.

Dijkstra:1985:MOU

[Dij85u] Edsger W. Dijkstra. Mainly on our use of the predicate calculus (Draft Ch.2). Circulated privately., Febru-

ary 1985. URL http://www.cs.utexas.edu/users/EWD/
ewd09xx/EWD905.PDF.

Dijkstra:1985:MPL

[Dij85v] Edsger W. Dijkstra. Misra’s proof of Lamport’s mutual exclusion. Circulated privately., November 1985. URL http:/

/www.cs.utexas.edu/users/EWD/ewd09xx/EWD948.PDF.

Dijkstra:1985:CG

[Dij85w] Edsger W. Dijkstra. On a cultural gap. Published as [Dij86x].,

September 1985. URL http://www.cs.utexas.edu/users/
EWD/ewd09xx/EWD924.PDF.

Dijkstra:1985:CGD

[Dij85x] Edsger W. Dijkstra. On a cultural gap (Draft). Circulated

privately., March 1985. URL http://www.cs.utexas.edu/
users/EWD/ewd09xx/EWD913.PDF.

Dijkstra:1985:PPM

[Dij85y] Edsger W. Dijkstra. On a problem posed by M. R. Khalil (for the record). Circulated privately., August 1985. URL http:

//www.cs.utexas.edu/users/EWD/ewd09xx/EWD933.PDF.

Dijkstra:1985:AS

[Dij85z] Edsger W. Dijkstra. On anthropomorphism in science. Circulated privately., September 1985. URL http://www.cs.

utexas.edu/users/EWD/ewd09xx/EWD936.PDF.

Dijkstra:1985:S

[Dij85-27] Edsger W. Dijkstra. On structures. early draft of Chapter 1; published as [?]., November 1985. URL http://www.cs.

utexas.edu/users/EWD/ewd09xx/EWD928.PDF.

Dijkstra:1985:NCS

[Dij85-28] Edsger W. Dijkstra. On the nature of computing science. In Control flow and data flow: concepts of distributed programming (Marktoberdorf, 1984), volume 14 of NATO Adv. Sci. Inst. Ser. F Comput. Systems Sci., pages 1–3. Springer, Berlin, 1985.

Dijkstra:1985:OT

[Dij85-29] Edsger W. Dijkstra. The operation “trickle”. Circulated pri-

vately., November 1985. URL http://www.cs.utexas.edu/
users/EWD/ewd09xx/EWD944.PDF.

Dijkstra:1985:PTD

[Dij85-30] Edsger W. Dijkstra. Predicate transformers (Draft Ch.3). appears to be an early draft of Chapter 6; circulated privately,

February 1985. URL http://www.cs.utexas.edu/users/
EWD/ewd09xx/EWD908.PDF.

Dijkstra:1985:PLR

[Dij85-31] Edsger W. Dijkstra. The problem of the Longest Ribbon. Circulated privately., October 1985. URL http://www.cs.

utexas.edu/users/EWD/ewd09xx/EWD943.PDF.

Dijkstra:1985:PP

[Dij85-32] Edsger W. Dijkstra. The proof of the pudding. Circulated privately., November 1985. URL http://www.cs.utexas.

edu/users/EWD/ewd09xx/EWD946.PDF.

Dijkstra:1985:RDC

[Dij85-33] Edsger W. Dijkstra. The repetition (Draft Ch.6). Circulated

privately., April 1985. URL http://www.cs.utexas.edu/
users/EWD/ewd09xx/EWD914.PDF.

Dijkstra:1985:RLA

[Dij85-34] Edsger W. Dijkstra. Rotating a linear array. Circulated pri-

vately., September 1985. URL http://www.cs.utexas.edu/
users/EWD/ewd09xx/EWD941.PDF.

Dijkstra:1985:SS

[Dij85-35] Edsger W. Dijkstra. The Saddleback Search. Circulated pri-

vately., September 1985. URL http://www.cs.utexas.edu/
users/EWD/ewd09xx/EWD934.PDF.

Dijkstra:1985:SSLa

[Dij85-36] Edsger W. Dijkstra. Semantics of straight-line programs (Draft Chap.4). Appears to be an early draft of Chapter 7; published as [Dij90r]., March 1985. URL http://www.cs.

utexas.edu/users/EWD/ewd09xx/EWD910.PDF.

Dijkstra:1985:SED

[Dij85-37] Edsger W. Dijkstra. A sequel to EWD912, i.e. to Draft Ch. 5. Circulated privately., April 1985. URL http://www.cs.

utexas.edu/users/EWD/ewd09xx/EWD915.PDF.

Dijkstra:1985:ST

[Dij85-38] Edsger W. Dijkstra. A simple theorem? Circulated privately.,

November 1985. URL http://www.cs.utexas.edu/users/
EWD/ewd09xx/EWD945.PDF.

Dijkstra:1985:SSLb

[Dij85-39] Edsger W. Dijkstra. Some simple lemmata on incremental sorting. Circulated privately., September 1985. URL http:/

/www.cs.utexas.edu/users/EWD/ewd09xx/EWD932b.PDF.

Dijkstra:1985:SST

[Dij85-40] Edsger W. Dijkstra. Some simple theorems on incremental sorting. Circulated privately., September 1985. URL http:/

/www.cs.utexas.edu/users/EWD/ewd09xx/EWD932.PDF.

Dijkstra:1985:SMAa

[Dij85-41] Edsger W. Dijkstra. The streamlining of the mathematical argument. Circulated privately., August 1985. URL http:/

/www.cs.utexas.edu/users/EWD/ewd09xx/EWD932a.PDF.

Dijkstra:1985:SMAb

[Dij85-42] Edsger W. Dijkstra. The streamlining of the mathematical argument. Circulated privately., 1985. URL http://www.cs.

utexas.edu/users/EWD/ewd09xx/EWD932c.PDF.

Dijkstra:1985:SYI

[Dij85-43] Edsger W. Dijkstra. A summary of a year’s impressions. Circulated privately., April 1985. URL http://www.cs.utexas.

edu/users/EWD/ewd09xx/EWD918.PDF.

Dijkstra:1985:ARA

[Dij85-44] Edsger W. Dijkstra. To an anonymous referee of AvG45/EWD901. Circulated privately., September 1985. URL http://www.cs.

utexas.edu/users/EWD/ewd09xx/EWD938.PDF.

Dijkstra:1985:TRWe

[Dij85-45] Edsger W. Dijkstra. Trip report E. W. Dijkstra, Ithaca, Newport, 30 May–13 June 1985. Circulated privately.,

June 1985. URL http://www.cs.utexas.edu/users/EWD/
ewd09xx/EWD923.PDF.

Dijkstra:1985:TRWf

[Dij85-46] Edsger W. Dijkstra. Trip report E. W. Dijkstra, London, 25– 27 June 1985. Circulated privately., June 1985. URL http:

//www.cs.utexas.edu/users/EWD/ewd09xx/EWD925.PDF.

Dijkstra:1985:TRWb

[Dij85-47] Edsger W. Dijkstra. Trip report E. W. Dijkstra, Netherlands, Austria, 23–31 Jan.1985. Circulated privately., Febru-

ary 1985. URL http://www.cs.utexas.edu/users/EWD/
ewd09xx/EWD907.PDF.

Dijkstra:1985:TRWa

[Dij85-48] Edsger W. Dijkstra. Trip report E. W. Dijkstra, Nuenen & London, 10 Dec.’84–10 J an.’85. Circulated privately., Jan-

uary 1985. URL http://www.cs.utexas.edu/users/EWD/
ewd09xx/EWD906.PDF.

Dijkstra:1985:TRWd

[Dij85-49] Edsger W. Dijkstra. Trip report E. W. Dijkstra, Tulsa, 1–6 March 1985. Circulated privately., March 1985. URL http:

//www.cs.utexas.edu/users/EWD/ewd09xx/EWD911.PDF.

Dijkstra:1985:TRWc

[Dij85-50] Edsger W. Dijkstra. Trip report E. W. Dijkstra, Westboro-Boston, 14–17 Feb. 1985. Circulated privately., Febru-

ary 1985. URL http://www.cs.utexas.edu/users/EWD/
ewd09xx/EWD909.PDF.

Dijkstra:1985:WMB

[Dij85-51] Edsger W. Dijkstra. What I missed in Batcher’s Baffler (EWD935a-0). Circulated privately., September 1985. URL

http://www.cs.utexas.edu/users/EWD/ewd09xx/EWD940.
PDF.

Dijkstra:1985:WRP

[Dij85-52] Edsger W. Dijkstra. Where is Russell’s “Paradox”? Circulated privately., May 1985. URL http://www.cs.utexas.

edu/users/EWD/ewd09xx/EWD923a.PDF.

Dijkstra:1986:W

[Dij86a] E. W. Dijkstra. Wiskundefreaks. NRC Handelsblad, 26 Juli, 1986. Ingezonden brief.

Dijkstra:1986:XXX

[Dij86b] Edsger W. Dijkstra. (∃y: : (∀x: : p.x.y)) ⇒ (∀x(∃y: : p.x.y)). Circulated privately., September 1986. URL http://www.

cs.utexas.edu/users/EWD/ewd09xx/EWD972.PDF.

Dijkstra:1986:AMSa

[Dij86c] Edsger W. Dijkstra. Address to my students. Circulated

privately., April 1986. URL http://www.cs.utexas.edu/
users/EWD/ewd09xx/EWD956.PDF.

Dijkstra:1986:AMSb

[Dij86d] Edsger W. Dijkstra. An address to my students (16.9.1986). Circulated privately., September 1986. URL http://www.cs.

utexas.edu/users/EWD/ewd09xx/EWD977.PDF.

Dijkstra:1986:AMSc

[Dij86e] Edsger W. Dijkstra. Address to my students (23.9.1986). Circulated privately., September 1986. URL http://www.

cs.utexas.edu/users/EWD/ewd09xx/EWD981.PDF.

Dijkstra:1986:BPSa

[Dij86f] Edsger W. Dijkstra. A belated proof of self-stabilization. Distributed Computing, 1(1):5–6, 1986.

Dijkstra:1986:CR

[Dij86g] Edsger W. Dijkstra. The cosine rule. Circulated privately.,

September 1986. URL http://www.cs.utexas.edu/users/
EWD/ewd09xx/EWD973.PDF.

Dijkstra:1986:CJM

[Dij86h] Edsger W. Dijkstra. Courtesy A. J. M. van Gasteren, C. S. Scholten and J. G. Wiltink. Circulated privately., Decem-

ber 1986. URL http://www.cs.utexas.edu/users/EWD/
ewd09xx/EWD996.PDF.

Dijkstra:1986:DE

[Dij86i] Edsger W. Dijkstra. Difficult is easy. Circulated privately.,

November 1986. URL http://www.cs.utexas.edu/users/
EWD/ewd09xx/EWD992.PDF.

Dijkstra:1986:ESE

[Dij86j] Edsger W. Dijkstra. Extreme solutions of equations. Chapter 8; see EWD 912 for an earlier version; published as [?].,

September 1986. URL http://www.cs.utexas.edu/users/
EWD/ewd09xx/EWD969.PDF.

Dijkstra:1986:FBC

[Dij86k] Edsger W. Dijkstra. F. L. Bauer’s Conjecture is F. L. Bauer’s Theorem. Circulated privately., August 1986. URL http://

www.cs.utexas.edu/users/EWD/ewd09xx/EWD970.PDF.

Dijkstra:1986:HEBa

[Dij86l] Edsger W. Dijkstra. A heuristic explanation of Batcher’s Baffler. Circulated privately., February 1986. URL http://www.

cs.utexas.edu/users/EWD/ewd09xx/EWD953.PDF.

Dijkstra:1986:HEBb

[Dij86m] Edsger W. Dijkstra. A heuristic explanation of Batcher’s Baffler. Published as [Dij87i]., 1986. URL http://www.cs.

utexas.edu/users/EWD/ewd09xx/EWD953a.PDF.

Dijkstra:1986:HDW

[Dij86n] Edsger W. Dijkstra. How do we contribute? (30.9.1986). Circulated privately., September 1986. URL http://www.

cs.utexas.edu/users/EWD/ewd09xx/EWD983.PDF.

Dijkstra:1986:HEC

[Dij86o] Edsger W. Dijkstra. How experimental is computing science? Circulated privately., October 1986. URL http://www.cs.

utexas.edu/users/EWD/ewd09xx/EWD988a.PDF.

Dijkstra:1986:IRA

[Dij86p] Edsger W. Dijkstra. Idempotence and reflexivity; associativity and transitivity; distribution and monotonicity. Circulated privately., February 1986. URL http://www.cs.

utexas.edu/users/EWD/ewd09xx/EWD954.PDF.

Dijkstra:1986:ICM

[Dij86q] Edsger W. Dijkstra. Introducing a course on mathematical methodology. Circulated privately., June 1986. URL http:/

/www.cs.utexas.edu/users/EWD/ewd09xx/EWD962.PDF.

Dijkstra:1986:LSO

[Dij86r] Edsger W. Dijkstra. The longer side is opposite to the greater angle. Circulated privately., September 1986. URL http://

www.cs.utexas.edu/users/EWD/ewd09xx/EWD974.PDF.

Dijkstra:1986:MM

[Dij86s] Edsger W. Dijkstra. Management and Mathemetics. Circulated privately., June 1986. URL http://www.cs.utexas.

edu/users/EWD/ewd09xx/EWD966.PDF.

Dijkstra:1986:NMR

[Dij86t] Edsger W. Dijkstra. The nature of my research and why I do it. Circulated privately., November 1986. URL http://www.

cs.utexas.edu/users/EWD/ewd09xx/EWD993.PDF.

Dijkstra:1986:NDS

[Dij86u] Edsger W. Dijkstra. Node degree and the size of disconnected subgraphs. Circulated privately., June 1986. URL http://

www.cs.utexas.edu/users/EWD/ewd09xx/EWD964.PDF.

Dijkstra:1986:CGMa

[Dij86v] Edsger W. Dijkstra. On a class of graphs with modest diameter. Circulated privately., October 1986. URL http://www.

cs.utexas.edu/users/EWD/ewd09xx/EWD987a.PDF.

Dijkstra:1986:CGMb

[Dij86w] Edsger W. Dijkstra. On a class of graphs with modest diameter. Circulated privately., October 1986. URL http://www.

cs.utexas.edu/users/EWD/ewd09xx/EWD987.PDF.

Dijkstra:1986:CGa

[Dij86x] Edsger W. Dijkstra. On a cultural gap. The Mathematical Intelligencer, 8(1):48–52, 1986. CODEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic).

Dijkstra:1986:PLP

[Dij86y] Edsger W. Dijkstra. On a proof I learned from prof. dr. J. Haantjes. Circulated privately., May 1986. URL http://

www.cs.utexas.edu/users/EWD/ewd09xx/EWD960.PDF.

Dijkstra:1986:SRN

[Dij86z] Edsger W. Dijkstra. On substitution, replacement, and the notion of a function. Chapters 2 and 3. Published as [?].,

April 1986. URL http://www.cs.utexas.edu/users/EWD/
ewd09xx/EWD949.PDF.

Dijkstra:1986:PSD

[Dij86-27] Edsger W. Dijkstra. On the phenomenon of scientific disciplines. Circulated privately., October 1986. URL http://

www.cs.utexas.edu/users/EWD/ewd09xx/EWD988.PDF.

Dijkstra:1986:QI

[Dij86-28] Edsger W. Dijkstra. On the “quadrangular” inequalities. Circulated privately., May 1986. URL http://www.cs.utexas.

edu/users/EWD/ewd09xx/EWD961.PDF.

Dijkstra:1986:TP

[Dij86-29] Edsger W. Dijkstra. On the theorem of Pythagoras. Circulated privately., 1986. URL http://www.cs.utexas.edu/

users/EWD/ewd09xx/EWD975.PDF.

Dijkstra:1986:WHJ

[Dij86-30] Edsger W. Dijkstra. On W. H. J. Feijen’s string equation. Circulated privately., December 1986. URL http://www.cs.

utexas.edu/users/EWD/ewd09xx/EWD995.PDF.

Dijkstra:1986:PPP

[Dij86-31] Edsger W. Dijkstra. Partitioning predicates and punctual functions (see EWD955). Circulated privately.,

May 1986. URL http://www.cs.utexas.edu/users/EWD/
ewd09xx/EWD957.PDF.

Dijkstra:1986:PPS

[Dij86-32] Edsger W. Dijkstra. Partitioning predicates and substitution; diabolical and angelical nondeterminism. Circulated

privately., March 1986. URL http://www.cs.utexas.edu/
users/EWD/ewd09xx/EWD955.PDF.

Dijkstra:1986:PGT

[Dij86-33] Edsger W. Dijkstra. Proving Gupta’s Theorem. Circulated privately., October 1986. URL http://www.cs.utexas.edu/

users/EWD/ewd09xx/EWD989.PDF.

Dijkstra:1986:PBB

[Dij86-34] Edsger W. Dijkstra. Ptolemaeus and Brahmagupta (or: Baffled by Symmetry). Circulated privately., August 1986. URL

http://www.cs.utexas.edu/users/EWD/ewd09xx/EWD965.
PDF.

Dijkstra:1986:RCA

[Dij86-35] Edsger W. Dijkstra. Relational Calculus according to ATAC. Circulated privately., September 1986. URL http://www.cs.

utexas.edu/users/EWD/ewd09xx/EWD982.PDF.

Dijkstra:1986:SFS

[Dij86-36] Edsger W. Dijkstra. Science fiction and science reality in computing. Circulated privately., 1986. URL http://www.

cs.utexas.edu/users/EWD/ewd09xx/EWD952.PDF.

Dijkstra:1986:SEa

[Dij86-37] Edsger W. Dijkstra. A sequel to EWD977 (18.9.1986). Circulated privately., September 1986. URL http://www.cs.

utexas.edu/users/EWD/ewd09xx/EWD978.PDF.

Dijkstra:1986:SNU

[Dij86-38] Edsger W. Dijkstra. A short note on universal quantification. Circulated privately., May 1986. URL http://www.cs.

utexas.edu/users/EWD/ewd09xx/EWD959.PDF.

Dijkstra:1986:SDB

[Dij86-39] Edsger W. Dijkstra. A solution designed by A. Blokhuis. Circulated privately., September 1986. URL http://www.

cs.utexas.edu/users/EWD/ewd09xx/EWD979.PDF.

Dijkstra:1986:SCP

[Dij86-40] Edsger W. Dijkstra. The strange case of The Pigeon-hole Principle. Circulated privately., September 1986. URL http:

//www.cs.utexas.edu/users/EWD/ewd09xx/EWD980.PDF.

Dijkstra:1986:SEb

[Dij86-41] Edsger W. Dijkstra. A supplement to EWD878. Circulated privately., October 1986. URL http://www.cs.utexas.edu/

users/EWD/ewd09xx/EWD985.PDF.

Dijkstra:1986:STD

[Dij86-42] Edsger W. Dijkstra. Suppose they differ! Circulated pri-

vately., October 1986. URL http://www.cs.utexas.edu/
users/EWD/ewd09xx/EWD986.PDF.

Dijkstra:1986:TCB

[Dij86-43] Edsger W. Dijkstra. A theorem of Charles Babbage’s extended. Circulated privately., October 1986. URL http://

www.cs.utexas.edu/users/EWD/ewd09xx/EWD990.PDF.

Dijkstra:1986:TRWa

[Dij86-44] Edsger W. Dijkstra. Trip report E. W. Dijkstra, Europe, 13.12.1985–10.1.1986. Circulated privately., Jan-

uary 1986. URL http://www.cs.utexas.edu/users/EWD/
ewd09xx/EWD951.PDF.

Dijkstra:1986:TRWb

[Dij86-45] Edsger W. Dijkstra. Trip report E. W. Dijkstra, Marktoberdorf, 29 July–10 Aug 1986. Circulated privately., Au-

gust 1986. URL http://www.cs.utexas.edu/users/EWD/
ewd09xx/EWD968.PDF.

Dijkstra:1986:VBV

[Dij86-46] Edsger W. Dijkstra. Visuals for BP’s Venture Research Conference. Circulated privately., June 1986. URL http://www.

cs.utexas.edu/users/EWD/ewd09xx/EWD963.PDF.

Dijkstra:1986:WYT

[Dij86-47] Edsger W. Dijkstra. Who is your “target audience”? Circulated privately., September 1986. URL http://www.cs.

utexas.edu/users/EWD/ewd09xx/EWD976.PDF.

Dijkstra:1986:WJC

[Dij86-48] Edsger W. Dijkstra. Why Johnny can’t understand. Circulated privately., November 1986. URL http://www.cs.

utexas.edu/users/EWD/ewd09xx/EWD991.PDF.

Dijkstra:1987:AE

[Dij87a] Edsger W. Dijkstra. Addendum to EWD969. Circulated

privately., April 1987. URL http://www.cs.utexas.edu/
users/EWD/ewd10xx/EWD1007.PDF.

Dijkstra:1987:CBSa

[Dij87b] Edsger W. Dijkstra. The calculus of boolean structures (Part 0). very early draft of part of Chapter 5; circulated privately,

February 1987. URL http://www.cs.utexas.edu/users/
EWD/ewd10xx/EWD1001.PDF.

Dijkstra:1987:CBSc

[Dij87c] Edsger W. Dijkstra. The calculus of boolean structures (Part 1). Early draft of the second part of Chapter 5; published as

[Dij90c]., March 1987. URL http://www.cs.utexas.edu/
users/EWD/ewd10xx/EWD1002.PDF.

Dijkstra:1987:CSU

[Dij87d] Edsger W. Dijkstra. Computing Science in the United Kingdom. Circulated privately., January 1987. URL http://www.

cs.utexas.edu/users/EWD/ewd09xx/EWD997.PDF.

Dijkstra:1987:ESE

[Dij87e] Edsger W. Dijkstra. Extreme solutions of equations. In Logic of programming and calculi of discrete design (Marktoberdorf, 1986), volume 36 of NATO Adv. Sci. Inst. Ser. F Comput. Systems Sci., pages 243–260. Springer, Berlin, 1987.

Dijkstra:1987:FYI

[Dij87f] Edsger W. Dijkstra. Fillers at a YoP-institute. Published as [Dij90h]., December 1987. URL http://www.cs.utexas.

edu/users/EWD/ewd10xx/EWD1017.PDF.

Dijkstra:1987:GMD

[Dij87g] Edsger W. Dijkstra. Graphs of modest diameter and degrees. Circulated privately., February 1987. URL http://www.cs.

utexas.edu/users/EWD/ewd10xx/EWD1003.PDF.

Dijkstra:1987:HEBa

[Dij87h] Edsger W. Dijkstra. A heuristic explanation of Batcher’s baffler. In Logic of programming and calculi of discrete design (Marktoberdorf, 1986), volume 36 of NATO Adv. Sci. Inst. Ser. F Comput. Systems Sci., pages 43–49. Springer, Berlin, 1987.

Dijkstra:1987:HEBb

[Dij87i] Edsger W. Dijkstra. A heuristic explanation of Batcher’s Baffler. Science of Computer Programming, 9(3):213–220, December 1987. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Dijkstra:1987:HP

[Dij87j] Edsger W. Dijkstra. The humble programmer. In Ashenhurst [Ash87], page ?? ISBN 0-201-07794-9. LCCN QA76.24 .A33 1987. ACM Turing Award lecture.

Dijkstra:1987:IMF

[Dij87k] Edsger W. Dijkstra. Introducing my Fall 1987 course on Mathematical Methodology. Circulated privately., Au-

gust 1987. URL http://www.cs.utexas.edu/users/EWD/
ewd10xx/EWD1011.PDF.

Dijkstra:1987:JMC

[Dij87l] Edsger W. Dijkstra. J. Misra’s comment on EWD964. Circulated privately., January 1987. URL http://www.cs.

utexas.edu/users/EWD/ewd09xx/EWD994.PDF.

Dijkstra:1987:MRM

[Dij87m] Edsger W. Dijkstra. A methodological remark on mathematical induction. Circulated privately., March 1987. URL http:

//www.cs.utexas.edu/users/EWD/ewd10xx/EWD1004.PDF.

Dijkstra:1987:MAR

[Dij87n] Edsger W. Dijkstra. A monotonicity argument revisited. Circulated privately., October 1987. URL http://www.cs.

utexas.edu/users/EWD/ewd10xx/EWD1014.PDF.

Dijkstra:1987:MNM

[Dij87o] Edsger W. Dijkstra. A much neglected mathematical object. Circulated privately., August 1987. URL http://www.cs.

utexas.edu/users/EWD/ewd10xx/EWD1010.PDF.

Dijkstra:1987:PTD

[Dij87p] Edsger W. Dijkstra. On a problem transmitted by Doug McIlroy. Circulated privately., November 1987. URL http://www.

cs.utexas.edu/users/EWD/ewd10xx/EWD1015.PDF.

Dijkstra:1987:SDC

[Dij87q] Edsger W. Dijkstra. On a somewhat disappointing correspondence. Circulated privately., May 1987. URL http://www.

cs.utexas.edu/users/EWD/ewd10xx/EWD1009.PDF.

Dijkstra:1987:BOT

[Dij87r] Edsger W. Dijkstra. On binary operators and their derived relations. Published as [Dij88l]., April 1987. URL http://

www.cs.utexas.edu/users/EWD/ewd09xx/EWD954a.PDF.

Dijkstra:1987:OPF

[Dij87s] Edsger W. Dijkstra. Our proof format. Early draft of Chapter 4. Published as [?]., January 1987. URL http://www.cs.

utexas.edu/users/EWD/ewd09xx/EWD999.PDF.

Dijkstra:1987:PPF

[Dij87t] Edsger W. Dijkstra. Position paper on “fairness”. Circulated privately., October 1987. URL http://www.cs.utexas.edu/

users/EWD/ewd10xx/EWD1013.PDF.

Dijkstra:1987:SSV

[Dij87u] Edsger W. Dijkstra. Shmuel Safra’s version of termination detection. Circulated privately., January 1987. URL http:/

/www.cs.utexas.edu/users/EWD/ewd09xx/EWD998.PDF.

Dijkstra:1987:SNV

[Dij87v] Edsger W. Dijkstra. The study of a notion, viz. that of residue classes. Circulated privately., January 1987. URL http://

www.cs.utexas.edu/users/EWD/ewd09xx/EWD984.PDF.

Dijkstra:1987:SE

[Dij87w] Edsger W. Dijkstra. A supplement to EWD1002. Circulated

privately., March 1987. URL http://www.cs.utexas.edu/
users/EWD/ewd10xx/EWD1005.PDF.

Dijkstra:1987:MBC

[Dij87x] Edsger W. Dijkstra. To the members of the Budget Council. Circulated privately., December 1987. URL http://www.cs.

utexas.edu/users/EWD/ewd10xx/EWD1018.PDF.

Dijkstra:1987:TRW

[Dij87y] Edsger W. Dijkstra. Trip report E. W. Dijkstra, Stanford/Palo Alto, 26–29 March 1987. Circulated privately.,

March 1987. URL http://www.cs.utexas.edu/users/EWD/
ewd10xx/EWD1006.PDF.

Dijkstra:1987:TEY

[Dij87z] Edsger W. Dijkstra. Twenty-eight years. Circulated pri-

vately., January 1987. URL http://www.cs.utexas.edu/
users/EWD/ewd10xx/EWD1000.PDF.

Dijkstra:1987:WCS

[Dij87-27] Edsger W. Dijkstra. What computing science is about. Circulated privately., May 1987. URL http://www.cs.utexas.

edu/users/EWD/ewd10xx/EWD1008.PDF.

Dijkstra:1988:AFY

[Dij88a] Edsger W. Dijkstra. Another filler of the YoP Institute. Published as [Dij90g]., January 1988. URL http://www.cs.

utexas.edu/users/EWD/ewd10xx/EWD1019.PDF.

Dijkstra:1988:AFW

[Dij88b] Edsger W. Dijkstra. Ascending functions and The Welfare Crook. Circulated privately., October 1988. URL http://

www.cs.utexas.edu/users/EWD/ewd10xx/EWD1031.PDF.

Dijkstra:1988:BIT

[Dij88c] Edsger W. Dijkstra. A bagatelle on implication’s transitivity. Circulated privately., April 1988. URL http://www.cs.

utexas.edu/users/EWD/ewd10xx/EWD1022.PDF.

Dijkstra:1988:CSA

[Dij88d] Edsger W. Dijkstra. A computing scientist’s approach to a once-deep theorem of Sylvester’s. Circulated privately.,

February 1988. URL http://www.cs.utexas.edu/users/
EWD/ewd10xx/EWD1016.PDF.

Dijkstra:1988:DIS

[Dij88e] Edsger W. Dijkstra. Decomposing an integer as sum of two squares. Circulated privately., October 1988. URL http://

www.cs.utexas.edu/users/EWD/ewd10xx/EWD1032.PDF.

Dijkstra:1988:DPJ

[Dij88f] Edsger W. Dijkstra. The derivation of a proof by J. C. S. P. van der Woude. Published as [Dij90d]., Jan-

uary 1988. URL http://www.cs.utexas.edu/users/EWD/
ewd10xx/EWD1020.PDF.

Dijkstra:1988:EVM

[Dij88g] Edsger W. Dijkstra. Equivalence versus mutual implication; junctivity and monotonicity. Circulated privately., Oc-

tober 1988. URL http://www.cs.utexas.edu/users/EWD/
ewd10xx/EWD1034.PDF.

Dijkstra:1988:ENP

[Dij88h] Edsger W. Dijkstra. Euclid, Netty, and the prime numbers. Circulated privately., September 1988. URL http://www.cs.

utexas.edu/users/EWD/ewd10xx/EWD1028.PDF.

Dijkstra:1988:MMB

[Dij88i] Edsger W. Dijkstra. My methodological blunder with grid polygons. Circulated privately., August 1988. URL http://

www.cs.utexas.edu/users/EWD/ewd10xx/EWD1027.PDF.

Dijkstra:1988:NSBa

[Dij88j] Edsger W. Dijkstra. A new science, from birth to maturity. Published as [Dij88k]., October 1988. URL http://www.cs.

utexas.edu/users/EWD/ewd10xx/EWD1024.PDF.

Dijkstra:1988:NSBb

[Dij88k] Edsger W. Dijkstra. A new science, from birth to maturity. In ETH Zürich: 20-year Anniversary of Institut für Informatik. ETH Zürich, Institut für Informatik, Zürich, Switzerland, October 1988.

Dijkstra:1988:BOTa

[Dij88l] Edsger W. Dijkstra. On binary operators and their derived relations. BIT (Nordisk tidskrift for informationsbehandling), 28(3):378–382, September 1988. CODEN BITTEL, NBITAB.

ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&
issn=0006-3835&volume=28&issue=3&spage=377.

Dijkstra:1988:CRT

[Dij88m] Edsger W. Dijkstra. On the cruelty of really teaching computing science. Circulated privately., December 1988. URL http:

//www.cs.utexas.edu/users/EWD/ewd10xx/EWD1036.PDF.

Dijkstra:1988:TBN

[Dij88n] Edsger W. Dijkstra. On triangles being nondegenerate. Circulated privately., June 1988. URL http://www.cs.utexas.

edu/users/EWD/ewd10xx/EWD1023.PDF.

Dijkstra:1988:PPF

[Dij88o] Edsger W. Dijkstra. Position paper on “fairness”. ACM SIG-SOFT Software Engineering Notes, 13(2):18–20, April 1988. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). See comments [SL88, CM88].

Dijkstra:1988:RMD

[Dij88p] Edsger W. Dijkstra. “Real mathematicians don’t prove”. Circulated privately., January 1988. URL http://www.cs.

utexas.edu/users/EWD/ewd10xx/EWD1012.PDF.

Dijkstra:1988:SST

[Dij88q] Edsger W. Dijkstra. A symmetry in the sorting task. Circulated privately., February 1988. URL http://www.cs.

utexas.edu/users/EWD/ewd10xx/EWD1021.PDF.

Dijkstra:1988:TRWa

[Dij88r] Edsger W. Dijkstra. Trip report E. W. Dijkstra, Marktoberdorf 1988. Circulated privately., August 1988. URL http://

www.cs.utexas.edu/users/EWD/ewd10xx/EWD1026.PDF.

Dijkstra:1988:TRWb

[Dij88s] Edsger W. Dijkstra. Trip report E. W. Dijkstra, Zürich, 15 – 20 Oct. 1988. Circulated privately., October 1988. URL http:

//www.cs.utexas.edu/users/EWD/ewd10xx/EWD1033.PDF.

Dijkstra:1988:UBS

[Dij88t] Edsger W. Dijkstra. Ulrich Berger’s solution to the rectangle problem. Circulated privately., August 1988. URL http://

www.cs.utexas.edu/users/EWD/ewd10xx/EWD1025.PDF.

Dijkstra:1988:ZJI

[Dij88u] Edsger W. Dijkstra. Zwanzig Jahre Institut für Informatik. (German) [Twenty years of the Institute for Informatics]: A new science, from birth to maturity. Report, ETH Zürich, Zürich, Switzerland, October 16, 1988. 25–38

pp. URL http://pascal.hansotten.com/uploads/wirth/
Zwanzig%20Jahre%20Institut%20Informatik%20095.pdf.

Dijkstra:1989:PMM

[Dij89a] Edsger W. Dijkstra. 0 Preface (Mathematical Methodology). Circulated privately., September 1989. URL http://www.cs.

utexas.edu/users/EWD/ewd10xx/EWD1059.PDF.

Dijkstra:1989:APE

[Dij89b] Edsger W. Dijkstra. Andrei P. Ershov in Nuenen. Circulated

privately., August 1989. URL http://www.cs.utexas.edu/
users/EWD/ewd10xx/EWD1057.PDF.

Dijkstra:1989:APS

[Dij89c] Edsger W. Dijkstra. ATAC’s proof of Scholten’s theorem. Circulated privately., March 1989. URL http://www.cs.

utexas.edu/users/EWD/ewd10xx/EWD1048.PDF.

Dijkstra:1989:BTO

[Dij89d] Edsger W. Dijkstra. A bagatelle on total orders. Circulated privately., January 1989. URL http://www.cs.utexas.edu/

users/EWD/ewd10xx/EWD1038.PDF.

Dijkstra:1989:WI

[Dij89e] Edsger W. Dijkstra. By way of introduction. Circulated pri-

vately., February 1989. URL http://www.cs.utexas.edu/
users/EWD/ewd10xx/EWD1041.PDF.

Dijkstra:1989:CLC

[Dij89f] Edsger W. Dijkstra. Composition, lambda-calculus, and some more. Circulated privately., September 1989. URL http://

www.cs.utexas.edu/users/EWD/ewd10xx/EWD1061.PDF.

Dijkstra:1989:CSA

[Dij89g] Edsger W. Dijkstra. A computing scientist’s approach to a once-deep theorem of Sylvester’s. In Constructive methods in computing science (Marktoberdorf, 1988), volume 55 of NATO Adv. Sci. Inst. Ser. F Comput. Systems Sci., pages 223–232. Springer, Berlin, 1989.

Dijkstra:1989:DPJ

[Dij89h] Edsger W. Dijkstra. The derivation of a proof by J. C. S. P. van der Woude. In Constructive methods in computing science (Marktoberdorf, 1988), volume 55 of NATO Adv. Sci. Inst. Ser. F Comput. Systems Sci., pages 233–241. Springer, Berlin, 1989.

Dijkstra:1989:DRC

[Dij89i] Edsger W. Dijkstra. Dijkstra’s reply to comments. Communications of the Association for Computing Machinery, 32(12): 1414, December 1989. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dijkstra:1989:EHL

[Dij89j] Edsger W. Dijkstra. An example of how to lengthen formal proofs. Circulated privately., October 1989. URL http://

www.cs.utexas.edu/users/EWD/ewd10xx/EWD1065.PDF.

Dijkstra:1989:ELC

[Dij89k] Edsger W. Dijkstra. Exploring the lexical coupling. Circulated privately., January 1989. URL http://www.cs.

utexas.edu/users/EWD/ewd10xx/EWD1037.PDF.

Dijkstra:1989:FF

[Dij89l] Edsger W. Dijkstra. Factorizing the factorial. Circulated

privately., 1989. URL http://www.cs.utexas.edu/users/
EWD/ewd10xx/EWD1041a.PDF.

Dijkstra:1989:FGB

[Dij89m] Edsger W. Dijkstra. Fair gambling with a biased coin. Circulated privately., October 1989. URL http://www.cs.

utexas.edu/users/EWD/ewd10xx/EWD1069.PDF.

Dijkstra:1989:FCC

[Dij89n] Edsger W. Dijkstra. A few comments on “Computing as a discipline”. Circulated privately., February 1989. URL http:

//www.cs.utexas.edu/users/EWD/ewd10xx/EWD1043.PDF.

Dijkstra:1989:BSM

[Dij89o] Edsger W. Dijkstra. For brevity’s sake (Mathematical Methodology). Circulated privately., November 1989. URL

http://www.cs.utexas.edu/users/EWD/ewd10xx/EWD1070.
PDF.

Dijkstra:1989:FDP

[Dij89p] Edsger W. Dijkstra, editor. Formal Development of Programs and Proofs. Addison-Wesley, Reading, MA, USA, 1989. ISBN 0-201-17237-2. xiv + 242 pp. LCCN QA76.6

.F657 1990. URL http://www.gbv.de/dms/bowker/toc/
9780201172379.pdf;   http://www.zentralblatt-math.
org/zmath/en/search/?an=0734.68063. Prepared with

LaTEX.

Dijkstra:1989:HEC

[Dij89q] Edsger W. Dijkstra. From the Hungarian Eötvös Competition, 1911. Circulated privately., June 1989. URL http://

www.cs.utexas.edu/users/EWD/ewd10xx/EWD1052.PDF.

Dijkstra:1989:HOD

[Dij89r] Edsger W. Dijkstra. Hoe onbelangkrijk het is of duikboten kunnen zwemmen. (Dutch) [how unimportant is it that submarines can swim?]. Circulated privately.,

July 1989. URL http://www.cs.utexas.edu/users/EWD/
ewd10xx/EWD1056.PDF.

Dijkstra:1989:HBL

[Dij89s] Edsger W. Dijkstra. Hungry? Have a byte . . .. Circulated privately., October 1989. URL http://www.cs.utexas.edu/

users/EWD/ewd10xx/EWD1066.PDF.

Dijkstra:1989:RC

[Dij89t] Edsger W. Dijkstra. In reply to comments. Circulated

privately., 1989. URL http://www.cs.utexas.edu/users/
EWD/ewd10xx/EWD1058.PDF.

Dijkstra:1989:IEM

[Dij89u] Edsger W. Dijkstra. An introductory example (Mathematical Methodology). Circulated privately., October 1989. URL

http://www.cs.utexas.edu/users/EWD/ewd10xx/EWD1063.
PDF.

Dijkstra:1989:MVA

[Dij89v] Edsger W. Dijkstra. The majority vote according to J. Gutknecht. Circulated privately., January 1989. URL http:

//www.cs.utexas.edu/users/EWD/ewd10xx/EWD1040.PDF.

Dijkstra:1989:MFRa

[Dij89w] Edsger W. Dijkstra. Making a fair roulette from a possibly biased coin. Published as [Dij90n]., Novem-

ber 1989. URL http://www.cs.utexas.edu/users/EWD/
ewd10xx/EWD1071a.PDF.

Dijkstra:1989:MFRb

[Dij89x] Edsger W. Dijkstra. Making a fair roulette from a possibly biased coin. Circulated privately., November 1989. URL http:

//www.cs.utexas.edu/users/EWD/ewd10xx/EWD1071.PDF.

Dijkstra:1989:MPT

[Dij89y] Edsger W. Dijkstra. Monochrome pairs in the three-coloured plane. Circulated privately., June 1989. URL http://www.

cs.utexas.edu/users/EWD/ewd10xx/EWD1053.PDF.

Dijkstra:1989:NFY

[Dij89z] Edsger W. Dijkstra. The next forty years. Circulated

privately., June 1989. URL http://www.cs.utexas.edu/
users/EWD/ewd10xx/EWD1051.PDF.

Dijkstra:1989:PTD

[Dij89-27] Edsger W. Dijkstra. On a problem transmitted by Doug McIlroy. In Constructive methods in computing science (Marktoberdorf, 1988), volume 55 of NATO Adv. Sci. Inst. Ser. F Comput. Systems Sci., pages 219–222. Springer, Berlin, 1989.

Dijkstra:1989:ETH

[Dij89-28] Edsger W. Dijkstra. On an exercise of Tony Hoare’s. Circulated privately., September 1989. URL http://www.cs.

utexas.edu/users/EWD/ewd10xx/EWD1062.PDF.

Dijkstra:1989:CFDa

[Dij89-29] Edsger W. Dijkstra. On covering a figure with diamonds. Circulated privately., October 1989. URL http://www.cs.

utexas.edu/users/EWD/ewd10xx/EWD1055c.PDF.

Dijkstra:1989:CFDb

[Dij89-30] Edsger W. Dijkstra. On covering a figure with diamonds. Circulated privately., December 1989. URL http://www.cs.

utexas.edu/users/EWD/ewd10xx/EWD1072.PDF.

Dijkstra:1989:HIO

[Dij89-31] Edsger W. Dijkstra. On hygiene, intellectual and otherwise. Circulated privately., June 1989. URL http://www.

cs.utexas.edu/users/EWD/ewd10xx/EWD1054.PDF.

Dijkstra:1989:CRT

[Dij89-32] Edsger W. Dijkstra. On the cruelty of really teaching computer science. Communications of the Association for Computing Machinery, 32(12):1398–1404, December 1989. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also responses in letters [BSE^{+}90, WWSD90].

Dijkstra:1989:DSP

[Dij89-33] Edsger W. Dijkstra. On the design of a simple proof for Morley’s Theorem. Circulated privately., April 1989. URL http:

//www.cs.utexas.edu/users/EWD/ewd10xx/EWD1050.PDF.

Dijkstra:1989:PC

[Dij89-34] Edsger W. Dijkstra. On the problem of the calissons. Circulated privately., July 1989. URL http://www.cs.utexas.

edu/users/EWD/ewd10xx/EWD1055.PDF.

Dijkstra:1989:QCM

[Dij89-35] Edsger W. Dijkstra. On the quality criteria for mathematical writing (Mathematical Methodology). Circulated

privately., 1989. URL http://www.cs.utexas.edu/users/
EWD/ewd10xx/EWD1068.PDF.

Dijkstra:1989:SDN

[Dij89-36] Edsger W. Dijkstra. On the selection of dummies: a notational freedom. Circulated privately., October 1989. URL

http://www.cs.utexas.edu/users/EWD/ewd10xx/EWD1064.
PDF.

Dijkstra:1989:OMB

[Dij89-37] Edsger W. Dijkstra. Once more the bichrome 3 × 7 grid (See EWD764). Circulated privately., February 1989. URL http:

//www.cs.utexas.edu/users/EWD/ewd10xx/EWD1042.PDF.

Dijkstra:1989:PCP

[Dij89-38] Edsger W. Dijkstra. “Predicate Calculus and Program Semantics”, Fall 1989. Circulated privately., Septem-

ber 1989. URL http://www.cs.utexas.edu/users/EWD/
ewd10xx/EWD1060.PDF.

Dijkstra:1989:PDD

[Dij89-39] Edsger W. Dijkstra. The problem of the difficult dartboard. Circulated privately., February 1989. URL http://www.cs.

utexas.edu/users/EWD/ewd10xx/EWD1045.PDF.

Dijkstra:1989:SPT

[Dij89-40] Edsger W. Dijkstra. Seemingly on a problem transmitted by Bengt Jonsson. Circulated privately., Jan-

uary 1989. URL http://www.cs.utexas.edu/users/EWD/
ewd10xx/EWD1039.PDF.

Dijkstra:1989:SOL

[Dij89-41] Edsger W. Dijkstra. A somewhat open letter to Uri Leron. Circulated privately., April 1989. URL http://www.cs.

utexas.edu/users/EWD/ewd10xx/EWD1049.PDF.

Dijkstra:1989:TTW

[Dij89-42] Edsger W. Dijkstra. A tale of two winters. Circulated

privately., March 1989. URL http://www.cs.utexas.edu/
users/EWD/ewd10xx/EWD1046.PDF.

Dijkstra:1989:HMI

[Dij89-43] Edsger W. Dijkstra. To hell with “meaningful identifiers”! Circulated privately., February 17, 1989. URL http://www.

cs.utexas.edu/users/EWD/ewd10xx/EWD1044.PDF.

Dijkstra:1989:UAY

[Dij89-44] Edsger W. Dijkstra. Untitled (advice to a young scientist). Circulated privately., 1989. URL http://www.cs.utexas.

edu/users/EWD/ewd10xx/EWD1055A.PDF.

Dijkstra:1989:VFI

[Dij89-45] Edsger W. Dijkstra. A very first introductory example (Mathematical Methodology). Circulated privately., Octo-

ber 1989. URL http://www.cs.utexas.edu/users/EWD/
ewd10xx/EWD1067.PDF.

Dijkstra:1990:MHE

[Dij90a] Albert J. Dijkstra. On the mechanism of the hydrogenation of edible oils. In Feijen et al. [FvGGM90], chapter 11, pages 102–111. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN

QA76 .B326 1990. URL http://www.zentralblatt-math.
org/zmath/en/search/?an=0718.68004. Contains impor-

tant treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

Dijkstra:1990:BC

[Dij90b] Edsger W. Dijkstra. The balance and the coins. Circulated privately., September 1990. URL http://www.cs.utexas.

edu/users/EWD/ewd10xx/EWD1083.PDF.

Dijkstra:1990:CBSb

[Dij90c] Edsger W. Dijkstra. The calculus of boolean structures (Part 1). In Predicate Calculus and Program Semantics [DS90g], chapter 5, pages 62–80. ISBN 0-387-96957-8. early draft of the second part of Chapter 5.

Dijkstra:1990:DPJ

[Dij90d] Edsger W. Dijkstra. The derivation of a proof by J. C. S. P. van der Woude. In Formal Development of Programs and Proofs, chapter 16, pages 201–207. Addison-Wesley, 1990.

Dijkstra:1990:EOO

[Dij90e] Edsger W. Dijkstra. The everywhere operator once more. Circulated privately., November 1990. URL http://www.cs.

utexas.edu/users/EWD/ewd10xx/EWD1086.PDF.

Dijkstra:1990:FGC

[Dij90f] Edsger W. Dijkstra. Fibonacci and the greatest common divisor. Circulated privately., April 1990. URL http://www.

cs.utexas.edu/users/EWD/ewd10xx/EWD1077.PDF.

Dijkstra:1990:FYIa

[Dij90g] Edsger W. Dijkstra. Fillers at a YoP-institute, chapter 17, pages 209–210. Addison-Wesley, 1990.

Dijkstra:1990:FYIb

[Dij90h] Edsger W. Dijkstra. Fillers at a YoP-institute, chapter 17, pages 211–227. Addison-Wesley, 1990.

Dijkstra:1990:BMA

[Dij90i] Edsger W. Dijkstra. From Discrete Mathematics with Applications by Susanna S. Epp. Circulated privately.,

March 1990. URL http://www.cs.utexas.edu/users/EWD/
ewd10xx/EWD1076.PDF.

Dijkstra:1990:GRB

[Dij90j] Edsger W. Dijkstra. A generalization of R. S. Bird’s problem. Circulated privately., February 1990. URL http://www.cs.

utexas.edu/users/EWD/ewd10xx/EWD1074.PDF.

Dijkstra:1990:HCS

[Dij90k] Edsger W. Dijkstra. How Computing Science created a new mathematical style. Published as [Dij91a, pages 897–905].,

March 1990. URL http://www.cs.utexas.edu/users/EWD/
ewd10xx/EWD1073.PDF.

Dijkstra:1990:LPM

[Dij90l] Edsger W. Dijkstra. Leibniz’s Principle (Mathematical Methodology). Circulated privately., April 1990. URL http:

//www.cs.utexas.edu/users/EWD/ewd10xx/EWD1078.PDF.

Dijkstra:1990:LAL

[Dij90m] Edsger W. Dijkstra. A logician’s anomaly or: Leibniz vindicated. Circulated privately., September 1990. URL http://

www.cs.utexas.edu/users/EWD/ewd10xx/EWD1084.PDF.

Dijkstra:1990:MFR

[Dij90n] Edsger W. Dijkstra. Making a fair roulette from a possibly biased coin. Information Processing Letters, 36(4):193, November 15, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dijkstra:1990:MIU

[Dij90o] Edsger W. Dijkstra. Mathematical induction and universal quantification. Circulated privately., July 1990. URL http:

//www.cs.utexas.edu/users/EWD/ewd10xx/EWD1081.PDF.

Dijkstra:1990:PTM

[Dij90p] Edsger W. Dijkstra. Proving the theorem of Menelaos. Circulated privately., October 1990. URL http://www.cs.

utexas.edu/users/EWD/ewd10xx/EWD1085.PDF.

Dijkstra:1990:RS

[Dij90q] Edsger W. Dijkstra. A relational summary. Circulated pri-

vately., November 1990. URL http://www.cs.utexas.edu/
users/EWD/ewd10xx/EWD1047.PDF.

Dijkstra:1990:SSLb

[Dij90r] Edsger W. Dijkstra. Semantics of straight-line programs (Draft Chap.4), chapter 7. Springer, 1990.

Dijkstra:1991:HCS

[Dij91a] E. W. Dijkstra. How computing science created a new mathematical style. Mitteilungen der Mathematischen Gesellschaft in Hamburg, 12(4):897–905 (1992), 1991. CODEN MNG-BAK. ISSN 0340-4358. Mathematische Wissenschaften gestern und heute. 300 Jahre Mathematische Gesellschaft in Hamburg, Teil 4 (Hamburg, 1990).

Dijkstra:1991:SPR

[Dij91b] Edsger W. Dijkstra. Are “systems people” really necessary? Circulated privately., February 1991. URL http://www.cs.

utexas.edu/users/EWD/ewd10xx/EWD1095.PDF.

Dijkstra:1991:CH

[Dij91c] Edsger W. Dijkstra. The computing habit. Circulated

privately., April 1991. URL http://www.cs.utexas.edu/
users/EWD/ewd11xx/EWD1100.PDF.

Dijkstra:1991:CRR

[Dij91d] Edsger W. Dijkstra. A conversion routine revisited. Circulated privately., January 1991. URL http://www.cs.

utexas.edu/users/EWD/ewd10xx/EWD1087.PDF.

Dijkstra:1991:CE

[Dij91e] Edsger W. Dijkstra. A correction of EWD1112. Circulated privately., December 1991. URL http://www.cs.utexas.

edu/users/EWD/ewd11xx/EWD1119.PDF.

Dijkstra:1991:CAE

[Dij91f] Edsger W. Dijkstra. Courtesy Apt, ETAC, Hoogerwoord, & Voermans. Circulated privately., August 1991. URL http:/

/www.cs.utexas.edu/users/EWD/ewd11xx/EWD1106.PDF.

Dijkstra:1991:DPE

[Dij91g] Edsger W. Dijkstra. The design of a proof of equality (Mathematical Methodology). Circulated privately., Jan-

uary 1991. URL http://www.cs.utexas.edu/users/EWD/
ewd10xx/EWD1090.PDF.

Dijkstra:1991:BWS

[Dij91h] Edsger W. Dijkstra. For Bill Walker a sequel to EWD1099. Circulated privately., July 1991. URL http://www.cs.

utexas.edu/users/EWD/ewd11xx/EWD1105.PDF.

Dijkstra:1991:REC

[Dij91i] Edsger W. Dijkstra. For the record: ETAC and the couples. Circulated privately., June 1991. URL http://www.cs.

utexas.edu/users/EWD/ewd11xx/EWD1103.PDF.

Dijkstra:1991:GN

[Dij91j] Edsger W. Dijkstra. Guided by necessity. Circulated pri-

vately., March 1991. URL http://www.cs.utexas.edu/
users/EWD/ewd10xx/EWD1098.PDF.

Dijkstra:1991:IPA

[Dij91k] Edsger W. Dijkstra. Inspired by Peter Auer’s proof. Circulated privately., March 1991. URL http://www.cs.utexas.

edu/users/EWD/ewd10xx/EWD1099.PDF.

Dijkstra:1991:MVA

[Dij91l] Edsger W. Dijkstra. The majority vote among three. Circulated privately., March 1991. URL http://www.cs.utexas.

edu/users/EWD/ewd10xx/EWD1097.PDF.

Dijkstra:1991:MMM

[Dij91m] Edsger W. Dijkstra. Manipulating mathematical macros. Circulated privately., December 1991. URL http://www.cs.

utexas.edu/users/EWD/ewd11xx/EWD1120.PDF.

Dijkstra:1991:MCS

[Dij91n] Edsger W. Dijkstra. A manuscript for the Coahuila student chapter of the ACM. Circulated privately., Novem-

ber 1991. URL http://www.cs.utexas.edu/users/EWD/
ewd11xx/EWD1115.PDF.

Dijkstra:1991:MDD

[Dij91o] Edsger W. Dijkstra. Monotonic demonstranda and dummy introduction. Circulated privately., December 1991. URL

http://www.cs.utexas.edu/users/EWD/ewd11xx/EWD1117.
PDF.

Dijkstra:1991:MEP

[Dij91p] Edsger W. Dijkstra. More equality proofs with instantiated definitions (Mathematical Methodology). Circulated pri-

vately., February 1991. URL http://www.cs.utexas.edu/
users/EWD/ewd10xx/EWD1091.PDF.

Dijkstra:1991:MPR

[Dij91q] Edsger W. Dijkstra. More pointless relational calculus: the transitive closure. Circulated privately., Novem-

ber 1991. URL http://www.cs.utexas.edu/users/EWD/
ewd11xx/EWD1116.PDF.

Dijkstra:1991:FDH

[Dij91r] Edsger W. Dijkstra. On a formula designed by Henk Doornbos. Circulated privately., December 1991. URL http://

www.cs.utexas.edu/users/EWD/ewd11xx/EWD1118.PDF.

Dijkstra:1991:ES

[Dij91s] Edsger W. Dijkstra. On extreme solutions. Circulated pri-

vately., October 1991. URL http://www.cs.utexas.edu/
users/EWD/ewd11xx/EWD1107.PDF.

Dijkstra:1991:OSE

[Dij91t] Edsger W. Dijkstra. An Oxford sequel to EWD1047. Circulated privately., January 1991. URL http://www.cs.

utexas.edu/users/EWD/ewd10xx/EWD1089.PDF.

Dijkstra:1991:PPD

[Dij91u] Edsger W. Dijkstra. Potter’s proof of disjunction’s symmetry. Circulated privately., March 1991. URL http://www.

cs.utexas.edu/users/EWD/ewd10xx/EWD1096.PDF.

Dijkstra:1991:PEI

[Dij91v] Edsger W. Dijkstra. Proving the equality of infinite sequences. Circulated privately., February 1991. URL http:/

/www.cs.utexas.edu/users/EWD/ewd10xx/EWD1093.PDF.

Dijkstra:1991:SPO

[Dij91w] Edsger W. Dijkstra. Simplifying a proof in our book. Circulated privately., September 1991. URL http://www.cs.

utexas.edu/users/EWD/ewd11xx/EWD1109.PDF.

Dijkstra:1991:MSV

[Dij91x] Edsger W. Dijkstra. So much for scientific visualization. Circulated privately., July 1991. URL http://www.cs.utexas.

edu/users/EWD/ewd11xx/EWD1104.PDF.

Dijkstra:1991:SOL

[Dij91y] Edsger W. Dijkstra. A somewhat open letter to Ben Kuipers. Circulated privately., November 1991. URL http://www.cs.

utexas.edu/users/EWD/ewd11xx/EWD1113.PDF.

Dijkstra:1991:TCK

[Dij91z] Edsger W. Dijkstra. A theorem communicated by Ken Calvert. Circulated privately., January 1991. URL http:/

/www.cs.utexas.edu/users/EWD/ewd10xx/EWD1088.PDF.

Dijkstra:1991:TMP

[Dij91-27] Edsger W. Dijkstra. Too much PSST in Texas. Circulated privately., September 1991. URL http://www.cs.utexas.

edu/users/EWD/ewd11xx/EWD1108.PDF.

Dijkstra:1991:TGS

[Dij91-28] Edsger W. Dijkstra. Triangles in graphs: a sharp bound. Circulated privately., May 1991. URL http://www.cs.utexas.

edu/users/EWD/ewd11xx/EWD1101.PDF.

Dijkstra:1991:THL

[Dij91-29] Edsger W. Dijkstra. Triggered by Higman’s Lemma. Circulated privately., October 1991. URL http://www.cs.

utexas.edu/users/EWD/ewd11xx/EWD1111.PDF.

Dijkstra:1991:TWD

[Dij91-30] Edsger W. Dijkstra. Two ways of determining an expectation value. Circulated privately., February 1991. URL http://

www.cs.utexas.edu/users/EWD/ewd10xx/EWD1092.PDF.

Dijkstra:1991:USP

[Dij91-31] Edsger W. Dijkstra. The undeserved status of the pigeon-hole principle (Mathematical Methodology). Circulated privately.,

March 1991. URL http://www.cs.utexas.edu/users/EWD/
ewd10xx/EWD1094.PDF.

Dijkstra:1991:WFR

[Dij91-32] Edsger W. Dijkstra. Well-foundedness and the relational calculus. Circulated privately., November 1991. URL http://

www.cs.utexas.edu/users/EWD/ewd11xx/EWD1114.PDF.

Dijkstra:1991:WPB

[Dij91-33] Edsger W. Dijkstra. Why preorders are beautiful. Circulated privately., June 1991. URL http://www.cs.utexas.

edu/users/EWD/ewd11xx/EWD1102.PDF.

Dijkstra:1991:WE

[Dij91-34] Edsger W. Dijkstra. Why “up to equivalence”. Circulated privately., November 1991. URL http://www.cs.utexas.

edu/users/EWD/ewd11xx/EWD1112.PDF.

Dijkstra:1992:AMG

[Dij92a] Edsger W. Dijkstra. The arithmetic mean and the geometric mean. Circulated privately., October 1992. URL http://

www.cs.utexas.edu/users/EWD/ewd11xx/EWD1140.PDF.

Dijkstra:1992:BTS

[Dij92b] Edsger W. Dijkstra. Bulterman’s theorem on shortest tree. Circulated privately., July 1992. URL http://www.cs.

utexas.edu/users/EWD/ewd11xx/EWD1131.PDF.

Dijkstra:1992:CPT

[Dij92c] Edsger W. Dijkstra. The checkers problem told to me by M. O. Rabin. Circulated privately., September 1992. URL http:

//www.cs.utexas.edu/users/EWD/ewd11xx/EWD1134.PDF.

Dijkstra:1992:CRP

[Dij92d] Edsger W. Dijkstra. A comparison of relational proofs. Circulated privately., October 1992. URL http://www.cs.

utexas.edu/users/EWD/ewd11xx/EWD1136.PDF.

Dijkstra:1992:CF

[Dij92e] Edsger W. Dijkstra. Computing the future? Circulated pri-

vately., October 1992. URL http://www.cs.utexas.edu/
users/EWD/ewd11xx/EWD1137.PDF.

Dijkstra:1992:CRB

[Dij92f] Edsger W. Dijkstra. Covering rectangles with bars of length 3. Circulated privately., January 1992. URL http://www.

cs.utexas.edu/users/EWD/ewd11xx/EWD1121.PDF.

Dijkstra:1992:IA

[Dij92g] Edsger W. Dijkstra. Interleaving is associative. Circulated privately., January 1992. URL http://www.cs.utexas.edu/

users/EWD/ewd11xx/EWD1122.PDF.

Dijkstra:1992:ICD

[Dij92h] Edsger W. Dijkstra. Introducing a course on the design and use of calculi. Circulated privately., August 1992. URL http:

//www.cs.utexas.edu/users/EWD/ewd11xx/EWD1132.PDF.

Dijkstra:1992:IAD

[Dij92i] Edsger W. Dijkstra. It is all distributivity. Circulated pri-

vately., November 1992. URL http://www.cs.utexas.edu/
users/EWD/ewd11xx/EWD1142.PDF.

Dijkstra:1992:KT

[Dij92j] Edsger W. Dijkstra. The knight’s tour. Circulated privately.,

September 1992. URL http://www.cs.utexas.edu/users/
EWD/ewd11xx/EWD1135.PDF.

Dijkstra:1992:MIS

[Dij92k] Edsger W. Dijkstra. Mathematical induction synthesized. Circulated privately., April 1992. URL http://www.cs.

utexas.edu/users/EWD/ewd11xx/EWD1128.PDF.

Dijkstra:1992:MAP

[Dij92l] Edsger W. Dijkstra. More annotated proofs (a sequel to 1143). Circulated privately., November 1992. URL http:

//www.cs.utexas.edu/users/EWD/ewd11xx/EWD1144.PDF.

Dijkstra:1992:MDC

[Dij92m] Edsger W. Dijkstra. A more disentangled characterization of extreme solutions. Circulated privately., Decem-

ber 1992. URL http://www.cs.utexas.edu/users/EWD/
ewd11xx/EWD1148.PDF.

Dijkstra:1992:MMP

[Dij92n] Edsger W. Dijkstra. More on monotonic predicate transformers and the introduction of dummies (see EWD1117). Circulated privately., December 1992. URL http://www.cs.

utexas.edu/users/EWD/ewd11xx/EWD1147.PDF.

Dijkstra:1992:NCR

[Dij92o] Edsger W. Dijkstra. Notational considerations and the relational calculus. Circulated privately., November 1992. URL

http://www.cs.utexas.edu/users/EWD/ewd11xx/EWD1141.
PDF.

Dijkstra:1992:PKT

[Dij92p] Edsger W. Dijkstra. On a proof of Kaplansky’s Theorem. Circulated privately., February 1992. URL http://www.cs.

utexas.edu/users/EWD/ewd11xx/EWD1124.PDF.

Dijkstra:1992:DSP

[Dij92q] Edsger W. Dijkstra. On the design of a simple proof for Morley’s theorem. In Programming and mathematical method (Marktoberdorf, 1990), volume 88 of NATO Adv. Sci. Inst. Ser. F Comput. Systems Sci., pages 3–9. Springer, Berlin, 1992.

Dijkstra:1992:EDMa

[Dij92r] Edsger W. Dijkstra. On the economy of doing mathematics. Circulated privately., June 1992. URL http://www.cs.

utexas.edu/users/EWD/ewd11xx/EWD1130.PDF.

Dijkstra:1992:EDMb

[Dij92s] Edsger W. Dijkstra. On the economy of doing mathematics. In Bird et al. [BMW93], pages 2–10. CODEN LNCSD9. ISBN 3-540-56625-2 (Berlin), 0-387-56625-2 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6

.M367 1993. DM72.00. URL http://link.springer-
ny.com/link/service/series/0558/tocs/t0669.htm;
http://www.springerlink.com/content/978-3-540-56625-
0;   http://www.springerlink.com/openurl.asp?genre=
issue&issn=0302-9743&volume=669. Papers . . . presented

at the Second International Conference on the Mathematics of Program Construction, held at St Catherine’s College, Oxford, during the week of 29 June–3 July, 1992.

Dijkstra:1992:PMP

[Dij92t] Edsger W. Dijkstra. Perry Moerland’s proof of Carel Scholten’s theorem. Circulated privately., December 1992.

URL http://www.cs.utexas.edu/users/EWD/ewd11xx/EWD1149.
PDF.

Dijkstra:1992:RRT

[Dij92u] Edsger W. Dijkstra. Referee’s Report on “A transformational approach to specifying recovery in asynchronous communicating systems” by Zhiming Liu and Mathai Joseph, University of Warwick. Paper No. 40. Circulated privately.,

April 1992. URL http://www.cs.utexas.edu/users/EWD/
ewd11xx/EWD1127.PDF.

Dijkstra:1992:RB

[Dij92v] Edsger W. Dijkstra. A relational bagatelle. Circulated pri-

vately., October 1992. URL http://www.cs.utexas.edu/
users/EWD/ewd11xx/EWD1139.PDF.

Dijkstra:1992:SAP

[Dij92w] Edsger W. Dijkstra. Some annotated proofs. Circulated pri-

vately., November 1992. URL http://www.cs.utexas.edu/
users/EWD/ewd11xx/EWD1143.PDF.

Dijkstra:1992:SOL

[Dij92x] Edsger W. Dijkstra. A somewhat open letter to Cathleen Synge Morawetz. Circulated privately., Novem-

ber 1992. URL http://www.cs.utexas.edu/users/EWD/
ewd11xx/EWD1145.PDF.

Dijkstra:1992:STC

[Dij92y] Edsger W. Dijkstra. A summary for Turner’s class (An extract from EWD1107). Circulated privately., Febru-

ary 1992. URL http://www.cs.utexas.edu/users/EWD/
ewd11xx/EWD1125.PDF.

Dijkstra:1992:TSO

[Dij92z] Edsger W. Dijkstra. A theorem of Scholten’s once more (see EWD1048). Circulated privately., August 1992. URL http:

//www.cs.utexas.edu/users/EWD/ewd11xx/EWD1133.PDF.

Dijkstra:1992:TOJ

[Dij92-27] Edsger W. Dijkstra. Total-order junctivity. Circulated pri-

vately., October 1992. URL http://www.cs.utexas.edu/
users/EWD/ewd11xx/EWD1138.PDF.

Dijkstra:1992:TI

[Dij92-28] Edsger W. Dijkstra. The transitivity of the implication. Circulated privately., December 1992. URL http://www.cs.

utexas.edu/users/EWD/ewd11xx/EWD1146.PDF.

Dijkstra:1992:UTC

[Dij92-29] Edsger W. Dijkstra. The unification of three calculi. Circulated privately., June 1992. URL http://www.cs.utexas.

edu/users/EWD/ewd11xx/EWD1123.PDF.

Dijkstra:1992:WF

[Dij92-30] Edsger W. Dijkstra. Who failed? Circulated privately.,

February 1992. URL http://www.cs.utexas.edu/users/
EWD/ewd11xx/EWD1126.PDF.

Dijkstra:1992:WHW

[Dij92-31] Edsger W. Dijkstra. Why “heavy-weight” bits are unavoidable. Circulated privately., April 1992. URL http://www.

cs.utexas.edu/users/EWD/ewd11xx/EWD1129.PDF.

Dijkstra:1993:BEA

[Dij93a] Edsger W. Dijkstra. A bagatelle on Euclid’s Algorithm. Circulated privately., September 1993. URL http://www.cs.

utexas.edu/users/EWD/ewd11xx/EWD1158.PDF.

Dijkstra:1993:DPD

[Dij93b] Edsger W. Dijkstra. A derivation of a proof by D. Zagier. Circulated privately., August 1993. URL http://www.cs.

utexas.edu/users/EWD/ewd11xx/EWD1154.PDF.

Dijkstra:1993:DDV

[Dij93c] Edsger W. Dijkstra. A detailed derivation of a very simple program. Circulated privately., October 1993. URL http://

www.cs.utexas.edu/users/EWD/ewd11xx/EWD1162.PDF.

Dijkstra:1993:ML

[Dij93d] Edsger W. Dijkstra. “From my life”. Circulated privately.,

November 1993. URL http://www.cs.utexas.edu/users/
EWD/ewd11xx/EWD1166.PDF.

Dijkstra:1993:IGV

[Dij93e] Edsger W. Dijkstra. Het intellectuele gehalte van informatica. (Dutch) [The intellectual content of computer science]. Circulated privately., February 1993. URL http://www.cs.

utexas.edu/users/EWD/ewd11xx/EWD1152.PDF.

Dijkstra:1993:HPF

[Dij93f] Edsger W. Dijkstra. How promiscuous are the French ? Circulated privately., October 1993. URL http://www.cs.

utexas.edu/users/EWD/ewd11xx/EWD1161.PDF.

Dijkstra:1993:HSS

[Dij93g] Edsger W. Dijkstra. How subtypes should enter the picture. Circulated privately., October 1993. URL http://www.cs.

utexas.edu/users/EWD/ewd11xx/EWD1160.PDF.

Dijkstra:1993:ICP

[Dij93h]  Edsger W. Dijkstra.  Introducing a course on pro-

gram design and presentation. Circulated privately., Au-

gust 1993. URL http://www.cs.utexas.edu/users/EWD/
ewd11xx/EWD1157.PDF.

Dijkstra:1993:NTI

[Dij93i] Edsger W. Dijkstra. A note on “tail invariants”. Circulated privately., October 1993. URL http://www.cs.utexas.edu/

users/EWD/ewd11xx/EWD1159.PDF.

Dijkstra:1993:BIT

[Dij93j] Edsger W. Dijkstra. On bags and identical twins. Circulated privately., February 1993. URL http://www.cs.utexas.

edu/users/EWD/ewd11xx/EWD1153.PDF.

Dijkstra:1993:DCP

[Dij93k] Edsger W. Dijkstra. On the design of calculational proofs. Circulated privately., April 1993. URL http://www.cs.

utexas.edu/users/EWD/ewd11xx/EWD1150.PDF.

Dijkstra:1993:EDM

[Dij93l] Edsger W. Dijkstra. On the economy of doing mathematics. In R. S. Bird, C. C. Morgan, and J. C. P. Woodcock, editors, Mathematics of Program Construction, Second International Conference, Oxford, U.K., volume 669 of Lecture Notes in Computer Science, pages 2–10. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., June/July 1993.

Dijkstra:1993:PMWa

[Dij93m] Edsger W. Dijkstra. A prime is in at most 1 way the sum of 2 squares. Circulated privately., August 1993. URL http://

www.cs.utexas.edu/users/EWD/ewd11xx/EWD1155a.PDF.

Dijkstra:1993:PMWb

[Dij93n] Edsger W. Dijkstra. A prime is in at most 1 way the sum of 2 squares. Circulated privately., June 1993. URL http://

www.cs.utexas.edu/users/EWD/ewd11xx/EWD1155.PDF.

Dijkstra:1993:RC

[Dij93o] Edsger W. Dijkstra. A regrettable cover. Circulated pri-

vately., December 1993. URL http://www.cs.utexas.edu/
users/EWD/ewd11xx/EWD1167.PDF.

Dijkstra:1993:TAA

[Dij93p] Edsger W. Dijkstra. A tentative axiomatization of ascending sequences. Circulated privately., October 1993. URL http:/

/www.cs.utexas.edu/users/EWD/ewd11xx/EWD1164.PDF.

Dijkstra:1993:TT

[Dij93q] Edsger W. Dijkstra. A terrible Thursday. Circulated pri-

vately., January 1993. URL http://www.cs.utexas.edu/
users/EWD/ewd11xx/EWD1151.PDF.

Dijkstra:1993:TSW

[Dij93r] Edsger W. Dijkstra. There is still a war going on. Circulated privately., December 1993. URL http://www.cs.utexas.

edu/users/EWD/ewd11xx/EWD1165.PDF.

Dijkstra:1993:TUC

[Dij93s] Edsger W. Dijkstra. Transitivity as an (unexpected?) consequence. Circulated privately., October 1993. URL http://

www.cs.utexas.edu/users/EWD/ewd11xx/EWD1163.PDF.

Dijkstra:1993:VVJ

[Dij93t] Edsger W. Dijkstra. Voorwoord (voor jaarboek studievereniging UT). (Dutch) [Foreword (for student yearbook UT [University of Texas?])]. Circulated privately., June 1993. URL

http://www.cs.utexas.edu/users/EWD/ewd11xx/EWD1156.
PDF.

Dijkstra:1994:AEE

[Dij94a] Edsger W. Dijkstra. An alternative of the ETAC to EWD1163. Circulated privately., January 1994. URL http:

//www.cs.utexas.edu/users/EWD/ewd11xx/EWD1169.PDF.

Dijkstra:1994:APP

[Dij94b] Edsger W. Dijkstra. Another ping-pong argument from Leibniz’s principle. Circulated privately., October 1994. URL

http://www.cs.utexas.edu/users/EWD/ewd11xx/EWD1189.
PDF.

Dijkstra:1994:AAA

[Dij94c] Edsger W. Dijkstra. The argument about the arithmetic mean and the geometric mean, heuristics included. Circulated privately., January 1994. URL http://www.cs.utexas.edu/

users/EWD/ewd11xx/EWD1171.PDF.

Dijkstra:1994:BCY

[Dij94d] Edsger W. Dijkstra. Boolean connectives yield punctual expressions. Circulated privately., September 1994. URL http:

//www.cs.utexas.edu/users/EWD/ewd11xx/EWD1187.PDF.

Dijkstra:1994:C

[Dij94e] Edsger W. Dijkstra. A chutzpah. Circulated privately.,

April 1994. URL http://www.cs.utexas.edu/users/EWD/
ewd11xx/EWD1176.PDF.

Dijkstra:1994:CC

[Dij94f] Edsger W. Dijkstra. Counting characters. Circulated

privately., May 1994. URL http://www.cs.utexas.edu/
users/EWD/ewd11xx/EWD1182.PDF.

Dijkstra:1994:DMU

[Dij94g] Edsger W. Dijkstra. Don’t mix unary pre- and postfix operators. Circulated privately., May 1994. URL http://www.cs.

utexas.edu/users/EWD/ewd11xx/EWD1181.PDF.

Dijkstra:1994:ETR

[Dij94h] Edsger W. Dijkstra. Equilateral triangles and rectangular grids. Circulated privately., January 1994. URL http://

www.cs.utexas.edu/users/EWD/ewd11xx/EWD1170.PDF.

Dijkstra:1994:HCP

[Dij94i] Edsger W. Dijkstra. Heuristics for a calculational proof. Circulated privately., July 1994. URL http://www.cs.utexas.

edu/users/EWD/ewd11xx/EWD1174a.PDF.

Dijkstra:1994:HVS

[Dij94j] Edsger W. Dijkstra. Heuristics for a very simple Euclidean proof. Circulated privately., May 1994. URL http://www.

cs.utexas.edu/users/EWD/ewd11xx/EWD1180.PDF.

Dijkstra:1994:HPJ

[Dij94k] Edsger W. Dijkstra. The heuristics of a proof by Jan L. A. van de Snepscheut and Richard S. Bird. Circulated privately.,

April 1994. URL http://www.cs.utexas.edu/users/EWD/
ewd11xx/EWD1174.PDF.

Dijkstra:1994:MJV

[Dij94l] Edsger W. Dijkstra. In memoriam Jan L. A. van de Snepscheut. Circulated privately., April 25, 1994. URL http://

www.cs.utexas.edu/users/EWD/ewd11xx/EWD1177.PDF.

Dijkstra:1994:JVS

[Dij94m] Edsger W. Dijkstra. Jan L. A. van de Snepscheut. Formal Aspects of Computing, 6(3):247–249, May 1994. CO-DEN FACME5. ISSN 0934-5043 (print), 1433-299X (elec-

tronic). URL http://link.springer.com/article/10.
1007/BF01215406.

Dijkstra:1994:JMQ

[Dij94n] Edsger W. Dijkstra. Junctivity and massaging quantification. Circulated privately., October 1994. URL http://www.cs.

utexas.edu/users/EWD/ewd11xx/EWD1191.PDF.

Dijkstra:1994:LTT

[Dij94o] Edsger W. Dijkstra. “Less than” in terms of “at most”. Circulated privately., August 1994. URL http://www.cs.utexas.

edu/users/EWD/ewd11xx/EWD1183.PDF.

Dijkstra:1994:LOG

[Dij94p] Edsger W. Dijkstra. Lifting, orders, and the Galois connection. Circulated privately., September 1994. URL http://

www.cs.utexas.edu/users/EWD/ewd11xx/EWD1185.PDF.

Dijkstra:1994:AJC

[Dij94q] Edsger W. Dijkstra. On an argument by J. A. Carruth & J. Misra. Circulated privately., April 1994. URL http://www.

cs.utexas.edu/users/EWD/ewd11xx/EWD1178.PDF.

Dijkstra:1994:AA

[Dij94r] Edsger W. Dijkstra. On arcs and angles. Circulated privately.,

December 1994. URL http://www.cs.utexas.edu/users/
EWD/ewd11xx/EWD1193.PDF.

Dijkstra:1994:TES

[Dij94s] Edsger W. Dijkstra. On two equations that have the same extreme solution. Circulated privately., January 1994. URL

http://www.cs.utexas.edu/users/EWD/ewd11xx/EWD1168.
PDF.

Dijkstra:1994:OBO

[Dij94t] Edsger W. Dijkstra. Our book’s omission on quantification over scalar subtypes. Circulated privately., Au-

gust 1994. URL http://www.cs.utexas.edu/users/EWD/
ewd11xx/EWD1184.PDF.

Dijkstra:1994:PPA

[Dij94u] Edsger W. Dijkstra. Ping-pong arguments and Leibniz’s principle. Circulated privately., September 1994. URL http://

www.cs.utexas.edu/users/EWD/ewd11xx/EWD1186.PDF.

Dijkstra:1994:PAM

[Dij94v] Edsger W. Dijkstra. Problem 10406 from The American Mathematical Monthly, Volume 101, Number 8 / October 1994. Circulated privately., October 1994. URL http://

www.cs.utexas.edu/users/EWD/ewd11xx/EWD1190.PDF.

Dijkstra:1994:PZP

[Dij94w] Edsger W. Dijkstra. A problem from Zhendong Patrick Su (including an alternative solution by R. Boute). Circulated privately., September 1994. URL http://www.cs.utexas.

edu/users/EWD/ewd11xx/EWD1188.PDF.

Dijkstra:1994:PTD

[Dij94x] Edsger W. Dijkstra. Pythagorean triples, or the design of a theorem. Circulated privately., February 1994. URL http:/

/www.cs.utexas.edu/users/EWD/ewd11xx/EWD1172.PDF.

Dijkstra:1994:SP

[Dij94y] Edsger W. Dijkstra. A sorry parade. Circulated privately.,

May 1994. URL http://www.cs.utexas.edu/users/EWD/
ewd11xx/EWD1179.PDF.

Dijkstra:1994:SAE

[Dij94z] Edsger W. Dijkstra. The strengths of the academic enterprise. Circulated privately., February 1994. URL http://www.cs.

utexas.edu/users/EWD/ewd11xx/EWD1175.PDF.

Dijkstra:1994:TPJ

[Dij94-27] Edsger W. Dijkstra. A theorem proved by John Segers. Circulated privately., February 1994. URL http://www.cs.

utexas.edu/users/EWD/ewd11xx/EWD1173.PDF.

Dijkstra:1994:VFB

[Dij94-28] Edsger W. Dijkstra. The very first beginnings of lattice theory. Circulated privately., November 1994. URL http://

www.cs.utexas.edu/users/EWD/ewd11xx/EWD1192.PDF.

Dijkstra:1994:ZPS

[Dij94-29] Edsger W. Dijkstra. Z. P. Su’s second problem. Circulated privately., December 1994. URL http://www.cs.utexas.

edu/users/EWD/ewd11xx/EWD1194.PDF.

Dijkstra:1995:AFP

[Dij95a] Edsger W. Dijkstra. Another forced ping-pong argument? Circulated privately., January 1995. URL http://www.cs.

utexas.edu/users/EWD/ewd11xx/EWD1195.PDF.

Dijkstra:1995:CPD

[Dij95b] Edsger W. Dijkstra. Calculational proof design; an experiment. Circulated privately., December 1995. URL http://

www.cs.utexas.edu/users/EWD/ewd12xx/EWD1221b.PDF.

Dijkstra:1995:CGD

[Dij95c] Edsger W. Dijkstra. The complete (n + 1)-graph in ndimensional space. Circulated privately., March 1995. URL

http://www.cs.utexas.edu/users/EWD/ewd12xx/EWD1202.
PDF.

Dijkstra:1995:CD

[Dij95d] Edsger W. Dijkstra. Complete DAGs. Circulated privately.,

March 1995. URL http://www.cs.utexas.edu/users/EWD/
ewd12xx/EWD1204.PDF.

Dijkstra:1995:CDB

[Dij95e] Edsger W. Dijkstra. Courtesy Dr. Birgit Schieder. Circulated privately., September 1995. URL http://www.cs.utexas.

edu/users/EWD/ewd12xx/EWD1215.PDF.

Dijkstra:1995:E

[Dij95f] Edsger W. Dijkstra. The equation x: [x]. Circulated privately.,

September 1995. URL http://www.cs.utexas.edu/users/
EWD/ewd12xx/EWD1216.PDF.

Dijkstra:1995:FAE

[Dij95g] Edsger W. Dijkstra. A few addenda to EWD1192. Circulated privately., October 1995. URL http://www.cs.utexas.edu/

users/EWD/ewd12xx/EWD1218.PDF.

Dijkstra:1995:RPS

[Dij95h] Edsger W. Dijkstra. For the record: painting the squared plane. Circulated privately., August 1995. URL http://

www.cs.utexas.edu/users/EWD/ewd12xx/EWD1212.PDF.

Dijkstra:1995:GEV

[Dij95i] Edsger W. Dijkstra. Generators of extreme values. Circulated privately., January 1995. URL http://www.cs.utexas.edu/

users/EWD/ewd11xx/EWD1196.PDF.

Dijkstra:1995:HCP

[Dij95j] Edsger W. Dijkstra. Heuristics for a calculational proof. Information Processing Letters, 53(3):141–143, February 10, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dijkstra:1995:HNP

[Dij95k] Edsger W. Dijkstra. Hm++, a new Professional Standard (From the desk of the chairman of the Board of Mathematics Inc.). Circulated privately., December 1995. URL http://

www.cs.utexas.edu/users/EWD/ewd12xx/EWD1224.PDF.

Dijkstra:1995:PL

[Dij95l] Edsger W. Dijkstra. “I have a proof that … ”. Circulated privately., October 1995. URL http://www.cs.utexas.edu/

users/EWD/ewd12xx/EWD1219.PDF.

Dijkstra:1995:I

[Dij95m] Edsger W. Dijkstra. Inspired by [x] ≡ [ x]. Circulated privately., November 1995. URL http://www.cs.utexas.

edu/users/EWD/ewd12xx/EWD1222.PDF.

Dijkstra:1995:ICC

[Dij95n] Edsger W. Dijkstra. Introducing a course on calculi. Circulated privately., August 1995. URL http://www.cs.utexas.

edu/users/EWD/ewd12xx/EWD1213.PDF.

Dijkstra:1995:JVS

[Dij95o] Edsger W. Dijkstra. Jan van de Snepscheut’s tiling problem. Circulated privately., February 1995. URL http://www.cs.

utexas.edu/users/EWD/ewd11xx/EWD1198.PDF.

Dijkstra:1995:MA

[Dij95p] Edsger W. Dijkstra. The marriage agency. Circulated pri-

vately., September 1995. URL http://www.cs.utexas.edu/
users/EWD/ewd12xx/EWD1214.PDF.

Dijkstra:1995:MWF

[Dij95q] Edsger W. Dijkstra. Misra’s weakest fair buffer. Circulated privately., December 1995. URL http://www.cs.utexas.

edu/users/EWD/ewd12xx/EWD1225.PDF.

Dijkstra:1995:NUN

[Dij95r] Edsger W. Dijkstra. The non-unique node multiplicity. Circulated privately., July 1995. URL http://www.cs.utexas.

edu/users/EWD/ewd12xx/EWD1210.PDF.

Dijkstra:1995:SVB

[Dij95s] Edsger W. Dijkstra. On some very binary patterns (dedicated to F. E. J. Kruseman Aretz). Circulated privately.,

September 1995. URL http://www.cs.utexas.edu/users/
EWD/ewd12xx/EWD1211a.PDF.

Dijkstra:1995:OMS

[Dij95t] Edsger W. Dijkstra. Only a matter of style? Circulated pri-

vately., February 1995. URL http://www.cs.utexas.edu/
users/EWD/ewd12xx/EWD1200.PDF.

Dijkstra:1995:PPU

[Dij95u] Edsger W. Dijkstra. Points in the plane (unfinished). Circulated privately., 1995. URL http://www.cs.utexas.edu/

users/EWD/ewd12xx/EWD1221a.PDF.

Dijkstra:1995:POM

[Dij95v] Edsger W. Dijkstra. Primes once more (re Kac & Ulam). Circulated privately., March 1995. URL http://www.cs.

utexas.edu/users/EWD/ewd12xx/EWD1203.PDF.

Dijkstra:1995:RRR

[Dij95w] Edsger W. Dijkstra. A result of rabbit removal. Circulated

privately., June 1995. URL http://www.cs.utexas.edu/
users/EWD/ewd12xx/EWD1208.PDF.

Dijkstra:1995:SGT

[Dij95x] Edsger W. Dijkstra. A simple geometrical theorem I did not know. Circulated privately., December 1995. URL http://

www.cs.utexas.edu/users/EWD/ewd12xx/EWD1226.PDF.

Dijkstra:1995:SEE

[Dij95y] Edsger W. Dijkstra. A supplement to EWD1140 and EWD1171. Circulated privately., February 1995. URL http:

//www.cs.utexas.edu/users/EWD/ewd11xx/EWD1199.PDF.

Dijkstra:1995:TA

[Dij95z] Edsger W. Dijkstra. A termination argument. Circulated privately., January 1995. URL http://www.cs.utexas.edu/

users/EWD/ewd11xx/EWD1197.PDF.

Dijkstra:1995:TAF

[Dij95-27] Edsger W. Dijkstra. A theorem about “factors” perhaps worth recording. Circulated privately., May 1995. URL http:

//www.cs.utexas.edu/users/EWD/ewd12xx/EWD1207.PDF.

Dijkstra:1995:TWF

[Dij95-28] Edsger W. Dijkstra. Triggered by Wim Feijen’s treatment of “∃∀ ⇒ ∀∃”. Circulated privately., February 1995. URL http:

//www.cs.utexas.edu/users/EWD/ewd12xx/EWD1201.PDF.

Dijkstra:1995:THL

[Dij95-29] Edsger W. Dijkstra. A trip to Harvard Law School, 10–11 April 1995. Circulated privately., April 1995. URL http://

www.cs.utexas.edu/users/EWD/ewd12xx/EWD1205.PDF.

Dijkstra:1995:UUS

[Dij95-30] Edsger W. Dijkstra. An unfortunate use of symbols. Circulated privately., May 1995. URL http://www.cs.utexas.

edu/users/EWD/ewd12xx/EWD1206.PDF.

Dijkstra:1995:WAC

[Dij95-31] Edsger W. Dijkstra. Why American Computing Science seems incurable. Circulated privately., August 1995. URL

http://www.cs.utexas.edu/users/EWD/ewd12xx/EWD1209.
PDF.

Dijkstra:1995:WMU

[Dij95-32] Edsger W. Dijkstra. WLOG, or the misery of the unordered pair. Circulated privately., November 1995. URL http://

www.cs.utexas.edu/users/EWD/ewd12xx/EWD1223.PDF.

Dijkstra:1996:AEE

[Dij96a] Edsger W. Dijkstra. An alternative of the ETAC to EWD1163. In Broy [Bro96], pages 27–28. ISBN 3-540-60947-4 (hardcover). LCCN QA76.9.D5 D38 1996.

URL http://www.loc.gov/catdir/enhancements/fy0812/
96010788-d.html.

Dijkstra:1996:ACO

[Dij96b] Edsger W. Dijkstra. Andrew’s challenge once more (see EWD1247). Circulated privately., October 1996. URL http:

//www.cs.utexas.edu/users/EWD/ewd12xx/EWD1249.PDF.

Dijkstra:1996:AAA

[Dij96c] Edsger W. Dijkstra. The argument about the arithmetic mean and the geometric mean, heuristics included. In Broy [Bro96], pages 29–32. ISBN 3-540-60947-4 (hardcover).

LCCN QA76.9.D5 D38 1996. URL http://www.loc.gov/

catdir/enhancements/fy0812/96010788-d.html.

Dijkstra:1996:AGM

[Dij96d] Edsger W. Dijkstra. The arithmetic and geometric means once more. Circulated privately., February 1996. URL http:

//www.cs.utexas.edu/users/EWD/ewd12xx/EWD1231.PDF.

Dijkstra:1996:BEA

[Dij96e] Edsger W. Dijkstra. A bagatelle on Euclid’s algorithm. In Broy [Bro96], pages 21–23. ISBN 3-540-60947-4 (hardcover). LCCN QA76.9.D5 D38 1996. URL http://www.loc.gov/

catdir/enhancements/fy0812/96010788-d.html.

Dijkstra:1996:BC

[Dij96f] Edsger W. Dijkstra. The balance and the coins. In Broy [Bro96], pages 11–13. ISBN 3-540-60947-4 (hardcover). LCCN QA76.9.D5 D38 1996. URL http://www.loc.gov/

catdir/enhancements/fy0812/96010788-d.html.

Dijkstra:1996:BER

[Dij96g] Edsger W. Dijkstra. Beware of the empty range. Circulated privately., September 1996. URL http://www.cs.utexas.

edu/users/EWD/ewd12xx/EWD1247.PDF.

Dijkstra:1996:BTS

[Dij96h] Edsger W. Dijkstra. Bulterman’s theorem on shortest trees. In Broy [Bro96], pages 15–16. ISBN 3-540-60947-4 (hardcover). LCCN QA76.9.D5 D38 1996. URL http://www.loc.

gov/catdir/enhancements/fy0812/96010788-d.html.

Dijkstra:1996:CSD

[Dij96i] Edsger W. Dijkstra. Convocation speech, December 8, 1996. Circulated privately., December 1996. URL http://www.cs.

utexas.edu/users/EWD/ewd12xx/EWD1252.PDF.

Dijkstra:1996:CSP

[Dij96j] Edsger W. Dijkstra. Cooperating sequential processes. In Laplante [Lap96], chapter 5.1, pages 321–377. ISBN 0-314-06365-X (paperback), 0-7803-1112-4 (hardcover). LCCN

QA76 .G686 1996. URL http://bit.csc.lsu.edu/~chen/
GreatPapers.html.

Dijkstra:1996:CRL

[Dij96k] Edsger W. Dijkstra. The couples, the river, and the little boat. Circulated privately., November 1996. URL http://

www.cs.utexas.edu/users/EWD/ewd12xx/EWD1250.PDF.

Dijkstra:1996:CRJ

[Dij96l] Edsger W. Dijkstra. Courtesy Rajeev Joshi. Circulated pri-

vately., September 1996. URL http://www.cs.utexas.edu/
users/EWD/ewd12xx/EWD1246.PDF.

Dijkstra:1996:SSD

[Dij96m] Edsger W. Dijkstra. Een scheve schaats. (Dutch) [A skewed skate]. Circulated privately., March 14, 1996. URL http://

www.cs.utexas.edu/users/EWD/ewd12xx/EWD1235.PDF.

Dijkstra:1996:EER

[Dij96n] Edsger W. Dijkstra. Elegance and effective reasoning (Fall 1996). Circulated privately., May 1996. URL http://www.

cs.utexas.edu/users/EWD/ewd12xx/EWD1237.PDF.

Dijkstra:1996:FGC

[Dij96o] Edsger W. Dijkstra. Fibonacci and the greatest common divisor. In Broy [Bro96], pages 7–10. ISBN 3-540-60947-4 (hardcover). LCCN QA76.9.D5 D38 1996. URL http://www.loc.

gov/catdir/enhancements/fy0812/96010788-d.html.

Dijkstra:1996:FEE

[Dij96p] Edsger W. Dijkstra. A first exploration of effective reasoning. Circulated privately., July 1996. URL http://www.cs.

utexas.edu/users/EWD/ewd12xx/EWD1239.PDF.

Dijkstra:1996:Fa

[Dij96q] Edsger W. Dijkstra. Foreword. Circulated privately.,

May 1996. URL http://www.cs.utexas.edu/users/EWD/
ewd12xx/EWD1238.PDF.

Dijkstra:1996:Fd

[Dij96r] Edsger W. Dijkstra. Foreword. In Teaching and Learning Formal Methods [DH96], pages vii–viii. ISBN 0-12-349040-5. LCCN QA76.76D47 T42. URL http://www.gbv.de/dms/

bowker/toc/9780123490407.pdf. Foreword by Edsger W.

Dijkstra.

Dijkstra:1996:Fb

[Dij96s] Edsger W. Dijkstra. The formula for sin.(α + β). Circulated privately., September 1996. URL http://www.cs.utexas.

edu/users/EWD/ewd12xx/EWD1244a.PDF.

Dijkstra:1996:Fc

[Dij96t] Edsger W. Dijkstra. The formula for sin.(α + β). Circulated privately., September 1996. URL http://www.cs.utexas.

edu/users/EWD/ewd12xx/EWD1244.PDF.

Dijkstra:1996:GSC

[Dij96u] Edsger W. Dijkstra. Go to statement considered harmful. In Laplante [Lap96], chapter 6.3, pages 420–422. ISBN 0-314-06365-X (paperback), 0-7803-1112-4 (hardcover). LCCN

QA76 .G686 1996. URL http://bit.csc.lsu.edu/~chen/
GreatPapers.html.

Dijkstra:1996:GCN

[Dij96v] Edsger W. Dijkstra. Guarded commands, nondeterminancy, and formal derivation of programs. In Laplante [Lap96], chapter 6.2, pages 412–419. ISBN 0-314-06365-X (paperback), 0-7803-1112-4 (hardcover). LCCN QA76 .G686 1996. URL

http://bit.csc.lsu.edu/~chen/GreatPapers.html.

Dijkstra:1996:HM

[Dij96w] Edsger W. Dijkstra. A hint from monotonicity? Circulated privately., February 1996. URL http://www.cs.utexas.

edu/users/EWD/ewd12xx/EWD1233.PDF.

Dijkstra:1996:HSE

[Dij96x] Edsger W. Dijkstra. Homework #1 (see EWD996). Circulated privately., October 1996. URL http://www.cs.

utexas.edu/users/EWD/ewd12xx/EWD1248.PDF.

Dijkstra:1996:HCM

[Dij96y] Edsger W. Dijkstra. Honors course “Mathematical Methodology”, Spring 1996. Circulated privately., 1996. URL http:

//www.cs.utexas.edu/users/EWD/ewd12xx/EWD1220.PDF.

Dijkstra:1996:KCL

[Dij96z] Edsger W. Dijkstra. A kind of converse of Leibniz’s Principle. Circulated privately., September 1996. URL http://www.cs.

utexas.edu/users/EWD/ewd12xx/EWD1245.PDF.

Dijkstra:1996:MIF

[Dij96-27] Edsger W. Dijkstra. Mathematical induction’s fixpoint. Circulated privately., December 1996. URL http://www.cs.

utexas.edu/users/EWD/ewd12xx/EWD1253.PDF.

Dijkstra:1996:MST

[Dij96-28] Edsger W. Dijkstra. My simplest theorem. Circulated pri-

vately., February 1996. URL http://www.cs.utexas.edu/
users/EWD/ewd12xx/EWD1232.PDF.

Dijkstra:1996:NFYa

[Dij96-29] Edsger W. Dijkstra. The next fifty years. Circulated

privately., 1996. URL http://www.cs.utexas.edu/users/
EWD/ewd12xx/EWD1243.PDF.

Dijkstra:1996:NFYb

[Dij96-30] Edsger W. Dijkstra. The next fifty years. Circulated

privately., 1996. URL http://www.cs.utexas.edu/users/
EWD/ewd12xx/EWD1243a.PDF.

Dijkstra:1996:NCA

[Dij96-31] Edsger W. Dijkstra. Nondeterministic construction of an arbitrary witness. Circulated privately., January 1996. URL

http://www.cs.utexas.edu/users/EWD/ewd12xx/EWD1229.
PDF.

Dijkstra:1996:TES

[Dij96-32] Edsger W. Dijkstra. On two equations that have the same extreme solution. In Broy [Bro96], pages 25–26. ISBN 3-540-60947-4 (hardcover). LCCN QA76.9.D5 D38 1996.

URL http://www.loc.gov/catdir/enhancements/fy0812/
96010788-d.html.

Dijkstra:1996:PMW

[Dij96-33] Edsger W. Dijkstra. A prime is in at most 1 way the sum of 2 squares. In Broy [Bro96], pages 17–20. ISBN 3-540-60947-4 (hardcover). LCCN QA76.9.D5 D38 1996.

URL http://www.loc.gov/catdir/enhancements/fy0812/
96010788-d.html.

Dijkstra:1996:PCL

[Dij96-34] Edsger W. Dijkstra. A problem communicated by Laurens de Vries. Circulated privately., March 1996. URL http://www.

cs.utexas.edu/users/EWD/ewd12xx/EWD1234.PDF.

Dijkstra:1996:PCH

[Dij96-35] Edsger W. Dijkstra. Programming considered as a human activity. In Laplante [Lap96], pages 648–?? ISBN 0-314-06365-X (paperback), 0-7803-1112-4 (hardcover). LCCN

QA76 .G686 1996. URL http://bit.csc.lsu.edu/~chen/
GreatPapers.html.

Dijkstra:1996:SEW

[Dij96-36] Edsger W. Dijkstra. A sequel to EWD1241 and WF 233/AvG 141. Circulated privately., December 1996. URL http://

www.cs.utexas.edu/users/EWD/ewd12xx/EWD1254.PDF.

Dijkstra:1996:SPC

[Dij96-37] Edsger W. Dijkstra. Solution of a problem in concurrent process control. In Laplante [Lap96], chapter 5.2, pages 378–380. ISBN 0-314-06365-X (paperback), 0-7803-1112-4 (hardcover).

LCCN QA76 .G686 1996. URL http://bit.csc.lsu.edu/~
chen/GreatPapers.html.

Dijkstra:1996:SOL

[Dij96-38] Edsger W. Dijkstra. A somewhat open letter to David Gries. Circulated privately., January 1996. URL http://www.cs.

utexas.edu/users/EWD/ewd12xx/EWD1227.PDF.

Dijkstra:1996:STU

[Dij96-39] Edsger W. Dijkstra. Sylvester’s theorem used (see EWD1016). Circulated privately., January 1996. URL http://www.cs.

utexas.edu/users/EWD/ewd12xx/EWD1228.PDF.

Dijkstra:1996:TVL

[Dij96-40] Edsger W. Dijkstra. Three very little problems from Eindhoven. Circulated privately., February 1996. URL http://

www.cs.utexas.edu/users/EWD/ewd12xx/EWD1230.PDF.

Dijkstra:1996:TCW

[Dij96-41] Edsger W. Dijkstra. The transitive closure of a wellfounded relation. Circulated privately., June 1996. URL http://www.

cs.utexas.edu/users/EWD/ewd12xx/EWD1241.PDF.

Dijkstra:1996:TPD

[Dij96-42] Edsger W. Dijkstra. Two problems derived from Hugo Steinhaus. Circulated privately., November 1996. URL http://

www.cs.utexas.edu/users/EWD/ewd12xx/EWD1251.PDF.

Dijkstra:1996:WSC

[Dij96-43] Edsger W. Dijkstra. The wife-swapping couples once more (see EWD1103 and EWD1115). Circulated privately.,

April 1996. URL http://www.cs.utexas.edu/users/EWD/
ewd12xx/EWD1236.PDF.

Dijkstra:1997:ABT

[Dij97a] Edsger W. Dijkstra. The angle bisectors of a triangle. Circulated privately., April 1997. URL http://www.cs.utexas.

edu/users/EWD/ewd12xx/EWD1261.PDF.

Dijkstra:1997:CE

[Dij97b] Edsger W. Dijkstra. A correction to EWD1240. Circulated

privately., May 1997. URL http://www.cs.utexas.edu/
users/EWD/ewd12xx/EWD1263.PDF.

Dijkstra:1997:DGC

[Dij97c] Edsger W. Dijkstra. Defining the greatest common divisor. Circulated privately., February 7, 1997. URL http://www.

cs.utexas.edu/users/EWD/ewd12xx/EWD1257.PDF.

Dijkstra:1997:LTa

[Dij97d] Edsger W. Dijkstra. The ladder theorem. Circulated pri-

vately., March 1997. URL http://www.cs.utexas.edu/
users/EWD/ewd12xx/EWD1258a.PDF.

Dijkstra:1997:LTb

[Dij97e] Edsger W. Dijkstra. The ladder theorem. Circulated pri-

vately., February 1997. URL http://www.cs.utexas.edu/
users/EWD/ewd12xx/EWD1258.PDF.

Dijkstra:1997:MCS

[Dij97f] Edsger W. Dijkstra. The marked coins and the scale. Circulated privately., March 1997. URL http://www.cs.utexas.

edu/users/EWD/ewd12xx/EWD1260.PDF.

Dijkstra:1997:MD

[Dij97g] Edsger W. Dijkstra. The Mathematical Divide. Circulated privately., December 1997. URL http://www.cs.utexas.

edu/users/EWD/ewd12xx/EWD1268.PDF.

Dijkstra:1997:MMS

[Dij97h] Edsger W. Dijkstra. Mathematical Methodology, Spring 1997. Circulated privately., January 1997. URL http://

www.cs.utexas.edu/users/EWD/ewd12xx/EWD1256.PDF.

Dijkstra:1997:DPT

[Dij97i] Edsger W. Dijkstra. On the determinant of the product of two square matrices. Circulated privately., Novem-

ber 1997. URL http://www.cs.utexas.edu/users/EWD/
ewd12xx/EWD1267.PDF.

Dijkstra:1997:PIC

[Dij97j] Edsger W. Dijkstra. Proving an implication via its converse. Circulated privately., 1997. URL http://www.cs.utexas.

edu/users/EWD/ewd12xx/EWD1266a.PDF.

Dijkstra:1997:PST

[Dij97k] Edsger W. Dijkstra. Pruning the search tree. Circulated privately., January 1997. URL http://www.cs.utexas.edu/

users/EWD/ewd12xx/EWD1255.PDF.

Dijkstra:1997:TCO

[Dij97l] Edsger W. Dijkstra. The theorem of de Ceva once more. Circulated privately., June 1997. URL http://www.cs.utexas.

edu/users/EWD/ewd12xx/EWD1265.PDF.

Dijkstra:1997:TW

[Dij97m] Edsger W. Dijkstra. The tide, not the waves. In Denning and Metcalfe [DM97], pages 59–64. ISBN 0-387-94932-1, 0-387-98588-3 (paperback), 1-4612-0685-5. LCCN QA76 .D348 1997.

Dijkstra:1998:APS

[Dij98a] Edsger W. Dijkstra. Appalling prose and the shortest path. In Shasha and Lazere [SL98], page ?? ISBN 0-387-98269-8. LCCN QA76.2.A2 S53 1998. US$16.00.

Dijkstra:1998:CE

[Dij98b] Edsger W. Dijkstra. A calculational example. Circulated privately., September 1998. URL http://www.cs.utexas.

edu/users/EWD/ewd12xx/EWD1278.PDF.

Dijkstra:1998:RYS

[Dij98c] Edsger W. Dijkstra. For the record: Yossi Shiloach’s Algorithm. Circulated privately., June 1998. URL http://www.

cs.utexas.edu/users/EWD/ewd12xx/EWD1274.PDF.

Dijkstra:1998:GPM

[Dij98d] Edsger W. Dijkstra. A geometry problem from “The Monthly”, March 1998. Circulated privately., April 1998.

URL http://www.cs.utexas.edu/users/EWD/ewd12xx/EWD1272.
PDF.

Dijkstra:1998:HCC

[Dij98e] Edsger W. Dijkstra. How closures could have been invented. Circulated privately., September 1998. URL http://www.cs.

utexas.edu/users/EWD/ewd12xx/EWD1270.PDF.

Dijkstra:1998:LBL

[Dij98f] Edsger W. Dijkstra. A little bit of lattice theory. Circulated privately., January 1998. URL http://www.cs.utexas.edu/

users/EWD/ewd12xx/EWD1240a.PDF.

Dijkstra:1998:CD

[Dij98g] Edsger W. Dijkstra. N cheers for determinants (N >> 0). Circulated privately., 1998. URL http://www.cs.utexas.

edu/users/EWD/ewd12xx/EWD1271.PDF.

Dijkstra:1998:DLK

[Dij98h] Edsger W. Dijkstra. On Dijkstra’s Lemma and Kruskal’s Algorithm. Circulated privately., May 1998. URL http://www.

cs.utexas.edu/users/EWD/ewd12xx/EWD1273.PDF.

Dijkstra:1998:SCD

[Dij98i] Edsger W. Dijkstra. A simple country dance for transitivity and shunting. Circulated privately., Septem-

ber 1998. URL http://www.cs.utexas.edu/users/EWD/
ewd12xx/EWD1279.PDF.

Dijkstra:1998:SPH

[Dij98j] Edsger W. Dijkstra. A simple proof of Hall’s Theorem. Circulated privately., January 1998. URL http://www.cs.

utexas.edu/users/EWD/ewd12xx/EWD1269.PDF.

Dijkstra:1998:SRM

[Dij98k] Edsger W. Dijkstra. Society’s role in mathematics. Circulated privately., September 1998. URL http://www.cs.utexas.

edu/users/EWD/ewd12xx/EWD1277.PDF.

Dijkstra:1999:CPR

[Dij99a] Edsger W. Dijkstra. A 2-colouring problem in the rational plane. Circulated privately., September 1999. URL http://

www.cs.utexas.edu/users/EWD/ewd12xx/EWD1289.PDF.

Dijkstra:1999:CSA

[Dij99b] Edsger W. Dijkstra. Computing Science: Achievements and Challenges. Circulated privately., March 1999. URL http:/

/www.cs.utexas.edu/users/EWD/ewd12xx/EWD1284.PDF.

Dijkstra:1999:CBS

[Dij99c] Edsger W. Dijkstra. Constructing the binary search once more. Circulated privately., November 1999. URL http://

www.cs.utexas.edu/users/EWD/ewd12xx/EWD1293.PDF.

Dijkstra:1999:CGC

[Dij99d] Edsger W. Dijkstra. Courtesy Georg Cantor and Jayadev Misra. Circulated privately., October 1999. URL http://

www.cs.utexas.edu/users/EWD/ewd12xx/EWD1291.PDF.

Dijkstra:1999:DTD

[Dij99e] Edsger W. Dijkstra. Dear Tony, dearest Jill, and other people, dear or not. Circulated privately., 1999. URL http://www.

cs.utexas.edu/users/EWD/ewd12xx/EWD1287.PDF.

Dijkstra:1999:ECC

[Dij99f] Edsger W. Dijkstra. Eliminating cascading carries. Circulated privately., September 1999. URL http://www.cs.utexas.

edu/users/EWD/ewd12xx/EWD1290.PDF.

Dijkstra:1999:EML

[Dij99g] Edsger W. Dijkstra. EWD1166: from my life. In People & ideas in theoretical computer science, Springer Ser. Discrete Math. Theor. Comput. Sci., pages 86–92. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999.

Dijkstra:1999:FWT

[Dij99h] Edsger W. Dijkstra. A formula is worth a thousand pictures (dedicated to Charles Antony Richard Hoare). Circulated

privately., June 1999. URL http://www.cs.utexas.edu/
users/EWD/ewd12xx/EWD1280a.PDF.

Dijkstra:1999:HTT

[Dij99i] Edsger W. Dijkstra. How “they” try to corrupt “us”. Circulated privately., February 1999. URL http://www.cs.

utexas.edu/users/EWD/ewd12xx/EWD1283.PDF.

Dijkstra:1999:IRC

[Dij99j] Edsger W. Dijkstra. An iteration for the k-th root with cubic convergence. Circulated privately., January 1999. URL http:

//www.cs.utexas.edu/users/EWD/ewd12xx/EWD1281.PDF.

Dijkstra:1999:GWN

[Dij99k] Edsger W. Dijkstra. On graphs whose nodes are Black or White. Circulated privately., February 1999. URL http://

www.cs.utexas.edu/users/EWD/ewd12xx/EWD1282.PDF.

Dijkstra:1999:PTE

[Dij99l] Edsger W. Dijkstra. A partition theorem of Euler’s generalized. Circulated privately., June 1999. URL http://www.cs.

utexas.edu/users/EWD/ewd12xx/EWD1286.PDF.

Dijkstra:1999:CMA

[Dij99m] Edsger W. Dijkstra. To Cambridge by mistake (13–18 April 1999). Circulated privately., April 1999. URL http://www.

cs.utexas.edu/users/EWD/ewd12xx/EWD1285.PDF.

Dijkstra:1999:TET

[Dij99n] Edsger W. Dijkstra. A tough experiment with the triangle calculus. Circulated privately., March 1999. URL http://

www.cs.utexas.edu/users/EWD/ewd12xx/EWD1280.PDF.

Dijkstra:1999:UBA

[Dij99o] Edsger W. Dijkstra. Ulrich Berger’s argument rephrased. Circulated privately., September 1999. URL http://www.cs.

utexas.edu/users/EWD/ewd12xx/EWD1288.PDF.

Dijkstra:1999:WSO

[Dij99p] Edsger W. Dijkstra. When a symmetric operator distributes over ↑ and ↓. Circulated privately., October 1999. URL http:

//www.cs.utexas.edu/users/EWD/ewd12xx/EWD1292.PDF.

Dijkstra:19xx:PCP

[Dijxxa] Edsger W. Dijkstra. Plotting a curve with a printer. Circulated privately., 19xx. URL http://www.cs.utexas.edu/

users/EWD/ewd02xx/EWD260.PDF.

Dijkstra:19xx:RR

[Dijxxb] Edsger W. Dijkstra. A rearranging routine. EWD316-89, 19xx.

Dijkstra:2000:AQS

[Dij00a] Edsger W. Dijkstra. Answers to questions from students of software engineering. Circulated privately., Novem-

ber 2000. URL http://www.cs.utexas.edu/users/EWD/
ewd13xx/EWD1305.PDF.

Dijkstra:2000:NCA

[Dij00b] Edsger W. Dijkstra. The notational conventions I adopted, and why. Circulated privately., July 2000. URL http://www.

cs.utexas.edu/users/EWD/ewd13xx/EWD1300.PDF.

Dijkstra:2000:DSS

[Dij00c] Edsger W. Dijkstra. On defining sets, suprema, and infima. Circulated privately., May 2000. URL http://www.

cs.utexas.edu/users/EWD/ewd12xx/EWD1299.PDF.

Dijkstra:2000:TCWa

[Dij00d] Edsger W. Dijkstra. On the transitive closure of a wellfounded relation. Circulated privately., January 2000. URL http://

www.cs.utexas.edu/users/EWD/ewd12xx/EWD1295.PDF.

Dijkstra:2000:TCWb

[Dij00e] Edsger W. Dijkstra. On the transitive closure of a wellfounded relation. In Pomberger et al. [PGB00], pages 31– 40. ISBN 1-55860-723-4 (Morgan Kaufmann), 3-932588-85-1

(dpunkt). LCCN ???? URL http://www.springerlink.
com/content/978-3-642-48354-7.

Dijkstra:2000:WMGa

[Dij00f] Edsger W. Dijkstra. On weighted means, geometric and arithmetic. Circulated privately., February 2000. URL http://

www.cs.utexas.edu/users/EWD/ewd12xx/EWD1296.PDF.

Dijkstra:2000:WMGb

[Dij00g] Edsger W. Dijkstra. On weighted means, geometric and arithmetic. Circulated privately., February 2000. URL http://

www.cs.utexas.edu/users/EWD/ewd12xx/EWD1296a.PDF.

Dijkstra:2000:OMB

[Dij00h] Edsger W. Dijkstra. Once more bichrome triangles in complete graphs. Circulated privately., October 2000. URL http:

//www.cs.utexas.edu/users/EWD/ewd13xx/EWD1302.PDF.

Dijkstra:2000:RIB

[Dij00i] Edsger W. Dijkstra. The river, the isles and the bridges. Circulated privately., July 2000. URL http://www.cs.utexas.

edu/users/EWD/ewd13xx/EWD1301.PDF.

Dijkstra:2000:THS

[Dij00j] Edsger W. Dijkstra. Triggered by a high-school exercise. Circulated privately., March 2000. URL http://www.cs.

utexas.edu/users/EWD/ewd12xx/EWD1297.PDF.

Dijkstra:2000:USL

[Dij00k] Edsger W. Dijkstra. Under the spell of Leibniz’s Dream. farewell lecture, to be submitted to Inf. Proc. Letters; circulated privately, April 2000. URL http://www.cs.utexas.

edu/users/EWD/ewd12xx/EWD1298.PDF.

Dijkstra:2001:ALP

[Dij01a] Edsger W. Dijkstra. Another look at a problem from Hugo Steinhaus. Circulated privately., October 2001. URL http:/

/www.cs.utexas.edu/users/EWD/ewd13xx/EWD1311.PDF.

Dijkstra:2001:CCD

[Dij01b] Edsger W. Dijkstra. The chessboard covered with dominoes. Circulated privately., February 2001. URL http://www.cs.

utexas.edu/users/EWD/ewd13xx/EWD1306a.PDF.

Dijkstra:2001:ECS

[Dij01c] Edsger W. Dijkstra. The end of computing science? Communications of the Association for Computing Machinery, 44(3):92, March 2001. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:/

/www.acm.org/pubs/citations/journals/cacm/2001-44-

3/p92-dijkstra/. Copyright held by author.

Dijkstra:2001:GM

[Dij01d] Edsger W. Dijkstra. The GCD and the minimum. Circulated privately., November 2001. URL http://www.cs.utexas.

edu/users/EWD/ewd13xx/EWD1313.PDF.

Dijkstra:2001:GSC

[Dij01e] Edsger W. Dijkstra. Go to statement considered harmful. In Broy and Denert [BD01b], pages 297–300. ISBN 3-540-42290-0 (print), 3-642-48354-2 (e-book). LCCN QA76.758.

URL http://link.springer.com/chapter/10.1007/978-
3-642-48354-7_12/.

Dijkstra:2001:IEE

[Dij01f] Edsger W. Dijkstra. Indirect equality enriched (and a proof by Netty). Circulated privately., December 2001. URL http:

//www.cs.utexas.edu/users/EWD/ewd13xx/EWD1315.PDF.

Dijkstra:2001:MRO

[Dij01g] Edsger W. Dijkstra. My recollections of operating system design. Circulated privately., April 2001. URL http://www.

cs.utexas.edu/users/EWD/ewd13xx/EWD1303.PDF.

Dijkstra:2001:DBN

[Dij01h] Edsger W. Dijkstra. On disjoint binary numbers. Circulated privately., November 2001. URL http://www.cs.utexas.

edu/users/EWD/ewd13xx/EWD1312.PDF.

Dijkstra:2001:PFS

[Dij01i] Edsger W. Dijkstra. Philips and I: a few snapshots. Circulated privately., December 2001. URL http://www.cs.utexas.

edu/users/EWD/ewd13xx/EWD1316.PDF.

Dijkstra:2001:SPC

[Dij01j] Edsger W. Dijkstra. Solution of a problem in concurrent programming control. In Broy and Denert [BD01b], pages 289–294. ISBN 3-540-42290-0 (print), 3-642-48354-2 (e-

book). LCCN QA76.758. URL http://link.springer.com/
chapter/10.1007/978-3-642-48354-7_10/.

Dijkstra:2001:SMS

[Dij01k] Edsger W. Dijkstra. The structure of the “THE” multiprogramming system. In Classic Operating Systems: from Batch Processing to Distributed Systems [Bri01], pages 223– 236. ISBN 0-387-95113-X. LCCN QA76.76.O63 B7425 2001.

URL https://link.springer.com/chapter/10.1007/978-
1-4757-3510-9_12.

Dijkstra:2001:TTR

[Dij01l] Edsger W. Dijkstra. Three trip reports rolled into one: 2001.04.16–2001.07.06. Circulated privately., Septem-

ber 2001. URL http://www.cs.utexas.edu/users/EWD/
ewd13xx/EWD1310.PDF.

Dijkstra:2001:UCA

[Dij01m] Edsger W. Dijkstra. An unavoidable case analysis. Circulated

privately., March 2001. URL http://www.cs.utexas.edu/
users/EWD/ewd13xx/EWD1307.PDF.

Dijkstra:2001:USL

[Dij01n] Edsger W. Dijkstra. Under the spell of Leibniz’s dream. Information Processing Letters, 77(2–4):53–61, February 28, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (elec-

tronic). URL http://www.elsevier.nl/gej-ng/10/23/20/
68/29/25/abstract.html;   http://www.elsevier.nl/
gej-ng/10/23/20/68/29/25/article.pdf. Special issue in

honor of Edsger W. Dijkstra (Austin, TX, 2000).

Dijkstra:2001:WLN

[Dij01o] Edsger W. Dijkstra. What led to “Notes on Structured Programming”. Circulated privately., June 2001. URL http://

www.cs.utexas.edu/users/EWD/ewd13xx/EWD1308.PDF.

Dijkstra:2001:ZPE

[Dij01p] Edsger W. Dijkstra. Zuckerman’s problem and the ETAC. Circulated privately., June 2001. URL http://www.cs.

utexas.edu/users/EWD/ewd13xx/EWD1309a.PDF.

Dijkstra:2002:DDP

[Dij02a] D. Dijkstra. Doubling the degree of precision without doubling the grid when solving a differential equation with a pseudo-spectral collocation method. Journal of Scientific Computing, 17(1-4):513–527, 2002. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

Dijkstra:2002:CSP

[Dij02b] Edsger W. Dijkstra. Cooperating sequential processes. In Brinch Hansen [Bri02], pages 65–138. ISBN 0-387-95401-5, 1-4419-2986-X, 1-4757-3472-7, 1-4757-3472-7. LCCN QA76.642

.O75 2002. URL http://link.springer.com/10.1007/
978-1-4757-3472-0.

Dijkstra:2002:CR

[Dij02c] Edsger W. Dijkstra. Coxeter’s rabbit. Circulated privately.,

April 2002. URL http://www.cs.utexas.edu/users/EWD/
ewd13xx/EWD1318.PDF.

Dijkstra:2002:EWL

[Dij02d] Edsger W. Dijkstra. EWD 1308: What led to “Notes on structured programming”. In Broy and Denert [BD02], pages 340–364. ISBN 3-540-43081-4, 3-642-48354-2 (e-book), 3-642-59412-3 (e-book), 3-642-59413-1, 3-642-63970-4 (print). LCCN QA76.2.A2 S62 2002. URL http://aleph.unisg.

ch/hsgscan/hm00048094.pdf;   http://link.springer.
com/10.1007/978-3-642-59412-0;   http://www.loc.gov/
catdir/enhancements/fy0817/2002070484-d.html;   http:
//www.springerlink.com/content/978-3-642-48354-7;
 z3950.loc.gov:7090/Voyager.

Dijkstra:2002:ENC

[Dij02e] Edsger W. Dijkstra. EWD1300: The notational conventions I adopted, and why. Formal Aspects of Computing, 14(2): 99–107, 2002. CODEN FACME5. ISSN 0934-5043 (print), 1433-299X (electronic).

Dijkstra:2002:VIC

[Dij02f] Edsger W. Dijkstra. From van IJzeren’s correspondence to my aunt & uncle. Circulated privately., Jan-

uary 2002. URL http://www.cs.utexas.edu/users/EWD/
ewd13xx/EWD1317.PDF.

Dijkstra:2002:GSC

[Dij02g] Edsger W. Dijkstra. Go to statement considered harmful. In Broy and Denert [BD02], pages 351–355. ISBN 3-540-43081-4, 3-642-48354-2 (e-book), 3-642-59412-3 (e-book), 3-642-59413-1, 3-642-63970-4 (print). LCCN QA76.2.A2 S62

2002. URL http://aleph.unisg.ch/hsgscan/hm00048094.
pdf;   http://link.springer.com/10.1007/978-3-642-
59412-0;   http://www.loc.gov/catdir/enhancements/
fy0817/2002070484-d.html;   http://www.springerlink.
com/content/978-3-642-48354-7;   z3950.loc.gov:7090/
Voyager.

Dijkstra:2002:HOS

[Dij02h] Edsger W. Dijkstra. Hierarchical ordering of sequential processes. In Brinch Hansen [Bri02], pages 198–227. ISBN 0-387-95401-5, 1-4419-2986-X, 1-4757-3472-7, 1-4757-3472-7. LCCN

QA76.642 .O75 2002. URL http://link.springer.com/10.
1007/978-1-4757-3472-0.

Dijkstra:2002:SPC

[Dij02i] Edsger W. Dijkstra. Solution of a problem in concurrent programming control. In Broy and Denert [BD02], pages 347–350. ISBN 3-540-43081-4, 3-642-48354-2 (e-book), 3-642-59412-3 (e-book), 3-642-59413-1, 3-642-63970-4 (print). LCCN QA76.2.A2 S62 2002. URL http://aleph.unisg.

ch/hsgscan/hm00048094.pdf;   http://link.springer.
com/10.1007/978-3-642-59412-0;   http://www.loc.gov/
catdir/enhancements/fy0817/2002070484-d.html;   http:
//www.springerlink.com/content/978-3-642-48354-7;
 z3950.loc.gov:7090/Voyager.

Dijkstra:2002:SMS

[Dij02j] Edsger W. Dijkstra. The structure of the “THE” multiprogramming system. In Brinch Hansen [Bri02], pages 139–152. ISBN 0-387-95401-5, 1-4419-2986-X, 1-4757-3472-7, 1-4757-3472-7. LCCN QA76.642 .O75 2002. URL http://link.

springer.com/10.1007/978-1-4757-3472-0.

Dijkstra:2002:SE

[Dij02k] Edsger W. Dijkstra. A synthesis emerging? In Brinch Hansen [Bri02], pages 397–412. ISBN 0-387-95401-5, 1-4419-2986-X, 1-4757-3472-7, 1-4757-3472-7. LCCN QA76.642 .O75

2002. URL http://link.springer.com/10.1007/978-1-
4757-3472-0.

Dijkstra:2004:MMD

[Dij04] Matthijs Dijkstra. A model for the management of dynamic web applications. In Janssen et al. [JSW04], pages 402–408. ISBN 1-58113-930-6. LCCN ????

Dijkstra:2005:MRO

[Dij05] Edsger W. Dijkstra. My recollections of operating system design. Operating Systems Review, 39(2):4–40, April 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Dijkstra:2006:TAN

[Dij06] Arie Dijkstra. Technology adds new principles to persuasive psychology: Evidence from health education. In IJsselsteijn et al. [IdKM06], pages 16–26. ISBN 3-540-34291-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

BF637.P4 I58 2006. URL http://www.springerlink.com/
content/978-3-540-34291-5.

Dijkstra:2008:HP

[Dij08] Edsger W. Dijkstra. Home page, 2008. URL http://www.

cs.utexas.edu/users/EWD/.

Dijkstra:2009:TP

[Dij09] E. W. Dijkstra. On the theorem of Pythagoras. Nieuw Archief voor Wiskunde (5), 10(2):94–99, 2009. ISSN 0028-9825. Photo reproduction of the handwritten 1986 manuscript, With an introduction in Dutch by Klaas Pieter Hart.

Dijkstra:2021:GSC

[Dij21a] E. W. Dijkstra. Go to statement considered harmful (1968). In Ideas That Created the Future: Classic Papers of Computer Science [Lew21], chapter 29, pages 289–292. ISBN 0-262-04530-3. LCCN Q124.6-127.2.

Dijkstra:2021:SPC

[Dij21b] E. W. Dijkstra. Solution of a problem in concurrent program control (1965). In Ideas That Created the Future: Classic Papers of Computer Science [Lew21], chapter 26, pages 267– 270. ISBN 0-262-04530-3. LCCN Q124.6-127.2.

Dijkstra:2021:SMS

[Dij21c] E. W. Dijkstra. The structure of the “THE”-multiprogramming system (1968). In Ideas That Created the Future: Classic Papers of Computer Science [Lew21], chapter 28, pages 279–288. ISBN 0-262-04530-3. LCCN Q124.6-127.2.

Dijkstra:20xx:ACW

[Dijxxa] Edsger W. Dijkstra. Aan de Commissie Wetenschapsbeleid. (Dutch) [To the Committee on Science Policy]. Circulated

privately., 20xx. URL http://www.cs.utexas.edu/users/
EWD/ewd02xx/EWD281.PDF.

Dijkstra:20xx:AKR

[Dijxxb] Edsger W. Dijkstra. Aan de keuzecommissie rekenmachine. (Dutch) [The Selection Committee computer]. Circulated

privately., 20xx. URL http://www.cs.utexas.edu/users/
EWD/ewd02xx/EWD270.PDF.

Dijkstra:20xx:ALVb

[Dijxxc] Edsger W. Dijkstra. Aan de leden van de Onderafdeling der Wiskunde. (Dutch) [To the members of the Sub-Section of Mathematics]. Circulated privately., 20xx. URL http://

www.cs.utexas.edu/users/EWD/OtherDocs/NN252.PDF.

Dijkstra:20xx:ALVa

[Dijxxd] Edsger W. Dijkstra. Aan de leden van de onderafdelingsvergadering. (Dutch) [To the members of the Division meeting]. Circulated privately., 20xx. URL http://www.cs.utexas.

edu/users/EWD/ewd01xx/EWD167.PDF.

Dijkstra:20xx:ASG

[Dijxxe] Edsger W. Dijkstra. ACM Symposium in Gatlinburg “Operating System Principles”. Circulated privately., 20xx. URL

http://www.cs.utexas.edu/users/EWD/ewd02xx/EWD214.
PDF.

Dijkstra:20xx:AVP

[Dijxxf] Edsger W. Dijkstra. Appreciatie van ponskaarten. (Dutch) [appreciation of punch cards]. Circulated privately., 20xx.

URL http://www.cs.utexas.edu/users/EWD/ewd01xx/EWD139.
PDF.

Dijkstra:20xx:AV

[Dijxxg] Edsger W. Dijkstra. Array variables. Circulated pri-

vately., 20xx. URL http://www.cs.utexas.edu/users/
EWD/ewd04xx/EWD428.PDF.

Dijkstra:20xx:BEE

[Dijxxh] Edsger W. Dijkstra. [Beantwoording ener enquete]. (Dutch) [[Answering listener survey]]. Circulated privately., 20xx.

URL http://www.cs.utexas.edu/users/EWD/ewd02xx/EWD247a.
PDF.

Dijkstra:20xx:BAV

[Dijxxi] Edsger W. Dijkstra. Bij de aanvang van het semester. (Dutch) [At the beginning of the semester]. Circulated

privately., 20xx. URL http://www.cs.utexas.edu/users/
EWD/ewd04xx/EWD471.PDF.

Dijkstra:20xx:CS

[Dijxxj] Edsger W. Dijkstra. The characterization of semantics. Circulated privately., 20xx. URL http://www.cs.utexas.edu/

users/EWD/ewd04xx/EWD401.PDF.

Dijkstra:20xx:CCH

[Dijxxk] Edsger W. Dijkstra. Complexity controlled by hierarchical ordering of function and variability. Circulated pri-

vately., 20xx. URL http://www.cs.utexas.edu/users/
EWD/ewd02xx/EWD236.PDF.

Dijkstra:20xx:CVP

[Dijxxl] Edsger W. Dijkstra. Computation versus program. incomplete; circulated privately, 20xx. URL http://www.cs.

utexas.edu/users/EWD/ewd02xx/EWD238.PDF.

Dijkstra:20xx:CSE

[Dijxxm] Edsger W. Dijkstra. Computer Science of enkel Software Engineering?. (Dutch) [Computer science or software engineering only?]. Circulated privately., 20xx. URL http://www.

cs.utexas.edu/users/EWD/ewd02xx/EWD254.PDF.

Dijkstra:20xx:CDN

[Dijxxn] Edsger W. Dijkstra. Context dependent names. Circulated

privately., 20xx. URL http://www.cs.utexas.edu/users/
EWD/ewd01xx/EWD155.PDF.

Dijkstra:20xx:BVD

[Dijxxo] Edsger W. Dijkstra. De bankiersalgorithme en verfijningen daarvan. (Dutch) [The Banker’s Algorithm and refinements thereof]. Circulated privately., 20xx. URL http://www.cs.

utexas.edu/users/EWD/ewd01xx/EWD116.PDF.

Dijkstra:20xx:LAA

[Dijxxp] Edsger W. Dijkstra. De logische automaat in academisch milieu. (Dutch) [The logical machine in an academic environment]. Circulated privately., 20xx. URL http://www.cs.

utexas.edu/users/EWD/OtherDocs/NN035.PDF.

Dijkstra:20xx:DOPa

[Dijxxq] Edsger W. Dijkstra. Description of the object program. Circulated privately., 20xx. URL http://www.cs.utexas.edu/

users/EWD/ewd01xx/EWD101.PDF.

Dijkstra:20xx:DOPb

[Dijxxr] Edsger W. Dijkstra. Description of the object program (a sequel to EWD102). Circulated privately., 20xx. URL http:

//www.cs.utexas.edu/users/EWD/ewd01xx/EWD105.PDF.

Dijkstra:20xx:DCM

[Dijxxs] Edsger W. Dijkstra. Design considerations in more detail. Circulated privately., 20xx. URL http://www.cs.utexas.

edu/users/EWD/ewd03xx/EWD302.PDF.

Dijkstra:20xx:DCA

[Dijxxt] Edsger W. Dijkstra. Documentatie over de communicatie apparatuur aan de EL X8. (Dutch) [Documentation of the communication equipment of the EL X8]. Circulated pri-

vately., 20xx. URL http://www.cs.utexas.edu/users/
EWD/ewd01xx/EWD140.PDF.

Dijkstra:20xx:ATV

[Dijxxu] Edsger W. Dijkstra. Een algorithme ter voorkoming van de dodelijke omarming. (Dutch) [An algorithm to prevent deadly embrace]. Circulated privately., 20xx. URL http://www.cs.

utexas.edu/users/EWD/ewd01xx/EWD108.PDF.

Dijkstra:20xx:PVX

[Dijxxv] Edsger W. Dijkstra. Een ponsbandorganisatie voor de X8. (Dutch) [A paper tape organization for the X8]. Circulated

privately., 20xx. URL http://www.cs.utexas.edu/users/
EWD/ewd00xx/EWD82.PDF.

Dijkstra:20xx:EVD

[Dijxxw] Edsger W. Dijkstra. Eerste verkenning over de dood van programma’s. (Dutch) [First exploration of the death of programs]. Circulated privately., 20xx. URL http://www.cs.

utexas.edu/users/EWD/ewd01xx/EWD163.PDF.

Dijkstra:20xx:ETS

[Dijxxx] Edsger W. Dijkstra. An effort towards structuring of programmed processes. Circulated privately., 20xx. URL http:

//www.cs.utexas.edu/users/EWD/ewd01xx/EWD198.PDF.

Dijkstra:20xx:EPL

[Dijxxy] Edsger W. Dijkstra. Exit “The Programming Laboratory”. Circulated privately., 20xx. URL http://www.cs.utexas.

edu/users/EWD/ewd03xx/EWD300.PDF.

Dijkstra:20xx:ERC

[Dijxxz] Edsger W. Dijkstra. An experiment with the “record class” as suggested by C. A. R. Hoare. Circulated pri-

vately., 20xx. URL http://www.cs.utexas.edu/users/
EWD/ewd01xx/EWD132.PDF.

Dijkstra:20xx:FMS

[Dijxx-27] Edsger W. Dijkstra. Finding the maximal strong components in a directed graph. Circulated privately., 20xx. URL http:

//www.cs.utexas.edu/users/EWD/ewd04xx/EWD453.PDF.

Dijkstra:20xx:FTS

[Dijxx-28] Edsger W. Dijkstra. The formal treatment of some small examples. Circulated privately., 20xx. URL http://www.cs.

utexas.edu/users/EWD/ewd04xx/EWD413.PDF.

Dijkstra:20xx:GSS

[Dijxx-29] Edsger W. Dijkstra. A generalization of the Sheffer Stroke for n-valued logic (by C. S. Scholten). Circulated pri-

vately., 20xx. URL http://www.cs.utexas.edu/users/
EWD/ewd04xx/EWD429.PDF.

Dijkstra:20xx:HVO

[Dijxx-30] Edsger W. Dijkstra. Heer, verlos ons van de charlatans!. (Dutch) [Lord, deliver us from charlatans!]. Circulated pri-

vately., 20xx. URL http://www.cs.utexas.edu/users/
EWD/ewd04xx/EWD451.PDF.

Dijkstra:20xx:CC

[Dijxx-31] Edsger W. Dijkstra. Het controlerende communicatieapparaat. (Dutch) [The communication controller device]. Circulated privately., 20xx. URL http://www.cs.utexas.edu/

users/EWD/ewd00xx/EWD77.PDF.

Dijkstra:20xx:CIA

[Dijxx-32] Edsger W. Dijkstra. Het curriculum informatica aan de THE. (Dutch) [The curriculum in computer science at the THE]. Circulated privately., 20xx. URL http://www.cs.utexas.

edu/users/EWD/ewd04xx/EWD449.PDF.

Dijkstra:20xx:EVA

[Dijxx-33] Edsger W. Dijkstra. Het einde van een ambacht. (Dutch) [The end of a craft]. Circulated privately., 20xx. URL http:

//www.cs.utexas.edu/users/EWD/ewd01xx/EWD194.PDF.

Dijkstra:20xx:V

[Dijxx-34] Edsger W. Dijkstra. Het vectorgeheugen. (Dutch) [The vector memory]. Circulated privately., 20xx. URL http://www.cs.

utexas.edu/users/EWD/ewd00xx/EWD41.PDF.

Dijkstra:20xx:HOSa

[Dijxx-35] Edsger W. Dijkstra. Hierarchical ordering of sequential processes. Published as [Dij72f]., 20xx. URL http://www.cs.

utexas.edu/users/EWD/ewd03xx/EWD310.PDF.

Dijkstra:20xx:HWP

[Dijxx-36] Edsger W. Dijkstra. Hoe wiskundig programmeren is. (Dutch) [What is mathematical programming?]. Circulated

privately., 20xx. URL http://www.cs.utexas.edu/users/
EWD/ewd02xx/EWD261.PDF.

Dijkstra:20xx:HOA

[Dijxx-37] Edsger W. Dijkstra. Hoger orde adressering. (Dutch) [Highorder addressing]. Circulated privately., 20xx. URL http://

www.cs.utexas.edu/users/EWD/ewd02xx/EWD254a.PDF.

Dijkstra:20xx:IWD

[Dijxx-38] Edsger W. Dijkstra. Informatica als wiskundige discipline. (Dutch) [Informatics as a mathematical discipline]. Circu-

lated privately., 20xx. URL http://www.cs.utexas.edu/
users/EWD/ewd02xx/EWD298.PDF.

Dijkstra:20xx:IMI

[Dijxx-39] Edsger W. Dijkstra. Ir. Medema, Ir. Hendriks, weledelgestrenge heren!. (Dutch) [Engineer Medema, Engineer Hendriks, noble gentlemen!]. Circulated privately., 20xx. URL http://www.

cs.utexas.edu/users/EWD/ewd01xx/EWD184.PDF.

Dijkstra:20xx:JHB

[Dijxx-40] Edsger W. Dijkstra. [“Journal for half-baked Ideas”]. Circulated privately., 20xx. URL http://www.cs.utexas.edu/

users/EWD/ewd02xx/EWD247.PDF.

Dijkstra:20xx:KTS

[Dijxx-41] Edsger W. Dijkstra. Keuze tussen symcharf en symchart. (Dutch) [Choice between symcharf and symchart]. Circulated

privately., 20xx. URL http://www.cs.utexas.edu/users/
EWD/ewd01xx/EWD164.PDF.

Dijkstra:20xx:LBL

[Dijxx-42] Edsger W. Dijkstra. A little bit of lattice theory. circulated privately; replaced by EWD1240a, 20xx. URL http://www.

cs.utexas.edu/users/EWD/ewd12xx/EWD1240.PDF.

Dijkstra:20xx:MLV

[Dijxx-43] Edsger W. Dijkstra. Mijn laatste verslag van een bijeenkomst van W.G.2.1. (Dutch) [My final report of a meeting of W.G.2.1]. Circulated privately., 20xx. URL http://www.

cs.utexas.edu/users/EWD/ewd02xx/EWD252.PDF.

Dijkstra:20xx:MEE

[Dijxx-44] Edsger W. Dijkstra. The moral of EWD237–EWD239. Circulated privately., 20xx. URL http://www.cs.utexas.edu/

users/EWD/ewd02xx/EWD240.PDF.

Dijkstra:20xx:MX

[Dijxx-45] Edsger W. Dijkstra. Multiprogrammering en de X8. (Dutch) [Multiprogramming and the X8]. Circulated pri-

vately., 20xx. URL http://www.cs.utexas.edu/users/
EWD/ewd00xx/EWD51.PDF.

Dijkstra:20xx:MXV

[Dijxx-46] Edsger W. Dijkstra. Multiprogrammering en de X8 (Vervolg van EWD51). (Dutch) [Multiprogramming and the X8 (continuation of EWD51)]. Circulated privately., 20xx. URL

http://www.cs.utexas.edu/users/EWD/ewd00xx/EWD54.
PDF.

Dijkstra:20xx:MD

[Dijxx-47] Edsger W. Dijkstra. On a methodology of design. Circulated

privately., 20xx. URL http://www.cs.utexas.edu/users/
EWD/ewd03xx/EWD317.PDF.

Dijkstra:20xx:AI

[Dijxx-48] Edsger W. Dijkstra. On avoiding the infinite. Circulated

privately., 20xx. URL http://www.cs.utexas.edu/users/
EWD/ewd04xx/EWD416.PDF.

Dijkstra:20xx:NDB

[Dijxx-49] Edsger W. Dijkstra. On non-determinacy being bounded. Circulated privately., 20xx. URL http://www.cs.utexas.

edu/users/EWD/ewd04xx/EWD458.PDF.

Dijkstra:20xx:RP

[Dijxx-50] Edsger W. Dijkstra. On the reliability of programs. Circulated privately., 20xx. URL http://www.cs.utexas.edu/

users/EWD/ewd03xx/EWD303.PDF.

Dijkstra:20xx:TSA

[Dijxx-51] Edsger W. Dijkstra. On trading storage against computation time. Circulated privately., 20xx. URL http://www.cs.

utexas.edu/users/EWD/ewd02xx/EWD239.PDF.

Dijkstra:20xx:UP

[Dijxx-52] Edsger W. Dijkstra. On understanding programs. Circulated

privately., 20xx. URL http://www.cs.utexas.edu/users/
EWD/ewd02xx/EWD264.PDF.

Dijkstra:20xx:US

[Dijxx-53] Edsger W. Dijkstra. On useful structuring. Circulated

privately., 20xx. URL http://www.cs.utexas.edu/users/
EWD/ewd02xx/EWD245.PDF.

Dijkstra:20xx:OCI

[Dijxx-54] Edsger W. Dijkstra. Ontwerp Collegebeschrijving “Inleiding tot de kunst van het programmeren”. (Dutch) [draft college description “Introduction to art of programming”]. Circu-

lated privately., 20xx. URL http://www.cs.utexas.edu/
users/EWD/ewd02xx/EWD253a.PDF.

Dijkstra:20xx:OVL

[Dijxx-55] Edsger W. Dijkstra. Ontwerp voor “Letter to the Manufacturers”. (Dutch) [draft for “Letter to the Manufacturers”]. Circulated privately., 20xx. URL http://www.cs.utexas.

edu/users/EWD/OtherDocs/NN034.PDF.

Dijkstra:20xx:OVA

[Dijxx-56] Edsger W. Dijkstra. Ontwerp voorwoord [afstudeerrichting Fundamentele Programmering]. (Dutch) [Draft foreword [specialization in fundamental programming]]. Circulated

privately., 20xx. URL http://www.cs.utexas.edu/users/
EWD/ewd02xx/EWD252a.PDF.

Dijkstra:20xx:I

[Dijxx-57] Edsger W. Dijkstra. Over de IBM360. (Dutch) [On the IBM360]. Circulated privately., 20xx. URL http://www.cs.

utexas.edu/users/EWD/ewd02xx/EWD255.PDF.

Dijkstra:20xx:SVP

[Dijxx-58] Edsger W. Dijkstra. Over de sequentialiteit van procesbeschrijvingen. (Dutch) [On the sequentiality of process descriptions]. Circulated privately., 20xx. URL http://www.

cs.utexas.edu/users/EWD/ewd00xx/EWD35.PDF.

Dijkstra:20xx:SIVa

[Dijxx-59] Edsger W. Dijkstra. Over de sequentiele interpretatie van een ponsband. (Dutch) [On the sequential interpretation of a punched tape]. Circulated privately., 20xx. URL http://

www.cs.utexas.edu/users/EWD/ewd01xx/EWD157.PDF.

Dijkstra:20xx:TVC

[Dijxx-60] Edsger W. Dijkstra. Over de toekomst van Computer Science. (Dutch) [On the future of Computer Science]. Circulated

privately., 20xx. URL http://www.cs.utexas.edu/users/
EWD/ewd02xx/EWD267.PDF.

Dijkstra:20xx:AOT

[Dijxx-61] Edsger W. Dijkstra. Over een academische opleiding tot “Computer Scientist”. (Dutch) [An academic program for “Computer Scientist”]. Circulated privately., 20xx. URL

http://www.cs.utexas.edu/users/EWD/ewd01xx/EWD188.
PDF.

Dijkstra:20xx:PA

[Dijxx-62] Edsger W. Dijkstra. Over pagina-administratie. (Dutch) [On page management]. Circulated privately., 20xx. URL http:

//www.cs.utexas.edu/users/EWD/ewd00xx/EWD71.PDF.

Dijkstra:20xx:Sa

[Dijxx-63] Edsger W. Dijkstra. Over seinpalen. (Dutch) [On semaphores]. Circulated privately., 20xx. URL http://www.cs.utexas.

edu/users/EWD/ewd00xx/EWD74.PDF.

Dijkstra:20xx:Sb

[Dijxx-64] Edsger W. Dijkstra. Over standaardroutines. (Dutch) [On standard routines]. Circulated privately., 20xx. URL http:/

/www.cs.utexas.edu/users/EWD/ewd00xx/EWD75.PDF.

Dijkstra:20xx:SI

[Dijxx-65] Edsger W. Dijkstra. Over stapeladministratie. II. (Dutch) [On stack administration. II]. Circulated privately., 20xx.

URL http://www.cs.utexas.edu/users/EWD/ewd00xx/EWD70.
PDF.

Dijkstra:20xx:P

[Dijxx-66] Edsger W. Dijkstra. A parabel. Circulated privately., 20xx.

URL http://www.cs.utexas.edu/users/EWD/ewd03xx/EWD365.
PDF.

Dijkstra:20xx:PMP

[Dijxx-67] Edsger W. Dijkstra. The pattern-matching problem. Circulated privately., 20xx. URL http://www.cs.utexas.edu/

users/EWD/ewd04xx/EWD459.PDF.

Dijkstra:20xx:PIC

[Dijxx-68] Edsger W. Dijkstra. A preliminary investigation into Computer Assisted Programming. Circulated privately., 20xx.

URL http://www.cs.utexas.edu/users/EWD/ewd02xx/EWD237.
PDF.

Dijkstra:20xx:PDV

[Dijxx-69] Edsger W. Dijkstra. Pretentie en doelstelling van het THE multiprogrammeringsproject. (Dutch) [Pretense and purpose of the THE multiprogramming project]. Circulated pri-

vately., 20xx. URL http://www.cs.utexas.edu/users/
EWD/ewd02xx/EWD232.PDF.

Dijkstra:20xx:PMI

[Dijxx-70] Edsger W. Dijkstra. The problem of the most isolated villages. Circulated privately., 20xx. URL http://www.cs.

utexas.edu/users/EWD/ewd04xx/EWD440.PDF.

Dijkstra:20xx:PS

[Dijxx-71] Edsger W. Dijkstra. Problem/solution. Circulated pri-

vately., 20xx. URL http://www.cs.utexas.edu/users/
EWD/OtherDocs/NN117.PDF.

Dijkstra:20xx:PLP

[Dijxx-72] Edsger W. Dijkstra. The programming laboratory project. Circulated privately., 20xx. URL http://www.cs.utexas.

edu/users/EWD/ewd02xx/EWD279.PDF.

Dijkstra:20xx:PMTa

[Dijxx-73] Edsger W. Dijkstra. Programming methodologies, their objectives and their nature. Published as [Dij72m]., 20xx. URL

http://www.cs.utexas.edu/users/EWD/ewd04xx/EWD469.
PDF.

Dijkstra:20xx:PIN

[Dijxx-74] Edsger W. Dijkstra. Proposal for the input and the notation of the system. Circulated privately., 20xx. URL http://www.

cs.utexas.edu/users/EWD/ewd01xx/EWD147.PDF.

Dijkstra:20xx:RCC

[Dijxx-75] Edsger W. Dijkstra. [Raw code for computing De Bruijnsequences]. Circulated privately., 20xx. URL http://www.

cs.utexas.edu/users/EWD/ewd02xx/EWD221.PDF.

Dijkstra:20xx:RBV

[Dijxx-76] Edsger W. Dijkstra. Reisverslag bezoek van E. W. Dijkstra aan IFIP W.G.2.3 te Warwick, Engeland. (Dutch) [Trip report by E. W. Dijkstra ti IFIP W.G.2.3 in Warwick, England].

Circulated privately., 20xx. URL http://www.cs.utexas.

edu/users/EWD/ewd03xx/EWD308.PDF.

Dijkstra:20xx:RVB

[Dijxx-77] Edsger W. Dijkstra. Reisverslag van bezoek aan het Seminar on the Teaching of Programming at University Level. (Dutch) [Travelogue of visits to the Seminar on the Teaching of Programming at University Level]. Circulated pri-

vately., 20xx. URL http://www.cs.utexas.edu/users/
EWD/ewd02xx/EWD295.PDF.

Dijkstra:20xx:RPT

[Dijxx-78] Edsger W. Dijkstra. Requirements of programming tools. Circulated privately., 20xx. URL http://www.cs.utexas.

edu/users/EWD/ewd02xx/EWD251.PDF.

Dijkstra:20xx:STM

[Dijxx-79] Edsger W. Dijkstra. Sans titre (but most definitely a predecessor to EWD316, “A short introduction into the art of programming”). Circulated privately., 20xx. URL http://

www.cs.utexas.edu/users/EWD/ewd02xx/EWD287.PDF.

Dijkstra:20xx:SE

[Dijxx-80] Edsger W. Dijkstra. A sequel to EWD126. Circulated

privately., 20xx. URL http://www.cs.utexas.edu/users/
EWD/ewd01xx/EWD130.PDF.

Dijkstra:20xx:SIVb

[Dijxx-81] Edsger W. Dijkstra. Sequentiele interpretatie van ponsband. (Dutch) [Sequential interpretation of a punched tape]. Circulated privately., 20xx. URL http://www.cs.utexas.edu/

users/EWD/ewd01xx/EWD161.PDF.

Dijkstra:20xx:SCA

[Dijxx-82] Edsger W. Dijkstra. Some comments on the aims of MIRFAC. Circulated privately., 20xx. URL http://www.cs.utexas.

edu/users/EWD/ewd00xx/EWD68.PDF.

Dijkstra:20xx:SMA

[Dijxx-83] Edsger W. Dijkstra. Some meditations on advanced programming. Circulated privately., 20xx. URL http://www.cs.

utexas.edu/users/EWD/ewd00xx/EWD32.PDF.

Dijkstra:20xx:SS

[Dijxx-84] Edsger W. Dijkstra. Synchronisatie en sequencing. (Dutch) [Synchronization and sequencing]. Circulated privately.,

20xx. URL http://www.cs.utexas.edu/users/EWD/ewd04xx/
EWD405.PDF.

Dijkstra:20xx:TCO

[Dijxx-85] Edsger W. Dijkstra. Tentamen “Co-operating Sequential Processes”. (Dutch) [Exam question “Co-operating Sequential Processes”]. Circulated privately., 20xx. URL http://www.

cs.utexas.edu/users/EWD/ewd01xx/EWD190.PDF.

Dijkstra:20xx:TCS

[Dijxx-86] Edsger W. Dijkstra. Tentamenopgave “Cooperating Sequential Processes”. (Dutch) [Exam question “Cooperating Sequential Processes”]. Circulated privately., 20xx. URL http:

//www.cs.utexas.edu/users/EWD/ewd01xx/EWD150.PDF.

Dijkstra:20xx:TWH

[Dijxx-87] Edsger W. Dijkstra. A time-wise hierarchy imposed upon the use of a two-level store. Circulated privately., 20xx. URL

http://www.cs.utexas.edu/users/EWD/ewd04xx/EWD408.
PDF.

Dijkstra:20xx:EA

[Dijxx-88] Edsger W. Dijkstra. To the EDITOR ALGOL 68. Circulated

privately., 20xx. URL http://www.cs.utexas.edu/users/
EWD/ewd02xx/EWD230.PDF.

Dijkstra:20xx:TFP

[Dijxx-89] Edsger W. Dijkstra. Toekomstverwachting Fundamentele Programmering. (Dutch) [Future outlook for basic programming]. Circulated privately., 20xx. URL http://www.cs.

utexas.edu/users/EWD/ewd02xx/EWD259.PDF.

Dijkstra:20xx:TNV

[Dijxx-90] Edsger W. Dijkstra. [Toespraak over de noodzaak van programmeermethodologie]. (Dutch) [[Speech on the need for programming methodology]]. Circulated privately., 20xx.

URL http://www.cs.utexas.edu/users/EWD/ewd03xx/EWD341.
PDF.

Dijkstra:20xx:TTA

[Dijxx-91] Edsger W. Dijkstra. [Toespraak tot een afstudeerder]. (Dutch) [[Addressing a graduate]]. Circulated privately.,

20xx. URL http://www.cs.utexas.edu/users/EWD/ewd04xx/
EWD493.PDF.

Dijkstra:20xx:TTS

[Dijxx-92] Edsger W. Dijkstra. [toespraak tot een student bij uitreiking ir-diploma]. (Dutch) [Speaking to students at IR-degree ceremony]. Circulated privately., 20xx. URL http://www.

cs.utexas.edu/users/EWD/ewd04xx/EWD499.PDF.

Dijkstra:20xx:TCP

[Dijxx-93] Edsger W. Dijkstra. Towards correct programs. Circulated

privately., 20xx. URL http://www.cs.utexas.edu/users/
EWD/ewd02xx/EWD241.PDF.

Dijkstra:20xx:TSS

[Dijxx-94] Edsger W. Dijkstra. A tree-structured system. Circulated

privately., 20xx. URL http://www.cs.utexas.edu/users/
EWD/ewd02xx/EWD282.PDF.

Dijkstra:20xx:TTT

[Dijxx-95] Edsger W. Dijkstra. Tweede toespraak tot mijn studenten. (Dutch) [Second lecture to my students]. Circulated

privately., 20xx. URL http://www.cs.utexas.edu/users/
EWD/ewd04xx/EWD477.PDF.

Dijkstra:20xx:VVT

[Dijxx-96] Edsger W. Dijkstra. Verslag van de tweede “Conference on Software Engineering”, georganiseerd door de NATO Science Committee te Rome, 27–31 oktober 1969. (Dutch) [Report of the second “Conference on Software Engineering”, organized by the NATO Science Committee in Rome, 27–31 October 1969]. Circulated privately., 20xx. URL http://www.cs.

utexas.edu/users/EWD/ewd02xx/EWD272.PDF.

Dijkstra:20xx:VVB

[Dijxx-97] Edsger W. Dijkstra. Verslag van het bezoek aan de NATO Conference on Software Engineering. (Dutch) [Report of a visit to the NATO Conference on Software Engineering]. Circulated privately., 20xx. URL http://www.cs.utexas.edu/

users/EWD/ewd02xx/EWD246.PDF.

Dijkstra:20xx:VTT

[Dijxx-98] Edsger W. Dijkstra. Vijfde toespraak tot mijn studenten. (Dutch) [Fifth lecture to my students]. Circulated

privately., 20xx. URL http://www.cs.utexas.edu/users/
EWD/ewd04xx/EWD486.PDF.

Dijkstra:20xx:WIN

[Dijxx-99] Edsger W. Dijkstra. Waarom ik niet gewoon hoogleraar wil blijven. (Dutch) [Why I want to stay not just as a professor]. Circulated privately., 20xx. URL http://www.cs.utexas.

edu/users/EWD/ewd03xx/EWD366.PDF.

Dijkstra:20xx:WHI

[Dijxx-100] Edsger W. Dijkstra. Weledelgestrenge Heer, Ingenieur Swinkels. (Dutch) [Noble gentleman, engineer Swinkels]. Circulated privately., 20xx. URL http://www.cs.utexas.edu/

users/EWD/ewd03xx/EWD364.PDF.

Dijkstra:1981:SASa

[Dij82] Edsger W. Dijkstra. Smoothsort, an alternative for sorting in situ. Science of Computer Programming, 1(3):223–233, 1981/82. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). See errata [Dij82o].

Djojo:2013:CLA

[DK13] M. A. Djojo and K. Karyono. Computational load analysis of Dijkstra, A*, and Floyd–Warshall algorithms in mesh network. In 2013 IEEE International Conference on Robotics, Biomimetics, and Intelligent Computational Systems (RO-BIONETICS), pages 104–108. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2013. ISBN 1-4799-1206-9. URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?arnumber=6743587.

Dijkstra:1975:FGCc

[DLM^{+}75a] Edsger W. Dijkstra, Leslie Lamport, A. J. Martin, C. S. Scholten, and E. F. M. Steffens. On-the-fly garbage collection: an exercise in cooperation. Published as [DLM^{+}78].,

June 1975. URL http://www.cs.utexas.edu/users/EWD/
ewd04xx/EWD496B.PDF.

Dijkstra:1975:FGCd

[DLM^{+}75b] Edsger W. Dijkstra, Leslie Lamport, A. J. Martin, C. S. Scholten, and E. F. M. Steffens. On-the-fly garbage collection: an exercise in cooperation. Circulated privately., Oc-

tober 1975. URL http://www.cs.utexas.edu/users/EWD/
ewd05xx/EWD520.PDF.

Dijkstra:1975:FGCb

[DLM^{+}75c] Edsger W. Dijkstra, Leslie Lamport, A. J. Martin, C. S. Scholten, and E. F. M. Steffens. On-the-fly garbage collection: an exercise in cooperation (preliminary version). marked “old”; circulated privately, May 1975. URL http://www.cs.

utexas.edu/users/EWD/ewd04xx/EWD496A.PDF.

Dijkstra:1976:FGC

[DLM^{+}76] E. W. Dijkstra, Leslie Lamport, A. J. Martin, C. S. Scholten, and E. F. M. Steffens. On-the-fly garbage collection: an exercise in cooperation. In Bauer and Samelson [BS76], pages 43–56. ISBN 0-387-07994-7 (New York), 3-540-07994-7 (Berlin), 3-540-37972-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

QA76.6 .L335.  URL http://link.springer-ny.com/
link/service/series/0558/tocs/t0046.htm;   http:/
/www.springerlink.com/content/978-0-387-07994-3;
http://www.springerlink.com/openurl.asp?genre=book&
isbn=978-3-540-07994-1;   http://www.springerlink.
com/openurl.asp?genre=issue&issn=0302-9743&volume=
46.

Dijkstra:1977:FGCa

[DLM^{+}77a] Edsger W. Dijkstra, Leslie Lamport, A. J. Martin, C. S. Scholten, and E. F. M. Steffens. On-the-fly garbage collection: an exercise in cooperation. Circulated privately., Jan-

uary 1977. URL http://www.cs.utexas.edu/users/EWD/
ewd05xx/EWD595.PDF.

Dijkstra:1977:FGCb

[DLM^{+}77b] Edsger W. Dijkstra, Leslie Lamport, A. J. Martin, C. S. Scholten, and E. F. M. Steffens. On-the-fly garbage collection: an exercise in cooperation. Circulated pri-

vately., 1977. URL http://www.cs.utexas.edu/users/
EWD/ewd06xx/EWD630.PDF.

Dijkstra:1978:FGCa

[DLM^{+}78] Edsger W. Dijkstra, Leslie Lamport, Alain J. Martin, C. S. Scholten, and E. F. M. Steffens. On-the-fly garbage collection: An exercise in cooperation. Communications of the Association for Computing Machinery, 21(11):966–975, November 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Also E. W. Dijkstra Note EWD496, June 1975.

Doherty:1999:CMU

[DLMB99] P. Doherty, W. Lukaszewicz, and E. Madalinska-Bugaj. Computing MPMA updates using Dijkstra’s semantics. Lecture Notes in Computer Science, 1609:271–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeMillo:1978:RRD

[DLP78] Richard A. DeMillo, Richard J. Lipton, and Alan J. Perlis. Response from R. A. DeMillo, R. J. Lipton, A. J. Perlis. ACM SIGSOFT Software Engineering Notes, 3(2):16–17, April 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). See [Dij78u].

DeMillo:1979:SPP

[DLP79] Richard A. De Millo, Richard J. Lipton, and Alan J. Perlis. Social processes and proofs of theorems and programs. Communications of the Association for Computing Machinery, 22 (5):271–280, May 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See negative review [Dij77w, Dij78u]. Original paper prepublished in Conference Record of the 4th Symposium on Principles of Programming Languages, Los Angeles, January 17–19, 1977.

Denning:1997:BCN

[DM97] Peter J. Denning and Robert M. Metcalfe, editors. Beyond Calculation: The Next Fifty Years of Computing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 0-387-94932-1, 0-387-98588-3 (paperback), 1-4612-0685-5. xviii + 313 pp. LCCN QA76 .D348 1997.

Dijkstra:2000:DCP

[DM00] Edsger W. Dijkstra and Jayadev Misra. Designing a calculational proof of Cantor’s theorem. circulated privately; to appear in Am. Math. Monthly, 2000. URL http://www.cs.

utexas.edu/users/EWD/ewd12xx/EWD1294a.PDF.

Dijkstra:2001:DCP

[DM01] Edsger W. Dijkstra and Jayadev Misra. Designing a calculational proof of Cantor’s theorem. American Mathematical Monthly, 108(5):440–443, 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/2695799.

Dewar:1982:SMA

[DMS83] Robert B. K. Dewar, Susan M. Merritt, and Micha Sharir. Some modified algorithms for Dijkstra’s longest upsequence problem. Acta Informatica, 18(1):1–15, November 1982/83. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Dahl:2001:CSD

[DN01] Ole-Johan Dahl and Kristen Nygaard. Class and subclass declarations. In Broy and Denert [BD01b], pages 235–253. ISBN 3-540-42290-0 (print), 3-642-48354-2 (e-book). LCCN

QA76.758. URL http://link.springer.com/chapter/10.
1007/978-3-642-48354-7_8/.

Dijkstra:1990:CPV

[DR90a] Edsger W. Dijkstra and J. R. Rao. Constructing the proof of Vizing’s Theorem. Circulated privately., Febru-

ary 1990. URL http://www.cs.utexas.edu/users/EWD/
ewd10xx/EWD1075.PDF.

Dijkstra:1990:DPV

[DR90b] Edsger W. Dijkstra and J. R. Rao. Designing the proof of Vizing’s Theorem (with J. R. Rao). Circulated privately., Oc-

tober 1990. URL http://www.cs.utexas.edu/users/EWD/
ewd10xx/EWD1082a.PDF.

Dijkstra:1972:PMRa

[DS72a] Edsger W. Dijkstra and C. S. Scholten. Parallelism in multirecord transactions. Published as [DS72b]., 1972. URL http:

//www.cs.utexas.edu/users/EWD/ewd03xx/EWD338.PDF.

Dijkstra:1972:PMRb

[DS72b] Edsger W. Dijkstra and C. S. Scholten. Parallelism in multirecord transactions. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 15–21. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982. Including a paper co-authored by C. S. Schölten.

DuWorks:1972:APD

[DS72c] Robert J. DuWorks and Steven W. Smoliar. The arrogant programmer: Dijkstra and Wegner considered harmful. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 4(4):19–21, December 1972. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dijkstra:1976:CAP

[DS76a] Edsger W. Dijkstra and C. S. Scholten. Comments on Arbeitsblatt 3 from o. Prof. Dr. F. L. Bauer. Circulated

privately., 1976. URL http://www.cs.utexas.edu/users/
EWD/ewd06xx/EWD602.PDF.

Dijkstra:1976:PTC

[DS76b] Edsger W. Dijkstra and C. S. Scholten. A proof of a theorem communicated to us by S. Ghosh. Published as [DS82i]., Au-

gust 1976. URL http://www.cs.utexas.edu/users/EWD/
ewd05xx/EWD582.PDF.

Dijkstra:1977:CSC

[DS77] Edsger W. Dijkstra and C. S. Scholten. A class of simple communication patterns. Published as [DS82d]., 1977. URL

http://www.cs.utexas.edu/users/EWD/ewd06xx/EWD643.
PDF.

Dijkstra:1978:TDDa

[DS78a] Edsger W. Dijkstra and C. S. Scholten. Termination detection for diffusing computations. Circulated privately., 1978. URL

http://www.cs.utexas.edu/users/EWD/ewd06xx/EWD684.
PDF.

Dijkstra:1978:TDDb

[DS78b] Edsger W. Dijkstra and C. S. Scholten. Termination detection for diffusing computations. Circulated privately., 1978. URL

http://www.cs.utexas.edu/users/EWD/ewd06xx/EWD687.
PDF.

Dijkstra:1979:TDD

[DS79] Edsger W. Dijkstra and C. S. Scholten. Termination detection for diffusing computations. Published as [DS80b]., Jan-

uary 1979. URL http://www.cs.utexas.edu/users/EWD/
ewd06xx/EWD687a.PDF.

Dijkstra:1980:DAD

[DS80a] Edsger W. Dijkstra and C. S. Scholten. Distributed arbitration (DRAFT). Circulated privately., 1980. URL http://

www.cs.utexas.edu/users/EWD/ewd07xx/EWD752.PDF.

Dijkstra:1980:TDDa

[DS80b] Edsger W. Dijkstra and C. S. Scholten. Termination detection for diffusing computations. Information Processing Letters, 11(1):1–4, August 29, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dijkstra:1981:DRS

[DS81] Edsger W. Dijkstra and C. S. Scholten. Distances from the root in skew trees. Circulated privately., August 1981. URL

http://www.cs.utexas.edu/users/EWD/ewd08xx/EWD801.
PDF.

Dijkstra:1982:CSCa

[DS82a] E. W. Dijkstra and C. S. Scholten. A class of simple communication patterns. In Selected Writings on Computing. EWD643. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982.

Dijkstra:1982:PMR

[DS82b] E. W. Dijkstra and C. S. Scholten. Parallelism in multirecord transactions. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 15–21. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24

.D54 1982. URL http://link.springer.com/chapter/10.

1007/978-1-4612-5695-3_2. Including a paper co-authored by C. S. Schölten.

Dijkstra:1982:APT

[DS82c] Edsger W. Dijkstra and C. S. Scholten. About predicate transformers in general. Circulated privately., March 1982.

URL http://www.cs.utexas.edu/users/EWD/ewd08xx/EWD813.
PDF.

Dijkstra:1982:CSCb

[DS82d] Edsger W. Dijkstra and C. S. Scholten. A class of simple communication patterns. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 334–337. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982. Including a paper co-authored by C. S. Schölten.

Dijkstra:1982:DNT

[DS82e] Edsger W. Dijkstra and C. S. Scholten. On different notions of termination. Circulated privately., October 1982. URL http:

//www.cs.utexas.edu/users/EWD/ewd08xx/EWD830.PDF.

Dijkstra:1982:SEc

[DS82f] Edsger W. Dijkstra and C. S. Scholten. A sequel to EWD822. Circulated privately., May 1982. URL http://www.cs.

utexas.edu/users/EWD/ewd08xx/EWD823.PDF.

Dijkstra:1982:UQR

[DS82g] Edsger W. Dijkstra and C. S. Scholten. A universal quantification revisited. Circulated privately., November 1982. URL

http://www.cs.utexas.edu/users/EWD/ewd08xx/EWD844.
PDF.

Dijkstra:1982:WPL

[DS82h] Edsger W. Dijkstra and C. S. Scholten. Weakest preconditions, liberal and not. Circulated privately., April 1982. URL

http://www.cs.utexas.edu/users/EWD/ewd08xx/EWD816.
PDF.

Dijkstra:1982:PTCa

[DS82i] Prof. Dr. Edsger W. Dijkstra and Drs. C. S. Scholten. A proof of a theorem communicated to us by S. Ghosh. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 233–234. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982. URL http:/

/link.springer.com/accesspage/chapter/10.1007/978-
1-4612-5695-3_42?coverImageUrl=/static%2F0.9313%2Fsites%
2Flink%2Fimages%2Fabstract_cover_placeholder.png. In-

cluding a paper co-authored by C. S. Schölten.

Dijkstra:1983:JESa

[DS83a] Edsger W. Dijkstra and C. S. Scholten. Junctivity of extreme solutions. Circulated privately., April 1983. URL http://

www.cs.utexas.edu/users/EWD/ewd08xx/EWD849.PDF.

Dijkstra:1983:JESb

[DS83b] Edsger W. Dijkstra and C. S. Scholten. Junctivity of extreme solutions. Circulated privately., April 1983. URL http://

www.cs.utexas.edu/users/EWD/ewd08xx/EWD849a.PDF.

Dijkstra:1983:SRP

[DS83c] Edsger W. Dijkstra and C. S. Scholten. Semantics of recursive procedures. Circulated privately., November 1983. URL

http://www.cs.utexas.edu/users/EWD/ewd08xx/EWD859.
PDF.

Dijkstra:1983:SE

[DS83d] Edsger W. Dijkstra and C. S. Scholten. A sequel to EWD859. Circulated privately., December 1983. URL http://www.cs.

utexas.edu/users/EWD/ewd08xx/EWD860.PDF.

Dijkstra:1984:MR

[DS84a] Edsger W. Dijkstra and C. S. Scholten. A miscellany of results. Circulated privately., February 1984. URL http://

www.cs.utexas.edu/users/EWD/ewd08xx/EWD874.PDF.

Dijkstra:1984:OIE

[DS84b] Edsger W. Dijkstra and C. S. Scholten. The operational interpretation of extreme solutions. Circulated privately.,

April 1984. URL http://www.cs.utexas.edu/users/EWD/
ewd08xx/EWD883.PDF.

Dijkstra:1984:SE

[DS84c] Edsger W. Dijkstra and C. S. Scholten. A sequel to EWD874. Circulated privately., April 1984. URL http://www.cs.

utexas.edu/users/EWD/ewd08xx/EWD884.PDF.

Dijkstra:1986:FME

[DS86] Edsger W. Dijkstra and C. S. Scholten. On finite models for the equivalence. Circulated privately., July 1986. URL http:

//www.cs.utexas.edu/users/EWD/ewd09xx/EWD967.PDF.

Dijkstra:1990:CBSa

[DS90a] Edsger W. Dijkstra and Carel S. Scholten. The calculus of boolean structures. In Predicate Calculus and Program Semantics [DS90g], pages 30–80. ISBN 0-387-96957-

8. URL http://link.springer.com/chapter/10.1007/
978-1-4612-3228-5_5.

Dijkstra:1990:CPT

[DS90b] Edsger W. Dijkstra and Carel S. Scholten. Converse predicate transformers. In Predicate Calculus and Program Semantics [DS90g], pages 201–208. ISBN 0-387-96957-

8. URL http://link.springer.com/chapter/10.1007/
978-1-4612-3228-5_11.

Dijkstra:1990:EPT

[DS90c] Edsger W. Dijkstra and Carel S. Scholten. Equations in predicates and their extreme solutions. In Predicate Calculus and Program Semantics [DS90g], pages 147–170. ISBN

0-387-96957-8. URL http://link.springer.com/chapter/
10.1007/978-1-4612-3228-5_8.

Dijkstra:1990:OPF

[DS90d] Edsger W. Dijkstra and Carel S. Scholten. On our proof format. In Predicate Calculus and Program Semantics [DS90g], pages 21–29. ISBN 0-387-96957-8. URL http://link.

springer.com/chapter/10.1007/978-1-4612-3228-5_4.

Dijkstra:1990:SRa

[DS90e] Edsger W. Dijkstra and Carel S. Scholten. On substitution and replacement. In Predicate Calculus and Program Semantics [DS90g], pages 11–16. ISBN 0-387-96957-

8. URL http://link.springer.com/chapter/10.1007/
978-1-4612-3228-5_2.

Dijkstra:1990:OC

[DS90f] Edsger W. Dijkstra and Carel S. Scholten. Operational considerations. In Predicate Calculus and Program Semantics [DS90g], pages 190–200. ISBN 0-387-96957-8.

URL http://link.springer.com/chapter/10.1007/978-
1-4612-3228-5_10.

Dijkstra:1990:PCP

[DS90g] Edsger W. Dijkstra and Carel S. Scholten. Predicate Calculus and Program Semantics. Texts and Monographs in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. ISBN 0-387-96957-8. xii + 220 pp.

Dijkstra:1990:SRb

[DS90h] Edsger W. Dijkstra and Carel S. Scholten. Semantics of repetitions. In Predicate Calculus and Program Semantics [DS90g], pages 170–189. ISBN 0-387-96957-8.

URL http://link.springer.com/chapter/10.1007/978-
1-4612-3228-5_9.

Dijkstra:1990:SSLa

[DS90i] Edsger W. Dijkstra and Carel S. Scholten. Semantics of straight-line programs. In Predicate Calculus and Program Semantics [DS90g], pages 121–146. ISBN 0-387-96957-

8. URL http://link.springer.com/chapter/10.1007/
978-1-4612-3228-5_7.

Dijkstra:1990:SPP

[DS90j] Edsger W. Dijkstra and Carel S. Scholten. Some properties of predicate transformers. In Predicate Calculus and Program Semantics [DS90g], pages 81–120. ISBN 0-387-

96957-8. URL http://link.springer.com/chapter/10.
1007/978-1-4612-3228-5_6.

Dijkstra:1990:SP

[DS90k] Edsger W. Dijkstra and Carel S. Scholten. The strongest postcondition. In Predicate Calculus and Program Semantics [DS90g], pages 209–215. ISBN 0-387-96957-8.

URL http://link.springer.com/chapter/10.1007/978-
1-4612-3228-5_12.

Dijkstra:1990:FE

[DS90l] Edsger W. Dijkstra and Carel S. Schotten. On functions and equality. In Predicate Calculus and Program Semantics [DS90g], pages 17–20. ISBN 0-387-96957-8.

URL http://link.springer.com/chapter/10.1007/978-
1-4612-3228-5_3.

Dijkstra:1990:S

[DS90m] Edsger W. Dijkstra and Carel S. Schotten. On structures. In Predicate Calculus and Program Semantics [DS90g], pages 1–9. ISBN 0-387-96957-8. URL http://link.

springer.com/chapter/10.1007/978-1-4612-3228-5_1;
 http://link.springer.com/content/pdf/bfm%3A978-1-
4612-3228-5/1.pdf.

Dijkstra:2004:THA

[DS04] Atze Dijkstra and S. Doaitse Swierstra. Typing Haskell with an attribute grammar. In Vene and Uustalu [VU05], pages 1–72. CODEN LNCSD9. ISBN 3-540-28540-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

????  URL http://www.springerlink.com/content/
978-3-540-28540-3;   http://www.springerlink.com/
openurl.asp?genre=issue&issn=0302-9743&volume=3622.

Dijkstra:2006:RPT

[DS06] Atze Dijkstra and S. Doaitse Swierstra. Ruler: Programming type rules. In Hagiya and Wadler [HW06], pages 30–46. CODEN LNCSD9. ISBN 3-540-33438-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

????  URL http://www.springerlink.com/content/
978-3-540-33438-5;   http://www.springerlink.com/
openurl.asp?genre=issue&issn=0302-9743&volume=3945.

Du:2010:NRR

[Du10] Yanwei Du. Notice of retraction: Research on path planning of materials supply of lean production based on the Dijkstra algorithm. In 2010 3rd IEEE International Conference on Computer Science and Information Technology (ICCSIT), volume 9, pages 692–695. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2010.

ISBN 1-4244-5537-5. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=5565082.

Dijkstra:2006:ITO

[DvdS06] Fokke Dijkstra and Aad J. van der Steen. Integration of two ocean models within Cactus. Concurrency and Computation: Practice and Experience, 18(2):193–202, 2006. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dijkstra:1981:APPb

[DvG81a] Edsger W. Dijkstra and A. J. M. van Gasteren. About the presentation of programs. Circulated privately.,

March 1981. URL http://www.cs.utexas.edu/users/EWD/
ewd07xx/EWD781.PDF.

Dijkstra:1981:OPK

[DvG81b] Edsger W. Dijkstra and A. J. M. van Gasteren. Our presentation of K. A. Post’s proof of R. Stanley’s theorem. Circulated

privately., April 1981. URL http://www.cs.utexas.edu/
users/EWD/ewd07xx/EWD785.PDF.

Dijkstra:1981:PEE

[DvG81c] Edsger W. Dijkstra and A. J. M. van Gasteren. Proving the existence of the Euler line. Circulated privately.,

March 1981. URL http://www.cs.utexas.edu/users/EWD/
ewd07xx/EWD780.PDF.

Dijkstra:1981:WWS

[DvG81d] Edsger W. Dijkstra and A. J. M. van Gasteren. What we seem to have learned. Circulated privately., May 1981. URL http:

//www.cs.utexas.edu/users/EWD/ewd07xx/EWD788.PDF.

Dijkstra:1982:EPE

[DvG82a] Edsger W. Dijkstra and A. J. M. van Gasteren. Elementary predicates and elimination. Circulated privately., Septem-

ber 1982. URL http://www.cs.utexas.edu/users/EWD/
ewd08xx/EWD834.PDF.

Dijkstra:1982:FICa

[DvG82b] Edsger W. Dijkstra and A. J. M. van Gasteren. Fast image construction in computerized axial tomography (CAT). manuscript version; circulated privately, February 1982. URL

http://www.cs.utexas.edu/users/EWD/ewd08xx/EWD810.
PDF.

Dijkstra:1982:FICb

[DvG82c] Edsger W. Dijkstra and A. J. M. van Gasteren. Fast image construction in computerized axial tomography (CAT). typescript version; circulated privately, Febru-

ary 1982. URL http://www.cs.utexas.edu/users/EWD/
ewd08xx/EWD810a.PDF.

Dijkstra:1982:ITAa

[DvG82d] Edsger W. Dijkstra and A. J. M. van Gasteren. An introduction to three algorithms for sorting in situ. manuscript version; circulated privately, April 1982. URL http://www.

cs.utexas.edu/users/EWD/ewd08xx/EWD817.PDF.

Dijkstra:1982:ITAb

[DvG82e] Edsger W. Dijkstra and A. J. M. van Gasteren. An introduction to three algorithms for sorting in situ. Typescript version; published as [DvG82f]., 1982. URL http://www.cs.

utexas.edu/users/EWD/ewd08xx/EWD817a.PDF.

Dijkstra:1982:ITAd

[DvG82f] Edsger W. Dijkstra and A. J. M. van Gasteren. An introduction to three algorithms for sorting in situ. Information Processing Letters, 15(3):129–134, October 11, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dijkstra:1982:IET

[DvG82g] Edsger W. Dijkstra and A. J. M. van Gasteren. An introductory essay on three algorithms for sorting in situ. Circulated privately., January 1982. URL http://www.cs.utexas.edu/

users/EWD/ewd08xx/EWD809.PDF.

Dijkstra:1982:RN

[DvG82h] Edsger W. Dijkstra and A. J. M. van Gasteren. Remarks on notation. Circulated privately., March 1982. URL http://

www.cs.utexas.edu/users/EWD/ewd08xx/EWD815.PDF.

Dijkstra:1983:MLS

[DvG83a] Edsger W. Dijkstra and A. J. M. van Gasteren. The maximum length of a segment satisfying a monotonic predicate. Circulated privately., February 1983. URL http://www.cs.

utexas.edu/users/EWD/ewd08xx/EWD846.PDF.

Dijkstra:1983:RPO

[DvG83b] Edsger W. Dijkstra and A. J. M. van Gasteren. Research proposal: Our plans for the years to come. Circulated

privately., 1983. URL http://www.cs.utexas.edu/users/
EWD/ewd08xx/EWD861.PDF.

Dijkstra:1984:MA

[DvG84a] Edsger W. Dijkstra and A. J. M. van Gasteren. A monotonicity argument. Circulated privately. Also issued as Eindhoven University of Technology technical report AvG36., Febru-

ary 1984. URL http://www.cs.utexas.edu/users/EWD/
ewd08xx/EWD878.PDF.

Dijkstra:1984:SFPa

[DvG84b] Edsger W. Dijkstra and A. J. M. van Gasteren. A simple fix-point argument without the restriction to continuity. Circulated privately., June 1984. URL http://www.cs.utexas.

edu/users/EWD/ewd08xx/EWD886.PDF.

Dijkstra:1984:SFPb

[DvG84c] Edsger W. Dijkstra and A. J. M. van Gasteren. A simple fix-point argument without the restriction to continuity. Circulated privately., July 1984. URL http://www.cs.utexas.

edu/users/EWD/ewd08xx/EWD886a.PDF.

Dijkstra:1984:SFPc

[DvG84d] Edsger W. Dijkstra and A. J. M. van Gasteren. A simple fix-point argument without the restriction to continuity. Circulated privately., November 1984. URL http://www.cs.

utexas.edu/users/EWD/ewd09xx/EWD901.PDF.

Dijkstra:1984:SFPd

[DvG84e] Edsger W. Dijkstra and A. J. M. van Gasteren. A simple fix-point argument without the restriction to continuity. Circulated privately., November 1984. URL http://www.cs.

utexas.edu/users/EWD/ewd09xx/EWD901a.PDF.

Dijkstra:1984:SUF

[DvG84f] Edsger W. Dijkstra and A. J. M. van Gasteren. Some useful formulae. Circulated privately., February 1984. URL http:/

/www.cs.utexas.edu/users/EWD/ewd08xx/EWD877.PDF.

Dijkstra:1984:TTU

[DvG84g] Edsger W. Dijkstra and A. J. M. van Gasteren. The termination theorem for unconstrained nondeterminacy. Circulated privately., February 1984. URL http://www.cs.utexas.

edu/users/EWD/ewd08xx/EWD880b.PDF.

Dijkstra:1984:WICa

[DvG84h] Edsger W. Dijkstra and A. J. M. van Gasteren. Why the importance of continuity seems to be overrated. Circulated pri-

vately., February 1984. URL http://www.cs.utexas.edu/
users/EWD/ewd08xx/EWD880.PDF.

Dijkstra:1984:WICb

[DvG84i] Edsger W. Dijkstra and A. J. M. van Gasteren. Why the importance of continuity seems to be overrated. Circulated pri-

vately., February 1984. URL http://www.cs.utexas.edu/
users/EWD/ewd08xx/EWD880a.PDF.

Dijkstra:1985:N

[DvG85a] Edsger W. Dijkstra and A. J. M. van Gasteren. On notation. Circulated privately., December 1985. URL http://www.cs.

utexas.edu/users/EWD/ewd09xx/EWD950.PDF.

Dijkstra:1985:SFP

[DvG85b] Edsger W. Dijkstra and A. J. M. van Gasteren. A simple fix-point argument without the restriction to continuity. In Control flow and data flow: concepts of distributed programming (Marktoberdorf, 1984), volume 14 of NATO Adv. Sci. Inst. Ser. F Comput. Systems Sci., pages 519–525. Springer, Berlin, 1985.

Dijkstra:1986:Nb

[DvG86a] Edsger W. Dijkstra and A. J. M. van Gasteren. On naming. Circulated privately., May 1986. URL http://www.cs.

utexas.edu/users/EWD/ewd09xx/EWD958.PDF.

Dijkstra:1986:Na

[DvG86b] Edsger W. Dijkstra and A. J. M. van Gasteren. On notation. Circulated privately., January 1986. URL http://www.cs.

utexas.edu/users/EWD/ewd09xx/EWD950a.PDF.

Dijkstra:1986:SFA

[DvG86c] Edsger W. Dijkstra and A. J. M. van Gasteren. A simple fixpoint argument without the restriction to continuity. Acta Informatica, 23(1):1–7, March 1986. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Dijkstra:1990:WFL

[DvG90a] Edsger W. Dijkstra and A. J. M. van Gasteren. Wellfoundedness and lexical coupling (with A. J. M. van Gasteren). Circulated privately., September 1990. URL http:

//www.cs.utexas.edu/users/EWD/ewd10xx/EWD1080.PDF.

Dijkstra:1990:WFT

[DvG90b] Edsger W. Dijkstra and A. J. M. van Gasteren. Wellfoundedness and the transitive closure. Circulated privately.,

April 1990. URL http://www.cs.utexas.edu/users/EWD/
ewd10xx/EWD1079.PDF.

Dijkstra:1981:H

[DvGF81] Edsger W. Dijkstra, A. J. M. van Gasteren, and W. H. J. Feijen. Heapsort. Circulated privately., June 1981. URL http:

//www.cs.utexas.edu/users/EWD/ewd07xx/EWD794.PDF.

Dijkstra:1955:TEI

[DvW55] E. W. Dijkstra and A. van Wijngaarden. Table of Everett’s interpolation coefficients. Technical Report R294, Computation Department of the Mathematical Centre, Amsterdam, Amsterdam, The Netherlands, 1955. 204 pp. URL http://

books.google.com/books?id=UflRAQAAIAAJ.

Everts:2009:VWM

[EBR09] M. H. Everts, H. Bekker, and J. B. T. M. Roerdink. Visualizing white matter structure of the brain using Dijkstra’s algorithm. In Proceedings of 6th International Symposium on Image and Signal Processing and Analysis, 2009. ISPA 2009, pages 569–574. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2009. ISBN 953-184-135-7. ISSN 1845-5921. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?arnumber=5297652.

Eckert:1980:E

[Eck80] J. Presper Eckert, Jr. The ENIAC. In Metropolis et al. [MHR80], pages 525–539. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5 .I63 1976.

URL http://www.sciencedirect.com/science/article/

pii/B9780124916500500381. Original versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Edsger:1982:PSE

[Eds82] Edsger. A post-scriptum to EWD501. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 141– 144. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982. URL http://link.springer.

com/chapter/10.1007/978-1-4612-5695-3_24. Including a

paper co-authored by C. S. Schölten.

EindhovenTuesdayAfternoonClub:1990:PMN

[Ein90] The Eindhoven Tuesday Afternoon Club. The problem of the majority network. In Feijen et al. [FvGGM90], chapter 12, pages 112–118. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X.

LCCN QA76 .B326 1990. URL http://www.zentralblatt-
math.org/zmath/en/search/?an=0718.68004. Contains

important treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

Eklund:1996:DMS

[EKP96] P. W. Eklund, S. Kirkby, and S. Pollitt. A dynamic multisource Dijkstra’s algorithm for vehicle routing. In Australian and New Zealand Conference on Intelligent Information Systems, 1996, pages 329–333. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-7803-3667-4. URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?arnumber=573976.

Eraghi:2014:PLC

[ELCdCG14] N. O. Eraghi, F. Lopez-Colino, A. de Castro, and J. Garrido. Path length comparison in grid maps of planning algorithms: HCTNav, A and Dijkstra. In 2014 Conference on Design of Circuits and Integrated Circuits (DCIS), pages 1–6. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2014. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?arnumber=7035557.

Eisenberg:1972:FCD

[EM72] Murray A. Eisenberg and Michael R. McGuire. Further comments on Dijkstra’s concurrent programming control problem. Communications of the Association for Computing Machinery, 15(11):999, November 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ershov:1980:EDP

[ESB80] Andrei P. Ershov and Mikhail R. Shura-Bura. The early development of programming in the USSR. In Metropolis et al. [MHR80], pages 137–196. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5 .I63 1976.

URL http://www.sciencedirect.com/science/article/

pii/B9780124916500500186. Original versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Everett:1980:W

[Eve80] Robert R. Everett. WHIRLWIND. In Metropolis et al. [MHR80], pages 365–384. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5 .I63 1976. URL http://www.

sciencedirect.com/science/article/pii/B9780124916500500253.

Original versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Fagan:2001:ASI

[Fag01a] Michael E. Fagan. Advances in software inspections. In Broy and Denert [BD01b], pages 335–360. ISBN 3-540-42290-0 (print), 3-642-48354-2 (e-book). LCCN QA76.758.

URL http://link.springer.com/chapter/10.1007/978-
3-642-48354-7_14/.

Fagan:2001:DCI

[Fag01b] Michael E. Fagan. Design and code inspections to reduce errors in program development. In Broy and Denert [BD01b], pages 301–334. ISBN 3-540-42290-0 (print), 3-642-48354-2 (e-book). LCCN QA76.758. URL http://link.springer.

com/chapter/10.1007/978-3-642-48354-7_13/.

Ferariu:2014:PGP

[FC14] L. Ferariu and C. Cimpanu. Pareto genetic path planning hybridized with multi-objective Dijkstra’s algorithm. In 2014 18th International Conference System Theory, Control and Computing (ICSTCC), pages 341–346. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD

20910, USA, 2014. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=6982439.

Fabbri:2008:EDT

[FCTB08] Ricardo Fabbri, Luciano Da F. Costa, Julio C. Torelli, and Odemir M. Bruno. 2D Euclidean distance transform algorithms: a comparative survey. ACM Computing Surveys, 40 (1):2:1–2:44, February 2008. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Feijen:1990:LED

[Fei90] W. H. J. Feijen. A little exercise in deriving multiprograms. In Feijen et al. [FvGGM90], chapter 13, pages 119–126.

ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN QA76 .B326

1990. URL http://www.zentralblatt-math.org/zmath/

en/search/?an=0718.68004. Contains important treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

Fuhao:2009:ASP

[FJ09] Zhang Fuhao and Liu Jiping. An algorithm of shortest path based on Dijkstra for huge data. In Sixth International Conference on Fuzzy Systems and Knowledge Discovery, 2009. FSKD ’09, volume 4, pages 244–247. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2009. ISBN 0-7695-3735-9. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?arnumber=5359145.

Floyd:1977:BRB

[Flo77] Robert W. Floyd. Book review: A Discipline of Programming. Series in Automatic Computation, by Edsger W. Dijkstra. American Scientist, 65(6):785, November/December 1977. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-

2786 (electronic). URL http://www.jstor.org/stable/
27848237.

Fujita:2003:DDS

[FNS03] Y. Fujita, Y. Nakamura, and Z. Shiller. Dual Dijkstra search for paths with different topologies. In ICRA ’03. IEEE International Conference on Robotics and Automation, 2003. Proceedings, volume 3, pages 3359–3364 vol.3. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2003. ISBN 0-7803-7736-2. ISSN 1050-

4729. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=1242109.

Fan:2010:IDA

[FS10] DongKai Fan and Ping Shi. Improvement of Dijkstra’s algorithm and its application in route planning. In 2010 Seventh International Conference on Fuzzy Systems and Knowledge Discovery (FSKD), volume 4, pages 1901–1904. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2010. ISBN 1-4244-5931-

1. URL http://ieeexplore.ieee.org/stamp/stamp.jsp?
arnumber=5569452.

Furculita:2013:IIM

[FURD13] A. G. Furculita, M.-V. Ulinic, A. B. Rus, and V. Dobrota. Implementation issues for modified Dijkstra’s and floyd-warshall algorithms in OpenFlow. In 2013 RoEduNet International Conference 12th Edition Networking in Education and Research, pages 1–6. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2013. ISBN 1-4799-2599-3. ISSN 2068-1038. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?arnumber=6714208.

Feijen:1990:BOB

[FvGGM90] W. H. J. Feijen, A. J. M. van Gasteren, David Gries, and J. Misra, editors. Beauty is our Business: a Birthday Salute to Edsger W. Dijkstra. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. xix + 453 pp. LCCN QA76 .B326

1990. URL http://www.zentralblatt-math.org/zmath/

en/search/?an=0718.68004. Contains important treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

Guzdial:2021:DDB

[GA21] Mark Guzdial and Joel C. Adams. Disputing Dijkstra, and birthdays in base 2. Communications of the Association for Computing Machinery, 64(3):12–13, March 2021. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/doi/10.1145/3446806.

Gallaher:1973:CDG

[Gal73] L. J. Gallaher. Correspondence: On Dijkstra and the GOTO statement. The Computer Journal, 16(3):284–285, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.uk/

computer_journal/hdb/Volume_16/Issue_03/tiff/284.
tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_16/Issue_03/tiff/285.tif.

Gawlik:1964:MRP

[Gaw64] H. J. Gawlik. MIRFAC: a reply to Professor Dijkstra. Communications of the Association for Computing Machinery, 7 (10):571, October 1964. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gritzner:1996:RAM

[GB96] Thomas F. Gritzner and Rudolf Berghammer. A relation algebraic model of robust correctness. Theoretical Computer Science, 159(2):245–270, June 03, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294

(electronic). URL http://www.elsevier.com/cgi-bin/
cas/tree/store/tcs/cas_sub/browse/browse.cgi?year=
1996&volume=159&issue=2&aid=2001.

Galan-Garcia:2015:NPE

[GGAVGGRC15] José L. Galán-García, Gabriel Aguilera-Venegas, María Á. Galán-García, and Pedro Rodríguez-Cielos. A new Probabilistic Extension of Dijkstra’s Algorithm to simulate more realistic traffic flow in a smart city. Applied Mathematics and Computation, 267(??):780–789, September 15, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0096300314016087.

Gamma:2001:DPA

[GHJV01] Erich Gamma, Richard Helm, Ralph Johnson, and John Vlissides. Design patterns: Abstraction and reuse of objectoriented design. In Broy and Denert [BD01b], pages 361–388. ISBN 3-540-42290-0 (print), 3-642-48354-2 (e-book). LCCN

QA76.758. URL http://link.springer.com/chapter/10.
1007/978-3-642-48354-7_15/.

Gao:1998:DAL

[GL98] Cheng Xiu Gao and Jun Liu. Dijkstra algorithms for largescale nonlinear network models. J. Math. (Wuhan), 18 (suppl.):37–40, 1998. ISSN 0255-7797.

Garcia:2007:PSP

[GLFM07] N. M. Garcia, P. Lenkiewicz, M. M. Freire, and P. P. Monteiro. On the performance of shortest path routing algorithms for modeling and simulation of static source routed networks — an extension to the Dijkstra algorithm. In Oana Dini, editor, Second International Conference on Systems and Networks Communications, 2007: ICSNC 2007; 25–31 August 2007, Cap Esterel, France; [including] HPC-Bio 2007, the First International Workshop on High Performance Computing Applied to Medical Data and Bioinformatics, page 60.

IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2007. ISBN 0-7695-2938-

0. URL http://ieeexplore.ieee.org/stamp/stamp.jsp?
arnumber=4300032.

Gold:2002:ATA

[Gol02] Virginia Gold. ACM Turing Award winner Edsger Dijkstra dies at 72: Internationally acclaimed computing pioneer known for wit and wisdom. Web site., Au-

gust 9, 2002. URL http://www.acm.org/announcements/
dijkstraACMobit.html.

Good:1980:PWC

[Goo80] I. J. Good. Pioneering work on computers at Bletchley. In Metropolis et al. [MHR80], pages 31–45. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5 .I63 1976.

URL http://www.sciencedirect.com/science/article/

pii/B9780124916500500125. Original versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Gouda:1990:SPP

[Gou90] Mohamed G. Gouda. Serializable programs, parallelizable assertions: a basis for interleaving. In Feijen et al. [FvGGM90], chapter 15, pages 135–140. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN QA76 .B326 1990. URL

http://www.zentralblatt-math.org/zmath/en/search/?

an=0718.68004. Contains important treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

Gries:1990:BDO

[Gri90] David Gries. Binary to decimal, one more time. In Feijen et al. [FvGGM90], chapter 16, pages 141–148. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN QA76 .B326

1990. URL http://www.zentralblatt-math.org/zmath/

en/search/?an=0718.68004. This paper presents an alternate proof of Knuth’s algorithm [Knu90] for conversion between decimal and fixed-point binary numbers.

Goldberg:1997:IDA

[GS97] Andrew V. Goldberg and Craig Silverstein. Implementations of Dijkstra’s algorithm based on multi-level buckets. In Network optimization (Gainesville, FL, 1996), volume 450 of Lecture Notes in Econom. and Math. Systems, pages 292–327. Springer, Berlin, 1997.

Habermann:1990:RD

[Hab90] A. N. Habermann. Rotate and double. In Feijen et al. [FvGGM90], chapter 17, pages 149–162. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN QA76 .B326 1990. URL

http://www.zentralblatt-math.org/zmath/en/search/?

an=0718.68004. Contains important treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

Hamming:1980:WWK

[Ham80] R. W. Hamming. We would know what they thought when they did it. In Metropolis et al. [MHR80], pages 3–9. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5

.I63 1976. URL http://www.sciencedirect.com/science/
article/pii/B9780124916500500095. Original versions of

these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Harel:1981:TCN

[Har81] D. Harel. On the total correctness of nondeterministic programs. Theoretical Computer Science, 13(2):175–192, February 1981. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Hart:1995:ELC

[Har95] Johnson M. Hart. Experience with logical code analysis in software maintenance. Software — Practice and Experience, 25(11):1243–1262, November 1995. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Harrison:1998:FD

[Har98] John Harrison. Formalizing Dijkstra. Lecture Notes in Computer Science, 1479:171–188, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hartley:1999:EAC

[Har99] David Hartley. EDSAC 1 and after — a compilation of personal reminiscences. Cambridge University Web site, July 21,

1999. URL https://www.cl.cam.ac.uk/events/EDSAC99/
reminiscences/.

Hoogendijk:1994:RPL

[HB94] Paul F. Hoogendijk and Roland C. Backhouse. Relational programming laws in the tree, list, bag, set hierarchy. Science of Computer Programming, 22(1–2):67–105, April 1994. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). Mathematics of program construction (Oxford, 1992).

Hu:2009:ASR

[HCSW09] Yi Hu, Zhiying Chang, Liying Sun, and Yi Wang. Analysis of the shortest repaired path of distribution network based on Dijkstra algorithm. In International Conference on Energy and Environment Technology, 2009. ICEET ’09, volume 2, pages 73–76. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2009.

ISBN 0-7695-3819-3. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=5362929.

Hehner:1990:BG

[Heh90] Eric C. R. Hehner. Beautifying Gödel. In Feijen et al. [FvGGM90], chapter 18, pages 163–172. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN QA76 .B326 1990. URL

http://www.zentralblatt-math.org/zmath/en/search/?

an=0718.68004. Contains important treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

Helmberg:1990:SE

[Hel90] G. Helmberg. A striptease of entropy. In Feijen et al. [FvGGM90], chapter 19, pages 173–175. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN QA76 .B326 1990. URL

http://www.zentralblatt-math.org/zmath/en/search/?

an=0718.68004. Contains important treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

Herman:1990:TJ

[Her90] Ted Herman. On a theorem of Jacobson. In Feijen et al. [FvGGM90], chapter 20, pages 176–181. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN QA76 .B326

1990. URL http://www.zentralblatt-math.org/zmath/

en/search/?an=0718.68004. Contains important treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

Hesselink:1990:MN

[Hes90] Wim H. Hesselink. Modalities of nondeterminacy. In Feijen et al. [FvGGM90], chapter 21, pages 182–192. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN QA76 .B326

1990. URL http://www.zentralblatt-math.org/zmath/

en/search/?an=0718.68004. Contains important treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

Hesselink:1995:ATD

[Hes95] Wim H. Hesselink. Angelic termination in Dijkstra’s calculus. Lecture Notes in Computer Science, 947:322–334, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Han:1994:FRD

[HHM94] Y. S. Han, C. R. P. Hartmann, and K. G. Mehrotra. Further results on decoding linear block codes using a generalized Dijkstra’s algorithm. In 1994 IEEE International Symposium on Information Theory, 1994. Proceedings, pages 342– ?? IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-7803-2015-

8. URL http://ieeexplore.ieee.org/stamp/stamp.jsp?
arnumber=394676.

Helgason:1993:OOS

[HKS93] Richard V. Helgason, Jeffery L. Kennington, and B. Douglas Stewart. The one-to-one shortest-path problem: an empirical analysis with the two-tree Dijkstra algorithm. Computational Optimization and Applications, 2(1):47–75, 1993. CODEN CPPPEF. ISSN 0926-6003 (print), 1573-2894 (electronic).

Hofner:2012:DFW

[HM12] Peter Höfner and Bernhard Möller. Dijkstra, Floyd and Warshall meet Kleene. Formal Aspects of Computing, 24(4–6): 459–476, 2012. CODEN FACME5. ISSN 0934-5043 (print), 1433-299x (electronic).

Hoare:1978:CSP

[Hoa78a] C. A. R. Hoare. Communicating sequential processes. Communications of the Association for Computing Machinery, 21(8):666–677, August 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See corrigendum [Hoa78b].

Hoare:1978:CCS

[Hoa78b] C. A. R. Hoare. Corrigendum: “Communicating Sequential Processes”. Communications of the Association for Computing Machinery, 21(11):958, November 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Hoa78a].

Hoare:1982:VTO

[Hoa82] Professor C. A. R. Hoare. Variations on a theme: An open letter to C. A. R. Hoare. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 132– 140. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982. URL http://link.springer.

com/chapter/10.1007/978-1-4612-5695-3_23. Including a

paper co-authored by C. S. Schölten.

Hoare:1985:CSP

[Hoa85] C. A. R. Hoare. Communicating Sequential Processes. Prentice Hall International Series in Computer Science. Prentice-Hall International, Englewood Cliffs, NJ, USA and London, UK, 1985. ISBN 0-13-153271-5 (hardcover), 0-13-153289-8 (paperback). viii + 256 pp. LCCN QA76.6 .H57 1985.

URL https://dl.acm.org/doi/book/10.5555/3921. With

a foreword by Edsger W. Dijkstra.

Hoare:1990:TDC

[Hoa90] C. A. R. Hoare. A theory for the derivation of C-mos circuit designs. In Feijen et al. [FvGGM90], chapter 22, pages 193–205. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN

QA76 .B326 1990. URL http://www.zentralblatt-math.
org/zmath/en/search/?an=0718.68004. Contains impor-

tant treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

Hoare:2001:ABC

[Hoa01a] C. A. R. Hoare. An axiomatic basis for computer programming. In Broy and Denert [BD01b], pages 419–438. ISBN 3-540-42290-0 (print), 3-642-48354-2 (e-book). LCCN

QA76.758. URL http://link.springer.com/chapter/10.
1007/978-3-642-48354-7_17/.

Hoare:2001:PCD

[Hoa01b] Prof. C. A. R. Hoare. Proof of correctness of data representations. In Broy and Denert [BD01b], pages 439–451. ISBN 3-540-42290-0 (print), 3-642-48354-2 (e-book). LCCN

QA76.758. URL http://link.springer.com/chapter/10.
1007/978-3-642-48354-7_18/.

Hoare:2003:EWD

[Hoa03] Tony Hoare. Edsger Wybe Dijkstra. Physics Today, 56(3):96, 98, March 2003. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Hoare:2022:FYE

[Hoa22] Tony Hoare. Forty years with Edsger. In Krzysztof R. Apt and Tony Hoare, editors, Edsger Wybe Dijkstra: His Life, Work, and Legacy, pages 411–422. ACM / Morgan & Claypool, ????, 2022.

Hoogerwoord:1990:MII

[Hoo90] Rob Hoogerwoord. On mathematical induction and the invariance theorem. In Feijen et al. [FvGGM90], chapter 23, pages 206–211. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X.

LCCN QA76 .B326 1990. URL http://www.zentralblatt-
math.org/zmath/en/search/?an=0718.68004. Contains

important treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

Horning:1990:FSC

[Hor90] J. J. Horning. Formalizing some classic synchronization primitives. In Feijen et al. [FvGGM90], chapter 24, pages 212–219. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN QA76 .B326

1990. URL http://www.zentralblatt-math.org/zmath/

en/search/?an=0718.68004. Contains important treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

Householder:1980:ROR

[Hou80] A. S. Householder. Reminiscences of Oak Ridge. In Metropolis et al. [MHR80], pages 385–388. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5 .I63 1976.

URL http://www.sciencedirect.com/science/article/

pii/B9780124916500500265. Original versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Hoare:1972:OST

[HP72] C. A. R. (Charles Antony Richard) Hoare and Ronald H. Perrott, editors. Operating systems techniques; proceedings, volume 9 of A.P.I.C. studies in data processing. Academic Press, New York, NY, USA, 1972. ISBN 0-12-350650-6. LCCN QA76.6 .I57 1971.

He:1997:PMG

[HSM97] Jifeng He, K. Seidel, and A. McIver. Probabilistic models for the guarded command language. Science of Computer Programming, 28(2–3):171–192, April 1997. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). Formal specifications: foundations, methods, tools and applications (Konstancin, 1995).

Holzer:2008:EMO

[HSW08] Martin Holzer, Frank Schulz, and Dorothea Wagner. Engineering multilevel overlay graphs for shortest-path queries. ACM Journal of Experimental Algorithmics, 13:5:1–5:??, September 2008. CODEN ???? ISSN 1084-6654.

Holzer:2005:CST

[HSWW05] Martin Holzer, Frank Schulz, Dorothea Wagner, and Thomas Willhalm. Combining speed-up techniques for shortest-path computations. ACM Journal of Experimental Algorithmics, 10:2.5:1–2.5:??, ???? 2005. CODEN ???? ISSN 1084-6654.

Hazel:2008:TCL

[HTVW08] Thomas Hazel, Laura Toma, Jan Vahrenhold, and Rajiv Wickremesinghe. Terracost: Computing least-cost-path surfaces for massive grid terrains. ACM Journal of Experimental Algorithmics, 12:1.9:1–1.9:??, June 2008. CODEN ???? ISSN 1084-6654.

Hurd:1980:CDI

[Hur80] Cuthbert C. Hurd. Computer development at IBM. In Metropolis et al. [MHR80], pages 389–418. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5 .I63 1976.

URL http://www.sciencedirect.com/science/article/

pii/B9780124916500500277. Original versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Huskey:1980:SNB

[Hus80] Harry D. Huskey. The SWAC: The National Bureau of Standards Western Automatic Computer. In Metropolis et al. [MHR80], pages 419–431. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5 .I63 1976.

URL http://www.sciencedirect.com/science/article/

pii/B9780124916500500289. Original versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Hagiya:2006:FLP

[HW06] Masami Hagiya and Philip Wadler, editors. Functional and Logic Programming: 8th International Symposium, FLOPS 2006, Fuji-Susono, Japan, April 24–26, 2006. Proceedings, volume 3945 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. CODEN LNCSD9. ISBN 3-540-33438-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

????  URL http://www.springerlink.com/content/
978-3-540-33438-5;   http://www.springerlink.com/
openurl.asp?genre=issue&issn=0302-9743&volume=3945.

Huang:2009:ODA

[HZW09] Yihu Huang, Genmin Zhang, and Jinli Wang. An optimization Dijkstra algorithm based on two-function limitation strategy. In Peihua Qiu et al., editors, Proceedings of the 2009 2nd International Congress on Image and Signal Processing: [October 17–19, 2009, Tianjin, China]; . CISP ’09, pages 1– 4. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2009. ISBN 1-4244-4117-X, 1-4244-4129-3, 1-4244-4131-5. LCCN TA1637 .I46

2009. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=5300828.

IJsselsteijn:2006:PTF

[IdKM06] Wijnand IJsselsteijn, Yvonne de Kort, and Cees Midden, editors. Persuasive Technology: First International Conference on Persuasive Technology for Human Well-Being, PERSUA-SIVE 2006, Eindhoven, The Netherlands, May 2006, Proceedings, volume 3962 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 3-540-34291-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

BF637.P4 I58 2006. URL http://www.springerlink.com/
content/978-3-540-34291-5.

IEEE:1979:ICS

[IEE79a] IEEE, editor. 4th International Conference on Software Engineering (ICSE ’79). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, September 1979.

IEEE:1979:PIC

[IEE79b] IEEE, editor. Proceedings: 4th International Conference on Software Engineering. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1979.

Ibrahim:2004:MCA

[ITOK04] Z. Ibrahim, Y. Tsuboi, O. Ono, and M. Khalid. Molecular computation approach to compete Dijkstra’s algorithm. In 5th Asian Control Conference, 2004, volume 1, pages 635– 642 Vol.1. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2004. ISBN 0-7803-

8873-9. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=1426021.

Jackson:2001:CMP

[Jac01] Michael A. Jackson. Constructive methods of program design. In Broy and Denert [BD01b], pages 453–478. ISBN 3-540-42290-0 (print), 3-642-48354-2 (e-book). LCCN

QA76.758. URL http://link.springer.com/chapter/10.
1007/978-3-642-48354-7_19/.

Jasika:2012:DSP

[JAE^{+}12] N. Jasika, N. Alispahic, A. Elma, K. Ilvana, L. Elma, and N. Nosovic. Dijkstra’s shortest path algorithm serial and parallel execution performance analysis. In 2012 Proceedings of the 35th International Convention MIPRO, pages 1811–1815. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2012. ISBN 1-4673-2577-

5. URL http://ieeexplore.ieee.org/stamp/stamp.jsp?
arnumber=6240942.

Jeffrey:1978:DPP

[Jef78] H. J. Jeffrey. On Dijkstra’s position paper on software reliability. ACM SIGSOFT Software Engineering Notes, 3(2): 18, April 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). See [Dij78u].

Jifeng:1993:AOS

[JH93] He Jifeng and C. A. R. Hoare. From algebra to operational semantics. Information Processing Letters, 45(2):75–80, February 26, 1993. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jiang:2014:EDS

[JHLC14] Jehn-Ruey Jiang, Hsin-Wen Huang, Ji-Hau Liao, and Szu-Yuan Chen. Extending Dijkstra’s shortest path algorithm for software defined networking. In 2014 16th Asia-Pacific Network Operations and Management Symposium (APNOMS), pages 1–4. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA,

2014. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=6996609.

Jones:1980:CFA

[JM80] Neil D. Jones and Steven S. Muchnick. Complexity of flow analysis, inductive assertion synthesis and a language due to Dijkstra. In 21st Annual Symposium on Foundations of Computer Science, 1980, pages 185–190. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1980. ISSN 0272-5428. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?arnumber=4567818.

Juzoji:2011:DNT

[JNK11] H. Juzoji, I. Nakajima, and T. Kitano. A development of network topology of wireless packet communications for disaster situation with genetic algorithms or with Dijkstra’s. In IEEE, editor, 2011 IEEE International Conference on Communications (ICC): ICC 2011: 5–9 June, 2010, Kyoto, Japan, pages 1–5. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2011. ISBN 1-61284-231-3, 1-61284-233-X, 1-61284-232-1.

LCCN TK5101.A1. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=5962439.

Johnson:1973:NDS

[Joh73] Donald B. Johnson. A note on Dijkstra’s shortest path algorithm. Journal of the Association for Computing Machinery, 20(3):385–388, July 1973. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Jones:1990:C

[Jon90] Cliff B. Jones. Consequences. In Feijen et al. [FvGGM90], chapter 25, pages 220–225. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN QA76 .B326 1990. URL

http://www.zentralblatt-math.org/zmath/en/search/?

an=0718.68004. Contains important treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

Janssen:2004:SIC

[JSW04] Marijn Janssen, Henk G. Sol, and René W. Wagenaar, editors. Sixth international conference on electronic commerce: ICEC04: towards a new service landscape, Delft, The Netherlands 25–27 October 2004. ACM Press, New York, NY 10036, USA, 2004. ISBN 1-58113-930-6. LCCN ????

Kaldewaij:1990:SLS

[Kal90] Anne Kaldewaij. Shortest and longest segments. In Feijen et al. [FvGGM90], chapter 26, pages 226–232. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN QA76 .B326

1990. URL http://www.zentralblatt-math.org/zmath/

en/search/?an=0718.68004. Contains important treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

Kessels:1988:EPS

[Kes88] J. L. W. Kessels. An exercise in proving self-stabilization with a variant function. Information Processing Letters, 29(1):39– 42, September 15, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kuznetsov:2008:DAI

[KF08] N. A. Kuznetsov and V. N. Fetisov. Dijkstra’s algorithm with improved robustness for the control of routing in IP networks. Automation and Remote Control, 2:80–85, 2008. CODEN AURCAT. ISSN 0005-2310 (print), 2413-9777 (electronic).

Khachiyan:2006:EDA

[KGZ06] Leonid Khachiyan, Vladimir Gurvich, and Jihui Zhao. Extending Dijkstra’s algorithm to maximize the shortest path by node-wise limited arc interdiction. In Computer science — theory and applications, volume 3967 of Lecture Notes in Comput. Sci., pages 221–234. Springer, Berlin, 2006.

Kilov:1979:LPV

[Kil79] H. Kilov. Letter on Parnas’ view of Dijkstra vs. deMillo, Lipton and Perlis. ACM SIGSOFT Software Engineering Notes, 4(1):19, January 1979. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Koutsopoulos:2014:DGS

[KKNI14] Iordanis Koutsopoulos, Iordanis Koutsopoulos, Eugenia Noutsi, and George Iosifidis. Dijkstra goes social: Socialgraph-assisted routing in next generation wireless networks. In Proceedings of European Wireless 2014; 20th European Wireless Conference, pages 1–7. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2014. ISBN 3-8007-3621-7. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?arnumber=6843155.

Kang:2008:PPA

[KLK08] Hwan Il Kang, Byunghee Lee, and Kabil Kim. Path planning algorithm using the particle swarm optimization and the improved Dijkstra algorithm. In Pacific-Asia Workshop on Computational Intelligence and Industrial Application, 2008. PACIIA ’08, volume 2, pages 1002–1004. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2008. ISBN 0-7695-3490-

2. URL http://ieeexplore.ieee.org/stamp/stamp.jsp?
arnumber=4756927.

Kong:2011:IRD

[KLMZ11] Dechuan Kong, Yunjuan Liang, Xiaoqin Ma, and Lijun Zhang. Improvement and realization of Dijkstra algorithm in GIS of depot. In Automation and Systems Engineering (CASE), 2011 International Conference on Control, pages 1– 4. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2011. ISBN 1-4577-0859-

0. URL http://ieeexplore.ieee.org/stamp/stamp.jsp?
arnumber=5997763.

Koparkar:1983:DPP

[KM83] P. A. Koparkar and S. P. Mudur. The development of programs for the processing of parametric curves. Computer Graphics Forum, 2(2-3):135–144, August 1983. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Knuth:1966:EAC

[Knu66a] Donald E. Knuth. Errata: “Additional comments on a problem in concurrent programming control”. Communications of the Association for Computing Machinery, 9(12):878, December 1966. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [?].

Knuth:1966:LEA

[Knu66b] Donald E. Knuth. Letter to the Editor: Additional comments on a problem in concurrent programming control. Communications of the Association for Computing Machinery, 9(5):321–322, May 1966. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Dij65h]. See errata [Knu66a]. Reprinted in Comm. ACM 26 (January 1983), 22.

Knuth:1974:SPG

[Knu74] Donald E. Knuth. Structured programming with go to statements. ACM Computing Surveys, 6(4):261–301, December 1974. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). Reprinted with revisions in Current Trends in Programming Methodology, Raymond T. Yeh, ed., 1 (Englewood Cliffs, NJ: Prentice-Hall, 1977), 140–194; Classics in Software Engineering, Edward Nash Yourdon, ed. (New York: Yourdon Press, 1979), 259–321. Reprinted with “final” revisions in [Knu92, pp. 17–89]. This paper is a response to [Dij68s].

Knuth:1977:GDA

[Knu77] Donald E. Knuth. A generalization of Dijkstra’s algorithm. Information Processing Letters, 6(1):1–5, February ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Knuth:1990:SPW

[Knu90] Donald E. Knuth. A simple program whose proof isn’t. In Feijen et al. [FvGGM90], chapter 27, pages 233–242. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN QA76 .B326

1990. URL http://www.zentralblatt-math.org/zmath/

en/search/?an=0718.68004. Contains important treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

Knuth:1992:LP

[Knu92] Donald E. Knuth. Literate Programming, volume 27 of CSLI Lecture Notes. Stanford University Center for the Study of Language and Information, Stanford, CA, USA, 1992. ISBN 0-937073-80-6 (paperback), 0-937073-81-4 (hardcover). xv + 368 pp. LCCN QA76.6.K644. URL http://csli-www.

stanford.edu/publications/literate.html.

Koh:1976:SPP

[Koh76] K. M. Koh. On the shortest-path problem and Dijkstra’s algorithm. Bull. Malaysian Math. Soc., 7(1):46–53. (1 loose figure), 1976. ISSN 0126-6705.

Kosaraju:1973:LDS

[Kos73] S. Rao Kosaraju. Limitations of Dijkstra’s Semaphore Primitives and Petri nets. Operating Systems Review, 7(4):122–126, October 1973. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kotov:1990:BSB

[Kot90] Vadim E. Kotov. Binding structure and behaviour in whole net concurrency semantics. In Feijen et al. [FvGGM90], chapter 28, pages 243–250. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X.

LCCN QA76 .B326 1990. URL http://www.zentralblatt-
math.org/zmath/en/search/?an=0718.68004. Contains

important treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

Knuth:1980:EDP

[KP80] Donald E. Knuth and Luis Trabb Pardo. The early development of programming languages. In Metropolis et al. [MHR80], pages 197–273. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5 .I63 1976.

URL http://www.sciencedirect.com/science/article/

pii/B9780124916500500198. Original versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Kim:2009:IHT

[KP09] Tae-Hwan Kim and In-Cheol Park. Implementation of a high-throughput and area-efficient MIMO detector based on modified Dijkstra’s search. In IEEE Global Telecommunications Conference, 2009. GLOBECOM 2009, pages 1–6. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2009. ISBN 1-4244-4148-

X. ISSN 1930-529X. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=5426187.

Kim:2010:HTA

[KP10] Tae-Hwan Kim and In-Cheol Park. High-throughput and area-efficient MIMO symbol detection based on modified Dijkstra’s search. IEEE Transactions on Circuits and Systems I: Regular Papers, 57(7):1756–1766, 2010. ISSN 1549-8328 (print), 1558-0806 (electronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?arnumber=5371849.

KrusemanAretz:2003:DZA

[Kru03] F. E. J. Kruseman Aretz. The Dijkstra–Zonneveld ALGOL 60 compiler for the Electrologica X1: historical note SEN, 2. Note SEN-N0301, Centrum voor Wiskunde en Informatica, Amsterdam, The Netherlands, June 30, 2003. vii + 328 pp.

URL http://archive.computerhistory.org/resources/
text/algol/algol_bulletin/EX/SN-N0301.PDF.

Kundu:1980:DLS

[Kun80] Sukhamay Kundu. A Dijkstra-like shortest path algorithm for certain cases of negative arc lengths. BIT (Nordisk tidskrift for informationsbehandling), 20(4):522– 524, December 1980. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&
issn=0006-3835&volume=20&issue=4&spage=522.

Kambayashi:2009:DBG

[KYTY09] Y. Kambayashi, H. Yamachi, Y. Tsujimura, and H. Yamamoto. Dijkstra beats genetic algorithm: Integrating uncomfortable intersection-turns to subjectively optimal route selection. In IEEE, editor, ICCC 2009: IEEE 7th International Conference on Computational Cybernetics: Tryp Bellver, Palma de Mallorca, Spain, November 26–29, 2009, pages 45–50. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2009. ISBN 1-4244-5310-0 (paperback), 1-4244-5311-

9. URL http://ieeexplore.ieee.org/stamp/stamp.jsp?
arnumber=5393932.

Lamport:1974:NSD

[Lam74] Leslie Lamport. A new solution of Dijkstra’s concurrent programming problem. Communications of the Association for Computing Machinery, 17(8):453–455, August 1974. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Lamport:2019:NSD

[Lam19] Leslie Lamport. A new solution of Dijkstra’s concurrent programming problem. In Malkhi [Mal19], pages 171– 178. ISBN 1-4503-7271-6 (paperback), 1-4503-7270-8 (hardcover), 1-4503-7273-2. LCCN ???? URL https://www.

morganclaypoolpublishers.com/catalog_Orig/product_
info.php?cPath=24&products_id=1468.

Lamport:2022:DBB

[Lam22] Leslie Lamport. Deconstructing the bakery to build a distributed state machine. Communications of the Association for Computing Machinery, 65(9):58–66, September 2022. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3513259.

Laplante:1996:GPC

[Lap96] Phillip Laplante, editor. Great Papers in Computer Science. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-314-06365-X (paperback), 0-7803-1112-4 (hardcover). iv + 717 pp. LCCN

QA76 .G686 1996. URL http://bit.csc.lsu.edu/~chen/
GreatPapers.html.

Lavington:1980:CDM

[Lav80] S. H. Lavington. Computer development at Manchester University. In Metropolis et al. [MHR80], pages 433–443. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5

.I63 1976. URL http://www.sciencedirect.com/science/
article/pii/B9780124916500500290. Original versions of

these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Liu:1994:ICB

[LCL^{+}94] Bing Liu, Siew-Hwee Choo, Shee-Ling Lok, Sing-Meng Leong, Soo-Chee Lee, Foong-Ping Poon, and Hwee-Har Tan. Integrating case-based reasoning, knowledge-based approach and Dijkstra algorithm for route finding. In Proceedings of the Tenth Conference on Artificial Intelligence for Applications, 1994, pages 149–155. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994.

ISBN 0-8186-5550-X. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=323680.

Liu:2014:MSO

[LCZW14] He Liu, Qimei Cui, Tong Zhang, and Jian Wang. Multistep offloading based on Dijkstra algorithm for 3D beamforming networks. In 2014 IEEE/CIC International Conference on Communications in China (ICCC), pages 873–878. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2014. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?arnumber=7008399.

Lee:2006:PMD

[Lee06] D. C. Lee. Proof of a modified Dijkstra’s algorithm for computing shortest bundle delay in networks with deterministically time-varying links. IEEE Communications Letters, 10(10):734–736, 2006. CODEN ICLEF6. ISSN 1089-7798 (print), 1558-2558 (electronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?arnumber=1705916.

Lehmann:1977:AST

[Leh77] D. J. Lehmann. Algebraic structures for transitive closure. Theoretical Computer Science, 4(1):59–76, February 1977. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Lehmer:1980:HSP

[Leh80] D. H. Lehmer. A history of the sieve process. In Metropolis et al. [MHR80], pages 445–456. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5 .I63 1976.

URL http://www.sciencedirect.com/science/article/

pii/B9780124916500500307. Original versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Lewis:2021:ICF

[Lew21] Harry R. Lewis. Ideas That Created the Future: Classic Papers of Computer Science. MIT Press, Cambridge, MA, USA, 2021. ISBN 0-262-04530-3. xxii + 495 pp. LCCN Q124.6-127.2.

Lengauer:1990:SPG

[LH90] Christian Lengauer and Duncan G. Hudson. A systolic program for Gauss–Jordan elimination. In Feijen et al. [FvGGM90], chapter 30, pages 262–273. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN QA76 .B326

1990. URL http://www.zentralblatt-math.org/zmath/

en/search/?an=0718.68004. Contains important treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

Lloyd:1974:BRB

[Llo74] Cliff Lloyd. Book review: Structured programming, by O.- J. Dahl, E. W. Dijkstra and C. A. R. Hoare. The Computer Journal, 17(2):186, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.

oup.co.uk/computer_journal/hdb/Volume_17/Issue_02/
tiff/186.tif.

Lukaszewicz:2002:RAA

[ŁMB02] Witold Łukaszewicz and Ewa Madalińska-Bugaj. Reasoning about action and change using Dijkstra’s semantics for programming languages. In Agent-based defeasible control in dynamic environments, volume 7 of Handb. Defeasible Reason. Uncertain. Manag. Syst., pages 383–398. Kluwer Acad. Publ., Dordrecht, 2002.

Li:2014:DAA

[LN14] Duanling Li and Kun Niu. Dijkstra’s algorithm in AGV. In 2014 IEEE 9th Conference on Industrial Electronics and Applications (ICIEA), pages 1867–1871. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2014. ISBN 1-4799-4316-

9. URL http://ieeexplore.ieee.org/stamp/stamp.jsp?
arnumber=6931472.

Liu:2012:FDH

[LSJT12] Hui Liu, N. Stoll, S. Junginger, and K. Thurow. A Floyd– Dijkstra hybrid application for mobile robot path planning in life science automation. In 2012 IEEE International Conference on Automation Science and Engineering (CASE), pages 279–284. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2012. ISBN 1-4673-0429-8. ISSN 2161-8070 (print), 2161-8089 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=6386325.

Lin:2012:DBH

[LSMS12] I-Te Lin, D. Sarkar, T. Murase, and I. Sasase. Dijkstrabased higher capacity route selection algorithm using bounded length and state change for automobiles. In IEEE, editor, IEEE 75th Vehicular Technology Conference (VTC Spring): proceedings: Yokohama, Japan, 6–9 May 2012, pages 1–5. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2012. ISBN 1-4673-0988-5, 1-4673-0989-3. ISSN 1550-

2252. LCCN TA1235. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=6239958. IEEE Catalog Num-

bers: CFP12VTC-USB and CFP12VTC-ART.

Lukkien:1994:OSG

[Luk94] Johan J. Lukkien. Operational semantics and generalized weakest preconditions. Science of Computer Programming, 22(1–2):137–155, April 1994. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). Mathematics of program construction (Oxford, 1992).

Lukkien:1990:TIA

[LvdS90] Johan J. Lukkien and Jan L. A. van de Snepscheut. Topologyindependent algorithms based on spanning trees. In Feijen et al. [FvGGM90], chapter 32, pages 280–288. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN QA76 .B326

1990. URL http://www.zentralblatt-math.org/zmath/

en/search/?an=0718.68004. Contains important treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

Liu:2011:ITA

[LW11] Shuiqiang Liu and Yanpeng Wu. Isomorphism testing algorithm based on Dijkstra algorithm for plan graphs. In 2011 International Conference on Information Technology Computer Engineering and Management Sciences (ICM), volume 4, pages 309–311. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2011.

ISBN 1-4577-1419-1. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=6113755.

Lu:2009:ALW

[LWQ09] He Lu, Chen Wei, and Gao Qiang. Analysis of the 3-layer workflow scheduling model based on improved Dijkstra algorithm. In 2009 1st International Conference on Information Science and Engineering (ICISE), pages 141–144. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2009. ISBN 1-4244-4909-

X. URL http://ieeexplore.ieee.org/stamp/stamp.jsp?
arnumber=5454827.

Li:2010:DAI

[LX10] Yuqiang Li and QianXing Xiong. Dijkstra algorithm implementing services composing way selection. In Chunzhi Wang and Hongwei Chen, editors, Proceedings, 2010 Second International Workshop on Intelligent Systems and Applications: ISA 2010: 22–23 May 2010, Wuhan, China, pages 1–4. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2010. ISBN 1-4244-3894-2, 1-4244-5872-2 (paperback), 1-4244-5874-9. LCCN

Q334. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=5473556.

Malkhi:2019:CWL

[Mal19] Dahlia Malkhi, editor. Concurrency: the Works of Leslie Lamport. Morgan and Claypool Publishers, Williston, VT 05495 USA, 2019. ISBN 1-4503-7271-6 (paperback), 1-4503-7270-8 (hardcover), 1-4503-7273-2. xx + 345 pp. LCCN

???? URL https://www.morganclaypoolpublishers.com/
catalog_Orig/product_info.php?cPath=24&products_id=
1468.

Martin:1990:LDI

[Mar90] Alain J. Martin. The limitations to delay-insensitivity in asynchronous circuits. In Feijen et al. [FvGGM90], chapter 34, pages 302–311. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X.

LCCN QA76 .B326 1990. URL http://www.zentralblatt-
math.org/zmath/en/search/?an=0718.68004. Contains

important treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

Markoff:2002:EDP

[Mar02] John Markoff. Edsger Dijkstra, 72, physicist who shaped computer era, dies. New York Times, ??(??):??, August 10, 2002. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-

667X, 1553-8095. URL http://www.nytimes.com/2002/08/
10/obituaries/10DIJK.html.

Maurer:1978:ALC

[Mau78] W. D. Maurer. Another letter on the correctness of real programs. ACM SIGSOFT Software Engineering Notes, 3(3): 22–24, July 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). See [Dij78u, DLP78, Jef78, Van78].

Mauchly:1980:E

[Mau80] John W. Mauchly. The ENIAC. In Metropolis et al. [MHR80], pages 541–550. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book).

LCCN QA75.5 .I63 1976. URL http://www.sciencedirect.
com/science/article/pii/B9780124916500500393. Origi-

nal versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

May:1980:HPC

[May80] Kenneth O. May. Historiography: A perspective for computer scientists. In Metropolis et al. [MHR80], pages 11–18. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5

.I63 1976. URL http://www.sciencedirect.com/science/
article/pii/B9780124916500500101. Original versions of

these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Mazurkiewicz:1976:MFC

[Maz76] Antoni W. Mazurkiewicz, editor. Mathematical foundations of computer science, 1976: proceedings, 5th Symposium, Gdansk, September 6–10, 1976, volume 45 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNCSD9. ISBN 0-387-07854-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.5.S4

S989 1976. URL http://link.springer-ny.com/link/
service/series/0558/tocs/t0045.htm;   http://www.
springerlink.com/content/978-0-387-07854-0;   http:
//www.springerlink.com/openurl.asp?genre=issue&issn=

0302-9743&volume=45. “The Symposium is being organized by the Computation Centre of the Polish Academy of Sciences in cooperation with the University of Gdansk.”.

Madalinska-Bugaj:1997:HSQ

[MB97] E. Madalinska-Bugaj. How to solve qualification and ramification using Dijkstra’s semantics for programming languages. Lecture Notes in Computer Science, 1321:381–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mao:2013:TBP

[MCX13] Xinyu Mao, Yuxin Cheng, and Haige Xiang. Two block partitioned Dijkstra algorithms. In 2013 IEEE 78th Vehicular Technology Conference (VTC Fall), pages 1–5. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2013. ISSN 1090-

3038. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=6692448.

Metropolis:1980:M

[Met80] N. Metropolis. The MANIAC. In Metropolis et al. [MHR80], pages 457–464. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book).

LCCN QA75.5 .I63 1976. URL http://www.sciencedirect.
com/science/article/pii/B9780124916500500319. Origi-

nal versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Meirosu:2005:NGE

[MGH^{+}05] Catalin Meirosu, Piotr Golonka, Andreas Hirstius, Stefan Stancu, Bob Dobinson, Erik Radius, Antony Antony, Freek Dijkstra, Johan Blom, and Cees de Laat. Native 10 Gigabit Ethernet experiments over long distances. Future Generation Computer Systems, 21(4):457–468, April 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Markoski:2010:DIA

[MHMO10] Branko Markoski, Petar Hotomski, Dušan Malbaški, and Danilo Obradović. Dijkstra’s interpretation of the approach to solving a problem of program correctness. Yugoslav Journal of Operations Research, 20(2):229–236, 2010. ISSN 0354-0243 (print), 1820-743x (electronic).

Metropolis:1980:HCT

[MHR80] Nicholas Metropolis, Jack Howlett, and Gian-Carlo Rota, editors. A History of Computing in the Twentieth Century: a Collection of Essays. Academic Press, New York, NY, USA, 1980. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5 .I63 1976. Original versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Mili:1989:BRB

[Mil89] Ali Mili. Book review: A method of programming, by Edsger W. Dijkstra and W. H. J. Feijen. Addison-Wesley, Wokingham, United Kingdom, 1988, Price £15.95 (paperback), ISBN 0-201-17536-3. Science of Computer Programming, 12(2):165–167, July 1989. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
0167642389900439.

Misra:1990:SPS

[Mis90] Jayadev Misra. A simple proof of a simple consensus algorithm. In Feijen et al. [FvGGM90], chapter 35, pages 312–318. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN QA76 .B326

1990. URL http://www.zentralblatt-math.org/zmath/

en/search/?an=0718.68004. Contains important treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

Misra:2001:WSP

[Mis01] Jayadev Misra. A walk over the shortest path: Dijkstra’s Algorithm viewed as fixed-point computation. Information Processing Letters, 77(2–4):197–200, February 28, 2001. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.elsevier.nl/gej-ng/10/23/20/68/29/
44/abstract.html;   http://www.elsevier.nl/gej-ng/
10/23/20/68/29/44/article.pdf. Special issue in honor of

Edsger W. Dijkstra (Austin, TX, 2000).

Misa:2010:IEW

[Mis10] Thomas J. Misa. An interview with Edsger W. Dijkstra. Communications of the Association for Computing Machinery, 53 (8):41–47, August 2010. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Myllyla:2007:LSD

[MJC07] M. Myllyla, M. Juntti, and J. R. Cavallaro. A list sphere detector based on Dijkstra’s algorithm for MIMO– OFDM systems. In IEEE, editor, IEEE 18th International Symposium on Personal, Indoor and Mobile Radio Communications, 2007: PIMRC 2007; 3-7 September 2007, Athens, Greece, pages 1–5. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2007. ISBN 1-4244-1144-0, 1-4244-1143-

2. LCCN TK5103.2. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=4394362.

Mitchell:1987:DGP

[MMP87] Joseph S. B. Mitchell, David M. Mount, and Christos H. Papadimitriou. The discrete geodesic problem. SIAM Journal on Computing, 16(4):647–668, August 1987. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Moore:2002:RFE

[Moo02] J. Strother Moore. Remarks at the funeral of Edsger W. Dijkstra. Edsger W. Dijkstra Archive document., August 10, 2002.

URL http://www.cs.utexas.edu/users/EWD/obituaries/
JSMremarks.html.

Morgan:1990:WC

[Mor90a] Carroll Morgan. Of wp and CSP. In Feijen et al. [FvGGM90], chapter 36, pages 319–326. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN QA76 .B326 1990. URL

http://www.zentralblatt-math.org/zmath/en/search/?

an=0718.68004. Contains important treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

Morris:1990:PER

[Mor90b] Joseph M. Morris. Programming by expression refinement: the KMP algorithm. In Feijen et al. [FvGGM90], chapter 37, pages 327–338. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X.

LCCN QA76 .B326 1990. URL http://www.zentralblatt-
math.org/zmath/en/search/?an=0718.68004. Contains

important treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

Moses:2010:BSK

[Mos10] Yoram Moses. Behind the scenes of K&CK: the undelivered speech for the 2009 Dijkstra Prize. SIGACT News (ACM Special Interest Group on Automata and Computability Theory), 41(1):58–62, March 2010. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Manna:1990:EVM

[MP90] Zohar Manna and Amir Pnueli. An exercise in the verification of multi-process programs. In Feijen et al. [FvGGM90], chapter 33, pages 289–301. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X.

LCCN QA76 .B326 1990. URL http://www.zentralblatt-
math.org/zmath/en/search/?an=0718.68004. Contains

important treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

Mitchell:1991:WRP

[MP91] Joseph S. B. Mitchell and Christos H. Papadimitriou. The weighted region problem: Finding shortest paths through a weighted planar subdivision. Journal of the Association for Computing Machinery, 38(1):18–73, January 1991. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (elec-

tronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/102784.html.

Metropolis:1980:IE

[MR80] N. Metropolis and Gian-Carlo Rota. Introductory essay. In Metropolis et al. [MHR80], pages xix–xx. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5 .I63 1976.

URL http://www.sciencedirect.com/science/article/

pii/B9780124916500500083. Original versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Misra:2003:MEW

[MR03] Jayadev Misra and Hamilton Richards. In memoriam: Edsger Wybe Dijkstra (1930–2002). Web document., August 25,

2003. URL www.cs.utexas.edu/users/EWD/MemRes(USLtr)
.pdf.

Mehlhorn:2001:HDA

[MS01] Kurt Mehlhorn and Guido Schäfer. A heuristic for Dijkstra’s algorithm with many targets and its use in weighted matching algorithms. Lecture Notes in Computer Science, 2161:242– 253, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer-ny.com/
link/service/series/0558/bibs/2161/21610242.htm;
http://link.springer-ny.com/link/service/series/0558/
papers/2161/21610242.pdf.

Murota:2014:DAC

[MS14] Kazuo Murota and Akiyoshi Shioura. Dijkstra’s algorithm and L-concave function maximization. Mathematical Programming, 145(1-2, Ser. A):163–177, 2014. CODEN MH-PGA4. ISSN 0025-5610.

Mohring:2006:PGS

[MSS^{+}06] Rolf H. Möhring, Heiko Schilling, Birk Schütz, Dorothea Wagner, and Thomas Willhalm. Partitioning graphs to speedup Dijkstra’s algorithm. ACM Journal of Experimental Algorithmics, 11:2.8:1–2.8:29, ???? 2006. CODEN ???? ISSN 1084-6654.

Moffat:1984:PQA

[MT84] Alistair Moffat and Tadao Takaoka. A priority queue for the all pairs shortest path problem. Information Processing Letters, 18(4):189–193, May 14, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Musznicki:2012:DBL

[MTZ12] B. Musznicki, M. Tomczak, and P. Zwierzykowski. Dijkstrabased localized multicast routing in wireless sensor networks. In 2012 8th International Symposium on Communication Systems, Networks & Digital Signal Processing (CSNDSP), pages 1–6. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2012. ISBN 1-4577-1472-

8. URL http://ieeexplore.ieee.org/stamp/stamp.jsp?
arnumber=6292692.

Mukherjee:2012:DAS

[Muk12] Sathi Mukherjee. Dijkstra’s algorithm for solving the shortest path problem on networks under intuitionistic fuzzy environment. J. Math. Model. Algorithms, 11(4):345–359, 2012. ISSN 1570-1166.

Mei:2002:NDS

[MZS02] Wei Feng Mei, Yu Zhang, and Fang Min Song. A note to Dijkstra’s study on quantification. Nanjing Daxue Xuebao Ziran Kexue Ban, 38(1):110–118, 2002. CODEN NCHPAZ. ISSN 0469-5097.

Nakayama:2013:DBA

[NA13] Akira Nakayama and Tsutomu Anazawa. Dijkstra-based algorithms for the shortest path problem with edges of negative length. Journal of the Operations Research Society of Japan, 56(2):137–154, 2013. CODEN JORJA5. ISSN 0453-4514 (print), 1878-6871 (electronic).

Nagappan:2018:RWG

[Nag18] M. Nagappan. Reconsidering whether GOTO is harmful. IEEE Software, 35(3):93–95, May 2018. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Nikas:2009:ETM

[NAGK09] K. Nikas, N. Anastopoulos, G. Goumas, and N. Koziris. Employing transactional memory and helper threads to speedup Dijkstra’s algorithm. In IEEE, editor, 2009, 38th international conference on parallel processing: (ICPP 2009) Vienna, Austria, 22–25 September 2009, pages 388– 395. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2009. ISBN 0-7695-3802-9, 1-4244-4961-8. ISSN 0190-3918. LCCN

QA76.58.  URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5362412.

Ng:1993:STP

[NB93] F. C. N. Ng and G. Butler. Specialised theorem-proving in an intelligent tutoring system for the Dijkstra–Gries programming methodology. In Proceedings ICCI ’93: Fifth International Conference on Computing and Information, 1993, pages 294–298. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN

0-8186-4212-2. URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=315360.

Naur:1960:RAL

[NBB^{+}60] Peter Naur, J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. McCarthy, A. J. Perlis, H. Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, and M. Woodger. Report on the algorithmic language ALGOL 60. Communications of the Association for Computing Machinery, 3(5):299–314, May 1960. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See letter [?]. Dedicated to the Memory of William Turanski [American group member who was killed by an automobile just prior to the January 1960 Conference].

Ng:1995:ITS

[NBK95] F. Ng, G. Butler, and Judy Kay. An intelligent tutoring system for the Dijkstra–Gries methodology. IEEE Transactions on Software Engineering, 21(5):415–428, May 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=387471.

Nepomniaschaya:2000:SID

[ND00] Anna S. Nepomniaschaya and Maria A. Dvoskina. A simple implementation of Dijkstra’s shortest path algorithm on associative parallel processors. Fundamenta Informaticae, 43 (1–4):227–243, August 2000. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nelson:1989:GDC

[Nel89] Greg Nelson. A generalization of Dijkstra’s calculus. ACM Transactions on Programming Languages and Systems, 11(4): 517–561, October 1989. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL http://www.acm.org/

pubs/toc/Abstracts/0164-0925/69559.html.

Nelson:1990:MCS

[Nel90] Greg Nelson. Methodical competitive snoopy-caching. In Feijen et al. [FvGGM90], chapter 38, pages 339–345. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN QA76 .B326

1990. URL http://www.zentralblatt-math.org/zmath/

en/search/?an=0718.68004. Contains important treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

Nelson:1992:SGA

[Nel92] Greg Nelson. Some generalizations and applications of Dijkstra’s guarded commands. In Programming and mathematical method (Marktoberdorf, 1990), volume 88 of NATO Adv. Sci. Inst. Ser. F Comput. Systems Sci., pages 157–191. Springer, Berlin, 1992.

Nepomniaschaya:1996:AVP

[Nep96] A. S. Nepomniaschaya. An associative version of the Prim– Dijkstra algorithm and its application to some graph problems. Lecture Notes in Computer Science, 1181:203–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nepomnyashchaya:2000:CPD

[Nep00] A. Sh. Nepomnyashchaya. Comparison of Prim–Dijkstra and Kraskal algorithms using an associative parallel processor. Kibernet. Sistem. Anal., 2:19–27, 187, 2000. ISSN 0023-1274. English translation in Cybernetics and Systems Analysis, 36(2) 162–169 (March 2000).

Neumann:1990:BBS

[Neu90] Peter G. Neumann. Beauty and the beast the software complexity elegance versus elephants. In Feijen et al. [FvGGM90], chapter 39, pages 346–351. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN QA76 .B326

1990. URL http://www.zentralblatt-math.org/zmath/

en/search/?an=0718.68004. Contains important treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

Natu:1995:PPC

[NF95] Madan Natu and Shu Cherng Fang. On the point-to-point connection problem. Information Processing Letters, 53(6): 333–336, March 24, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Noshita:1978:EBD

[NMM78] Kohei Noshita, Etsuo Masuda, and Hajime Machida. On the expected behaviors of the Dijkstra’s shortest path algorithm for complete graphs. Information Processing Letters, 7(5):

237–243, August ??, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Noshita:1985:TEC

[Nos85] Kohei Noshita. A theorem on the expected complexity of Dijkstra’s shortest path algorithm. Journal of Algorithms, 6(3):400–408, September 1985. CODEN JOALDV. ISSN 0196-6774 (print), 1090-2678 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0196677485900094.

Naur:1969:SER

[NR69] Peter Naur and Brian Randell, editors. Software Engineering: Report on a conference sponsored by the NATO Science Committee, Garmisch, Germany, 7 to 11 October, 1968. NATO Scientific Affairs Division, Brussels, Belgium, January 1969. LCCN QA 264 W675 1968.

Naur:1976:SEC

[NRB76] Peter Naur, Brian Randell, and J. N. Buxton, editors. Software Engineering Concepts and Techniques: Proceedings of the NATO Conferences. Petrocelli/Charter, New York, NY, USA, 1976. ISBN 0-88405-334-2. LCCN QA76.6 S618 1968. Conferences sponsored by the NATO Science Committee, held at Garmisch, Germany, 7–11 October 1968, and at Rome, Italy, 27–31 October 1969.

Noto:2000:MSP

[NS00] M. Noto and H. Sato. A method for the shortest path search by extended Dijkstra algorithm. In 2000 IEEE International Conference on Systems, Man, and Cybernetics, volume 3, pages 2316–2320 vol.3. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2000. ISBN 0-7803-6583-6. ISSN 1062-

922X. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=886462.

Obruca:1978:AM

[Obr78] A. Obruča. Algorithm 1: Mintree. The Computer Journal, 21(3):276, August 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.

oup.co.uk/computer_journal/hdb/Volume_21/Issue_03/

tiff/276.tif. Reprinted from The Computer Bulletin, September 1964, p. 67. See notes [?, ?, ?, ?, ?, WT79].

Oh:2010:HMD

[OC10] Keunhyun Oh and Sung-Bae Cho. A hybrid method of Dijkstra algorithm and evolutionary neural network for optimal Ms. Pac-Man agent. In 2010 Second World Congress on Nature and Biologically Inspired Computing (NaBIC), pages 239–243. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2010. ISBN 1-4244-

7377-2. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=5716312.

Padlipsky:1985:ENS

[Pad85] M. A. Padlipsky. The Elements of Networking Style and Other Essays and Animadversions on the Art of Intercomputer Networking. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1985. ISBN 0-13-268129-3 (hardcover), 0-13-268111-0 (paperback). 236 pp. LCCN TK5105.5 .P34 1985. US$34.95, CDN$44.95. Mike was the Multics system-programmer at Project Mac and is personally responsible for putting two systems on the ARPAnet back when that meant hacking the OS designed and implemented a number of protocols including being working-group chair, and has been a commentator on Internet RFCs since the beginning. (He was also Dennis Ritchie’s officemate before Bell Labs pulled out of Multics and Dennis had to invent UNIX, and has a personal exemption from Dijkstra to use a GOTO when, in his professional opinion as a protocol implementor, it is necessary.) This is a book on style in protocol design and the great Bromdinagian TCP vs. ISO debates. It’s a nice piece of history and explains where TCP, TELNET, FTP came from.

Parnas:1978:AVD

[Par78] David L. Parnas. Another view of the Dijkstra–dMLP controversy. ACM SIGSOFT Software Engineering Notes, 3(4): 20–21, October 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Parnas:2001:BHS

[Par01a] David L. Parnas. On a ‘buzzword’: Hierarchical structure. In Broy and Denert [BD01b], pages 499–513. ISBN 3-540-42290-0 (print), 3-642-48354-2 (e-book). LCCN QA76.758.

URL http://link.springer.com/chapter/10.1007/978-
3-642-48354-7_21/.

Parnas:2001:CUD

[Par01b] David L. Parnas. On the criteria to be used in decomposing systems into modules. In Broy and Denert [BD01b], pages 479–498. ISBN 3-540-42290-0 (print), 3-642-48354-2 (e-book). LCCN QA76.758. URL http://link.springer.

com/chapter/10.1007/978-3-642-48354-7_20/.

Parnas:2018:SSC

[Par18] David Lorge Parnas. Software structures: A careful look. IEEE Software, 35(6):68–71, November/December 2018. CO-DEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Payette:2014:HDC

[Pay14] S. Payette. Hopper and Dijkstra: Crisis, revolution, and the future of programming. IEEE Annals of the History of Computing, 36(4):64–73, October/December 2014. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=6982140.

Peremans:1990:NF

[Per90] W. Peremans. A note on feasibility. In Feijen et al. [FvGGM90], chapter 40, pages 352–355. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN QA76 .B326 1990. URL

http://www.zentralblatt-math.org/zmath/en/search/?

an=0718.68004. Contains important treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

Petitfrere:1971:ADP

[Pet71] Marianne Petitfrère. Sur l’algorithme de Dijkstra pour l’obtention des plus courts chemins dans un graphe. (French) [On Dijkstra’s algorithm for finding the shortest paths in a graph]. Cahiers Centre Études Recherche Opér., 13:111–123, 1971. ISSN 0774-3068.

Pereida:2013:HDP

[PG13] K. Pereida and J. Guivant. Hybrid Dijkstra–PSO algorithm for motion planning of non-holonomic multiple-trailer platforms in dense contexts. In 2013 IEEE/ASME International Conference on Advanced Intelligent Mechatronics (AIM), pages 13–18. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2013. ISBN 1-4673-5319-1. ISSN 2159-6247 (print), 2159-6255 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=6584061.

Pomberger:2000:SNW

[PGB00] Gustav Pomberger, Jürg Gutknecht, and László Böszörmenyi, editors. The School of Niklaus Wirth: the art of simplicity. dpunkt.verlag, Heidelberg, Germany, 2000. ISBN 1-55860-723-4 (Morgan Kaufmann), 3-932588-85-1 (dpunkt). LCCN

???? URL http://www.springerlink.com/content/978-
3-642-48354-7.

Page:1989:SME

[PJ89] I. P. Page and R. T. Jacob. The solution of mutual exclusion problems which can be described graphically. The Computer Journal, 32(1):45–54, February 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/32/1/45.full.pdf+html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_32/Issue_01/
tiff/45.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_01/tiff/46.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_32/Issue_01/
tiff/47.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_01/tiff/48.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_32/Issue_01/
tiff/49.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_01/tiff/50.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_32/Issue_01/
tiff/51.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_01/tiff/52.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_32/Issue_01/
tiff/53.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_32/Issue_01/tiff/54.tif.

Parra:2011:DAM

[PMR11] O. J. S. Parra, C. Manta, and G. L. Rubio. Dijkstra’s algorithm model over MPLS / GMPLS. In 2011 7th International Conference on Wireless Communications, Networking and Mobile Computing (WiCOM), pages 1–4. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2011. ISBN 1-4244-6250-9. ISSN 2161-

9646. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=6040536.

Popa:1995:CFP

[Pop95] Marin Popa. Concurrently firing in Petri-nets. Application in solving Dijkstra’s problem. Studii şi cercetări Matematice, 47 (3-4):337–347, 1995. CODEN APTMAD. ISSN 0039-4068, 0567-6401.

Post:1990:CPP

[Pos90] Karel A. Post. A curious property of points and circles in the plane. In Feijen et al. [FvGGM90], chapter 41, pages 356–357. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN QA76 .B326

1990. URL http://www.zentralblatt-math.org/zmath/

en/search/?an=0718.68004. Contains important treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

Parulekar:2013:AVN

[PPS^{+}13] M. Parulekar, V. Padte, T. Shah, K. Shroff, and R. Shetty. Automatic vehicle navigation using Dijkstra’s algorithm. In 2013 International Conference on Advances in Technology and Engineering (ICATE), pages 1–5. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2013. ISBN 1-4673-5618-

2. URL http://ieeexplore.ieee.org/stamp/stamp.jsp?
arnumber=6524721.

Parra:2012:DAB

[PRC12] O. J. S. Parra, G. L. Rubio, and L. Castellanos. Dijkstra algorithm based reliability model. In 2012 IEEE/ACM 16th International Symposium on Distributed Simulation and Real Time Applications (DS-RT), pages 157–160. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2012. ISBN 1-4673-2954-1. ISSN 1550-

6525. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=6365079.

Parra:2014:DAB

[PRD14] O. J. Salcedo Parra, G. Lopez Rubio, and B. S. Reyes Daza. Dijkstra algorithm based on a reliability model. In 2014 International Conference on Information Society (i-Society), pages 45–49. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA,

2014. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=7009008.

Pritchard:1990:PIS

[Pri90] Paul Pritchard. A problem involving subsequences. In Feijen et al. [FvGGM90], chapter 42, pages 358–364. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN QA76 .B326

1990. URL http://www.zentralblatt-math.org/zmath/

en/search/?an=0718.68004. Contains important treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

Parmar:2011:WBR

[PRM11] H. Parmar, S. Ruparelia, and U. Mehta. WTM based reordering of combine test vector & output response using Dijkstra algorithm for scan power reduction. In 2011 Nirma University International Conference on Engineering (NUiCONE), pages 1–6. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2011. ISBN 1-4577-2169-

4. URL http://ieeexplore.ieee.org/stamp/stamp.jsp?
arnumber=6153276.

Peyer:2009:GDS

[PRV09] Sven Peyer, Dieter Rautenbach, and Jens Vygen. A generalization of Dijkstra’s shortest path algorithm with applications to VLSI routing. Journal of Discrete Algorithms, 7(4):377– 390, 2009. ISSN 1570-8667 (print), 1570-8675 (electronic).

Perdana:2014:MRP

[PS14] D. Perdana and R. F. Sari. Multicast routing performance analysis for different mobility models on the IEEE 1609.4 standard using random Dijkstra algorithm. In 2014 International Conference on Intelligent Green Building and Smart Grid (IGBSG), pages 1–7. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA,

2014. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=6835255.

Parnas:1989:CRD

[PSvE^{+}89] David L. Parnas, W. L. Scherlis, M. H. van Emden, Jacques Cohen, R. W. Hamming, Richard M. Karp, and Terry Winograd. Colleagues respond to Dijkstra’s comments. Communications of the Association for Computing Machinery, 32(12): 1405–1413, December 1989. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rajchman:1980:ERC

[Raj80] Jan Rajchman. Early research on computers at RCA. In Metropolis et al. [MHR80], pages 465–469. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5 .I63 1976.

URL http://www.sciencedirect.com/science/article/

pii/B9780124916500500320. Original versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Randell:1980:C

[Ran80a] B. Randell. The COLOSSUS. In Metropolis et al. [MHR80], pages 47–92. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book).

LCCN QA75.5 .I63 1976. URL http://www.sciencedirect.
com/science/article/pii/B9780124916500500137. Origi-

nal versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Randell:1980:ODC

[Ran80b] B. Randell. The origins of digital computers: Supplementary bibliography. In Metropolis et al. [MHR80], pages 629–659. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5

.I63 1976. URL http://www.sciencedirect.com/science/
article/pii/B9780124916500500460. Original versions of

these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Randell:2002:ASE

[Ran02] Brian Randell. From Algol 60 to software engineering. Talk at Dijkstra Tribute, Sun Microsystems, London, UK., November 14, 2002.

Randell:2003:ED

[Ran03] B. Randell. Edsger Dijkstra. In The Ninth IEEE International Workshop on Object-Oriented Real-Time Dependable Systems, 2003. WORDS 2003 Fall, pages 3–8. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2003. ISBN 0-17-952054-

5. URL http://ieeexplore.ieee.org/stamp/stamp.jsp?
arnumber=1410938.

Rice:1968:LEG

[RD68] John R. Rice and Edsger W. Dijkstra. Letter to the Editor: The go to statement reconsidered. Communications of the Association for Computing Machinery, 11(8):538, August 1968. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rao:1992:DPV

[RD92] Josyula R. Rao and Edsger W. Dijkstra. Designing the proof of Vizing’s theorem. In Programming and mathematical method (Marktoberdorf, 1990), volume 88 of NATO Adv. Sci. Inst. Ser. F Comput. Systems Sci., pages 17–25. Springer, Berlin, 1992.

Reid:1983:MCA

[Rei83a] E. Reid. More comments after Dijkstra and Tompkins. ACM SIGPLAN Notices, 18(10):16–21, October 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Reid:1983:FDM

[Rei83b] Edward Reid. Fighting the disease: more comments after Dijkstra and Tompkins. ACM SIGPLAN Notices, 18(10):16– 21, October 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Reisig:2008:SDP

[Rei08] Wolfgang Reisig. The Scholten/Dijkstra pebble game played straightly, distributedly, online and reversed. In Pillars of computer science, volume 4800 of Lecture Notes in Comput. Sci., pages 589–595. Springer, Berlin, 2008.

Rem:1990:PPA

[Rem90] Martin Rem. A personal perspective of Alphern–Schneider characterization of safety and liveness. In Feijen et al. [FvGGM90], chapter 43, pages 365–372. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN QA76 .B326 1990. URL

http://www.zentralblatt-math.org/zmath/en/search/?

an=0718.68004. Contains important treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

Robertson:1980:OI

[Rob80] James E. Robertson. The ORDVAC and the ILLIAC. In Metropolis et al. [MHR80], pages 347–364. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5 .I63 1976.

URL http://www.sciencedirect.com/science/article/

pii/B9780124916500500241. Original versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Rosyidi:2014:TDW

[RPGS14] L. Rosyidi, H. P. Pradityo, D. Gunawan, and R. F. Sari. Timebase dynamic weight for Dijkstra Algorithm implementation in route planning software. In 2014 International Conference on Intelligent Green Building and Smart Grid (IGBSG), pages 1–4. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA,

2014. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=6835261.

Randell:1964:AIT

[RR64] Brian Randell and L. J. (Lawford John) Russell. ALGOL 60 implementation; the translation and use of ALGOL 60 programs on a computer, volume 5 of A.P.I.C. studies in data processing. Academic Press, New York, NY, USA, 1964. xiv + 418 pp. LCCN QA76.5 .R35 1964. With a foreword by E. W. Dijkstra. Published for the Automatic Programming Information Centre, Brighton College of Technology, England. Russian translation in 1967.

Ruane:1990:PSU

[Rua90] Lawrence M. Ruane. Process synchronization in the UTS kernel. Computing systems: the journal of the USENIX Association, 3(3):387–421, Summer 1990. CODEN CMSYE2. ISSN 0895-6340.

Svetlana:2013:DAC

[SAID13] C. Svetlana, G. Alexander, C. Ihor, and S. Dmytriy. Dijkstra’s algorithm for cyber structures analysis. In 2013 12th International Conference on the Experience of Designing and Application of CAD Systems in Microelectronics (CADSM), pages 61–65. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2013. ISBN

1-4673-6461-4. URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6543192.

Samelson:1976:ECP

[Sam76] Klaus Samelson, editor. ECI Conference 1976, Proceedings of the 1st European Cooperation in Informatics, Amsterdam, The Netherlands, August 9–12, 1976, Proceedings, volume 44 of Lecture Notes in Computer Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. ISBN 3-540-07804-5.

Santini:2005:WSI

[San05] Simone Santini. We are sorry to inform you . . .. Computer, 38 (12):128, 126–127, December 2005. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Samelson:2001:SFG

[SB01] K. Samelson and F. L. Bauer. Sequentielle Formelübersetzung. (German) [Sequential formula translation]. In Broy and Denert [BD01b], pages 1–27. ISBN 3-540-42290-0 (print), 3-642-48354-2 (e-book). LCCN QA76.758. URL http://

link.springer.com/chapter/10.1007/978-3-642-48354-
7_1/;   http://link.springer.com/content/pdf/bfm:
978-3-642-48354-7/1.pdf.

Singh:2004:ALO

[SB04] Awadhesh Kumar Singh and Anup Kumar Bandyopadhyay. Adding the leads-to operator to Dijkstra’s calculus. ACM SIGPLAN Notices, 39(2):12–17, February 2004. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schneider:1990:SPC

[Sch90a] Fred B. Schneider. Simpler proofs for concurrent reading and writing. In Feijen et al. [FvGGM90], chapter 44, pages 373–379. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN

QA76 .B326 1990. URL http://www.zentralblatt-math.
org/zmath/en/search/?an=0718.68004. Contains impor-

tant treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

Scholten:1990:GJ

[Sch90b]  Carel S. Scholten. Goodbye junctivity?  In Feijen

et al. [FvGGM90], chapter 45, pages 380–385. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN QA76 .B326

1990. URL http://www.zentralblatt-math.org/zmath/

en/search/?an=0718.68004. Contains important treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

Schneider:2018:HCD

[Sch18] Fred Schneider. History and context for defining liveness: Winner 2018 Edsger W. Dijkstra Prize. SIGACT News (ACM Special Interest Group on Automata and Computability Theory), 49(4):60–63, December 2018. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Scholten:1982:CSC

[SD82] C. S. Scholten and Prof. Dr. Edsger W. Dijkstra. A class of simple communication patterns. In Selected Writings on Computing: a Personal Perspective [Dij82-69], pages 334– 337. ISBN 0-387-90652-5 (New York), 3-540-90652-5 (Berlin). LCCN QA76.24 .D54 1982. URL http://link.springer.

com/chapter/10.1007/978-1-4612-5695-3_60. Including a

paper co-authored by C. S. Schölten.

Stamp:2001:RHD

[SEEM01] Mark Stamp, Brad Engel, McIntosh Ewell, and Victor Morrow. Rush hour(R) and Dijkstra’s algorithm. Graph Theory Notes N.Y., 40:23–30, 2001. ISSN 1040-8118. Dedicated to the memory of Robert E. Sacks (Purchase, NY, 2000).

Sharma:1981:MPE

[Sha81] D. K. Sharma. McDonald’s problem — an example of using Dijkstra’s programming method. The Bell System Technical Journal, 60(9):2157–2165, November 1981. CODEN BSTJAN. ISSN 0005-8580. URL http://bstj.bell-labs.

com/BSTJ/images/Vol60/bstj60-9-2157.pdf;   http:/
/www.alcatel-lucent.com/bstj/vol60-1981/articles/
bstj60-9-2157.pdf.

Shearing:2003:TEW

[She03a] B. Shearing. Tribute to Edsger W. Dijkstra. IEEE Annals of the History of Computing, 25(1):67–69, 2003. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=1179892.

Shearing:2003:EDR

[She03b] Brian Shearing. Edsger Dijkstra remembered. Computer Resurrection: The Bulletin of the Computer Conservation Society, 30:??, Spring 2003. ISSN 0958-7403. URL http://www.

cs.man.ac.uk/CCS/res/res30.htm#i.

Schneider:1988:WDP

[SL88] F. B. Schneider and L. Lamport. On E. W. Dijkstra’s position paper on “fairness”. ACM SIGSOFT Software Engineering Notes, 13(3):18–19, July 1988. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). See [Dij88o, CM88].

Shasha:1998:TML

[SL98] Dennis Shasha and Cathy Lazere, editors. Out of Their Minds: The Lives and Discoveries of 15 Great Computer Scientists. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 0-387-98269-8. xi + 293 pp. LCCN QA76.2.A2 S53 1998. US$16.00.

Slutz:1980:MBS

[Slu80] Ralph J. Slutz. Memories of the Bureau of Standards’ SEAC. In Metropolis et al. [MHR80], pages 471–477. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5

.I63 1976. URL http://www.sciencedirect.com/science/
article/pii/B9780124916500500332. Original versions of

these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Smith:2011:UDC

[Smi11] B. R. Smith. Using Dijkstra to compute hop-by-hop QoS paths. In 2011 Proceedings of 20th International Conference on Computer Communications and Networks (ICCCN), pages 1–6. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2011. ISBN 1-4577-0637-7. ISSN 1095-2055. URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?arnumber=6006055.

Sniedovich:2006:DAR

[Sni06] Moshe Sniedovich. Dijkstra’s algorithm revisited: the dynamic programming connection. Control and cybernetics, 35 (3):599–620, 2006. CODEN CCYBAP. ISSN 0324-8569.

Sedeno-Noda:2015:DLM

[SNR15] Antonio Sedeño-Noda and Andrea Raith. A Dijkstra-like method computing all extreme supported non-dominated solutions of the biobjective shortest path problem. Computers and Operations Research, 57:83–94, 2015. CODEN CMORAP. ISSN 0305-0548 (print), 1873-765X (electronic).

Solka:1992:AUP

[SPPR92] J. L. Solka, J. C. Perry, B. R. Poellinger, and G. W. Rogers. Autorouting using a parallel Dijkstra algorithm with embedded constraints. In International Joint Conference on Neural Networks, 1992. IJCNN, volume 2, pages 239–247. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. ISBN 0-7803-0559-

0. URL http://ieeexplore.ieee.org/stamp/stamp.jsp?
arnumber=227001.

Shilova:2004:AET

[SS04a] S. O. Shilova and Nikolay V. Shilov. Addendum to Etude on theme of Dijkstra. SIGACT News (ACM Special Interest Group on Automata and Computability Theory), 35(4):118, December 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Shilova:2004:ETD

[SS04b] S. O. Shilova and Nikolay V. Shilov. Etude on theme of Dijkstra. SIGACT News (ACM Special Interest Group on Automata and Computability Theory), 35(3):102–108, September 2004. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Shankar:2010:UMD

[SS10] N. Ravi Shankar and V. Sireesha. Using modified Dijkstra’s algorithm for critical path method in a project network. Int. J. Comput. Appl. Math., 5(2):217–225, 2010. ISSN 1819-4966.

Schewe:1999:GDC

[ST99] Klaus-Dieter Schewe and Bernhard Thalheim. A generalization of Dijkstra’s calculus to typed program specifications. In Fundamentals of computation theory (Iaşi, 1999), volume 1684 of Lecture Notes in Comput. Sci., pages 463–474. Springer, Berlin, 1999.

Stibitz:1980:EC

[Sti80] George R. Stibitz. Early computers. In Metropolis et al. [MHR80], pages 479–483. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5 .I63 1976.

URL http://www.sciencedirect.com/science/article/

pii/B9780124916500500344. Original versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Streitferdt:1972:BPD

[Str72] Lothar Streitferdt. Eine Bemerkung zur Programmierung des Dijkstra-Verfahrens. (German) [A remark on the programming of Dijkstra’s algorithm]. Z. Operations Res. Ser. A-B, 16:B253–B256, 1972. ISSN 0340-9422.

Suekane:1980:EHC

[Sue80] Ryota Suekane. Early history of computing in Japan. In Metropolis et al. [MHR80], pages 575–578. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5 .I63 1976.

URL http://www.sciencedirect.com/science/article/

pii/B9780124916500500423. Original versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Svoboda:1980:MLE

[Svo80] Antonin Svoboda. From mechanical linkages to electronic computers: Recollections from Czechoslovakia. In Metropolis et al. [MHR80], pages 579–586. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5 .I63 1976.

URL http://www.sciencedirect.com/science/article/

pii/B9780124916500500435. Original versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Sudhakar:2004:SRD

[SVSR04] T. D. Sudhakar, N. S. Vadivoo, S. M. R. Slochanal, and S. Ravichandran. Supply restoration in distribution networks using Dijkstra’s algorithm. In 2004 International Conference on Power System Technology, 2004. PowerCon 2004, volume 1, pages 640–645 Vol.1. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2004. ISBN 0-7803-8610-8. URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?arnumber=1460073.

Swierstra:2011:SVP

[Swi11] Wouter Swierstra. Sorted: Verifying the problem of the Dutch National Flag in Agda. Journal of Functional Programming, 21(6):573–583, November 2011. CO-DEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (elec-

tronic). URL https://www.cambridge.org/core/product/
B94802514DC2786CCF323154AEB91953.

Swamy:2013:VHO

[SWS^{+}13] Nikhil Swamy, Joel Weinberger, Cole Schlesinger, Juan Chen, and Benjamin Livshits. Verifying higher-order programs with the Dijkstra monad. ACM SIGPLAN Notices, 48(6):387–398, June 2013. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schulz:1999:DAL

[SWW99] Frank Schulz, Dorothea Wagner, and Karsten Weihe. Dijkstra’s algorithm on-line: An empirical case study from public railroad transport. Lecture Notes in Computer Science, 1668:

110–??, 1999. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer-ny.

com/link/service/series/0558/bibs/1668/16680110.htm;
 http://link.springer-ny.com/link/service/series/
0558/papers/1668/16680110.pdf.

Schulz:2000:DAL

[SWW00] Frank Schulz, Dorothea Wagner, and Karsten Weihe. Dijkstra’s algorithm on-line: an empirical case study from public railroad transport. ACM Journal of Experimental Algorithmics, 5(Special Issue 2):12:1–12:??, ???? 2000. CODEN ???? ISSN 1084-6654. 3rd Workshop on Algorithm Engineering (London, 1999).

Sen:2010:SCA

[SWX10] Ma Sen, Chang Wei, and Cui Xiaoxian. Service-correlation aware service selection for composite service based on the improved Dijkstra algorithm. In 2010 International Conference on E-Business and E-Government (ICEE), pages 2330–2334. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2010. ISBN 0-7695-3997-

1. URL http://ieeexplore.ieee.org/stamp/stamp.jsp?
arnumber=5592401.

Shu-xi:2011:IDS

[SxXq11] Wang Shu-xi and Zhao Xing-qiu. The improved Dijkstra’s shortest path algorithm. In 2011 Seventh International Conference on Natural Computation (ICNC), volume 4, pages 2313–2316. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2011. ISBN 1-4244-9950-X. ISSN 2157-9555. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?arnumber=6022589.

Sghaier:2010:DDA

[SZHT10] M. Sghaier, H. Zgaya, S. Hammadi, and C. Tahon. A distributed Dijkstra’s algorithm for the implementation of a real time carpooling service with an optimized aspect on siblings. In 2010 13th International IEEE Conference on Intelligent Transportation Systems (ITSC), pages 795–800. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2010. ISBN 1-4244-7657-

7. ISSN 2153-0009. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=5625288.

Tatarchenko:2023:ACC

[Tat23] Ksenia Tatarchenko. Algorithm’s cradle: Commemorating al-Khwarizmi in the Soviet history of mathematics and Cold War computer science. Osiris (Series 2), 38:286–304, ???? 2023. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287 (electronic).

Terwilliger:1993:TTS

[Ter93] Robert B. Terwilliger. Toward tools to support the Gries/ Dijkstra design process. ACM SIGSOFT Software Engineering Notes, 18(2):50–59, April 1993. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Tomayko:1991:AFC

[TLR^{+}91] James E. Tomayko, J. A. N. Lee, Martin Rem, David Gries, Andrzej Blikle, and Robert L. Glass. Anecdotes: Frontiers of Computing: A Tribute to Edsger W. Dijkstra on the occasion of his 60th birthday University of Texas (Austin) 10–11 May 1990. Annals of the History of Computing, 13(1):91–100, January/March 1991. CODEN AHCOE5.

ISSN 0164-1239. URL http://dlib.computer.org/an/
books/an1991/pdf/a1091.pdf;   http://www.computer.
org/annals/an1991/a1091abs.htm.

Tomash:1980:SEE

[Tom80] Erwin Tomash. The start of an ERA: Engineering Research Associates, Inc., 1946–1955. In Metropolis et al. [MHR80], pages 485–495. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book).

LCCN QA75.5 .I63 1976. URL http://www.sciencedirect.
com/science/article/pii/B9780124916500500356. Origi-

nal versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Traff:1995:ECT

[Trä95] Jesper Larsson Träff. An experimental comparison of two distributed single-source shortest path algorithms. Parallel Computing, 21(9):1505–1532, September 12, 1995. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.elsevier.com/cgi-bin/cas/
tree/store/parco/cas_sub/browse/browse.cgi?year=1995&
volume=21&issue=9&aid=1004.

Tropp:1980:SCH

[Tro80] Henry S. Tropp. The Smithsonian Computer History Project and some personal recollections. In Metropolis et al. [MHR80], pages 115–122. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5 .I63 1976.

URL http://www.sciencedirect.com/science/article/

pii/B9780124916500500162. Original versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Tommiska:2001:DSP

[TS01] Matti Tommiska and Jorma Skyttä. Dijkstra’s shortest path routing algorithm in reconfigurable hardware. Lecture Notes in Computer Science, 2147:653–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer-ny.com/link/
service/series/0558/bibs/2147/21470653.htm;   http:
//link.springer-ny.com/link/service/series/0558/papers/
2147/21470653.pdf.

Tseitlin:1995:CFC

[Tse95] G. E. Tseĭtlin. A criterion for functional completeness in Dijkstra’s algebra. Kibernet. Sistem. Anal., 5:28–40, 186, 1995. ISSN 0023-1274. English translation in Cybernetics and Systems Analysis, 31(5) 659–668 (September 1995).

Tseitlin:1997:CLS

[Tse97] G. E. Tseĭtlin. Construction of the lattice of subalgebras of the Dijkstra algebra. Kibernet. Sistem. Anal., 1:27–45, 188, 1997. ISSN 0023-1274. English translation in Cybernetics and Systems Analysis, 33(1) 21–34 (January 1997).

Turner:1990:DMP

[Tur90a] D. A. Turner. Duality and De Morgan principles for lists. In Feijen et al. [FvGGM90], chapter 47, pages 390–398. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN QA76 .B326

1990. URL http://www.zentralblatt-math.org/zmath/

en/search/?an=0718.68004. Contains important treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

Turski:1990:QTS

[Tur90b] W. M. Turski. The quest for timeless specifications leads to non-stepping automata. In Feijen et al. [FvGGM90], chapter 48, pages 399–409. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X.

LCCN QA76 .B326 1990. URL http://www.zentralblatt-
math.org/zmath/en/search/?an=0718.68004. Contains

important treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

Tang:2010:TFC

[TZDM10] Yan Tang, Gang Zhao, P. De Baer, and R. Meersman. Towards freely and correctly adjusted Dijkstra’s algorithm with semantic decision tables for ontology based data matching. In Venkatesh Mahadevan and Jianhong Zhou, editors, Proceedings, 2010 the 2nd International Conference on Computer and Automation Engineering (ICCAE 2010): February 26– 28, 2010, Singapore, volume 1, pages 345–348. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2010. ISBN 1-4244-5585-5, 1-4244-5586-3. LCCN TK7885.A1. URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?arnumber=5451937.

Uchida:2014:ORT

[Uch14] K. Uchida. Optical ray tracing based on Dijkstra algorithm in inhomogeneous medium. In 2014 Ninth International Conference on Broadband and Wireless Computing, Communication and Applications (BWCCA), pages 371–376. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2014. ISBN 1-4799-4174-

3. URL http://ieeexplore.ieee.org/stamp/stamp.jsp?
arnumber=7016099.

Udding:1990:MLP

[Udd90] Jan Tijmen Udding. The maximum length of a palindrome in a sequence. In Feijen et al. [FvGGM90], chapter 49, pages 410–416. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN

QA76 .B326 1990. URL http://www.zentralblatt-math.
org/zmath/en/search/?an=0718.68004. Contains impor-

tant treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

Ulam:1980:NIM

[Ula80] S. M. Ulam. Von Neumann: The interaction of mathematics and computing. In Metropolis et al. [MHR80], pages 93–99. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5

.I63 1976. URL http://www.sciencedirect.com/science/
article/pii/B9780124916500500149. Original versions of

these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Uchida:2014:TSB

[UNTH14] K. Uchida, S. Nogami, M. Takematsu, and J. Honda. Tsunami simulation based on Dijkstra algorithm. In 2014 17th International Conference on Network-Based Information Systems (NBiS), pages 114–119. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2014. ISBN 1-4799-4226-X. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?arnumber=7023942.

Ustun:2011:IDA

[UOZ11] T. S. Ustun, C. Ozansoy, and A. Zayegh. Implementation of Dijkstra’s algorithm in a dynamic microgrid for relay hierarchy detection. In 2011 IEEE International Conference on Smart Grid Communications (SmartGrid-Comm 2011): Brussels, Belgium, 17–20 October 2011, pages 481–486. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2011. ISBN 1-4577-1702-6, 1-4577-1704-2 (paperback). LCCN

TK5105.5. URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=6102370.

Urbiha:2001:SPF

[Urb01] Igor Urbiha. Some properties of a function studied by de Rham, Carlitz and Dijkstra and its relation to the (Eisenstein–)Stern’s diatomic sequence. Math. Commun., 6 (2):181–198, 2001. ISSN 1331-0623.

VanGhent:1978:LRS

[Van78] Roger Van Ghent. Letter on real software, regarding the commentary by Dijkstra and the reply by DeMillo, Lipton and Perlis. ACM SIGSOFT Software Engineering Notes, 3(3):20– 21, July 1978. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). See [Dij78u, DLP78, Jef78].

vandenHovedErtsenryck:2019:NIO

[van19] Gauthier Marie Claude Jean Thérèse Gérald van den Hove d’Ertsenryck. New Insights from Old Programs: The Structure of The First ALGOL 60 System. Ph.D. thesis, Centrum Wiskunde & Informatica (CWI), Informatics Institute, Faculty of Science, Universit of Amsterdam, Amsterdam, The Netherlands, February 2019. ???? + 302 pp. URL

https://hdl.handle.net/11245.1/8ecd6e9b-767d-44ab-
84a7-8eed35822da5;   https://ir.cwi.nl/pub/28427/.

Preface by Donald E. Knuth.

vandenHove:2017:DHC

[vdH17] Gauthier van den Hove. Dissolving a half century old problem about the implementation of procedures. Science of Computer Programming, 150(??):75–86, December 15, 2017. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0167642317301582. See [Dij60e].

vanDriel:2001:NAC

[vDO01] Kees van Driel and Paul Oosterveld. Nonoptimal alignment: a comment on “Measuring the agreement between sequences” by Dijkstra and Taris. Sociological Methods & Research, 29 (4):524–531, 2001. ISSN 0049-1241 (print), 1552-8294 (electronic).

vanderWoude:1990:CRE

[vdW90] Jaap van der Woude. Calculations with relations, an example. In Feijen et al. [FvGGM90], chapter 52, pages 435–441. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN QA76 .B326

1990. URL http://www.zentralblatt-math.org/zmath/

en/search/?an=0718.68004. Contains important treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

vanGasteren:1990:ERC

[vG90] A. J. M. van Gasteren. Experimenting with a refinement calculus. In Feijen et al. [FvGGM90], chapter 14, pages 127–134. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN QA76 .B326

1990. URL http://www.zentralblatt-math.org/zmath/

en/search/?an=0718.68004. Contains important treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

vanGasteren:1986:N

[vGD86] A. J. M. van Gasteren and E. W. Dijkstra. On notation. Technical Report AvG65a/EWD950a, Department of Mathematics and Computing Science, Eindhoven University of Technology, January 1986.

vanGasteren:1992:WFT

[vGD92] A. J. M. van Gasteren and Edsger W. Dijkstra. Wellfoundedness and the transitive closure. In Programming and mathematical method (Marktoberdorf, 1990), volume 88 of NATO Adv. Sci. Inst. Ser. F Comput. Systems Sci., pages 11–15. Springer, Berlin, 1992.

vanGasteren:20xx:RCF

[vGDxx] A. van Gasteren and Edsger W. Dijkstra. The regularity calculus: a first trial. Circulated privately., 20xx. URL http:

//www.cs.utexas.edu/users/EWD/ewd08xx/EWD882.PDF.

vanLint:1990:CCL

[vL90] J. H. van Lint. Coding for channels with localized errors. In Feijen et al. [FvGGM90], chapter 31, pages 274–279. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN QA76 .B326

1990. URL http://www.zentralblatt-math.org/zmath/

en/search/?an=0718.68004. Contains important treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

vanLint:2004:EWD

[vL04] J. H. van Lint. Edsger Wybe Dijkstra. In Levensberichten en Herdenkingen 2004 (Dutch) [Biographies and Commemorations 2004, Royal Dutch Academy of Sciences, Amsterdam], pages 32–36. Koninklijke Nederlandse Akademie van Wetenschappen, Amsterdam, The Netherlands, 2004.

Verscheure:2010:DAA

[VPM^{+}10] L. Verscheure, L. Peyrodie, N. Makni, N. Betrouni, S. Maouche, and M. Vermandel. Dijkstra’s algorithm applied to 3D skeletonization of the brain vascular tree: Evaluation and application to symbolic. In 2010 Annual International Conference of the IEEE Engineering in Medicine and Biology Society (EMBC), pages 3081–3084. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2010. ISBN 1-4244-4123-4. ISSN 1557-

170X. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=5626112.

vanTilborg:1990:APV

[vT90] Henk C. A. van Tilborg. An assignment problem for the vertices of a cycle. In Feijen et al. [FvGGM90], chapter 46, pages 386–389. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN

QA76 .B326 1990. URL http://www.zentralblatt-math.
org/zmath/en/search/?an=0718.68004. Contains impor-

tant treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

Vene:2005:AFP

[VU05] Varmo Vene and Tarmo Uustalu, editors. Advanced functional programming: 5th international school, AFP 2004, Tartu, Estonia, August 14–21, 2004, revised lectures, volume 3622 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. CODEN LNCSD9. ISBN 3-540-28540-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

????  URL http://www.springerlink.com/content/
978-3-540-28540-3;   http://www.springerlink.com/
openurl.asp?genre=issue&issn=0302-9743&volume=3622.

Vantilborgh:1972:EDS

[Vv72] H. Vantilborgh and A. van Lamsweerde. On an extension of Dijkstra’s semaphore primitives. Information Processing Letters, 1(5):181–186, October ??, 1972. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

vanWijngaarden:1956:CUN

[vWBD56] A. van Wijngaarden, J. Berghuis, and Edsger W. Dijkstra. Capita uit de numerieke wiskunde. (Dutch) [Capitalisation of numerical mathematics]. Technical Report 8, Mathematisch Centrum, Amsterdam, 1956. URL http://www.cs.utexas.

edu/users/EWD/MCReps/CR1956-008.PDF.

vanWijngaarden:1969:RAL

[vWMPK69] A. van Wijngaarden, B. J. Mailloux, J. E. L. Peck, and C. H. A. Koster. Report on the algorithmic language ALGOL 68. Numerische Mathematik, 14(2):79–218, December 1969. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http://www.fh-jena.

de/~kleine/history/languages/Algol68-Report.pdf;
http://www.fh-jena.de/~kleine/history/languages/Algol68-
RevisedReport.pdf.

Wallen:1990:FFE

[Wal90] Lincoln A. Wallen. On form, formalism and equivalence. In Feijen et al. [FvGGM90], chapter 50, pages 417–426. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN QA76 .B326

1990. URL http://www.zentralblatt-math.org/zmath/

en/search/?an=0718.68004. Contains important treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

Weiss:1988:BOP

[Wei88] Eric A. Weiss. Biographies: Oh, pioneers! Annals of the History of Computing, 10(4):348–361, October/December 1988. CODEN AHCOE5. ISSN 0164-1239. URL http://

dlib.computer.org/an/books/an1988/pdf/a4348.pdf;
http://www.computer.org/annals/an1988/a4348abs.htm.

Wells:1980:REA

[Wel80] Mark B. Wells. Reflections on the evolution of algorithmic language. In Metropolis et al. [MHR80], pages 275–287. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5

.I63 1976. URL http://www.sciencedirect.com/science/
article/pii/B9780124916500500204. Original versions of

these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Wegbreit:2001:ADT

[WG01] B. Wegbreit and John Guttag. Abstract data types and the development of data structures. In Broy and Denert [BD01b], pages 389–418. ISBN 3-540-42290-0 (print), 3-642-48354-2 (e-book). LCCN QA76.758. URL http://link.springer.

com/chapter/10.1007/978-3-642-48354-7_16/.

Wilkes:1980:EPD

[Wil80a] M. V. Wilkes. Early programming developments in Cambridge. In Metropolis et al. [MHR80], pages 497–501.

ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5

.I63 1976. URL http://www.sciencedirect.com/science/
article/pii/B9780124916500500368. Original versions of

these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Wilkinson:1980:TWN

[Wil80b] J. H. Wilkinson. Turing’s work at the National Physical Laboratory and the construction of Pilot ACE, DEUCE, and ACE. In Metropolis et al. [MHR80], pages 101–114. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5

.I63 1976. URL http://www.sciencedirect.com/science/
article/pii/B9780124916500500150. Original versions of

these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Winkowski:1978:MFC

[Win78] Józef Winkowski, editor. Mathematical foundations of computer science, 1978: proceedings, 7th Symposium, Zakopane, Poland, September 4–8, 1978, volume 64 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. CODEN LNCSD9. ISBN 0-387-08917-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6

.S9194 1978. URL http://link.springer-ny.com/link/
service/series/0558/tocs/t0064.htm;   http://www.
springerlink.com/content/978-0-387-08917-1;   http:
//www.springerlink.com/openurl.asp?genre=issue&issn=
0302-9743&volume=64.

Wirth:1990:DLC

[Wir90] N. Wirth. Drawing lines, circles, and ellipses in a raster. In Feijen et al. [FvGGM90], chapter 51, pages 427–434. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN QA76 .B326

1990. URL http://www.zentralblatt-math.org/zmath/

en/search/?an=0718.68004. Contains important treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

Wirth:2001:PDS

[Wir01a] Niklaus Wirth. Program development by stepwise refinement. In Broy and Denert [BD01b], pages 545–569.

ISBN 3-540-42290-0 (print), 3-642-48354-2 (e-book). LCCN

QA76.758. URL http://link.springer.com/chapter/10.
1007/978-3-642-48354-7_23/.

Wirth:2001:PLP

[Wir01b] Prof. Dr. Niklaus Wirth. The programming language Pascal. In Broy and Denert [BD01b], pages 515–543. ISBN 3-540-42290-0 (print), 3-642-48354-2 (e-book). LCCN

QA76.758. URL http://link.springer.com/chapter/10.
1007/978-3-642-48354-7_22/.

Wengui:2009:AOD

[WLXX09] Su Wengui, Li Li, Zhou Xiaorong, and Liao Xiaoping. Application of optimized Dijkstra algorithm in storage/retrieval routes scheduling of a stacker crane. In International Conference on Information Engineering and Computer Science, 2009. ICIECS 2009, pages 1–4. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2009. ISBN 1-4244-4994-4. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?arnumber=5364990.

Wei:2014:ROR

[WM14] Mingjun Wei and Yu Meng. Research on the optimal route choice based on improved Dijkstra. In 2014 IEEE Workshop on Advanced Research and Technology in Industry Applications (WARTIA), pages 303–306. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD

20910, USA, 2014. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=6976257.

Walker:1979:AAM

[WT79] A. N. Walker and The Editor. Algorithms: On “Algorithm 1: Mintree”. The Computer Journal, 22(3):288, August 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_03/tiff/288.tif. See [Obr78].

Woodger:1969:MRA

[WTD^{+}69] Mike Woodger, Władisław Marek Turski, Edsger W. Dijkstra, C. A. R. Hoare, Brian Randell, Jan V. Garwick, Fraser G. Duncan, and Gerhard Seegmueller. Minority report on Algol 68. The draft and the final text. ALGOL Bulletin (Amsterdam: Mathematisch Centrum), ??(??):??, January 17, 1969.

URL http://ershov.iis.nsk.su/en/node/805785.

Wielgus:1990:AFM

[WWSD90] Martin Wielgus, Marvin L. Watkins, Edward Sacharuk, and Srinivasarao Damerla. ACM Forum: More debatable. Communications of the Association for Computing Machinery, 33 (4):396–397, April 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See [Dij89-32, BSE^{+}90].

Wagner:2005:GCE

[WWZ05] Dorothea Wagner, Thomas Willhalm, and Christos Zaroliagis. Geometric containers for efficient shortest-path computation. ACM Journal of Experimental Algorithmics, 10: 1.3:1–1.3:??, ???? 2005. CODEN ???? ISSN 1084-6654.

Wang:2011:ADA

[WYY11] Huijuan Wang, Yuan Yu, and Q. Yuan. Application of Dijkstra algorithm in robot path-planning. In 2011 Second International Conference on Mechanic Automation and Control Engineering (MACE), pages 1067–1069. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2011. ISBN 1-4244-9436-

2. URL http://ieeexplore.ieee.org/stamp/stamp.jsp?
arnumber=5987118.

Xu:2007:IDS

[XLH^{+}07] M. H. Xu, Y. Q. Liu, Q. L. Huang, Y. X. Zhang, and G. F. Luan. An improved Dijkstra’s shortest path algorithm for sparse network. Applied Mathematics and Computation, 185(1):247–254, February 1, 2007. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0096300306008563.

Xu:2012:ADA

[XWZH12] Yuanzhe Xu, Zixun Wang, Qingqing Zheng, and Zhiyong Han. The application of Dijkstra’s algorithm in the intelligent fire evacuation system. In 2012 4th International Conference on Intelligent Human-Machine Systems and Cybernetics (IHMSC), volume 1, pages 3–6. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2012. ISBN 1-4673-1902-3. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?arnumber=6305611.

Xiao:2010:ISP

[xXL10] Ji xian Xiao and Fang-Ling Lu. An improvement of the shortest path algorithm based on Dijkstra algorithm. In Venkatesh Mahadevan and Jianhong Zhou, editors, Proceedings, 2010 the 2nd International Conference on Computer and Automation Engineering (ICCAE 2010): February 26– 28, 2010, Singapore, volume 2, pages 383–385. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2010. ISBN 1-4244-5585-5, 1-4244-5586-3. LCCN TK7885.A1. URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?arnumber=5451564.

Xie:2012:IDA

[XZY^{+}12] Dexiang Xie, Haibo Zhu, Lin Yan, Si Yuan, and Junqiao Zhang. An improved Dijkstra algorithm in GIS application. In 2012 World Automation Congress (WAC), pages 167–169. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2012. ISBN 1-4673-4497-

4. ISSN 2154-4824. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=6321059.

Yap:1983:HAS

[Yap83] Chee K. Yap. A hybrid algorithm for the shortest path between two nodes in the presence of few negative arcs. Information Processing Letters, 16(4):181–182, May 13, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yan:2010:DAB

[YJ10] Zhang Yan and Zhang Jun. Dijkstra’s algorithm based robust optimization to airline network planning. In 2010 International Conference on Mechanic Automation and Control Engineering (MACE), pages 2783–2786. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2010. ISBN 1-4244-7737-

9. URL http://ieeexplore.ieee.org/stamp/stamp.jsp?
arnumber=5536824.

Yang:2010:ERM

[YL10] Shu Yang and Chunhua Li. An enhanced routing method with Dijkstra algorithm and AHP analysis in GIS-based emergency plan. In 2010 18th International Conference on Geoinformatics, pages 1–6. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2010. ISBN

1-4244-7301-2. URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5567840.

Yershov:2011:SDA

[YL11] Dmitry S. Yershov and S. M. LaValle. Simplicial Dijkstra and A* algorithms for optimal feedback planning. In 2011 IEEE/RSJ International Conference on Intelligent Robots and Systems (IROS), pages 3862–3867. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2011. ISBN 1-61284-454-5. ISSN 2153-

0858. URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=6095032.

Li:2009:RSI

[yLsLtZ09] Xiang ying Li, Guo shun Li, and Sheng tian Zhang. Routing space Internet based on Dijkstra’s algorithm. In IEEE, editor, First Asian Himalayas International Conference on Internet, 2009: AH-ICI 2009 ; Kathmandu, Nepal, 3 - 5 Nov. 2009, pages 1–4. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2009. ISBN 1-4244-4569-8 (paperback), 1-4244-4570-

1. URL http://ieeexplore.ieee.org/stamp/stamp.jsp?
arnumber=5340354.

Li:2010:RSI

[yLsLtZ10] Xiang ying Li, Guo shun Li, and Sheng tian Zhang. Routing space Internet based on Dijkstra’s algorithm. In Zhitang Li and Zhengbing Hu, editors, Second International Conference on Networks Security, Wireless Communications and Trusted Computing (NSWCTC), 2010: 24– 25 April 2010, Wuhan, Hubei, China; proceedings, volume 2, pages 118–121. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2010. ISBN 0-7695-4011-2, 1-4244-6597-4, 1-4244-6598-2.

LCCN TK5105.5. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=5480671.

Zemanek:1980:CEP

[Zem80] H. Zemanek. Central European prehistory of computing. In Metropolis et al. [MHR80], pages 587–609. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5 .I63 1976.

URL http://www.sciencedirect.com/science/article/

pii/B9780124916500500447. Original versions of these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Zemanek:1990:TPP

[Zem90] Heinz Zemanek. Two proofs for Pythagoras. In Feijen et al. [FvGGM90], chapter 53, pages 442–447. ISBN 0-387-97299-4, 3-540-97299-4, 1-4612-8792-8 (print), 1-4612-4476-5 (online). ISSN 0172-603X. LCCN QA76 .B326

1990. URL http://www.zentralblatt-math.org/zmath/

en/search/?an=0718.68004. Contains important treatment of accurate binary-to-decimal conversion [Gri90, Knu90].

Fan:2010:NRI

[zFmLcWcJ10] Yue zhen Fan, Dun min Lu, Qing chun Wang, and Fa chao Jiang. Notice of retraction: An improved Dijkstra algorithm used on vehicle optimization route planning. In 2010 2nd International Conference on Computer Engineering and Technology (ICCET), volume 3, pages V3–693–V3–696. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2010. ISBN 1-4244-6347-

5. URL http://ieeexplore.ieee.org/stamp/stamp.jsp?
arnumber=5485737.

Zhang:2012:IDA

[ZJM12] Wei Zhang, Chong Jiang, and Yunxiang Ma. An improved Dijkstra algorithm based on pairing heap. In 2012 Fifth International Symposium on Computational Intelligence and Design (ISCID), volume 2, pages 419–422. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2012. ISBN 1-4673-2646-1. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?arnumber=6406028.

Zhu:2013:OPA

[ZLY13] Yong Zhu, Xiaohuan Liu, and Xiaohong Yu. An optimal path algorithm of high security based on Dijkstra algorithm.

In 2013 International Conference on Sensor Network Security Technology and Privacy Communication System (SNS & PCS), pages 93–96. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2013.

ISBN 1-4673-6452-5. URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?arnumber=6553842.

Zhou:2012:NBD

[ZTZY12] Huifeng Zhou, Ling Tian, Xiangjun Zeng, and Jun Yuan. Network-based Dijkstra algorithm for fault location with traveling wave. In IEEE, editor, China International Conference on Electricity Distribution: (CICED): 10–14 September 2012, Shanghai, China, pages 1–5. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2012. ISBN 1-4673-6065-1, 1-4673-6063-5. ISSN 2161-7481 (print), 2161-749X (electronic). LCCN TK3001. URL

http://ieeexplore.ieee.org/servlet/opac?punumber=
6504849;   http://ieeexplore.ieee.org/stamp/stamp.
jsp?arnumber=6508488.

Zuse:1980:SRH

[Zus80] Konrad Zuse. Some remarks on the history of computing in Germany. In Metropolis et al. [MHR80], pages 611–627. ISBN 0-12-491650-3, 1-4832-9668-7 (e-book). LCCN QA75.5

.I63 1976. URL http://www.sciencedirect.com/science/
article/pii/B9780124916500500459. Original versions of

these papers were presented at the International Research Conference on the History of Computing, held at the Los Alamos Scientific Laboratory, 10–15 June 1976.

Zhang:2010:ODA

[ZWKT10] Dong Zhang, ZuKuan Wei, Jae-Hong Kim, and ShuGuang Tang. An optimized Dijkstra algorithm for Embedded-GIS. In Jinkuan Wang and Bin Wang, editors, ICCDA 2010: 2010 International Conference on Computer Design and Applications: Qinhuangdao, Hebei, China: June 25-27, 2010, volume 1, pages V1–147–V1–150. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2010. ISBN 1-4244-7163-X, 1-4244-7164-8 (e-book). LCCN

TK7885.A1. URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5541160.

Zou:2009:DOC

[ZXZ^{+}09] Qinghua Zou, Naixue Xiong, Shengzhong Zou, Li Chen, and Chuanbing Wang. Dijkstra optimization and CAD based on VC lowest cost process. In 2009 1st International Conference on Information Science and Engineering (ICISE), pages 3760–3763. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2009. ISBN

1-4244-4909-X. URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?arnumber=5455164.